A Bibliography of Publications about the R, S, and S-Plus Statistics Programming Languages

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

22 May 2024 Version 4.150

Title word cross-reference

$ [Sha23b]. (Q, r) [OOSM18]. $109.95 [Has18]. $135.00 [Lip21a]. $139.95 [Sta21]. $150.00 [Cho22a]. 2 [EN11, Grö14b]. 2 × 2 [ILS11]. 3 [GGK10, LPLPD14]. $34.95 [Ree19]. $38.95 [Joh20]. 4 [Asq14, HWY18]. $44.95 [Hou07]. $47.96 [Peb21]. $49.95 [Ree19]. $50.99 [Sha21c]. $54.95 [Ben21]. $59.95 [Cho22a]. $63.96 [Lip22]. $69.95 [Put06, Sta05, Vie11]. $74.21 [Sha23a]. $74.94 [Cox05]. $79.95 [Ben21, Lip20b, Pet02, Sab19]. $85 [Ano03b]. $88.00 [Asz23]. $89.95 [Kel10, Lum08]. $94.94 [Mal21]. $95.00 [Aji17]. $99.95 [Lee21, Len20, Lip20a, Lip21c, Liu18, Ree19, Sha23b, Lip21b]. = [Sar06]. ^{2} [Sar06]. α [LPLPD14, dVSWAL17]. · [Sar06]. h [HBA19]. K [KSBZ16, PG15, Lei10]. L [Asq14]. μ [CBHLG21]. n [HMR^{+}13]. p [MF15]. R [QZLP21]. s [HJM08]. t [AHvD09, HJM08, ML13]. × [Sta08]. z [HJM08].

-function [PG15]. -Gram [HMR^{+}13]. -Level [Grö14b]. -Means [KSBZ16]. -Modes [Lei10]. -moments [Asq14]. -Parameter [HWY18, Asq14]. -Shape [LPLPD14]. -squared [QZLP21]. -Stable [dVSWAL17].

. [Lip22].

//www.amherst.edu/ [Grö15a].

//www.crcpress.com/9781498709576 [Grö16]. //www.crcpress.com/9781498711548 [Gle16a]. //www.crcpress.com/9781498712361 [Kha16]. //www.imperial.ac.uk/bio/research/crawley/statistics/ [Hel15]. //www.r [Iac15]. //www.r-datacollection.com/ [Iac15]. //www.springer.com/9783319140926 [Zei16]. //www.springer.com/9783319234458 [How16b]. //www.springer.com/9783319238821 [Mat16a]. //www.springer.com/book/9781493917013 [Mat15]. //www.springer.com/book/9783319144351 [Rus15]. //www.springer.com/book/9788132223399 [Hof15]. /2021 [Lee21].

0 [Abr97, Cox05, Dal98, Har03, Orm17, Pet02, Plu01, Til96]. 0-387-40081-8 [Cox05]. 0-387-94350-1 [Abr97]. 0-387-94727-2 [Dal98]. 0-387-98503-4. [Plu01]. 0-387-98814-9 [Har03]. 0-412-56310-X [Til96]. 0-8493-7168-6 [Pet02].

1 [Abr97, Hou07, Kal24a, Lip20b, Lum08, Mai21, Pod18, Put06, SS18a, Sta05]. 1-58488-408-8 [Sta05]. 1-58488-425-8 [Put06]. 1-58488-450-9 [Hou07]. 1-58488-619-6 [Lum08]. 1-58488-962-4 [Kel10]. 1013 [AE21]. 1017 [SM21]. 125.00 [Laz24b]. 13ARIMA [SE18]. 1969 [MN69]. 1985 [All86]. 1st [Bia94, Kal24a, Kal24b, Piñ24, Sau10, Sau11, Sau12].

2 [Cho22a, Dal98, Laz24a, Num20, Sab19, SS18b]. 2.0 [GK16]. 2.0.0 [HL04]. 2.1.0 [HL05]. 2011-04-13 [R D11a, R D11c, R D11d, R D11e, R D11f, R D11g]. 2019 [For20, Pet19]. 20th [Ano95a]. 21 [MN03]. 224pp [Ree19]. 242 [IDE15]. 256pp [Ree19]. 25th [DO94]. 2608 [MN03]. 294pp [Laz24b]. 2nd [Ano09a, Ano10, Ano12a, Ano12b, Bos09, CH11b, Cox05, Dem17, DN17, Edd09a, Gal17, GR18, Grö15a, Hel15, Hel18, Hew05, Joh09, KK99, Kha17, Len20, Lor18, Lu18, Man03, Pfa12, Qia10a, Ree18, Rob22, SCD07, Sau21, Sch08].

3 [Len20, Rui17]. 33487 [Sha23a]. 368pp [AS24]. 390 [Wan22]. 3rd [Ano08, Lum08, McG20, San10a].

4 [Cap19, Sta97a, Sta97b, Sta97c]. 4.

[Plu01]. 4.0 [DCB^{+}17]. 4.5

[Ano98, Gen98, Mat98, Sta99]. 423pp [Ree19]. 426 [Esk21]. 460 [Fus22]. 4th [Sta21].

5 [Ano16, Dan18, Die21, GC23, Kel10, Pod21, Sta98a, Sta98b, Vie11]. ‘50 [HJ17]. 531 [Rao14].

6 [Har23, Has18, Lip20a, Lip21a, Lum08, Peb21, Pet02]. 68 [IDE15].

7 [Che23, Hac17, Liu18, Rob17, Sha21c]. 711 [Das21]. 77 [GH18, GHN19].

8 [Cox05, Joh20, Put06, Sta05]. 80.99/$108 [Liu23].

9 [Cur18, Har03, Hou07, Mal21, Sta21, Ree19, Hu21]. 95 [LS05]. 978 [Ano16, Cap19, Che23, Cho22a, Cur18, Die21, GC23, Hac17, Har23, Has18, Joh20, Kal24a, Kel10, Laz24a, Len20, Lip20a, Lip20b, Lip21a, Liu18, Mai21, Mal21, Num20, Orm17, Peb21, Pod18, Pod21, Rob17, Rui17, Sab19, Sha21c, Sta21, Vie11, Ree19, Hu21]. 978-0-12-820788-8 [Tol23]. 978-0-19-872906-8 [Ree19]. 978-0-19-878784-6 [Ree19].

978-0-367-13991-9 [Mal21]. 978-0-367-17387-6 [Liu23, Har23]. 978-0-367-22109-6 [Nun20]. 978-0-367-25538-1 [AS24, Lip21b]. 978-0-367-27798-7 [Sha23a]. 978-0-367-35798-6 [Lip20a]. 978-0-367-36547-9 [Laz24b]. 978-0-367-40982-1 [Lip20b]. 978-0-367-56995-2 [Cho22a]. 978-0-367-73432-9 [Hu21]. 978-0-367-81889-0 [Lip21c]. 978-0-429-05646-8 [Liu23]. 978-0-429-06666-5 [Lee21]. 978-1-003-10027-0 [Cho22a]. 978-1-003-27921-1 [Kal24a]. 978-1-032-04174-2 [Lip22, Piñ24]. 978-1-032-06536-6 [Asz23]. 978-1-032-15373-5 [Sha23b]. 978-1-032-24407-5 [GC23]. 978-1-032-25775-4 [Kal24b]. 978-1-032-25948-2 [Laz24a]. 978-1-032-37674-5 [Liu23]. 978-1-107-15075-1 [Pod18]. 978-1-118-38798-6 [Lip21a]. 978-1-118-83481-7 [Iac15]. 978-1-118-94109-6 [Hel15]. 978-1-138-19745-9 [Sta21]. 978-1-138-35933-8 [Joh20]. 978-1-138-59963-5 [Die21]. 978-1-138-62691-1 [Ben21]. 978-1-138-63197-7 [Hac17]. 978-1-138-71113-6 [Peb21]. 978-1-315-21266-1 [Ben21]. 978-1-4398-7885-9 [Cur18]. 978-1-4398-8538-3 [Lee21]. 978-1-4422-4992-9 [Aji17]. 978-1-4665-0225-3 [Rao14]. 978-1-4665-1585-7 [Sha21c]. 978-1-4665-5332-3 [Len20]. 978-1-4822-3736-8 [Grö15a]. 978-1-4822-4936-1 [Oli17]. 978-1-4822-5344-3 [Rui17]. 978-1-4822-5383-2 [Lu18]. 978-1-4939-1701-3 [Mat15].

978-1-4939-8853-2 [Num20]. 978-1-4987-0957-6 [Grö16]. 978-1-4987-1154-8 [Gle16a]. 978-1-4987-1236-1 [Kha16]. 978-1-4987-2096-0 [Orm17]. 978-1-4987-2448-7 [Liu18]. 978-1-4987-3422-6 [Has18]. 978-1-4987-7013-2 [Sab19]. 978-1-4987-7571-7 [Kle17, Rob17]. 978-1-4987-8185-5 [Pod21]. 978-1-4987-8905-9 [Ree19]. 978-1-58488-087-5 [Vie11]. 978-1-58488-962-5 [Kel10]. 978-1-59327-651-5 [Ano16]. 978-3-030-56484-1 [Mai21]. 978-3-030-73791-7 [Che23]. 978-3-319-14092-6 [Zei16]. 978-3-319-14435-1 [Rus15]. 978-3-319-23446-5 [How16b]. 978-3-319-23882-1 [Mat16a]. 978-3-319-72346-4 [Cap19]. 978-81-322-2339-9 [Hof15]. 9781493917013 [Mat15]. 9781498709576 [Grö16]. 9781498711548 [Gle16a]. 9781498712361 [Kha16]. 9783319140926 [Zei16]. 9783319144351 [Rus15]. 9783319234458 [How16b]. 9783319238821 [Mat16a]. 9788132223399 [Hof15]. 99.99 [Tu21]. 9th [HI97].

= [CCP^{+}11].

A. [Ano08, Ano09c, Ber10]. Absence [FM08]. Absolute [Bel19]. abstract [KMMV14]. Abstraction [BLM^{+}15]. Abundance [Fie12, FC11]. Academic [Tol23]. Accelerated [CKY14, HXY12]. Accelerating [ES14]. Acceptance [CC11, Cho22a, Kie08]. AcceptanceSampling [Kie08]. accessibility [Aji17, Fri16]. accessible [CPB^{+}23, FAM^{+}20]. Accessing [DC09]. accounting [Cad13]. accumulated [Alb19]. Accuracy [NAA17, AY22, AGM07].

Accurate [AGM07]. Achim [Mye09, Sco09]. ACM [The99]. Acoustic [OS95, HCSH15]. Acquiring [Grö18a]. across [GW18]. ACS [Gle16b]. action [Kab11, Edd12b]. Active [dLHM09, HCSH15]. actuar [DGP08]. Actuarial [Cha15b, Dur15, DGP08, Spe13, NB13]. acyclic [Ste24]. ada [CJM06]. adabag [AGG13]. Adams [AS24, Lip21b]. Adaptive [AHvD09, Cha08b, DHM11, MR12, MMB15, MJGM10, PT07, PT09, WLK08, ZS17, Ros07, SA15, SGPSzC22, Hu09, Kel10, O’B08, Rob09]. Add [SH17]. Additional [CFSR15, Den16]. Additive [AC07, BKL05, Chi07, GR18, Han06a, HC05, KV13, LC10, Mai06, Rob07, Sch11, Sta07b, SR07, UAK^{+}15, Woo06, PW18]. ade4 [BD18, Sha20, TD07]. ade4TkGUI [TD07]. adehabitat [Cal06]. Adelchi [Ros00]. adimpro [PT07]. Adjacent [dLHM09]. Adjustment [KK15, KMTS14, SE18]. Administration [R D11d]. AdMit [AHvD09]. Adrian [Ros00]. ads [PG15]. Advanced [Dem20, DJS^{+}18, Lan17a, Lan14b, Wic19, ACG^{+}16, Ano13, Lan14a, Gep21, Lip21a, Sau21]. Advances [Vin10, Str10]. Advection [SKS15]. Advection-Diffusion [SKS15]. Africans [Tyn16]. AFS [Ano99, GHG19]. After [Mar06]. aftgee [CKY14]. ag5Tools [BdSvE23]. Age [Jon07]. Age-structured [Jon07]. Agency [WRD23]. AgERA5 [BdSvE23]. Agglomerative [Mül13]. aggressive [SWHJ17]. agree [FBB20]. agricultural [dSBSvE23]. Agriculture [Pla12, Pla19, For20]. agrometeorological [BdSvE23]. Aid [NV11]. Air [Buc09, PD08, YCLL20]. Airbnb [BCG18]. Ajay [Mat15]. Akalin [Pod21]. Akers [Mor03]. al [Lud13]. Alain [Bai11]. Alan [Bat08b, Has18, Ng09]. Albert [Cho09, Fus22, Lip20b, Pol11a, Bau14]. Albrecht [AA12]. Aldo [Ros10b]. Alejandro [Wan22]. Alexandria [Ano99].

Algebra [Fie15, Kha16, Liu16b, ES14].

Algorithm

[AE21, CLL17, HC05, KSBZ16, ML13, PK08, dLHM09, LAJJ18, AE21, SM21]. Algorithms [Bas18, Cic15, Cur18, Das21, GFS14, How18, Iri19, JPH21, Lab12, Lip20a, MJR93, MYK07, MCM12, NV11, Scr13, YEL18, ZFZ10, Che22, FGZ14, MGHR16, SS18a, Val09, WML14]. Aligning [CS12]. Alignment [KSS^{+}07, DS20]. Alignments [Gio09]. Allaire [Joh20, Kim20b]. Allerhand [Sen14]. Alley [ZMS21]. Alley-oop [ZMS21]. allocation [RK20]. allowing [Kom09]. Alon [Aji17]. Alonso [Oli17]. alphahull [PLRC10]. altered [MAY23]. alternative [HBA19]. Altuna [Pod21]. Amazon [HT19]. AMCMC [Ros07]. amei [MJGM10]. American [IDE15, Gle16b]. amplitude [LBL^{+}21]. amsrpm [KSS^{+}07]. Ana [Bat08b, Ng09, Sco13b]. anacor [dLM09c]. Analyses [Ano06a, Eve94, Eve02, EH06, HM18, Hil06, Kim95, Ko95a, Lun07b, Man03, Pen08, Pfa12, Rob23, TMCE22, Vie10, Zie02a, KKL^{+}15, MÇSD14, SCK^{+}23, Til96]. Analysing [CPA20, KWE^{+}17, MAY23]. Analysis [Agr16, AL16, Ano06b, Ano09d, Ano12b, Arm19, BL14a, BBG16, BP12, Bea17, Beb13, BC76, BC77, BC81, BC84a, BC84b, BW88, BCW88, BCW18, BBG12, BK17, BL15, BPGR08, Blo14, BPGC14, BDdM11, BPB09, BD18, Bow10, Bur10, Bur12, Cap22, Car10, CHT98, CFHBK11, Cha86, Cha08a, Cha22, CKY14, Cho22b, CFSR15, CS07, CO16, CsC08, Das21, Dav95, Den98, Dia06, Dia05, Dis23, Duo07, DHF15, Eme08, EE07, EL09, Eva11, Fal12, FBdlF12, Fel12, Fer11, FDGD16, Fle11, FC12, FM08, FKP17, Fri11, FM16, FGEMI12, GP12, GK16, GKD14, GZ11, GK14, GCA12, GDMB08, GT10, GVM16, Grö15a, GV12, HD18, HSL11, Han05, Han98, Har01, Har23, Has18, Hec07, HH15, HP09, Hla16, Hoe09, HK15, How16b, HSG12, IG96, IRCA21]. Analysis [Iri19, JM15, JD15, Joc14, Joh09, Jon07, Kal24b, Kav15, KK15, Kee18, Kha18, KE14, KH10, Kos15, Kos16, Kus03, Lal17, LRGTA12, Laz24b, LJH08, Lei10, Lep14, LFF17, LZHC17, LL11, Lip20a, LC10, Liu17, Liu23, LRN18b, LT19, dUJ13, Lor18, Lun07a, LCSC14, MF14, Mai08, Mai11, Mai14a, MdL10, MC97, Sta93e, Mat16a, MF15, MCAP19, MYK07, Meu13, NGBK12, Nor15, Num13, Nun20, Obe14, Oli10b, OK14, PB15, PH16, PPGD15, Par06a, Par12, PP18, Pfa06, Pfa08a, Pla12, Pla19, Pol11c, Pol13, Pri05b, Qia10a, R D11a, RHG09, RG96, Ree18, RJH14, RS05, Riz06, Rob19, RFKM12, RMG12, Ros00, RGD12, RD92, San10a, SMM^{+}15, SCK95, Sco10b, Sha16, She11, SC07, Sho18, SS19b, SS92, SR16, SP10, SVCB18, Sta93f, Ste11, Sun15]. Analysis [Sun16, TV11, TMN16, TM05, TD07, TYH^{+}14, Tod22, Tsa10, ÜS10, Unw11, Unw12a, Unw15, Utl05, VM09a, VR99, Ver18, WLH^{+}18, WHK21, WTB^{+}15, WS11, WM14, Wic09, WG10, Wil14a, WGE17, Yu12, YL17, ZQS16, Zie05, dL05b, dLM09c, dWFP11, vdA07, vdA12, Agr13, ACG^{+}16, Ano03b, ATCA20, BFA14, BC85a, Bee13, Beh04, Beh05, Beh12, Bos09, Bos12, BA97, Bro12, BG96, Cal06, Car04, Car14, CCKT83, Cha99, CP11b, CP13, Che22, CRP13, CPB^{+}23, Cot13, Cra02, CCP^{+}11, DJS^{+}18, Dră12, DTDd19, Eve05, EH11, For20, GCS^{+}14, GS19, GIM^{+}24, HKS08, HH04, HN09, HK11, How17, HRC20, HLP11, Kab11, Kee10, KC14a, Kon04, KHJ^{+}23, LB12a, LLM^{+}20, LBL^{+}21, MB03, Mir14, MPV12, MCSAGB20, Mur05, Nie14, PL23, Pal15b, Par15a, PCAS09, PG20]. analysis [PW18, RFGD08, RD93, Rup11, SVM^{+}17, Sch08, Sch17b, Sha22c, Sha12, Sha19, Sha20, Shi16, SS06, SS11, SA20, Sta05, Tei22, Tie09, Tsa13, Tsa14b, Tuk77, Tur12, VS02, Ven04, Vie11, WWDS18, Wil18, WD18, Wil14b, Wün13, YCLL20, ZY19, dSBSvE23, Alb15, Ano06c, CH17, GAE23, Lüt11, Par15b, Lun06, Wil11a, Ano08, Ano10, Ano12a, Edd09a, Fri12, Har07, How16a, Liu16a, Lum08, O’B12a, Sta21, Suc07, dL05a]. analysts [Tav17]. Analytical [MBM18]. Analytics [Cap19, Lan17a, Lip22, Mye12a, Otn17, Piñ24, Pod18, Sco13a, Cha15a, Lan14a, Mil15a, PK12, RN17, ZMS21, Rus15]. Analyze [EBO^{+}13, GW18, MdUÁC10, BMOF17]. AnalyzeFMRI [BDdM11]. Analyzing [Arm14, BT05, BCHY09, BKL05, CB17, DHM11, Dév09, Eva14, ERH01, GRMS11, GR16, GRD13, Gos11, Grö14b, KHR21, LZ10, LX12, MA14, Mug10, RR11, SZ11, Sch16, SM07, Vis10, BBN10, PDH16, SGPSzC22, Bau14, Cha21, Daw03, Pal15a, Zie02b]. André [Unw13b]. Andreas [Ano03c, Ano06d, Eas03, Har03, Zie98, Zie01c]. Andrés [Lip21c]. Andrew [Han13a, Han98, Mai10, Mye12a, Ng11, O’B09a, Peb21, Pet19, Ree19]. Andrzej [Mai14b]. Andy [Ree19, Unw13a]. Angewandte [DK09]. Angoff [MF14]. animals [Cal06]. Animated [NL12, SVC^{+}19]. animation [Xie13a]. Animations [Xie13a]. anisotropic [Alb19]. Anleitung [Ano97]. Anne [Cox05, Dal98]. Annie [Cox05, Dal98]. Annual [Ano95a]. ANOVA [Ber10, Ros10b, BPSS09, DW17, Gu14, RV20]. ANOXic [KP23]. Antagonism [BN10]. Anwendung [Süs93]. API [Smi17]. app2web [VFV13]. Appendix [TRM16b]. Application [BFRP13, DCB^{+}17, MH07, DH97]. Applications [BPSS09, Ber10, Cap19, Cha22, CsC08, Dem20, DMS14, Gep21, Har01, Hos22b, Joh09, Jon13, Lip21a, Lun07a, Mar12, Mül16, Nor14, O’B13, Pol13, Qia10a, dREP12, Rao14, Ros10b, Sch14, Sch17a, Tro09, VFV13, Ven10, VL21, Bia94, Bos09, BCMR19, Car16, CL11, CCP^{+}11, Dem13, HKS08, Hor09, HFY22, Hos22a, Hu21, JWHT13, Riz12, Sch08, SS06, SS11, Tsa14b, VP16, SS18b]. Applied [Abr97, Arg98, BPGR08, BA97, BG96, CP13, Cow03, Eyl22, Fer02, Fou09, GF09, Har23, Hos22a, Jam96, KK99, KZ08, KM01, Lee21, Liu23, Lu18, Mar07, Par06b, Sco13a, Unw12a, VR94, VR02, Wei14, WGE17, Yu12, Zie00, Zie01b, Zie04, CC08b, EH11, Fox02, PK12, RFGD08, Sel98, VR97, VR99, VR00a, DK09, Has18, Krä20, Par15a, PG20, Bra03, Cho22b, Den98, Eas03, Edd09b, Han98, Har03, Hoe09, Lor18, Mye09, Oli17, Pal15b, Ros00, SL09, Sco09, Sha19, Zei16, Zha11]. Applix [Ano99]. Applixware [Ano99]. apply [HRC20, WPW15]. Applying [ZS17]. Approach [BBG16, Den98, EN11, FGEMI12, Gou10, Hub11, Lal17, Mar11, Nun20, Oja10, RLWP16, RJH14, Ros00, San10a, Sha22a, She09, SS19b, SCS13, BA97, CC08b, GB13, Gav10, GAE23, Lan09a, MB03, Ohr14, OOSM18, PH16, Par06b, RN17, XYC22, Otn17]. Approaches [LZ10, PU13, Vis10, De 16]. approximant [LF15]. Approximate [Gra16, Ros08]. Approximation [HH14, LPLPD14, RRSPTR12]. April [Ano95a, MN69]. aquatic [KP23]. Arbeiten [Ano97]. ARC [Ger94, Mat94]. ARC/INFO [Ger94, Mat94]. Archetypal [EL09]. Archimedean [HM11]. Archive [Hor12b, MCSAGB20, MCSGBSA20]. archivist [BK17]. area [GZ21a]. Areal [LRN18a]. Argo [KHR21]. argoFloats [KHR21]. Arguments [Hof11]. Ariel [Oli17]. Arithmetic [Wol94, Wol99]. Arm [IDE15, JPM19, KWE^{+}17, KF14, LZHC17]. Arm-Based [LZHC17]. Arnaud [Unw13b]. Arnholt [Bat08b, Ng09]. Array [BS18, LBW18]. arrgh [Kor18]. Art [IR12, Mat11, SME^{+}09, Wil11b]. Artifacts [BML19]. ASMap [TB17]. aspect [MdL10]. Assemblage [Yua07]. Assertive [Tho18]. Assessing [BN10, Kie08, LQC^{+}12, CC23, HBA19]. Assessment [CB17, RG96, SlRC16, Smi07, USHH18, SO23]. Assessments [ZUL14, LB12b]. Association [Fou09, HSG12, MF15, SL09, ZFZ10, Zha11, Sha21d, Wan22]. associations [HTWT23]. asymmetric [KK21, Asq14]. Asymptotic [FS10a, ZFZ10]. asymptotics [YWL02]. Atlantic [Tyn16]. atmospheric [IESdF18]. Attaining [Gav10]. AUC [GZ21a]. Aufgaben [Kra97]. Augmentation [TDRD15]. augSIMEX [ZY19]. August [HI97]. Australian [SPPP17]. Authoring [Mor18]. Automated [CdM10, Mun14a, Sek11, HCSH15, Ste24, Sel17, Iac15]. Automatic [GZ09, HK08, Tan18]. Automatically [HL09]. autoplotly [Tan18]. Autoregression [MZ08]. Autoregressive [ABC19, Lee13, LRN18a, Wan13, CRP13]. available [AFGH22]. Average [HPWdL15, CRP13]. Averaging [ZF15]. Award [The99]. AZIAD [MAY23]. Azzalini [Den98, Ros00].

B [Abr97, Ano06a, Ano06c, Arg98, Dav95, Fer02, Few09, Jam96, Kim95, Yu12, Zie01b, Zie02b, Lip21b, Lip21c, The99]. b/w [Lip21b, Lip21c]. B2Z [MBR11]. BACCO [Han05]. Back [WDT^{+}12]. Back-End [WDT^{+}12]. Baclawski [Pic09]. Bagging [AGG13]. Balance [Sek11]. balanced [CLK21]. bamdit [Ver18]. Bandwidth [dREP12]. Bank [Van18]. BANOVA [DW17]. Baptiste [Tat18]. Barely [Bro14]. Baron [Mai14a]. BARS [WLK08]. Base [Hel16, Rob22, Wei12, R D04]. Baseball [Bau14, Pal15a, MA14]. Based [AR14, AWBM18, ABGF17, BBG16, BBG12, Bra15, CQZ^{+}10, Cul11, DCB^{+}17, FHH17, Fuj17, GFS14, HD12, HMR^{+}13, Hub11, KK14, Lal17, LIL^{+}15, LSPvdL17, LZHC17, LBC^{+}16, Mel16, MVS13, NAA17, Oja10, RBHB15, RGD12, SL09, San10a, SR18, Sha21a, SLS^{+}12, TM05, TFR16, WMS17, WHK21, Wee10, AFGH22, BCMR19, CPD^{+}20, Cha99, De 16, ELS17, Fou09, GZ21b, HLT09, HBA19, HBNS14, HT19, HMRS14, HRC20, KF10, KK22, LB21, LF15, MB03, NWH21, RRSPTR12, SA15, SA20, SADU23, WRV23, YCLL20, Zha11, McN22]. Basic [Aiz12, Fox05b, GK14, Lof21, PLR^{+}16, Tol23]. Basic-Statistics [Fox05b]. Basics [Ano03c, Ano06d, Bar02, Bro07, Fie15, Fox05a, Mac98, Sis03, Zie98, Zie01c, KO97, KO00, Kra05, Liu16b, Kha16]. Basketball [ZMS21]. Basso [Ber10, Ros10b]. Batch [BLM^{+}15, LBS17, Hof11]. BatchExperiments [BLM^{+}15]. Batching [Hof11]. BatchJobs [BLM^{+}15]. batchtools [LBS17]. Bates [Woo01, Zie01a]. Baumer [Liu18]. Bayes [JS05, KG17, Rob23, Woo05]. bayesclust [GFC12]. Bayesian [Gre22a, Har23, Hew16, Liu23, Mal21, Rui17, Alb07, Alb09, AA20, ACG^{+}16, AKL^{+}21, BS15, BGH^{+}17, BKL05, Bür17, Bur12, CF14, CPA20, CCJMR16, Cho09, DW17, EE07, GCS^{+}14, GG16, Gra07, Han05, Hel22, JHQ^{+}11, Kom09, Kur19, Lee13, LBC^{+}16, MR14, McE16, MdN22, MBR11, Mwi11, ND12, NAA17, OK14, PLGHG^{+}23, Pol11a, RAM05, Sah22, SP14a, Sav16, Sch11, Sch17c, Scu10, SD15, SWCP20, SCS13, SVCB18, SR17, Tat18, TDRD15, TNM17, Ver18, WLK08, WLH^{+}18, WM14, XYC22, ZF15, Cha14a, Cho15, Dem17, Sha23a, VSS^{+}17]. Bayesianbetareg [CCJMR16]. BayesLCA [WM14]. bayesTFR [SAR11]. BayesTwin [Sch17c]. BayesX [BKL05, UAK^{+}15]. Baylac [Mai09]. BB [VG09]. bccp [Tei22]. bclust [ND12]. bcp [EE07]. BDEsize [CLK21]. be [GJK^{+}21]. beam [OS95]. beamplots [HBA19]. beanz [WLH^{+}18]. Beautiful [BY18]. Beckerman [Ree19]. Bécue [Ben21, Rie19, Sán19]. Bécue-Bertaut [Ben21]. Beginner [Bai11, ZIM09]. Beginners [Bel19].

Behavioral [LB12a, SD18, Mai14a].

Behavioural [Bee13]. Benjamin [Liu18]. Bennet [Pod18]. Bergm [CF14]. Bernhard [Ano10, Lüt11]. Bernoulli [BS18]. Bertaut [Ben21, Rie19, Sán19]. Bespoke [NRD16]. Best [Nas14]. Beta [CNZ10, GKZ12, MP14, NAA17, CCJMR16, CHB14]. Beta-Binomial [NAA17]. Between [SBMG06, De 16, LS16, Urb09, WDT^{+}12]. Beyond [Ald20, Eyl22, RG96, Var14, ZUL14, Lee21]. bgmm [BSVT12]. Bi [TV11]. Bi-Parental [TV11]. bias [AY22]. BibLaTeX [McL17]. BibTeX [McL17]. Bidimensional [Car13]. BiDimRegression [Car13]. Biennial [Bia94]. big [HTWT23, Kut13, SKD22]. Bigirumurame [Oli17]. bild [GCA12]. Bilder [Par15b]. Bin [Laz24b]. Binary [AVS20, GCA12, IDE15, KF14, ZGZvB19]. Binomial [DM18, KV13, NAA17, SR17]. Binseqtest [IDE15, KF14]. bio [Hel15, RB11]. bio-ontologies [RB11]. bio.infer [Yua07]. Bioassay [RS05]. Bioconductor [Ano12b, Dră12, GIH^{+}05, LM14a, Sha21d, Doe06, Wan22, Dav07]. biodata [Wün13]. Biodiversity [Han07b]. Bioenergetics [DCB^{+}17]. Bioinformatics [AVS20, GIH^{+}05, SL05, Gen09, Ber09, Dav07, Doe06, dL09b]. Biological [DK09, KHLF^{+}10, CC23]. biologischen [DK09]. biologists [BP17b, Hec15, LŠ20, Ree19]. Biology [BFRP13, Dav07, Doe06, HH18, HH19b, Lab12, GIH^{+}05, Shi16, Val09, Wün13]. BioMark [WF12]. Biomedical [Cap19]. Biometry [Esk21]. Biostatics [Oli17]. Biostatistical [Zie00, Sel98]. Biostatistics [HGG08, LŠ20, H^{+}09]. biotic [GSJC22]. BiplotGUI [lGlRGL09]. Biplots [lGlRGL09]. Biquandle [FW18]. Birds [EBO^{+}13]. Birnbaum [BPL09]. birthday [Bor16b]. bittrex [Kan17]. Bivand [Hoe09]. Bivariate [KN05, Nag18, NAA17, PZK^{+}12, SSV14, Hos22a, Ros08, Sha22c].

BivRegBLS [FBB20]. Black [SGHY11, AGdSC20]. black-box [AGdSC20]. Blind [GL14]. blm [KV13]. BLNN [SWCP20]. bluffs [PL23]. bluffs/cliffs [PL23]. BMS [ZF15]. bnlearn [Scu10]. BNPdensity [AKL^{+}21]. boa [Smi07]. Boca [Agr16, Asz23, Ben21, Cho22a, Cho20, Cur18, Das21, Die21, Edd18, Esk21, For20, Fus22, Gal17, GC23, Gle16a, Grö15a, Grö16, Hou07, Joh20, Kel10, Kha16, Laz24a, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Liu17, Lum08, Mal21, Nun20, Orm17, Par15b, Pet19, Put06, Rob17, Sab19, Sán19, Sha21c, Sha23a, Sha23b, Sta21, Sta05, Tat18, Tu21, Vie11, Wan22]. Bocard [Mye12b]. Bolin [Kal24b, Sha21c]. Bolker [Few09]. bomrang [SPPP17]. Book [AS24, Abr97, AA12, Agr16, Aji17, All11b, AC04, AC07, Ano03a, Ano03c, Ano03b, Ano06b, Ano06a, Ano06c, Ano08, Ano09a, Ano09b, Ano09c, Ano09d, Ano10, Ano12a, Ano12b, Ano16, Arg98, Arm19, Asz23, Bai11, Bar02, Bar18, Bas18, Bat04, Bat08a, Bat08b, Bau14, Beb13, Ben21, Ber10, Ber09, Boo10, Bos09, Bos10, Bow09, Bow10, Bra03, Bro07, Bro10, Buc09, Bul06, Bur07, Bur09, Bur10, BL11, Cam09, Cap19, Cap22, Car10, Car16, Cha14a, Cha21, Cha22, Che23, Che11, Chi07, Cho22a, Cho15, Cho09, Cho22b, Cho20, Chr09, CH11b, Cow03, Cox05, Cur18, The99, Dal98, Das21, Dav95, Dav07, Daw03, Dem17, DN17, Dem18, Den98, Dia06, Dia18, Dic12, Die21, Doe06, Doe10, Dow17, Dre19, Dur14, Dur15, Eas03]. Book [Edd09c, Edd09a, Edd09b, Edd11, Edd12b, Edd12a, Edd18, Eme08, Esk21, Esp15, Eva11, Eva14, Fer02, Few09, Fin11, Fis06, Fle11, For20, Fot07, Fox05a, Fri12, Fri11, Fus22, Gal17, GC23, Gep21, Gil14, Gle16a, GR18, Gou05, Gou10, Gre22a, Gre22b, Grö11, Grö13, Grö15b, Grö15a, Grö16, Grö18a, Gro18c, Gro08, Gun06, Gut11, Hac17, Häg12, Han13b, Han13a, Han98, Har03, Har19, Har23, Har07, Has18, Hec07, Hel15, Hel16, Hel18, Hew05, Hew16, Hil06, Hil10, Hly09, Hoe09, Hof15, Hor12c, Hos22b, Hou07, How11, How16a, How16b, Hu09, Hu21, Hub11, Hüs18, IR12, Iac15, Jam96, Joh09, Joh20, Jon13, KK99, Kal24a, Kal24b, Kau13, Kel10, Kha16, Kha17, Kha18, Kim20b, Kim95, Kin20, Kle17, Ko95b, Ko95a, Kos15, Kos16]. Book [Krä20, Kuh10, Kum07, Kum10, Kus03, Lal17, Lan17b, Law02, Laz11a, Laz11b, Laz24a, Laz24b, Lee21, Len20, Leo10, Lep14, Lew16, Lha14, Li11, Lig09, Lip20a, Lip20b, Lip21a, Lip21b, Lip21c, Lip22, Liu15, Liu16a, Liu16b, Liu17, Liu18, Liu23, Lor18, Lu18, Lud13, Lum01, Lum08, Lun02, Lun06, Lun07b, Lun07a, Luo09, Lüt11, Mac98, Mai06, Mai08, Mai09, Mai10, Mai11, Mai12, Mai14a, Mai14b, Mai21, Mal13, Mal21, Mal09, Man03, Mar11, Mar12, Mat13, Mat15, Mat16a, McG20, McN14, McN04, McN22, Mig10, Mil12, Mil10, Mor18, Mor03, Mor09, Mül16, Mun14b, Mur14, Mye09, Mye12b, Mye12a, Neu11, Neu12, Ng06, Ng09, Ng11, Nie11, Nie14, Nor08, Nor09, Nor14, Nor15, Num13, Num20, Nun20, O’B08, O’B09c, O’B09a, O’B09b, O’B10]. Book [O’B12b, O’B12a, O’B13, O’B14, Oli07, Oli10b, Oli10a, Oli17, Oom10, Orm17, Otn17, Pal15a, Pan15, Par15b, Par15a, Pav16, PG20, Peb21, Pet02, Pet19, Pfa12, Pic09, Pie09, Piñ24, Plu01, Pod15, Pod18, Pod21, Pol09, Pol11a, Pol11c, Pol13, Put06, Qia10a, Rao14, Ree09, Ree18, Ric11, Rie19, Rip11, Rob12, Rob13, Rob19, Rob05, Rob07, Rob08, Rob09, Rob17, Ros09, Ros10b, Ros00, Rui16, Rui17, Rus15, SL09, Sab19, San10b, San10a, San11, Sán19, San03, SCD07, Sau10, Sau11, Sau12, Sau21, Sch08, Sch09, Sco09, Sco10b, Sco10a, Sco11, Sco12, Sco13b, Sco13a, Sel17, Sen14, Sha16, Sha21a, Sha21b, Sha22a, Sha11, Sha19, Sha21c, Sha23a, Sha23b, She11, Sis03, Smi06, Soe10, Sól10, Sta21, Sta05, Sta07a, Sta07b, Ste11].

Book [Sto11, Str10, Suc07, Sun15, Sun16, Tat18, Til96, Tod22, Tol23, Tol03, Tsa14a, Tu21, Tur12, Tus05, Tyl07, Unw11, Unw12b, Unw12a, Unw13b, Unw13a, Utl05, VM09b, VM09a, Veh13, Ven10, Vey14, Vie11, VSS^{+}17, VL21, Vis10, Voe09, Wan06, Wan16, Wan22, Wic08b, Wic08a, WG10, Wil14a, Wil11a, Woo01, Woo11, Woo98, Yal10, Yu12, Zei16, Zha11, Zie98, Zie99, Zie00, Zie01c, Zie01b, Zie01a, Zie02a, Zie02b, Zie04, Zie05, dL05b, dL05a, dL06, dL09a, dL09c, dL09b, Cra07, Cra12, Dav15, MN03, Laz11b, Lig09, Sta07a, Gil14, Sha23b]. bookdown [Mor18]. Books [Mor18]. Boosting [AGG13, CJM06, De 16, GAE23, HMRS14]. boostingDEA [GAE23]. Bootstrap [DH97, HH14, Han13a, Mar12, O’B13, VW13, VdL09, CL11, ZHL11]. bor [SD18]. Borcard [O’B12b]. Boundary [DLN17, HS18]. Bounds [Hla16]. Bouvier [AC04, Cox05, Dal98, Woo98]. Bowman [Den98, Ros00]. Box [SGHY11, AGdSC20]. bracelets [DG23]. Bradley [Fir05, TF12]. BradleyTerry2 [TF12]. Brain [RR11]. branch [GL15]. Braun [Hly09, Nor09]. Bretz [Che11, Dic12, Ric11]. Brian [Ano06b, Daw03, Kau13, Ko95a, Lun06, Pfa12, Til96, Unw12a, Zie02a]. Brief [Bec94, EB18, RL15, Edd21, FM18b]. brms [Bür17]. Broken [Sha21c, Sha23a, GJK^{+}21]. Broman [Doe10]. Browser [BP17a]. Brq [AA20]. Bruce [Han98, Laz11a]. Bruno [O’B12a]. BSP [Röh00]. bspmma [Bur12]. BugsXLA [Woo05]. Building [BCHR15, CG15, HMS16, Kuh08, SSV14, GZ21a, GZ21b, SMHBR06]. builds [Ano96]. built [FGG^{+}94]. Bulut [Sab19]. Bundle [Han05]. Bureau [SPPP17]. bursts [DTDd19]. Burt [Utl05]. Burzykowski [Oli17, Mai14b]. Business [RN17, Sco13a, GF09, HTWT23, Par06b, PK12, Otn17]. Buyse [Oli17].

C [Ano09b, Cur18, Edd18, Has18, Rip11, Sco12, Sco13b, Tol23, Woo01, Zie01a, Ano99, BS13, EF11, ES14, EB18, EK12, GH18, GHN19, GMF18, Lan09b, Mat16b, Mwi13b, dVSWAL17, SCS13, Smi18, Sta08, Ste00, Urb09, WLK08]. C- [BS13]. c-tree [Ano99]. C. [Ano09b]. C/C [Lan09b, dVSWAL17]. Caceres [Wan22]. Caching [Pen08, SNR18]. CADF [Lup09]. CAinterprTools [Alb15]. Calculate [TFR16, ZFZ10]. calculating [Alb19]. Calculation [FW18]. Calculations [LQC^{+}12, PLZ^{+}15, Spe13, LLM^{+}20]. Call [PLLC11]. Calling [But05]. Cambridge [Bar95, Pod18]. Campagnoli [Mig10, O’B10]. can [KP23]. cancer [Hu21]. cancerclass [JKvT^{+}14]. Candidate [AHvD09]. Canonical [GDMB08, LM18, dLM09c]. Can’t [The99]. Capabilities [KK14]. Capability [SFS12]. Capture [BR07]. Capture-Recapture [BR07]. captured [CBHLG21]. CARBayes [Lee13]. CARBayesST [LRN18a]. caret [Kuh08]. Carey [Doe06]. Carl [O’B09b]. Carlo [Neu11, Rip11, Sco10a, CGC11, DK18, LT16, MQP11, RC10, SP10, TT21]. CARMA [IM15]. Carmona [Dia06, Zie05]. Carpal [The99]. Carson [LB21]. Cartograms [Pan18]. cartography [GL16b]. Case [ADH11, Bra03, Buc09, Eas03, HSL11, PD08, ZFZ10, BVFB19, Har03, HBNS14, KM01, TNM17]. Case-Control [HSL11, ZFZ10, HBNS14]. Casella [Rip11, Neu11, Sco10a]. Cat5 [Ano99]. Catastrophe [GvdMW09]. Categorical [Agr13, Agr16, BPP17, BL15, FM16, GT10, Har23, Liu16a, Liu23, MH18, Mel16, Par15b, PU13]. Categorization [HMR^{+}13]. catR [MR12]. Causal [HPWdL15, KMC^{+}12, RJH14, SH17, SvdLN17, TYH^{+}14, Shi16, SMM^{+}22]. Cause [Shi16]. CCA [GDMB08]. CDF [SS19a]. CDF-Quantile [SS19a]. cdfquantreg [SS19a]. CDM [GRK^{+}16]. CDVine [BS13]. Cell [BFRP13, OSH^{+}21].

Censored [AB17a, FS10a, OL17, TGJ17, Kom09]. Census [Alm10, Smi17]. Center [IEE93, IEE94]. centered [JPOJ12]. Central [Phi10]. centric [SKD22]. CGIwithR [Fir03]. Chain [MQP11, VV16]. Chained [vBGO11]. Chains [Sch16]. challenges [TLH11]. Chambers [Ano09d, Kle17, Plu01, Rob17, Oli10b]. Chan [Joh09, Qia10a]. Chang [Hu09, Kel10, O’B08]. Change [EE07, JM15, ZLHK02, RHW21, XYC22]. Changepoint [KE14, ACG^{+}16, KE14]. channel [DTDd19]. Chapman [AS24, Ben21, Die21, Esk21, Fus22, GC23, Gle16a, Grö16, Hac17, Has18, Hou07, Joh20, Kal24a, Kel10, Kha16, Kle17, Laz24a, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Liu18, Liu23, Lum08, Oli17, Orm17, Par15b, Peb21, Pod21, Put06, Rao14, Ree19, Rob17, Rus15, Sab19, Sha23a, Sha23b, Sta05, Til96, Tu21, Vie11, Wan22, Cho22a]. Characteristics [GG16]. Characterization [BLY18, SADU23]. Characterize [MTPL15]. Chart [ZP13]. Charts [Anh18, KK21]. Chaudhary [Ano08, Lum08]. checkr [Tho18]. Chemical [Guh07]. Chen [Num13, Par15a]. Chernick [Mar12, O’B13]. Chester [Fus22, Lip20b]. chi [CM14]. chi-square [CM14]. Chichester [Hel15, Iac15]. CHICOM [GHN19]. Chihara [Sco12]. Childs [Ree19]. CHIWEI [GH18]. Choice [Aiz12, Eva14, dREP12, Sar16, Sch11, Arm14]. Chollet [Kim20b]. Chris [Rus15]. Christian [Cha14a, Iac15, Mye09, Neu11, O’B09b, Sco09, Sco10a, Sel17]. Christopher [AS24, Lip21b, Par15b, Pet19, Sab19, McG20, Peb21]. Chromatograms [CS12]. Churchill [Bar95]. Cichosz [Bas18]. CircNNTSR [FDGD16]. CircSpaceTime [LSM20]. Circulant [DB13]. Circular [BFRP13, FDGD16, Mul15, OCRC14, PNR13, LSM20, Lew16, Liu15, Pan15]. ck [HTWT23]. ck-FARM [HTWT23]. CL [Spa17]. Class [Bea17, Lei04, LHS08, LL11, PC11, Sho13, WM14, ATYK20, BBG14]. Classes [KR10]. classical [SS18b]. Classification [AVS20, AGG13, AWBM18, Arc10, CC08a, GSD12, GZP14, LIL^{+}15, Lee18, RBHB15, BCMR19, GZ21a, GZ21b, SO23, ZGZvB19]. Claude [Kum10, Mai09]. Claus [Häg12, Neu12, Hac17]. Cleaning [Grö18a]. ClickClust [Mel16]. Clickstream [Sch16, Sch16]. Client [Ano99, AB12, CH17, Kan17]. cliffs [PL23]. Climate [HCM^{+}18, AFGH22]. Climatology [Dia18, EJ13, Spa17]. ClimDown [HCM^{+}18]. Clinical [CP11b, JPM19, LQC^{+}12, USHH18, Hu21, PFT^{+}12, TPAM07, Dre19, Num13, She11]. Closing [WDT^{+}12]. Cloud [HT19, Mat15, Ohr14, YCLL20]. Cloud-based [HT19, YCLL20]. clues [CQZ^{+}10]. Cluster [BPDD08, FGEMI12, GASA15, Hof11]. Cluster/Multicore [Hof11]. Clustered [HC05]. Clustering [AWBM18, AM14, BBG12, Bra15, CQZ^{+}10, CKSLS12, FM18a, HD10, KK14, KSBZ16, LH12, MCM12, Mel16, MV14, Mül13, ND12, SR18, YEL18, BCMR19, CKSLS18, MT20, SKD22, SHR97]. Clusters [CGBN14, GFC12, FBNRG21]. ClustGeo [CKSLS18]. ClustOfVar [CKSLS12]. clustvarsel [SR18]. clValid [BPDD08]. cmaRs [YÖYB23]. CMIP6 [AFGH22]. CMIP6-D&A [AFGH22]. Co [GVM16, Har07]. Co-Integrated [Har07]. Co-Occurrence [GVM16]. Coarse [MBGK18]. Coarse-Grain [MBGK18]. coastal [PL23]. Code [Aiz12, AD15, BDMP15, BK11, BS18, CG15, Car13, CS12, CO16, DMB18, FKP17, GHN19, GVM16, Han05, Han06a, HW07, Han07a, HXY12, Hof11, Hoh18, Jal19, LW16, LQC^{+}12, LRN18b, MF14, MHJS16, MTPL15, MBGK18, MF15, MP14, NK06, NAA17, Pan18, Pap16, PZK^{+}12, Phi10, Rec10, RBB18, Rob18, RD92, Ruf09, SBMG06, TR14, Tie09, VSV09, VYD^{+}12, Wei12, XMW10, dSJdSF14, BVFB19, Beh04, Beh05, GH18, GMF18, Lan09b, RD93, SMM^{+}22, Gro18c]. CodeMeta [Boe17]. codes [ABEY18, VP16]. Coding [Hec07, dL05b, Mur05, SS18a]. Coefficients [TR14, Hel22]. Cognitive [GRK^{+}16]. Cohorts [BBGL17]. Cointegrated [Ano10, Edd09a, Lüt11, Pfa06, Pfa08a]. Colin [Cho20]. Collaboration [BBGL17]. collected [DO94]. Collection [Iac15, Mun14a, RG07, Sel17]. College [Bar95]. Collocation [Sch17a]. color [Lip22]. column [YH14]. COM [BN07]. Combination [YEL18]. Combinator [Siu17]. Combinatorial [Val09, Lab12]. Combinatorics [Han07a]. Combining [BMGT15]. Command [HL09, Hof11]. Command-Line [HL09]. Commander [Fox05b, FC12, Wil12]. Comment [Lub91, Pen17]. Common [MBT^{+}20, AFGH22, Woo05]. Communication [Kin20, LA19]. Communications [RL15, Gav10]. Community [BCHR15, Gle16b, Wu24, ZGS^{+}18]. comorbidity [Gas18, Gas18]. Companion [Ano06c, Cow03, Lun06, McN22, Zie04, dL05a, Den13, Eve05, Fox02, Kau13, Mat13, Mur14]. Comparative [ADH11]. compareGroups [SSV14]. Comparing [LA19, ZHL11, GHN19, Han13a]. Comparison [Lan95, OL17, Sch14, TM15, De 16, FBB20, Las97]. Comparisons [BHW11, Che11, Dic12, HD12, KPSH15, MVS13, Ric11]. Competing [BAS12, MN17, SZ11, dWFP11, BVFB19, Mal13]. Compilation [Lan14b]. compiled [GMF18]. complement [OS95]. Complete [OK14]. Completely [JJJ14]. Complex [AT13, BY18, GP12, MN14, Obe14, SvdLN17, TMKD17, Kut13, MCSAGB20, SK17, SGPSzC22]. Complexity [Bra15]. Complexity-Based [Bra15]. Component [BDdM11, GKD14, MW07]. Components [BdMM15, Kom09]. Composition [Yua07]. compositional [PG20, Sha19]. compounding [NPR13, NPR13]. CompRandFld [PB15]. Comprehensive [Kal24b, RFKM12, SR17, Hor12b, MCSAGB20, MCSGBSA20]. Computation [AB17b, ABGF17, Dem17, KC14b, MN69, dVSWAL17, Alb07, Alb09, Cho09, Kut13, Mwi11, Pol11a, Röh00]. Computational [Cha15b, Dav07, Din06, DO94, Doe06, How17, Lab12, LZ10, Saw09, Vis10, WDT^{+}12, Wün13, DG22, GIH^{+}05, SMM^{+}22, Val09, Mil12, Boo10, Dur15, Edd09c, Kha18, Pod21, Rob19]. Computationally [BG18]. computations [SR17]. Compute [Phi10, dSJdSF14]. Computer [All86, Ano99, DHF15, Han05, PPGD15, RGD12, Hor09, SS18a, SA15, All86]. Computerized [CPA20, MR12]. Computing [Ald20, AGM07, Cha10, CQ95, Den16, DO94, Edd18, Esp15, FBdlF12, FT08, GI00, Gio09, Kus03, Len20, Mat16b, R D11b, RTL07, SFS12, SME^{+}09, Sco13b, Tem97, TRL09, WSZ12, ACG^{+}16, Ano03b, Cra02, Dan18, DS20, Day15, Edd21, EK12, EKP^{+}11, Gas18, GSJC22, HLT09, JPOJ12, Lan18, Lon15, Mwi13b, NPR13, Ohr14, Riz08, WWDS18, Wil12, YS13, Mat15]. Concave [CGS09]. Concepts [BVFB19, Lun02, Mor03, SA01]. Concerto [SA15]. Concurrent [AD15, KHJ^{+}23]. Conditional [Lee13, LRN18a, OHD17]. Conditioning [Mar06]. Conditions [Yua07]. Conducting [KWE^{+}17, SvdLN17, Vie10]. condvis [OHD17]. Cone [LM14b]. coneproj [LM14b]. Conference [DO94, HI97, Ano95a, Bia94, IEE93, IEE94, MN69]. Confidence [KPSH15, dCCM^{+}22]. Confounding [Laz24b]. Conjugate [Sho13]. connect [CBL^{+}19]. connectivity [MS23, Urb09]. Conservation [VL21, MS23]. Consistency [LQC^{+}12]. Constrained [BFRP13, LM14b, LC10, Mug10]. Constraints [Grö10, CKSLS18]. Constructing [CSNF18, GGS23, MB15]. Construction [Jon07, Kav15, TB17]. Consumed [PZK^{+}12]. consumption [KKL^{+}15]. Contact [GW18]. Contaminated [PMM18]. ContaminatedMixt [PMM18]. contemporary [SP14b]. Context [RG07, BVFB19]. conting [OK14]. Contingency [OK14]. Continuous [AE21, BHS00, CC08a, Hoh18, KY10, KK14, Mul14, PJSPC17, Wan13, NPR13, TPAM07]. Contrast [WS11]. Contrast-Enhanced [WS11]. contribution [Sta05]. Control [ADH11, BCS96, HSL11, MG09, TKM15, ZFZ10, ZP13, ELS17, HBNS14, KK21]. Controls [BCAB07, KHJ^{+}23]. controversy [CM14]. Convention [IEE93, IEE94]. Convergence [Smi07]. Conversion [Hil96, Lei13]. convert [ABEY18]. Convex [KM14, PLRC10]. convoSPAT [RC17]. cooccur [GVM16]. Cook [Eme08, Mai08]. Cookbook [Cha18, LT19, XDR21, Tee11, How11, Sha21b]. cooperative [CBGGV17]. coordination [Amo21]. Copula [Lyu21, Nag18, NAA17, WMR16, ELS17, KY10]. Copula-Based [NAA17, ELS17]. CopulaDTA [NAA17]. copulaedas [GFS14]. Copulas [BS13, GFS14, HD18, HM11]. Copy [SIR^{+}11]. Corel [Ano99]. Cornillon [Unw13b]. corporate [LPR21]. Corporation [Ano99]. Corrected [BK11, HS18]. Correcting [AY22]. Correlated [AD15, Ano08, RMG12, Sho18, BF17, Lum08, SC07, Sta21]. Correlation [BKT14, FM19, GDMB08, HM18, CPB^{+}23, LAJJ18, PMLM23, Ros08, Shi16]. Correlations [LH12]. Correspondence [Mur05, dLM09c, Beh04, Beh05, Alb15, Hec07, dL05b]. Cosenza [Bia94]. Cost [PN97, Alb19]. cost-surfaces [Alb19]. Count [Agr16, FM16, HD18, Hil14, JD15, LFF17, ZKJ08, WM18]. Counterfactuals [SKZ05]. country [ABEY18]. countrycode [ABEY18]. Counts [Cam12]. Course [Gre22a, HH15, Hew16, Hly09, Mal21, Sta93a, Nor09, Pri05b, Rui17, Sha16, Sta93b, Utl05, Wic08a, BM07, BM16, BM21, Dav15, HH04, McE16, TRM16a]. Covariable [KK15]. Covariance [PP17, RC17]. Covariate [HPWdL15, HXY12, Sta08]. Covariates [VKVC15, ZY19]. cover [Das21, Tu21, WRV23]. CovSel [HPWdL15]. Cowpertwait [Li11, Mai10]. Cox [CG15, De 16, DPSH18, Gan15a, TDRD13, TDRD15, TR14]. CoxBoost [De 16]. coxphw [DPSH18]. COZIGAM [LC10]. Cramér [AFM22]. CRAN [MCSAGB20, MCSGBSA20]. Crash [Sta93a, Sta93b]. Crawley [Ano03b, Hel15, Dur14, Kus03, Laz11b, Lig09, Ng06, Sta07a]. crawling [KF17]. CRC [AS24, Agr16, Asz23, Ben21, Cho20, Cur18, Das21, Die21, Edd18, Esk21, For20, Fus22, Gal17, GC23, Gle16a, Grö15a, Grö16, Hac17, Har23, Has18, Hou07, Hu21, Joh20, Kal24a, Kal24b, Kel10, Kha16, Kle17, Laz24a, Laz24b, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Liu17, Liu18, Liu23, Lu18, Lum08, Mal21, Oli17, Orm17, Par15a, Peb21, Pet02, Pet19, Piñ24, Put06, Rao14, Ree19, Rob17, Rui17, Sab19, Sán19, Sha21c, Sha23a, Sha23b, Sta21, Sta05, Tat18, Tu21, Vie11, Cho22a, Par15b, Pod21, Wan22]. Create [GL16b]. creates [ZGS^{+}18]. Creating [BY18, GHG19, Grö14b, HL09, Ver12, Xie13a]. Creation [NRD16]. Creator [The99]. credit [LB12b]. Cressie [Peb21, Pet19]. Crime [GC23, Kal24a]. Criminologist [GC23, Kal24a]. critical [Las97]. crlmm [SIR^{+}11]. crop [CPB^{+}23]. Cross [CH18, ESO16, KO06]. Cross-Language [ESO16].

Cross-Validation [CH18, KO06]. CRS [HHJC14]. CRU [Spa17]. Cruising [The99]. Crunching [Gen98]. Cruz [Lip21c]. Cryer [Joh09, Qia10a]. Crypto [Kan17]. Crypto-Currency [Kan17]. CSS [LeB18]. cts [Wan13]. CUDA [Edd18, Mat16b]. curates [ZGS^{+}18]. Currency [Kan17]. Curve [Wil14a, DCMCPF20, GZ21a, Mir14]. Curves [KHLF^{+}10, OL17, TFR16, PDH16]. Cusp [GvdMW09, GvdMW09]. Customer [PK12, Sco13a]. Customizable [MH09]. Customizing [IRCA21]. Cutpoints [LRRÁCSGS14]. cvmgof [AFM22]. CVTresh [KO06]. CXXR [Run13]. cyanoFilter [OSH^{+}21]. cycle [SGPSzC22]. Cycling [FK17]. cytometry [OSH^{+}21].

D [Abr97, Arg98, Ber10, Buc09, Cap19, Die21, Eme08, Fer02, Han13a, Jam96, Joh09, Mye12a, Qia10a, Sab19, Woo01, Zie01b, BS13, EN11, GGK10, LPLPD14]. D&A [AFGH22]. D-Vine [BS13]. D. [Eme08]. D3GB [BP17a]. DAD [RG07]. Daily [SHN17, RVHG23]. DAKS [ÜS10]. Dalgaard [Ano09a, San03]. Dandelion [MÇSD14]. Daniel [Liu18, Mye12b, O’B12b, Sco13a, Zei16]. Daniela [Nor14]. Dario [Ros10b]. Data [Agr16, AL16, Alm10, AD15, AB17a, Ano08, Ano09d, Ano12b, ARC04, BL14a, BFRP13, BY18, BPP17, Bau14, BC76, BC77, BC81, BC84a, BC84b, BW88, BCW88, BCW18, BML19, BB18, Ben21, BBG12, BK17, BL15, BCG18, BPGR08, Bol08, BDdM11, Bow10, Bur10, But08, CPA20, Cap22, Car10, CB17, Cha69, Cha86, Cha98, Cha08a, CGBN14, Cic15, CFZ09, Cle93, CS07, Das21, Daw03, Den98, Dia06, Dia05, Don17, Dow17, Duo07, ESO16, Edd18, Eme08, FS10a, FBdlF12, Fel12, Fer11, FDGD16, Few09, Fle11, FM18a, FK17, Fri12, Fri11, FM16, Fuj17, GASA15, GP12, Ger20, GKZ16, GGK10, GCA12, Gos11, Gre22b, Grö15a, Grö18a, Gro08, HD10, HD18, HBBP18, Har19, HM18, Hec07, HH15, HH19a, Hil14, HC05, Hoe09, HJ17, HK15]. Data [IP08, Iac15, IG96, Iri19, IK20, Jac11, Jal19, JM15, JKvT^{+}14, JD15, Jon13, Kee18, KHR21, KSS^{+}07, KH10, KK14, Kos16, Kuh10, Kus03, Kut13, LRGTA12, Laz24a, Lei10, LSPvdL17, Lep14, LM03, LZ10, Lip20a, Lip20b, Lip21c, Liu16a, Liu17, Liu18, LT19, MB03, Mai08, Mai14a, MBGK18, MC97, MBM18, Mat16b, MH18, MCM12, Mil17a, Mil17b, MP12, MdUÁC10, MN14, Mun14a, Mun14b, Mur03, Mwi13a, NGBK12, NL12, Nor08, Nor15, NRD16, Num13, Nun20, O’B12a, Oli10b, OL17, Pal15a, PP17, Par15b, Pav16, Peb12, Pen17, Pie15, PU13, Pla12, Pla19, Plu01, Pod15, Pol11c, Pri05b, R D11a, R D11c, Rah17, RHG09, Ree18, Rie19, Riz10, Riz16, RMG12, Ros00, SP14a, Sav16, SZ11, Sch16, SCK95, SO13, Sch17c, Sco10b, Sco13a, Sel17, Sha16, Sha21a]. Data [Sha22a, She11, Sho18, SS19b, SS92, SR16, SCS13, SYC08, SvdLN17, Sól09, SD18, SPPP17, SHN17, Spe08, Ste11, Su07, Sun16, TDRD15, TKM15, TMKD17, TD07, Tur12, ÜS10, Unw11, Unw15, UC21, Utl05, VM09a, Van18, VR99, Ver18, VYD^{+}12, Vis10, WGSL12, WDM^{+}11, WST11, Wic09, WG10, WCHB11, Wil11a, Wil11b, WsT18, ZKJ08, ZQS16, Zie02b, Zie05, dL05b, AFGH22, Agr13, AD17, Ano03b, AB90, BSS^{+}24, BKH17, BC85a, Bee13, BLW12, Bos12, BCMR19, BA97, BdSvE23, Car04, Car14, CCKT83, Cha20, Cha99, Cha21, CP11b, Che22, Cho20, Cle85, Cle94, CC08b, CPB^{+}23, Cot13, Cra02, DG22, Dév09, Dir18, Dră12, Esk21, ERH01, Fal12, For20, FM18b, Fus22, GCS^{+}14, GF09, GGS23, GS19, HCSH15, HH04, HN09, HTWT23, HK11, JPOJ12, Kab11]. data [Kee10, Kle09, KC14a, Kom09, Kon04, Kub23, Lan09b, LSM20, LB12a, LB21, LF15, Lof21, Lum08, MA14, Mil10, Mil15a, MT20, MZ18, MAY23, Mur05, NKHZ21, Nie14, NL14, Ohr14, OSH^{+}21, PCAS09, PG20, Per14, PK12, RFGD08, Riz12, RL15, Rup11, Sah22, Sán19, Sar08, Sha22b, Sha22c, Sha12, Sha19, Sha20, Sha23a, SKD22, SC07, SA20, Smi17, Sta21, Tag23, Tsa13, Tuk77, VS02, Ven04, Vie11, WWDS18, WG17, WM18, XLX^{+}19, YÖYB23, ZY19, ZM14, dSBSvE23, Agr16, Bas18, GAE23, Hüs18, Die21, Mil17b, Sta05, Arm19, Bat08a, Bro10, Cap19, San10a, Sau10, Sau12, Wic08b]. data-centered [JPOJ12]. data-centric [SKD22]. Data-driven [Kut13]. Database [FGG^{+}94, HI97, MC97, BdSvE23]. DataBlade [MC97]. datacollection.com/ [Iac15]. Dataset [SVCB18]. datasets [FAM^{+}20, GZ21a, GZ21b, WRV23]. datastructures [Dir18]. Datendesign [Rah14]. DATforDCEMRI [Fer11]. David [Agr16, Car16, Num20, Nun20, Pol13, Rao14, Sch08]. Davies [Ano16]. dawai [CFSR15]. Dayal [Hof15]. dbEmpLikeGOF [MVS13]. DBKGrad [MP14]. dbmss [MTPL15]. DCE [Fer11]. DCE-MRI [Fer11]. DCEM [SKD22]. Dealing [Kas16, Pap16]. Death [TGJ17]. Deborah [Mai08]. December [Bar95]. Decision [Kim20a, Rob23, Sco13a, MM22, PK12]. Decision-Theoretic [Kim20a]. decisions [Lof21]. Decomposition [DMB18, Lei10, LBW18, BFA14, LHR16]. decon [WW11]. Deconvolution [Fer11, SSH^{+}20, WW11]. Deducer [Fel12]. Deep [Arn17, CA18, Kim20b]. Deepayan [Kuh10, Nor08]. Definition [R D11f]. Definitive [Joh20, XAG19]. Degradation [CP12]. Dehli [Hof15]. delivery [SA15]. Delta [MF14]. deltaPlotR [MF14]. Demidenko [Lip21a]. demogR [Jon07]. Demographic [Alm10, Jon07, SM07]. Demonstrating [Xie13a]. Denis [Tat18]. Dennis [Kau13]. Densities [Nag18, RRSPTR12]. Density [AM14, CGS09, DHM11, Duo07, HS18, Kle09, LF15]. DEoptim [MAG^{+}11]. Dependence [BDMP15, BS13, CFHBK11]. Dependent [KO06, Alb19, DCMCPF20]. Deployment [VFV13]. depmixS4 [VS10]. Derivatives [MS11]. Derive [Die21, Ger20]. Deriving [Arc10, Cad13]. Description [SMHBR06]. Descriptive [Ruf09]. Design [AA12, BC84a, BC85a, BCS96, BPB09, Cha08b, DHF15, GG16, Gra07, Grö11, HSL11, HB92, Hu09, KSC^{+}00, Mat11, MHOV12, O’B08, Ros10a, ZP13, BHH05, CLK21, GV19, Hu21, Kel10, RPVG11, Rob09, You12, IR12]. design-With [Hu21]. Designed [FBF14]. Designing [JPM19, KK21, TGKV20]. Designs [FKP17, Grö14b, HWY18, SlRC16, Was15]. Desjardins [Sab19]. deSolve [SPS10]. Detecting [CGC11, FBNRG21]. Detection [RG07, XYC22]. Determination [RV20, CLK21]. Determining [CGBN14]. Deterministic [MRC15]. Developed [RG07]. Developer [BG18]. Developer-Friendly [BG18]. Developers [LA19]. Developing [LS05, LB12b]. Development [JKvT^{+}14, PN97, RFKM12, SA15, SGPSzC22, TLH11]. Developments [vdA12]. Devices [FK17]. dglars [AMW14]. Diagnosis [GRK^{+}16, TB17, CPD^{+}20, CPD^{+}20]. Diagnostic [JKvT^{+}14, LRRÁCSGS14, LX12, NAA17, PJSPC17, RG07, Ver18, AY22]. Diagnostics [LH14, SGHY11]. Diagonal [KN05]. Diagrams [BDMP15]. DiagTest3Grp [LX12]. dialog [Cha99]. Dianne [Mai08]. DiaThor [GSJC22]. diatom [GSJC22]. DiceDesign [DHF15]. DiceEval [DHF15]. DiceKriging [RGD12]. DiceOptim [RGD12]. DICOM [WST11]. Dieter [AA12, Veh13, Krä20]. Difference [KM08]. different [Ros08]. Differential [Bur09, Cam09, CGC11, Iac08, Laz24a, Lud13, Luo09, MF14, MAG^{+}11, PLGHG^{+}23, Sch14, Sch17a, SPS10, SCM12, Sto11, SS18a, SS18b, Wu24]. Differentially [WHK21]. diffIRT [MTvdM15]. Diffusion [MTvdM15, PT09, SKS15, WS11, RRSPTR12, RRSPTR14]. Diffusion-Weighted [WS11]. Digital [PT07, SSH^{+}20]. Dimension [AR14, Wei02]. Dimensional [ALE10, BBG12, JKvT^{+}14, ND12, SF18, VG09, Cha21, Che23, KB11, Pen03, SA20, TKM16]. Din [Num13]. Ding [Num13, Par15a]. Ding-Geng [Num13, Par15a]. Dinov [Cap19]. directed [Ste24]. Dirichlet [CHB14, LHA^{+}15]. Dirk [Pod18]. Disclosure [TKM15]. discover [HTWT23]. Discoverability [BML19]. Discovering [FMF12, LAF^{+}17, Unw13a]. Discovery [Asz23, DK18, BHH05, Wil11b]. Discrete [FM16, KK14, MP14, NS94, Sar16, Wol99, BF17, LS16, NPR13, Agr16]. discretely [BI13]. Discretized [Su07]. Discriminant [BBG12, CFSR15, Duo07]. discriminative [PLGHG^{+}23]. Discussion [HJ17]. Disease [BVE^{+}15, JGM18]. Diseases [Kub23, Höh07]. Disequilibria [SBMG06]. diskette [Abr97]. Display [BCS96, HH15, Pri05b, Sha16, SBMG06, Utl05, HH04]. Displaying [FT08, Per14, Tag23, Pod15]. Displays [Fox03, FH09, Gru95, NL12, Sta95e]. Distance [Dro18]. distances [LS16]. Distributed [Gas11, Mug10, TFH12, HHJC11, MGHR16, SNR18]. Distributing [Pen08]. Distribution [AWBM18, AHvD09, GFS14, Phi10, dREP12, ZQS16, Asq14, LS16, NPR13, OOSM18]. Distributions [AHvD09, BS18, DMD15, Dem18, Den16, HG14, KY10, LM18, LHS08, ML13, MvSB^{+}07, NK06, NR16, PMM18, Rob12, dVSWAL17, Su07, ZFZ10, BPL09, Fog17, KD11, KK21, NPR13, TN22]. distrMod [KR10]. Dive [Lip20b, Fus22]. Diversity [MH15]. dlmap [HSG12]. dlnm [Gas11]. dmapply [MGHR16]. DNA [Cha21, TNM17]. Do [SVCB18]. document [Fil08]. documentation [Vid22, VC23]. Documents [LeB18, McN14, Mor18, Vey14, Xie13b].

DoE.base [Grö18b]. Doing [SO13, Ano97]. Dolores [Bat08b, Ng09]. Dominic [Iac15, Sel17]. Dominici [Buc09]. doRedis [HT19]. Dose [BPB09, CO16]. Dose-Finding [BPB09]. Dose-Response [CO16]. dosresmeta [CO16]. dotCall64 [GMF18]. Douc [Car16, Rao14]. Douglas [Mat16a, Zie01a]. downloading [BdSvE23]. downloads [HBA19]. downscale [MBGK18]. Downscaling [HCM^{+}18, MBGK18]. dozens [Vin08]. dplbnDE [PLGHG^{+}23]. DPpackage [JHQ^{+}11]. Dräghici [Ano12b]. drake [Lan18]. Drift [Han07b, BSG20]. driven [Kut13]. dropout [XLX^{+}19]. dropouts [SOD^{+}16]. Drug [BN10]. DTD [SSH^{+}20]. DTDA [MdUÁC10]. dti [PT09]. DTR [TM15]. DTRreg [SMWP20, WMS17]. dtw [Gio09]. dtwSat [MCAP19]. Dual [LM14b]. Ducks [Kur19]. Dudewicz [Rob12]. Dudoit [Doe06]. dummies [Pie15, Sch17b, dVM12, Edd12a, Rob13]. Dumont [Gle16a]. Duncan [Hly09, Nor09]. DWSurv [SMWP20]. Dylan [Ree19]. Dynameta [SCK^{+}23]. Dynamic [CQ95, CS12, CS07, Eme08, Gio09, KKEM15, Mai08, MCAP19, Oom13, PPC09, Pet10, TM15, VM09a, Vey14, WMS17, WS11, Xie13b, Dra23, SMWP20, SCK^{+}23, McN14, Mig10, O’B10, Sha11]. dynamical [NWH21]. dynamics [DJS^{+}18, Wu24]. dynmix [Dra23].

e-book [Ben21, Cho22a, Lee21, The99]. E-Learning [ZUL14]. early [PFT^{+}12]. Earthquake [Jal19]. Ease [LT16]. Easy [CdM10, HM16, DP13]. EbayesThresh [JS05]. ebk [Liu23]. eBook [Gle16a, Kal24a, Kal24b, Kle17, Orm17, Zei16, Num20]. echosounder [OS95]. Echoview [HCSH15]. EchoviewR [HCSH15]. eco [ILS11]. ecode [Wu24]. Ecological [AB12, Bol08, Fin11, Han07b, ILS11, PR07, Qia10b, Sól09, Woo11, Qia16, Sha20, SCK^{+}23, Few09]. Ecology [Chr09, Hor12c, Mye12b, O’B12b, O’B14, Oli10a, Pla12, Pla19, Sch09, Soe10, Sól10, Ste09, VL21, ZIW^{+}09, Swe13, For20]. Econometric [CN97, CNZ99, MR09]. Econometrics [Edd09b, KZ08, Mye09, Sco09, Yal10, RH02, Vin08]. Economic [ZP13]. economics [Day15, Hof15]. ecosystem [KP23, MCSAGB20, MCSGBSA20, PSZ17, SO23]. ecp [JM15]. ed [Ano09a, Ano10, Ano12a, Hel18, Kal24b, Len20, Man03, Pfa12, Qia10a, Sau10, Sau11, Sau12, Sau21, Sta21]. edcc [ZP13]. edgar [LPR21]. Edge [HL07]. Edited [Ano06a, Wan06, Doe06, Lip21c, Mal21, Vis10]. Edition [Cha22, CH11b, Dem17, DN17, Edd09a, GR18, Grö15a, Hac17, Has18, Hel15, Kal24a, Kha17, Lor18, Lu18, Mai12, Pol13, San10a, SCD07, Unw12b, VM09b, Ano08, Gal17, Joh09, KK99, McG20, Orm17, Piñ24, Ree19, Rob22, AC04, Bos09, Gil14, Tsa14a, Ano12b, Dur14, Fox05a]. Editor [MN03]. edn [Cox05, Lum08, Hew05, Ree18]. edn. [Sch08]. eds [Har03]. Educational [DB18, Sab19, HFY22]. Edward [Rob12]. Edwards [Han13b]. Effect [CLL17, Fox03, FH09, SvdLN17]. effectiveness [KK22]. Effects [Bea17, Chr09, Far06, Far16, Gal17, Gan15a, HPWdL15, LQC^{+}12, Mai14b, Mug10, Oli10a, SP14a, Sch09, Wan06, WLH^{+}18, ZIW^{+}09, GB13, Orm17, PB00, Ros09, Yan95, FH09, Woo01, Zie01a]. Efficient [AHvD09, BG18, BGH^{+}17, ESO16, GL16a, Hel22, TB17, CLK21, GMF18, LAJJ18, RRSPTR12, Lan17b]. Effort [PN97]. Einführung [Luh13, Süs93, Kra97]. Einsteiger [Luh13]. eixport [IESdF18]. Ekstrøm [Hac17, Neu12, Häg12]. Elea [Rus15]. Electronic [Kra95]. Elegant [Wic09, Wil11a]. Elements [Lyu21, Spa17, Cle85, Cle94]. Elena [Bai11]. Elliott [Has18]. Elliptic [Han06b, Han06b]. else [Wil12]. Elsevier [Tol23]. Elst [Oli17]. Emails [LA19]. Embarrassingly [DP13].

Embedded [BDdM11]. Embedding [DB13]. emil [BG18]. EmiR [PS22]. emissions [IESdF18]. Emmanuel [Ano12a, Cox05, Dal98, Mai09]. EMMIXcskew [LM18]. EMMIXuskew [ML13]. Emphasis [FS10b]. Empirical [ADN15, Cha10, JS05, KG17, LHR16, YS13]. empirically [TGKV20]. Employing [ZF15]. Enabled [FK17]. Enables [Wei12]. Encounter [Car17]. End [HM16, WDT^{+}12, WRV23]. End-to-End [HM16, WRV23]. End-Users [WDT^{+}12]. endemism [GL15]. endpoint [Oli17]. endpoints [SMWP20, TPAM07]. Energy [PZK^{+}12]. Enforcing [Oom13]. Engine [MC97]. Engineering [Blo14, Bor16a, Rup11, Kos15]. Enhanced [WS11]. Enhancing [BBGL17, BML19, Lan14b]. Enjoying [CBGGV17]. Enrichment [IRCA21]. Ensemble [AVS20, KOC21, LHR16, VWDB16]. entropart [MH15]. Entropy [MVS13, VdL09, GZ21b]. entropy-based [GZ21b]. Enumerative [Han07a]. Envelopment [GAE23]. ENVI [Ano99]. Environment [BC84b, BW88, BCW88, BCW18, Cha12, Cha86, Mug10, R D11a, R D11b, RFKM12, Tem97, Wei12, CNZ99, HLT09, VS02, Ven04]. Environmental [Buc09, MN01, PD08, Qia10b, Qia16, Yua07, dL09c, CBHLG21, Esk21, RFGD08, SO23, Fin11, Law02, Lum01, MN03, Pet02, Woo11, Zie99]. EnvironmentalStats [Mil98, Mil00, Mil02]. Environments [BLM^{+}15, Lub91, CN97]. epidemic [HF07]. Epidemics [GW18]. Epidemiological [MJGM10, PC11, Kub23]. Epidemiology [Buc09, PD08, dL09c]. Epilogue [TRM16b]. EpiModel [JGM18]. epinet [GW18]. Episodically [PZK^{+}12]. EPT [KOC21]. Equality [Grö10]. equate [Alb16]. equateIRT [Bat15]. Equating [Alb16, Bat15]. Equation [Obe14, QZLP21, Ros12, Tu21, Wu24]. Equations [Bur09, Cam09, HH07, HHY05, Iac08, Laz24a, Luo09, Sch14, Sch17a, SPS10, SCM12, Sto11, Tou15, VG09, SOD^{+}16, SS18a, SS18b, Shi16, vBGO11, Lud13]. Equivalent [LAF^{+}17]. Eric [Ano03a, Bur07, Car16, Fis06, Fot07, Laz11a, McN04, Rao14, Unw13b]. Erich [Fle11, Sco10b]. Erik [Bai11]. eRm [MH07]. Error [HXY12, PZK^{+}12, WW11, ZY19]. Errors [AB17b, BK11]. Essentials [Cha14a, MR14]. Estimate [AMW14, PZK^{+}12, Zag18]. Estimates [Gan15a]. Estimating [BKT14, Bra17, Fie12, HPWdL15, HH07, HHY05, KS14, SM07, TMCE22, Tou15, You10, PMLM23, SOD^{+}16]. Estimation [AMYR16, AT13, AM14, CLL17, CGS09, Cul11, DHM11, DBZ^{+}11, Duo07, FDB12, GFS14, GvdL12, Har19, HS18, Joh07, KL14, KSP15, LSPvdL17, dUJ13, MMB15, Nag18, PJSPC17, PN97, dREP12, RC17, RMG12, Rov17, Sav16, SIR^{+}11, SVMMRP17, Sho13, SvdLN17, TM15, TR14, VV16, WMS17, WW11, Wan11, Asq14, BI13, DCMCPF20, Dra23, ELS17, HT11, HHJC11, HHJC14, KK22, Kle09, Kom09, SO23, War23]. Estimator [SP05]. ETAS [Jal19]. ethnicity [Xie22]. EU [Lon15, Esp15]. EU-SILC [Lon15, Esp15]. Eubank [Sco13b]. Euclidean [AE21]. Eugene [Lip21a]. EUR [Hof15]. EUR55.99 [Pod18]. EUR74.95 [Cox05]. EUR85.39 [Num20]. Europe [Esp15, Lon15]. eval [GDBK^{+}21]. Evaluating [GG16, MHOV12, SKZ05, LB12b, MS23]. Evaluation [HSL11, DS20, Oli17, TPAM07]. Event [Bro12, Jon13, Riz10, Riz16, Riz12, Sta05, Beb13]. events [LS16]. Everitt [Ano06b, Ano06a, Daw03, Pfa12, Til96, Unw12a, Yu12, Zie02b, Ano06c, Dav95, Kim95, Ko95a, Lun06, Zie02a]. Everyone [Lan17a, Lan14a]. evidence [BL14b]. evmix [HS18]. Evolution [Ano12a, MCSGBSA20, MAG^{+}11, NdSL16, Par12, PSZ17, Suc07, Par06a, PLGHG^{+}23]. Evolutionary [GZP14, PS22]. evtree [GZP14]. Exact [FS10a]. Example [Eva11, Lud13, Mai11, San10a, AR12, Beh12, Fil08, HLP11, MB03]. Example-Based [San10a, MB03]. Examples [Cap22, DMS14, Edd18, Gre22a, HH15, Hew16, Iac08, Lip22, Mal21, Mat16b, Piñ24, Pri05b, Rah17, Rui17, SD15, Sha16, Tat18, Utl05, VSS^{+}17, Cho15, CC08b, Esk21, HH04, H^{+}96, HBPJ04, McE16, PG20, Sha19, SS06, SS11, Vin08, AC04, Rao14, Sch08, Bur09, Cam09, Hew05, Luo09, Pol13, Sto11, Woo98, Car16, Cox05, Dal98, Har19, Lun07a]. Exams [GZ09, ZUL14]. excavating [Wil11b]. Excel [Fle11, Sco10b, BN07, But05, HN09, Pol09, Tav17]. Exchange [Kan17]. Exclusion [AT13]. execution [SWHJ17]. exercises [Che23]. Expectation [SKD22]. Expected [CNA16]. experiment [Bar95]. Experimental [AA12, AL16, Grö11, RV20, SlRC16, RPVG11, You12]. experimentation [dSBSvE23]. experimenters [BHH05]. Experiments [Aiz12, DHF15, FO15, FBF14, Grö18b, LW16, NGBK12, RGD12, TP11, CLK21]. Explained [Cic15, Bas18, RFGD08]. explicit [BSG20]. Exploration [HBQ04, RVHG23]. Exploratory [Beh12, HLP11, LRGTA12, Tuk77, FBNRG21, GIM^{+}24, MÇSD14, Eva11, Mai11]. Explore [Spa17, Dan18]. Exploring [ARC04, GLC^{+}15, WCHB11, BLW12, Laz24a]. Exponential [Asq14, CF14]. Export [R D11c, IESdF18]. ExPosition [BFA14]. exposure [BSG20]. express [MGHR16]. expression [CPD^{+}20, Cha21]. Extend [GDMB08]. Extendable [BY18]. Extended [GKZ12, MH07, ZC10, ZGS^{+}18]. Extending [BC85b, Cha16, EB18, Far06, Far16, FC12, LWC^{+}09, SVC^{+}19, Vin08, Gal17, Kle17, Ros09, Orm17, Rob17, Wan06]. Extensibility [Lan95]. Extensible [HD10, Lan09a, Run13]. Extension [Tem97, YEL18, VP16]. Extensions [Chr09, FH09, Oli10a, R D11g, Sch09, ZIW^{+}09, MdN22]. Extensive [KS14]. extents [GL15]. extracat [PU13]. extracting [BdSvE23]. extraction [SA20, WRV23]. Extreme [GK16, Hla16, HS18]. ExtremeBounds [Hla16]. extremes [MdN22, GK16].

F [Bai11, Bat08b, Ber10, Dic12, Mai08, Ng09, Oli10a, Sch09]. F. [Buc09]. Facilities [Fir03]. FactoMineR [LJH08]. Factor [BP12, CFHBK11, GIM^{+}24, MÇSD14]. Factorial [FKP17, Grö14b, Grö18b, NGBK12]. Factors [Rob23, SO23]. Failure [CKY14, HXY12]. FairCom [Ano99]. Faithful [AB90]. Falissard [O’B12a]. Families [NR16, TN22]. Family [PG15]. FAMT [CFHBK11]. Far [RBB18]. Faraway [Orm17, Put06, Ros09, Wan06, Gal17, Rob05]. FARM [HTWT23]. Fast [BS16, LH12, Mül13, PG15, dVSWAL17, VW13, KMMV14, KF10]. fastcluster [Mül13]. fastWKendall [LAJJ18]. Fay [IDE15]. fda.usc [FBdlF12]. fdANOVA [GS19]. Feature [DK18, LAF^{+}17, Meu13, SA20]. features [Tag23]. February [Ree19]. fechner [ÜKD09]. Fechnerian [ÜKD09]. Feit [Rus15]. fen [CCP^{+}11]. Fertility [SAR11]. Fetch [SPPP17]. fgui [HL09]. FHtest [OL17]. FIAR [RR11]. Field [Bar95, Unw13a, LŠ20, Unw13a]. Fields [AE21, DB13, FS10b, PB15, SMM^{+}15]. Fieller [Kha16, Liu16b]. Figures [GHG19]. File [Hil96, OS95, NPP17]. Files [EW16, SVM^{+}17]. filesstrings [NPP17]. filings [LPR21]. Filter [Wan13]. Filtering [Tus11, VYD^{+}12]. Filzmoser [PG20]. finance [Rup04, Mül16]. Financial [Ano03a, Bur07, Cap22, Dia06, Dia05, Fis06, Fot07, Kum07, Lep14, McN04, Pfa13, SCD07, Spe13, Tol03, Zie05, Car04, Car14, Kon04, Nie14, Rup11, Tsa10, Tsa13, Tsa14b, ZW03, Ziv05, Pod18]. Finch [Sha21c]. Finding [BPB09, HWY18, Was15]. Fine [MBGK18]. Fine-Grain [MBGK18]. Finite [Lei04, Dra23, GL07]. FinMetrics [Kon04]. Fionn [Hec07]. firm [BMOF17]. First [Hly09, Nor09, Sta95b, Wic08a, BM07, BM16, BM21, Dav15, RRSPTR12, RRSPTR14]. first-passage [RRSPTR14]. first-passage-time [RRSPTR12]. Fish [DCB^{+}17]. FishBase [BLW12]. Fisher [HG14]. Fisheries [Ald20, WRD23]. fishery [KK22]. Fit [MVS13, TMN16, AFM22, GH18, GEV18, JV14, LT16, LB12b]. fitdistrplus [DMD15]. Fitting [BS18, Cad13, CAA15, CKY14, DMD15, Dem18, FC11, GvdMW09, GL07, GH11, HG14, IV05, Jal19, KHLF^{+}10, KV13, LM18, MRC15, ML13, MTvdM15, PZK^{+}12, PMM18, Riz16, Rob12, SM21, Su07, TGJ17, KD11, XLX^{+}19]. Fixed [CC11, DHM11, RV20, ZF15, LS16]. fixedTimeEvents [LS16]. FL [Agr16, Asz23, Ben21, Cho20, Das21, Esk21, For20, Fus22, GC23, Hou07, Joh20, Laz24a, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Mal21, Par15b, Pet19, Sab19, Sha21c, Sha23a, Sha23b, Sta21, Tu21, Wan22, Ano95a, Cho22a]. FlashR [ZMV^{+}18]. fleets [HT19]. Flexible [AHvD09, SlRC16, TPAM07, ZUL14, ZF15, Bar18]. FlexMix [Lei04]. flexsurv [Jac16]. Flickr [FAM^{+}20]. Florida [Cur18, Har23]. flow [OSH^{+}21]. FluxSimulator [MvSB^{+}07]. Flyby [KSC^{+}00]. FME [SP10]. fMRI [BDdM11, WDM^{+}11, TP11]. Focus [DLC06, De 16]. Follow [PC11]. Follow-Up [PC11]. Following [LFF17]. Foods [PZK^{+}12]. forbidden [VV19]. forecast [HK08]. Forecasting [HK08, GK14]. forecasts [RVHG23]. Forensic [Rob23]. forest [BSS^{+}24, Mye12a]. forestry [Esk21].

Forests [Mai21, BSS^{+}24]. Form [KS14]. formalization [BDT18]. Format [Wee10, AFGH22]. Formatting [LeB18]. forms [Fir03]. Formulas [ZC10]. Formulating [DLC06]. FORTRAN [MJR93, Yan95, GH18, GHN19, GMF18, Las97, LS05]. Fortran-77 [GH18, GHN19]. Fortunato [Ros10b]. Forward [ARC04]. Foulkes [Zha11]. Foundations [WML14, Cur18]. Fourteen [Bro14]. Fourth [Fox05a]. Fox [Zie04]. FPGas [Röh00]. fptdApprox [RRSPTR14]. FPTL [RRSPTR12]. FracSim [DC05]. Fractional [Grö14b, BI13, SS18a, SS18b, SMHBR06]. fractionally [CRP13]. Frailty [MRL12, RMG12]. frailtypack [RMG12]. Framework [ATF10, BML19, CH18, HMS16, Lei04, MdL10, NAA17, PR07, TRL09, Bor16a, CPB^{+}23, MR09, SGPSzC22]. Francis [Asz23, GC23, Har23, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Mal21, Nun20]. Francisco [Lip21c]. Francois [Kim20b, Mai11, Mye12b, Unw13b, O’B12b]. Frank [Che11, Ric11]. FRB [VW13]. frbs [RBHB15]. Free [IP08]. Frequencies [Meu13]. Frequency [GKZ16, CHT98]. Frequentist [Har23, Liu23]. Fresh [Sha22a]. FrF2 [Grö14b]. Friedman [Aji17]. Friendly [BG18, Agr16]. Friends [MBM18]. FTIR [CBHLG21]. FTLinuxCourse [Ano99]. Fun [Kur19]. Function [DBZ^{+}11, HB92, MBT^{+}20, PZK^{+}12, dREP12, Rov17, SBMG06, VSV09, VG09, Cot13, PG15, PSM^{+}11, RRSPTR12]. Functional [Swe13, Cha14b, FBdlF12, Lub91, RHG09, RR11, Sav16, SCS13, TP11, YEL18, GS19, MGHR16, Bow10, Bur10, Car10, O’B14]. Functionalities [BMGT15]. Functionality [Guh07]. Functioning [CGC11, MF14]. Functions [Aiz12, BCAB07, Han06b, LH12, LQC^{+}12, MJR93, Phi10, RC17, Ruf09, BH94, Las97]. funcy [YEL18]. Fundamental [LM18]. Fundamentals [Che23]. Further [BL14b]. fuse [VWDB16]. Fusion [SYC08]. Future [Ano99, Iha98]. Fuzzy [RBHB15, SM21, VV16]. fuzzyreg [SM21]. FuzzyStatProb [VV16].

G [Bai11, Bur10, BL11, Car10, Mig10, Mor09, O’B09c, O’B14, Oli10a, Sch09]. G. [Rip11]. GA [Scr13]. Gabor [CHT98]. Gabriel [Lab12]. Galecki [Mai14b]. gamboostLSS [HMS16]. game [FVVL23]. Games [KS14, CBGGV17]. GameTheory [CBGGV17]. GAMLSS [Bar18, HMS16, SR07]. gamma [Ros08]. gammSlice [PW18]. Gandrud [McG20]. Gao [Han98]. Gap [WDT^{+}12]. Gareth [Nor14]. Garrett [Joh20, Mil17b]. GAS [ABC19]. GAUSS [CN97]. Gaussian [TDRD13, AE21, BKT14, DB13, FS10b, Gra07, GT10, Gra16, HD18, HL07, LHS08, MRC15, PLZ^{+}15, SR18, TDRD15, VRP23]. GBP [Die21, Iac15, Kle17, Oli17, Rob17]. gcKrig [HD18]. Gebhardt [AA12]. geepack [HHY05]. Geert [Oli17]. GEEs [SOD^{+}16]. gems [BVE^{+}15]. Gene [IRCA21, WHK21, Cha21, HBNS14]. Gene-Based [WHK21]. GeneDMRs [WHK21]. General [FKP17, Lei04, LHS08, MdL10, RRSPTR14]. Generalised [Fox03, PW18]. Generalization [YEL18]. Generalized [AMYR16, ABC19, Arm19, AMW14, BdMM15, CdM10, Cha10, CH18, Eyl22, Far06, Far16, Gal17, Had10, HC05, HHY05, ICL16, Lee21, LFF17, LC10, MH09, ML11, Sch11, SR07, Su07, Tou15, Wan06, Woo06, dSJdSF14, BPL09, BC11, Er12, GL07, Hel22, LB12b, Orm17, PH16, Ros09, SOD^{+}16, AC07, Chi07, GR18, Mai06, Rob07, Sta07b]. Generalizing [PLRC10]. Generate [Ruf09]. Generating [Boe17, CHB14, Fuj17, WDM^{+}11]. Generation [AD15, GZ09, MR12, Thi18, VYD^{+}12, ZUL14, AD17, DG22, Dev86, DG23, Tan18]. generator [Ros08]. generators [Fog17]. Genetic [LZ10, MF15, MS11, Scr13, TB17, Vis10, ZFZ10, HBNS14]. Genetics [Fou09, SL09, Zha11]. geneXplain [SKW17]. geneXplainR [SKW17]. Geng [Num13, Par15a]. Genome [BP17a]. Genomic [Hic16]. Genomics [LM14a, Pod21]. genotype [SADU23]. Genotyping [SIR^{+}11]. Gentle [Sta93c, Sta93d]. Gentleman [Doe06, Ber09]. Genuer [Mai21]. geoCount [JD15]. Geographical [Kav15]. Geographically [GLC^{+}15]. geomnet [Tyn16]. geomorphic [PL23]. georeferenced [GL15]. George [Neu11, Sco10a]. Geospatial [CH17]. Geostatistical [HD18, JD15, SYC08]. geostatistics [LKD08]. GeoXp [LRGTA12]. Gerbing [Die21]. German [DK09]. Germany [Har03]. Get [Cun12]. getCRUCLdata [Spa17]. Getting [BP17b, HHB08, Mat98, Ree19]. Geyser [AB90]. GFD [FKP17]. ggm [Mar06]. GGobi [CS07, LWC^{+}09, Eme08, Mai08, VM09a]. ggplot2 [SVC^{+}19, Ste23, Wic09, Wic11b, Wil11a]. Gibbs [Laz11a, ST10]. Gifi [dLM09a]. Gillespie [PK08]. GillespieSSA [PK08]. Gillet [Mye12b, O’B12b]. Giovanni [O’B10]. GIS [RG96, VEF^{+}23]. GLDEX [Su07]. glmulti [CdM10]. Global [MAG^{+}11, Mul14, SHN17, Sta95c, Wol94, HRC20]. Globally [GZP14]. Glotaran [SLS^{+}12]. glow [PDH16]. gMWT [FO15]. Go [Kor18]. Godalming [Plu01]. goft [GEV18]. Gomez [Hoe09]. Gomez-Rubio [Hoe09]. Gonzalez [Wan22]. Good [Pol09]. Goodness [LT16, MVS13, AFM22, GH18, GEV18, JV14]. Goodness-of-Fit [MVS13, LT16, AFM22, JV14]. Gosset [HJM08, dSBSvE23]. Government [SPPP17]. GPfit [MRC15]. GPS [FK17]. GPS-Enabled [FK17]. GR [BC76, BC77]. GR-Z [BC76, BC77]. gradient [NWH21]. Graduation [MP14]. Graham [Sau12]. Grain [MBGK18, Høj12]. Gram [HMR^{+}13]. gramEvol [NdSL16]. Grammatical [NdSL16]. granularity [OSH^{+}21]. Graph [Cul11]. Graph-Based [Cul11]. gRapHD [ALE10]. Graphical [ALE10, AB12, BC76, BC77, CCKT83, Fox05b, HL09, Høj04, HL07, Høj12, HEL12, KMC^{+}12, LL10, Mar06, RFKM12, SlRC16, SBMG06, SLS^{+}12, TMW18, Unw15, VML12, WLH^{+}18, Han13b, Kos16, Sun16]. Graphics [BC84b, BW88, BCW88, BCW18, Cha86, Cha18, CS07, Eme08, Esp15, HK15, IG96, Kee10, Kee18, KH10, Lan17a, LBC^{+}16, LR15, LT19, Mai08, Mur06, Mur09a, MG09, Mur09b, Mur18, NL12, R D11a, San10a, Sha12, SVC^{+}19, VM09a, WC07, Wic09, Wil11a, Bos12, HN09, HK11, Kab11, Lan14a, Lon15, MB03, Mur11, Ree19, Ste23, Tak12, VS02, Ven04, Liu17, Ree19, CH11b, Fle11, Grö15a, Gun06, Pol11c, Sco10b, Ste11, Unw12b, Bul06, Fri11, Lha14, Ree18, Unw11, Vie11, WG10, Tur12]. graphicsQC [MG09]. graphing [Cle85, Cle94, Day15]. Graphs [CF14, Ste24]. Grassmann [ACW12]. GrassmannOptim [ACW12]. Graves [Car10, Bur10]. Gray [Has18]. gRc [HL07]. Great [The99]. Greek [Van18]. grImport [Mur09a]. groc [BdMM15]. grofit [KHLF^{+}10]. Grolemund [Joh20, Mil17b]. Group [Asz23, GC23, GG16, Har23, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Nun20, Was15, War23]. Group-Sequential [Was15]. Groups [Han13a, LX12, ZHL11]. groupTesting [War23]. Growth [KHLF^{+}10, Mir14, Wil14a]. Gruet [Dal98, Woo98]. gsbDesign [GG16]. GSEAplot [IRCA21]. GSODR [SHN17]. gss [Gu14]. GUI [AFGH22, Fel12]. Guide [AC04, Bai11, Bel19, BS09, Cha98, Dal98, Dia18, Doe10, EJ13, GC23, Grö16, Grö18a, Hew05, Iac15, Kal24a, Mat85, Sta93e, Mat16a, Mun14a, Plu01, Sel17, Sta93f, Sta97a, Sta97b, Sta97c, Sta98a, Sta98b, UC21, Woo98, ZIM09, Cot13, Cox05, GL16a, Gle16b, H^{+}96, HBPJ04, LŠ20, OS95, Shi16, XAG19, GHG19, Joh20, Lip21c]. Guidebook [Tak12]. guidelines [SOD^{+}16]. guides [Ree19]. Gunther [Boo10, Mil12]. Guyader [Unw13b]. GWAS [Tur18]. gWidgetsWWW [Ver12]. GWmodel [GLC^{+}15].

H [Asz23, Bai11, Ben21, Cho22a, Grö15a, Hor12c, Iac15, Kal24b, Kha16, Lee21, Len20, MN03, Nun20, Pol13, Sab19, Sch08, Sól10]. H. [Hub11, Sól10, Vis10, Wil11a]. H/e [Lee21]. H/e-book [Lee21]. H1 [The99]. H1-B [The99]. habitat [Cal06, MS23]. habitats [BSG20]. hackers [CW12]. Hadley [Mil17b]. Hadoop [FM18a]. HAFU [OS95]. Hall [AS24, Ben21, Die21, Esk21, Fus22, GC23, Gle16a, Grö16, Hac17, Has18, Hou07, Joh20, Kal24a, Kel10, Kha16, Kle17, Laz24a, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Liu18, Liu23, Lum08, Oli17, Orm17, Peb21, Put06, Rao14, Ree19, Rob17, Sab19, Sha23a, Sha23b, Sta05, Til96, Tu21, Vie11, Cho22a, Par15b, Pod21, Wan22]. Hall/CRC [Ben21, Die21, Esk21, Fus22, GC23, Gle16a, Grö16, Hac17, Has18, Hou07, Joh20, Kal24a, Kel10, Kha16, Kle17, Laz24a, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Liu18, Liu23, Lum08, Oli17, Orm17, Peb21, Put06, Ree19, Rob17, Sab19, Sta05, Tu21, Vie11, Cho22a, Par15b, Pod21, Wan22]. Hamann [Mye12a]. Hamilton [Ano99]. Hamiltonian [TT21]. Handbook [Ano06b, Ano06a, Dav95, DB18, Eve94, Eve02, EH06, Hil06, KD11, Kim95, Ko95a, LZ10, Lun07b, Man03, Pfa12, Sen14, Zie02a, All11a, Til96, Cha22, Dem18, Lip22, Piñ24, Rob12, Sab19, Vis10]. Hands [EKP^{+}11, Gro14a, Vin08, Bor16a, HMRS14, Yal10]. Hands-on [EKP^{+}11, Gro14a, Vin08, Bor16a, HMRS14, Yal10]. Hard [Das21, Tu21]. Hardback [Kal24a, Kal24b, Lip21a, Lip21c, Lu18, Mal21, Nun20, Oli17, Aji17, Cur18, Sha23a, Sha23b, Sta21]. Hardcover [Laz24a, Peb21, Zei16, Cap19, Liu18, Liu23, Pod18, Pod21, Rui17]. Harezlak [Num20]. harnessing [FVVL23]. Harring [Han13a]. harvesting [FAM^{+}20]. Hastie [Nor14]. Hazard [Gan15a]. hbk [AS24, Lip22, Liu23, Piñ24]. HDclassif [BBG12]. Health [Cap19, Gre22b, PD08, Buc09]. Heatmaps [BY18, SK17]. Hector [Ree19]. Heiberger [Fle11, Sco10b, Utl05]. Heidelberg [Har03]. held [MN69]. help [Alb15]. Helps [BCG18]. Henry [Has18, Hor12c]. heplots [FFM09]. Hero [EL09]. Hesketh [Daw03, Zie02b]. Hessian [Bra17]. Hessians [Bra14]. Hesterberg [Sco12]. Heterogeneity [GLC^{+}15]. Heterogeneous [BL14a, WLH^{+}18]. Heteroskedastic [Pir10]. Heteroskedasticity [SYC08]. HGLMMM [ML11]. hhsmm [ABS23]. Hidden [BHS00, HH19a, OH11, VS10, ZM09, ZML16, ABS23, BBN10, TNM17, All11b, Gut11, Kha17, Lu18, San11]. Hierarchical [Arm19, CGC14, DW17, FBC07, FC11, LH12, LH14, ML11, Mül13, Rec10, CKSLS18]. High [ALE10, BBG12, JKvT^{+}14, ND12, PSS^{+}17, dVSWAL17, Tem97, VG09, Cha21, Che23, ES14, Gav10, Lan18, SA20, TKM16]. High-Dimensional [ALE10, BBG12, JKvT^{+}14, ND12, VG09, Cha21, Che23, TKM16]. high-performance [ES14, Lan18]. High-Precision [dVSWAL17]. High-Throughput [PSS^{+}17]. higher [YWL02]. higher-order [YWL02]. highlightHTML [LeB18]. Hilbe [Grö16, Gol11]. histograms [GH18, GHN19]. Historical [SVCB18]. Histories [Sun15, Wil14b]. History [Beb13, Bec94, Iha98, Bro12]. HLMdiag [LH14]. hmmm [CGC14]. Hoboken [Bas18, Lip21a, Sel17]. Højsgaard [Han13b]. Holland [Utl05]. Holmes [Sha21c]. homals [dLM09a]. homogeneity [ATYK20]. Homology [FW18, WWDS18]. Homoscedasticity [JJJ14]. Hong [Han98]. Hong-Ye [Han98]. Hooker [Bur10, Car10]. Horton [Grö15a, Liu17, Liu18, Pol11c, Ste11]. Hotel [All86]. Hothorn [Ano06a, Che11, Dic12, Ric11, Pfa12, Unw12a, Yu12]. House [Mil15b, BL14b]. HPLC [RG07]. HPLC-DAD [RG07]. Hron [PG20]. hsmm [BBN10]. HTI [OS95]. HTML [Lei13]. htmlwidget [Coe18]. httk [PSS^{+}17]. http [Grö15a, Hel15, Hof15, How16b, Iac15, Mat15, Mat16a, Rus15, Zei16]. https [Gle16a, Grö16, Kha16]. Huber [Doe06]. Huet [AC04, Cox05, Dal98, Woo98]. huge [Tag23]. Hugen [Pod18]. Hull [PLRC10]. Human [LZ10, Vis10]. Humanities [Pav16]. humanleague [Smi18]. Hurricane [EJ13, Dia18]. Husson [Mai11, Unw13b]. Hybrid [CGC11, ABS23]. Hydro [OS95]. Hydro-Acoustic [OS95]. Hydrographic [FGG^{+}94]. Hydrological [Van18, VWDB16]. hydroscoper [Van18]. hyperspectral [CBHLG21]. Hypothesis [Kim20a, FFM09, Tae14].

I/O [NPP18]. Iacus [Cam09, Bur09, Luo09, Sto11]. Iain [Lu18]. iBluff [PL23]. ic.infer [Grö10]. icenReg [AB17a]. IDE [Rac12]. iDEMO [CP12]. identification [AGdSC20]. identify [OSH^{+}21]. idm [DMB18]. Ieno [Bai11, Oli10a, Sch09]. if [CM14, HJM08]. iheatmapr [SK17]. II [KWE^{+}17, Lei03, PFT^{+}12]. II/III [PFT^{+}12]. III [How16b, PFT^{+}12]. IISC [Ano99]. ijtiff [NPP18]. Illness [TGJ17]. Illness-Death [TGJ17]. Illustrating [Gan15a]. illustrations [BA97, Lip21b, Lip21c, Lip22, Den98, Ros00].

Image [BLY18, MCAP19]. ImageJ [NPP18]. Images [PT07, CBHLG21]. Imaging [CMS^{+}11, PT09, TW11, WS11]. imbalanced [GZ21a, GZ21b, ZGZvB19]. ImbTreeAUC [GZ21a]. ImbTreeEntropy [GZ21b]. iml [MCB18]. impact [BSS^{+}24, PSZ17]. Implement [BMGT15, XWHL15]. Implementation [Aiz12, AE21, BHS00, CHT98, Cha08b, CLL17, Här91, Hu09, IM15, LPLPD14, Lo 93, Lo 94, O’B08, PG15, Pfa08b, USHH18, WLK08, WsT18, YWL02, BPL09, Cho20, GV19, HFY22, JPH21, Kel10, KB11, Rob09, SOD^{+}16, SA20, YCLL20, YH14]. implemented [JV14]. Implementing [BK11, EN11, LSPvdL17, PK08, SR18, Wel18, BMOF17, PFT^{+}12]. Import [McL17, Mil17b, R D11c, VYD^{+}12, WG17]. Import/Export [R D11c]. Importance [GT10, RJH14]. Importing [Mur09a]. Improper [Dee99]. Improved [MZ08, dSJdSF14]. Improvement [Anh18]. ImpuR [RG07]. Impurity [RG07]. Imputation [CF08, KT16, SGHY11, vBGO11]. inch [Abr97]. Including [Gan15a]. Income [Esp15, Lon15]. Incomplete [OK14]. Incorporating [Sta08]. Incremental [DMB18]. Independence [Høj12, SF18]. Independencies [Mar06]. Independent [BDdM11, Rac12]. Index [R D11b, Sta95c, HBA19]. Indexed [Har10, Sav16]. Indicators [AT13]. Indices [SFS12, GSJC22]. Individual [TMCE22]. Induced [Mar06]. Industry [Bra03, Eas03, GF09, Har03, KM01]. Inequality [Esp15, Grö10, Lon15]. Infectious [JGM18, Höh07]. Inference [Asz23, BFRP13, BGH^{+}17, Bur09, Cam09, Grö10, GV12, HL07, Iac08, ILS11, IK20, KMC^{+}12, KNI16, Lip20b, Luo09, RJH14, SH17, Sch17c, SVMMRP17, Smi07, Sto11, TDRD13, TDRD15, Tro09, VW13, Ven10, BNB96, BNB07, BNB08, Fus22, MdN22, NWH21, SWCP20, Shi16, SMM^{+}22, SR17]. Inferences [IDE15, KF14]. inferference [SH17]. inferring [Amo21]. infinite [KB11]. infinite-dimensional [KB11]. Inflated [KN05, LC10, MAY23]. INFO [Ger94, Mat94]. Informatics [Guh07]. Information [BG18, CFSR15, Dro18, Sta08, Aji17, Fri16, Van18]. INFORMIX [MC97]. INFORMIX-Universal [MC97]. informR [MB15]. Infrastructure [FHM08, FK17]. INLA [BRF^{+}18]. innovation [BHH05]. Innovations [Pir10]. Inputs [GT10]. Installation [Sta93g, R D11d, Sta95d]. instructions [Bor16a]. instrumental [Zag18]. Intake [PZK^{+}12]. Intake epis food [PZK^{+}12]. Integer [Han06a, SS18b]. Integrate [GL16b]. Integrated [Ano10, CP12, Edd09a, Har07, Lüt11, RFKM12, CRP13, Pfa06, Pfa08a, SO23]. Integrating [SHR97]. Integration [EF11, Kuo03, RR11, SS18a, SS18b]. intelligence [HTWT23]. Intensity [RG07]. Interaction [QZLP21, YH14]. Interactive [BP17a, BC84b, BCAB07, Cha99, CS07, Gan15a, LB21, LLS15, NL12, SK17, Tem97, TD07, Ver12, VYD^{+}12, lGlRGL09, Dan18, Tan18, Eme08, Mai08, Sha21a, VM09a]. Interdependency [DK18]. Interface [All86, Arn17, AB12, Den16, Fle11, Fox05b, FC12, KSC^{+}00, LL10, MJR93, RFKM12, Sco10b, SLS^{+}12, UAK^{+}15, WLH^{+}18, Coe18, EK23, GMF18, HN09, Kor18, Ros07, SKW17, Van18]. Interfaces [HL09, VML12, Smi17, Smi18]. Interference [SH17]. Intermediate [HH15, Pri05b, Sha16, Utl05, Yal10, HH04, Vin08]. Internal [WRD23, Lan09a]. Internals [R D11e]. International [CB17, Che23, HI97]. Interpretable [MCB18]. interpretation [WPW15]. Interpreter [DC09, KMMV14, Run13]. interpreting [Alb15]. Interrupt [The99]. Interval [AB17a, FS10a, TGJ17, Wol94, Wol99, Kom09, FS10a]. Interval-Censored [TGJ17, Kom09]. Intervals [KPSH15, You10, FBB20]. Interventions [MJGM10]. intra [ATYK20]. intra-class [ATYK20]. Introducing [BKT14, Han05, Han06b, Han07b, Han20, LC10, LHR16, RC10, WMS17, SMWP20, Neu11, Rip11, Sco10a]. Introduction [All11b, AC07, Bac08, Boo10, Che22, Chi07, Cho22a, Day15, Edd09c, EB18, GR18, Goo05, Goo13, Gut11, HHB08, Har06, Hel15, Hof15, IDE15, Iri19, JMR14, Kha17, KF14, Ko95b, Kor22, Kus03, Lep14, Lu18, Mäc07, Mai06, Mar12, Sta93c, MPV12, Ng06, Nor14, O’B13, OHD17, OJMR09, R D11a, Ree19, Rob07, RB11, Rom07, San11, SM05, Sha23b, SMM^{+}22, Spe94, Sta07b, Sta93d, ST10, Thi14, Tro09, Unw12a, Ven10, Woo06, Yu12, ZML16, Ano03b, BP17b, CL11, Cra02, Cra05, Cra15, Dév09, EH11, Hec15, Hor09, JWHT13, Mil12, Nie14, Rup04, Saw09, Tav17, Tsa13, VS02, Ven04, Wün13, ZM09, Pol09, SS18a, Sha23b, Das21, DN17, Hil10, Laz11a, Lip20a, Mal09, Ng11, O’B09a, Pic09, Sco11]. Introductory [Ano06c, CM09, Dal02, Dal08, Gou05, McN22, Sch14, Smi06, Wan16, Hou07, LŠ20, Ver05, Ver14, Ano09a, Bat04, Bos10, Li11, Mai10, Nie11, San10b, San03, VM09b]. intsvy [CB17]. Invariant [IP08]. inventory [BSS^{+}24, OOSM18]. Inverse [LHS08, SP10, VSV09, vdWG11]. investigate [Wu24]. IP [KSC^{+}00]. ipw [vdWG11]. iqLearn [LLS15]. IRIC [ZGZvB19]. Irizarry [Doe06, Lip20a]. Irregular [PG15]. Irregular-Shaped [PG15]. Irrizarry [Das21]. IRT [Bat15, MH07, TMN16, Wee10]. IRT-Based [Wee10]. ISBN [AS24, Abr97, Aji17, Ano16, Asz23, Ben21, Cap19, Che23, Cox05, Cur18, Dal98, Die21, GC23, Gle16a, Grö15a, Grö16, Hac17, Har03, Har23, Has18, Hel15, Hof15, Hou07, How16b, Hu21, Iac15, Joh20, Kal24a, Kal24b, Kel10, Kha16, Kle17, Laz24a, Laz24b, Lee21, Len20, Lip20a, Lip20b, Lip21a, Lip21b, Lip21c, Lip22, Liu18, Liu23, Lu18, Lum08, Mai21, Mal21, Mat15, Mat16a, Num20, Nun20, Oli17, Orm17, Peb21, Pet02, Piñ24, Plu01, Pod18, Pod21, Put06, Rao14, Ree19, Rob17, Rui17, Rus15, Sab19, Sha21c, Sha23a, Sha23b, Sta21, Sta05, Til96, Tol23, Vie11, Zei16, Cho22a]. Island [Kim20b]. Ismay [Fus22, Lip20b]. isocir [BFRP13]. Isosurfaces [FT08]. Isotone [dLHM09]. Isotonic [BFRP13]. Isotopomer [MvSB^{+}07]. Isotropy [Wel18]. Issue [PLR^{+}16, FGZ14]. Italian [Bia94]. Italy [Bia94]. Item [Cha12, CGC11, DBZ^{+}11, HFY22, Joh07, MF14, MPM14, MTvdM15, Riz06, Sch17c, BBG14]. Item-Level [Sch17c]. Iterative [CGC11]. Iteratively [HB92]. Ivo [Cap19]. ivporbit [Zag18]. ix [Dal98, Sha23b].

J [Ano03b, Ano06c, Ano09b, Ano09d, Bur10, Car10, Dur14, Gal17, Grö15a, Hel15, Hly09, Hoe09, IDE15, Joh09, Joh20, Kim20b, Kum07, Kus03, Laz11b, Lig09, Liu17, Liu18, MN03, Ng06, Nor09, Oli07, Oli10b, Oli10a, Orm17, Pod18, Pol11c, Put06, Rob12, Rob05, Ros09, Sch09, Sta07a, Tol03, Unw11, Vie11, Wan06, Woo01]. J. [Fri11, Kum10, Mar11, Ros09, Wan06]. Jackknife [SP05]. Jacob [GC23, Kal24a]. JAGS [Den16]. James [Nor14]. January [Ree19]. Jaroslaw [Num20]. Java [Hec07, dL05b, Dan18, HLT09, Mur05, Nar05, SLS^{+}12, Urb09]. Java-Based [SLS^{+}12]. Java/R [HLT09]. Java/R-based [HLT09]. Javascript [CH17, Dan18]. JavaStat [HLT09]. Jean [Cha14a, Mai21, Tat18]. Jean-Baptiste [Tat18]. Jean-Michel [Cha14a, Mai21]. Jeff [Mye12a]. Jeffrey [Han13a]. Jégou [Unw13b]. Jens [O’B09b]. Jérôme [Mai11]. Jeremy [Unw13a]. ji [CCP^{+}11]. Jiahui [Ano03a, Bur07, Fis06, Fot07, McN04]. jian [CCP^{+}11]. Jianrong [Hu21]. Jim [Bau14, Cho09, Pol11a]. JM [Riz10]. JMbayes [Riz16]. jmcm [PP17]. job [Wil12]. Jocelyn [Kal24b, Sha21c]. Johannes [Che23]. John [Ano03b, Hel15, Hly09, Hou07, Iac15, Laz24a, Lip21a, Mil17a, Nor09, Plu01, Rob17, Smi06, Zie04, Cho22a, Kle17]. Joint [PP17, Riz10, Riz12, Riz16, Jon13]. Jolivet [Cox05, Dal98, AC04, Woo98]. Jonathan [Mai14a, Qia10a]. Jones [Ng11, O’B09a, Sco11]. Jong [Sta05]. JOP [KR13]. José [Zie01a]. Joseph [Gol11, Grö16, Kim20b]. Josse [Unw13b]. Journal [FL16]. Journalism [Kin20]. Joy [Oli17]. Juan [Wan22]. Judge [The99]. Judgment [Eva14, Arm14]. Juha [Esk21]. Julia [Dan18]. Julian [Gal17, Orm17, Put06, Rob05]. Julie [Lee21, Unw13b]. Julien [Mai09]. jump [GSB19]. Jun [Har23, Liu23]. Jupyter [Dan18]. JupyterLab [Dan18]. Jureä [Oli07]. Jurgen [AA12, Krä20]. Just [Bro14, The99, MBT^{+}20].

K. [Joh09]. K.-S. [Joh09]. Kalman [Tus11, Wan13]. kamila [FM18a]. Kao [Asz23]. Kao-Tai [Asz23]. kaphom [ATYK20]. Kaplan [GC23, Kal24a, Liu18]. kappa [ATYK20]. Karel [PG20]. Karian [Rob12]. Karl [Num13, Par15a]. kdecopula [Nag18]. Keen [Fri11, Vie11, Unw11]. Keith [Lip22, Piñ24]. Kelley [Sha21c]. Ken [Grö15a, Liu17, Pol11c, Sha21c]. Kendall [LAJJ18]. Kentucky [All86]. Kenward [HH14]. Keras [Arn17]. kerasR [Arn17]. kerdiest [dREP12]. Kernel [DHM11, Den98, Duo07, FHH17, HS18, KSHZ04, KF10, MP14, MPM14, Nag18, dREP12, Ros00, BA97, HHJC11, NWH21]. Kernel-Based [FHH17, KF10]. kernlab [KSHZ04]. KernSmoothIRT [MPM14]. Kevin [Unw11, Vie11]. KGode [NWH21]. Kim [Lip20b, Sta05, Fus22]. kinetic [HRC20]. Kirk [IDE15]. Klaus [Veh13].

Kleiber [Sco09]. Kleinman [Grö15a, Ste11, Liu17, Pol11c]. Klieber [Mye09]. Kloareg [Unw13b]. kml [GASA15]. kml3d [GASA15]. knitr [Vey14, Xie13b, McN14]. kNN [CF08]. Knowledge [Asz23, BSVT12, ÜS10, Wil11b, ZGS^{+}18]. Kolmogorov [ZS17]. Korosteleva [Sha23b]. Kraljic [BMOF17]. KraljicMatrix [BMOF17]. Krause [Ano03c, Ano06d, Bar02, Bro07, Har03, Mac98, Sis03, Zie98, Zie01c, Bra03, Eas03]. Krider [Sco13a]. Kriging [RGD12]. Kriging-Based [RGD12]. KRLS [FHH17]. ks [Duo07]. Kubinger [Veh13]. Kung [Qia10a]. Kung-Sik [Qia10a]. Kupresanin [Sco13b].

L [AA12, All11b, Ber10, Has18, IDE15, Len20, Lu18, Pod18, Sco13b]. L. [Gut11]. Lê [Mai11]. Label [Pap16]. label.switching [Pap16]. Labs [Che23]. LaBudde [Mar12, O’B13]. laeken [AT13]. Lag [Gas11, Mug10]. laGP [Gra16]. Lambda [Su07]. land [WRV23]. landings [KK22]. landsat [Gos11]. landscape [BSG20, MS23]. Langevin [RLWP16]. Langrock [Lu18]. Language [BC81, BW88, BCW88, BCW18, Cha69, Cha98, ESO16, IG96, KSC^{+}00, MHOV12, R D11f, R D11b, Ste00, KKEM15, LMY^{+}11, Röh00, RL15, SWHJ17, Plu01]. Lanham [Aji17]. Lappi [Esk21]. Large [CB17, CFHBK11, Gra16, VG09, FAM^{+}20, GDBK^{+}21, LBL^{+}21, SNR18]. Large-Scale [CB17, CFHBK11, Gra16, GDBK^{+}21, SNR18]. Latent [BPP17, Bea17, Lei04, LL11, Meu13, QZLP21, Riz06, WM14, BBG14]. LaTeX [Ruf09]. Lattice [Sar08, Kuh10, Nor08, Gro08]. Laura [Sco12]. Laurent [Unw13b]. Lauri [Esk21]. Lauritzen [Han13b]. lavaan [Ros12]. lavaan.survey [Obe14]. Law [HGG08]. lawn [CH17]. Lawson [Cho22a]. lawstat [HGG08]. laziness [GV19]. LC [KSS^{+}07]. LC/MS [KSS^{+}07]. lconnect [MS23]. LDheatmap [SBMG06]. ldr [AR14]. Leacky [Oli17]. leadership [Amo21]. Leakage [BG18]. Learning [AVS20, Arn17, Asz23, CFZ09, Cot13, Dan18, GZP14, Lan19, LLS15, MCB18, Nor14, Scu10, TT21, Tod22, ZUL14, Ano97, CBL^{+}19, CA18, CW12, HBZ09, JWHT13, KF10, Kim20b, KKL^{+}15, MGHR16, PLGHG^{+}23, ZMV^{+}18, AS24, Hel16, Lip21b, Rob22]. Least [FHH17, HB92, Len16, MW07, Alb19, Tu21]. least-cost [Alb19]. Least-Squares [Len16]. Lederer [Che23]. Legendre [Mye12b, O’B12b]. Legler [Lee21]. LEGO [Kur19]. Lessons [BCHR15]. Letter [MN03]. Letters [KSC^{+}00]. Level [Grö14b, KO06, Rec10, Sch17c, Tem97, HBNS14]. Level-Dependent [KO06]. leverage [Urb09]. Lévy [DC05, IM15]. Lexical [GI00]. Lexington [All86]. Lexis [CP11a, PC11]. lgcp [TDRD13]. Li [Laz24b, Mai14a]. libeemd [LHR16]. Library [CH17, LIL^{+}15, MJR93, Aji17, AGdSC20, Coe18, Fri16, ZGZvB19, Arn17]. libstable [dVSWAL17]. lie [CCP^{+}11]. Life [SVMMRP17, Sun15, Tei22, Wil14b]. life-testing [Tei22]. lifecontingencies [Spe13]. Ligges [Cur18]. Light [TFR16]. Likelihood [AR14, Cha10, CGS09, GvdL12, GV12, Joh07, KM08, MVS13, RC17, RMG12, Su07, Yua07, De 16, HT11, YS13]. Likelihood-Based [AR14, De 16]. Limited [Har23, Liu23]. Lin [Vis10]. linbin [WTB^{+}15]. Line [HL09, Hof11]. Linear [Arm19, AMW14, But05, CdM10, Eyl22, Far05, Far06, Far16, Fox03, FBF14, GB13, Gal17, Gas11, Grö10, Had10, HH14, Har01, KL14, Lee21, LFF17, LH14, MYK07, Mig10, ML11, O’B10, PPC09, Pet10, Pri05a, Rec10, Sta09, TV11, TM05, VSV09, Wan06, ZLHK02, dSJdSF14, BC11, BH94, Cox05, ES14, Er12, FFM09, GL07, Hel22, MPV12, Orm17, Ros09, Sha11, Wei14, Hos22b, Mai14b, Put06, Rob05, Tus05]. linguistic [Dév09]. link [Ger94, Mat94]. Linkage [SBMG06, TB17]. Linked [PMW^{+}15, SVC^{+}19]. Linking [Alb16, Wee10]. Linux [Ano99, Oom13, Wün13]. Lisp [Lub91, Nar05]. Lisp-Stat [Lub91]. Lists [LA19, WD18, ZGS^{+}18]. Literature [Joc14]. Littlefield [Aji17]. Live [SVCB18]. LLC [Asz23, Nun20, Pet02]. lmdme [FBF14]. lme4 [DBZ^{+}11]. lmer [DBZ^{+}11]. LMest [BPP17]. loa [LR15]. Løber [Unw13b]. Local [BKT14, CQZ^{+}10, Gra16, RC17]. localgauss [BKT14]. Locally [EN11]. Location [SR07, dCCM^{+}22]. Loftus [Tol23]. Log [CGS09, DM18, TDRD13, TDRD15, VP16]. Log-Binomial [DM18]. Log-Concave [CGS09]. Log-Gaussian [TDRD13, TDRD15]. log-symmetric [VP16]. logbin [DM18]. LogConcDEAD [CGS09]. Loggamma [AMYR16]. logging [BSS^{+}24]. LoggingLab [BSS^{+}24]. Logic [Ano99]. Logistic [CGC11, Grö16, Har01, HWY18, PCAS09]. Logit [FH09]. Loglinear [BR07]. Logrank [FS10a]. London [Nun20, Pet02]. Long [PC11, SVCB18, GMF18, Han13a]. Long-Term [PC11]. Longitudinal [BPP17, DLC06, GASA15, GCA12, Jon13, KK14, LSPvdL17, NGBK12, PP17, Riz10, Riz16, SP14a, SvdLN17, WsT18, Cho20, Riz12, SOD^{+}16, XLX^{+}19]. look [AB90]. lordif [CGC11]. Loss [LSPvdL17]. Loss-Based [LSPvdL17]. Lösungen [Kra97]. Loughin [Par15b]. Lovelace [Gle16a]. lp solve [But05]. LS2W [EN11]. lsmeans [Len16]. ltm [Riz06]. ltmle [LSPvdL17]. Luigi [Ros10b]. Luke [Mat16a].

M [AC04, Ano08, Bar02, Bro07, Fle11, Gol11, Grö16, IDE15, Kle17, Laz11b, Lig09, Lum08, Luo09, Mac98, Mar12, Oli10b, Oli10a, Par15b, Plu01, Sco10b, Sco12, Sis03, Sól10, Sta21, Sto11, Utl05, Ven10, Woo01, Woo98, Zie01a]. M. [Ano09d, Cam09, Few09, Hu09, Sch09]. M.-A [AC04, Woo98]. MacDonald [Gut11, Lu18, All11b]. Machine [CFZ09, CW12, Lan19, MCB18, Tod22, CBL^{+}19, HBZ09, KF10, KKL^{+}15, MGHR16, ZMV^{+}18, Asz23]. Machines [KMH06]. Macro [Wei12]. Made [HM16, GJK^{+}21]. Madison [MN69]. Maela [Unw13b]. Magnetic [CMS^{+}11, TW11, WS11]. Mailing [LA19, ZGS^{+}18]. Maillardet [Ng11, O’B09a, Sco11]. Maintenance [Sta93g, Sta95d]. Majorization [dLM09b]. Make [BCG18]. Making [Sco13a, Lan09a, MM22, PK12]. Malaccha [WRV23]. Mallows [ICL16]. MAMS [JPM19]. Man [EL09, Woo05]. Manage [McL17]. Management [BBGL17, Grö15a, HI97, HK15, KH10, Liu17, MJGM10, Pol11c, Ste11, Bos12, HK11, Tur12]. Managing [BK17, But08, Grö18a]. Mangion [Peb21, Pet19]. Manhattan [Tur18]. Manifold [ACW12]. manipulating [BLW12]. Manipulation [Bat08a, Bro10, Hic16, Sau10, Spe08, Wic08b, NPP17]. Mann [FO15]. Manning [Kim20b]. Manual [Sta92a, Sta92b, Sta93g, Sta93h, Sta93j, Mil00, Sta93i, Sta93k, Sta95d, Sta95e, KVCS98, Mil98, Mil02, R D04, Sta95f]. Map [LR15, TB17, Sho13]. Mapping [BS09, Doe10, GL15, RG96, TKM06]. Maps [GL16b, TP11, Ten18, Kub23]. Mara [Sta05]. Marc [Oli17]. March [All86]. Marchi [Bau14]. Marco [Tat18]. Marginal [CGC14, Joh07, Sta08, Ros08]. Marginalization [Mar06]. marginals [AD17]. Margins [KY10]. Maria [Bat08b, Len20, Ng09]. Marie [Cox05, Dal98]. Marie-Anne [Cox05, Dal98]. Marin [Cha14a]. marine [SO23]. Mark [Kel10, O’B08, Pod18]. Markdown [Joh20, Sha21b, XAG19, XDR21, BU15, LeB18, Mor18]. Marked [Har10]. marker [TPAM07]. Marketing [Mil15a, Rus15, Cha15a, RAM05]. Markov [All11b, Gut11, Kha17, Lu18, San11, ABS23, BPP17, BBN10, BHS00, FS10b, HD10, HH19a, KNI16, KSP15, Mar06, MQP11, OH11, RLWP16, Sch16, TNM17, VV16, VS10, WRV23, ZM09, ZML16]. Markov/semi [ABS23]. Markov/semi-Markov [ABS23]. Marries [Thi14]. Martin [Hor12c]. MaskDensity14 [LF15]. Mass [CS12, Kin20]. MATCH [Wie04]. Matching [IP08, KSS^{+}07, Lab12, RBB18, RL15, Sek11, Wie04, NWH21, Val09]. Math [Sta99]. Mathematical [CH11a, HBQ04, JGM18, Sta93e, Sta93f, Bia94, Sco12]. Mathematics [Spe13]. MathSoft [Ano99]. MATLAB [EK23, Mar07, RHG09, dVSWAL17, VFV13, Sch14, SS15, Bur10, Car10, Bow10]. Matloff [Edd18, IR12]. Matrices [Bra17, RG07, SD18]. Matrix [Fie15, FS10b, Kha16, KN03, Liu16b, WRV23, BMOF17]. Matt [Num20]. Matthias [PG20]. Matzner [Unw13b]. Matzner-Løber [Unw13b]. Max [Bau14]. Maximization [SKD22]. Maximum [CGS09, GvdL12, Joh07, KM08, RV20, Su07, VdL09, Yua07, HT11]. maxLik [HT11]. mboost [De 16, HMRS14]. MCAR [JJJ14]. McDonnell [Rus15]. McElreath [Mal21, Rui17]. mcglm [Bon18]. MCMC [Den16, FS10b, GSB19, Had10, Pap16, Riz16, Ros07, Smi07]. MCMC4Extremes [MdN22]. MCMCglmm [Had10]. MCMCpack [MQP11]. McNulty [Lip22, Piñ24]. MCPMod [BPB09]. MD [Aji17]. mdscore [dSJdSF14]. Me [Sta95b]. Mean [PP17]. Mean-Covariance [PP17]. Meaning [Die21, Ger20]. Means [KSBZ16, Len16, TMCE22, PH16]. Measure [MH15]. Measurement [DB18, HXY12, PZK^{+}12, Sab19, WW11, HFY22, SA15, ZY19]. Measurements [MvS07, SADU23]. Measures [Rec10, PCAS09]. Measuring [LA19]. meboot [VdL09]. Mechanisms [BLM^{+}15]. mechanized [BDT18]. median [KK21]. Mediation [Laz24b, TYH^{+}14, YL17, TYH^{+}14]. Mediators [YL17]. Medical [Daw03, Zie02b, ERH01]. Medicine [MN03]. Medline [RL15]. Medline/PubMed [RL15]. Meesters [Bai11]. Meet [GHG19, HM11]. meets [HBZ09]. mefa [Sól09]. Mehmet [Tu21]. Mehmetoglu [Tu21]. Mehtätalo [Esk21]. Meissner [Sel17, Iac15]. Melvin [Ano03c, Ano06d, Zie98, Zie01c]. Membership [SP14a]. memory [KKL^{+}15]. Menu [BP12]. Mersmann [Cur18]. Message [The99]. Meta [Bur12, Cho22b, CO16, GV12, LZHC17, LCSC14, MF15, TMCE22, Ver18, Vie10, WF12, CP13, HKS08, Lan09b, Pal15b, Par15a, SCK^{+}23, How16a]. Meta-Analyses [TMCE22, Vie10, SCK^{+}23]. Meta-Analysis [Bur12, Cho22b, CO16, GV12, LZHC17, LCSC14, MF15, Ver18, CP13, HKS08, Pal15b, Par15a, How16a]. meta-data [Lan09b]. Meta-Statistics [WF12]. meta.shrinkage [TMCE22]. Metabolic [MvSB^{+}07]. Metadata [Boe17]. metafor [Vie10]. metaLik [GV12]. Metamodeling [RGD12]. metaphor [SGPSzC22]. Metcalfe [Li11, Mai10]. Meteorological [Van18]. Meteorology [SPPP17]. Method [CC08a, Cha10, HL09, TKM16, Yua07, Asq14, FAM^{+}20, FBB20, MÇSD14]. Methodologies [Asz23]. Methodologists [WDT^{+}12]. Methods [ADH11, Buc09, BL11, DMB18, Den16, DMS14, FS10b, Had10, HH14, Han13a, Har06, Hub11, KSHZ04, Kha18, KG17, KR10, Lan14b, Laz24b, Len09, Mar12, Mor09, Nas14, Nas08, Neu11, O’B09c, O’B13, Oja10, Oli07, OCRC14, PD08, Pol09, Rao14, Rip11, Rob19, Sch17a, Sco10a, Su07, Tyl07, ÜS10, Wee10, Xie13a, Zie00, dL06, dL09c, dLM09a, dLHM09, Car16, CCKT83, CL11, DH97, Goo05, Goo13, H^{+}09, How17, Hu21, JP06, Kut13, Oli17, RC10, Rom07, Sel98, Tsa14a, YS13, ZHL11, Bia94]. Methylated [WHK21]. methylation [Cha21]. Metric [IP08]. Metrics [PSZ17, GSJC22]. metrix [CC23]. Meyer [Agr16]. mFLICA [Amo21]. MFSAS [CC11]. mgpr [VRP23]. mhsmm [OH11]. mi [SGHY11]. mice [vBGO11]. Michael [Agr16, Ano03b, Dur14, Hel15, Kus03, Ng06, O’B13, Sta07a]. Michel [Cha14a, Mai09, Mai21]. MICR [KP23]. Micro [TKM15]. Micro-Data [TKM15]. Microarray [FO15]. Microarrays [Ano12b, Dră12]. Microdata [Kav15]. Microeconometrics [AS24, Lip21b]. micromap [PMW^{+}15]. Micromaps [PMW^{+}15]. MicroPEM [SVM^{+}17]. microscopes [CBHLG21]. Microsimulation [Gle16a]. Microsoft [Pol09, Tak12]. Microstate [Ano99]. microsynthesis [Smi18]. Microxanox [KP23]. midasr [GKZ16]. migrating [GJK^{+}21]. Mike [Sen14]. Miles [Unw13a]. Militino [Bat08b, Ng09]. Millard [Bra03, Eas03, Har03, Law02, Lum01, Pet02, Zie99]. mimR [Høj04]. Minimax [Wol99]. minimization [JPH21, PS22]. Minimum [LSPvdL17, RV20]. Mining [Bas18, Cic15, CFZ09, FHM08, Fuj17, Mun14a, Sau12, Sco13a, SR16, TFH12, GF09, KF10, Mwi13a, PK12, Wil11b, YÖYB23, Sel17, Iac15]. Minitab [GL14]. minPtest [HBNS14]. mipfp [BS18]. MiRNA [CPD^{+}20, CPD^{+}20]. MiRNA-QC-and-Diagnosis [CPD^{+}20]. mirt [Cha12]. misclassification [ZY19]. Mises [HG14, AFM22]. Missing [JJJ14, Har19]. missingness [XLX^{+}19]. MissMech [JJJ14]. MitISEM [BGH^{+}17]. Mix [SH17]. mixAK [KK14]. Mixed [CLL17, Dem13, DG22, Far06, Far16, FM18a, Gal17, GKZ16, GKZ12, Had10, HH14, HSG12, Mai14b, MF15, PB00, RV20, Sch11, TV11, Wan06, Wee10, ZIW^{+}09, BC11, GB13, Orm17, PW18, Ros09, ZY19, Chr09, Oli10a, Sch09, Woo01, Zie01a]. Mixed-Effects [Mai14b, PB00, GB13, Woo01, Zie01a]. Mixed-Format [Wee10]. Mixed-Type [FM18a]. mixexp [LW16]. Mixing [Lei02]. MixMAP [MF15]. Mixmod [LIL^{+}15]. MixSim [MCM12]. mixtools [BCHY09]. Mixture [AHvD09, BCHY09, BSVT12, BMGT15, HH19a, HS18, KG17, LW16, LM18, Lei04, LHA^{+}15, Sav16, Su07, AKL^{+}21, MC18]. Mixtures [HG14, ML13, PMM18, Dra23, GL07, Kom09]. MLDS [KM08]. MLwiN [ZPC^{+}16]. mma [YL17]. mmeta [LCSC14]. MNP [IV05]. Mode [Meu13, LHR16, MÇSD14]. Model [ALE10, AMYR16, AWBM18, BBG12, BCMR19, CdM10, Den16, Dis23, DM18, Far06, Far16, Gal17, HD10, HMRS14, HMS16, HSG12, HWY18, Jal19, LIL^{+}15, Lei13, LC10, MRC15, Mar06, Mel16, Mil17b, Mil15b, PZK^{+}12, PN97, QZLP21, Sch11, SR18, Sta08, TMW18, Wan06, ZC10, ZF15, BL14b, Cad13, De 16, IM15, KB11, LB12b, OOSM18, OS95, Orm17, Ros09, SMHBR06, WG17, XLX^{+}19, Yan95, Zag18, IV05]. Model-Based [AWBM18, BBG12, LIL^{+}15, Mel16, SR18, BCMR19, HMRS14, De 16]. Modeling [Agr16, Asz23, AB12, BG18, BS15, BSVT12, BS13, Car13, DCB^{+}17, FM16, Gra16, Grö15b, Har01, HD12, Hil14, HPCS14, HS18, Jac16, JHQ^{+}11, JGM18, KY10, KC14b, Laz24a, Lee13, LRN18a, Lip22, Lyu21, MH07, MF15, MvS07, PP17, PR07, Piñ24, RLWP16, Riz06, Ros12, SKS15, SYC08, SS15, VKM18, VL21, ZW03, Ziv05, BRF^{+}18, Fin14, Hos22c, HRC20, Kut13, MC18, Mil15a, MZ18, Par06b, RHW21, Sah22, Sha21c, Sha23a, Tu21, Fis06, Ano03a, Bur07, Fot07, McN04, SCD07, Tol03]. Modelling [Car17, Har10, HC05, Høj04, Riz10, SP10, TvEK20, VWDB16, AFHD09, AKL^{+}21, LSM20, MT20, Pfa13, TN22, Kum07]. Models [All11b, AB17a, AC07, ABC19, Arm19, AMW14, BR07, BPP17, BCHY09, BVE^{+}15, Bol08, Bon18, BKL05, BMGT15, BHS00, Bür17, Bur12, CPA20, CdM10, CG15, CP11a, CH93, CP12, Chi07, CKY14, Chr09, CGC14, CLL17, DBZ^{+}11, Den16, DLC06, Eva14, Eyl22, Far06, Far16, FDB12, Few09, Fie12, FBC07, Fir05, FC11, Fox03, FH09, FBF14, Gan15a, Gas11, GRK^{+}16, GKZ16, GLC^{+}15, GR18, Gra07, GT10, GH11, Gu14, Gut11, Had10, HH14, Han13b, Har01, HXY12, HH19a, Hoh18, HL07, Hos22b, ICL16, Jac11, Joh07, Jon07, Jon13, KMC^{+}12, KN05, KL14, Kha17, KR10, KV13, KSP15, Kuh08, Lee21, Lei04, LFF17, LHA^{+}15, LH14, Lu18, Mai06, Mai14b, MH07, Mal13, Mig10, MP12, ML11, MTvdM15, MBR11, MRL12, NAA17, O’B10, OH11, OHD17]. Models [Obe14, Oli10a, PPGD15, Pen03, PPC09, Pet10, PP11, Pfa08b, Pir10, Pri05a, RV20, Rec10, Riz16, Rob05, Rob07, RMG12, SF18, San11, Sar16, SP14a, Sch11, Sch09, Sho13, SM21, Sta07b, SR07, SM07, TV11, TM05, TR14, TGJ17, TF12, Tus05, UAK^{+}15, VS10, WW11, Wan13, WMR16, Woo01, Woo06, WsT18, XWHL15, ZLHK02, ZKJ08, Zie01a, ZML16, ZIW^{+}09, dSJdSF14, dWFP11, ABS23, Arm14, AFM22, BBG14, BC11, BAS12, BBN10, CH92, Cho20, CRP13, Dee99, Dem13, ELS17, Er12, Far05, FFM09, GB13, Hel22, HF07, HEL12, IESdF18, MC18, NB13, PW18, PB00, Put06, Riz12, Sha11, Sta09, TNM17, VP16, WM18, YH14, ZM09, Gal17, Wan06, Orm17, Ros09]. Moderation [Laz24b]. Modern [BKH17, Dia18, EJ13, Gou10, HH18, HH19b, Lip20b, Mar11, Sel98, Sha22b, She09, VR94, VR97, VR99, VR00a, VR02, Fus22, SM05, Abr97, Arg98, Dow17, Fer02, Har23, Jam96, KK99, Liu18, Liu23, Rui16, Zie00, Zie01b].

ModernDive [IK20]. Modes [Lei10]. modest [Lan09a]. modification [SKD22]. Modified [Har19, KK21]. Modifying [MB15]. Modular [Han06b]. MODULE [Lan95]. Mohamed [Lum08, Sta21]. Mohammad [Lum08]. Mokken [vdA07, vdA12]. Molecular [JKvT^{+}14, DJS^{+}18]. Molenberghs [Oli17]. Moments [Cha10, Phi10, Asq14]. Momocs [BPGC14]. Mónica [Ben21, Sán19, Rie19]. monitoring [Höh07]. Monogan [How16b]. Monographs [Lu18]. Montage [FGG^{+}94]. Monte [Neu11, Rip11, Sco10a, TT21, CGC11, DK18, LT16, MQP11, RC10, SP10]. Moodle [ZUL14]. Morgane [Gle16a]. Morphometrics [Cla08, Bow09, Kum10, Mai09]. Morse [GP12]. Mortality [MP14, Mug10, VKM18]. MortalitySmooth [Cam12]. Most [BCG18]. Motions [DC05]. Moulines [Car16, Rao14]. Moult [EBO^{+}13, EBO^{+}13]. Mount [Mil17a]. movecost [Alb19]. Moving [LRN18b, CRP13]. movMF [HG14]. MPCI [SFS12]. MPI [Röh00]. mplot [TMW18]. MPS [TN22]. MRI [Fer11, TP11]. MS [KSS^{+}07]. msm [Jac11]. msr [GP12]. msSurv [FDB12]. MST [CSNF18]. mstate [dWFP11]. MSY [Cad13]. Muchene [Oli17]. Muenchen [Ano09c, Gol11, Mai12, Ree09, Sau11, Voe09]. multgee [Tou15]. Multi [CP11a, Had10, Jac11, JPM19, KSP15, LQC^{+}12, MvS07, Pen03, Tem97, dWFP11, MUM16]. Multi-Arm [JPM19]. Multi-dimensional [Pen03]. Multi-Regional [LQC^{+}12]. Multi-Response [Had10]. Multi-Stage [JPM19, MUM16]. Multi-State [CP11a, Jac11, KSP15, dWFP11]. Multi-way [MvS07]. Multiblock [BD18]. Multiclass [CC08a]. Multicore [Hof11, JPOJ12]. Multidimensional [BS18, Cha12, LBW18, dLM09b, BBG14]. Multifractional [DC05]. MultiLCIRT [BBG14]. Multilevel [Bür17, CC11, Eyl22, Fin14, Grö15b, Lee21, Sha21c]. Multinomial [CGC14, FH09, Tou15, IV05]. multiPIM [RJH14]. Multiple [Bon18, BHW11, CFHBK11, Eyl22, FM19, HWY18, JM15, KPSH15, KR13, Lee21, OH11, SP14a, SGHY11, WGSL12, YL17, ZC10, Beh05, Las97, Che11, Dic12, Ric11]. Multiple-Objective [HWY18]. Multiple-Table [WGSL12]. Multiplication [LBW18]. multiplied [LF15]. Multiresolution [SCS13]. Multiscale [WTB^{+}15]. MultispeQ [SADU23]. Multistate [FDB12, Mal13, Wil14b, BAS12, Sun15]. multitable [WGSL12]. Multivariable [SMHBR06]. Multivariate [AWBM18, Ano06c, ARC04, BS18, BC11, CSNF18, CO16, CGS09, Duo07, Er12, Eva11, FDGD16, FBC07, FBF14, Gro08, JM15, JJJ14, KY10, KK14, Kuh10, LJH08, LM03, LBC^{+}16, Lun06, LCSC14, Mai11, MdL10, ML13, Nor08, Oja10, Phi10, PMM18, SFS12, SMM^{+}15, Sek11, Sha20, TDRD15, TD07, Tsa14b, Unw12a, VW13, WCHB11, Yu12, Zei16, dL05a, AD17, BFA14, Beh12, Eve05, EH11, FFM09, GS19, Hos22a, HLP11, JV14, Kle09, Kom09, Sar08, Sha22c, SA20, VRP23, vBGO11, Hub11, Mül16]. Multiway [Lei10]. Munzert [Iac15, Sel17]. Murdoch [Hly09, Nor09]. Murrell [Ree19, Gun06, Unw12b]. Murtagh [Hec07]. Mutable [Wic11a]. MXM [LAF^{+}17]. MySQL [Wün13].

N [Abr97, Arg98, Bai11, Chi07, Fer02, Jam96, Law02, Liu16b, Oli10a, Rob07, Sch09, Sta07b, Zie01b]. N. [Ste11]. nacopula [HM11]. Nagaraj [Lum01, Pet02]. names [ABEY18, Xie22]. Narrow [Mil15b]. Nash [CI17]. Nason [Mor09, BL11, O’B09c]. Nathan [O’B14]. National [RG96, Van18, RVHG23]. Native [Wei12]. Natural [MC97, Hos22a]. NbClust [CGBN14]. NCC [KHJ^{+}23]. NCIFD [WRD23]. Near [RBB18, Was15]. Near-Far [RBB18]. Near-Optimal [Was15]. nearfar [RBB18]. necklaces [DG23]. Neerchal [Pet02, Law02, Lum01]. Negative [LA19]. Nested [HM11, WD18]. NetLogo [Thi14]. Network [BB12, LZHC17, Mat16a, MCSGBSA20, Nor15, AFGH22, KC14a, MCSAGB20, Hor12b, But08]. Networks [BB12, GW18, HPCS14, Høj12, JGM18, MvSB^{+}07, PLGHG^{+}23, Scu10, SD15, Tat18, VSS^{+}17, WTB^{+}15, Cho15, SWCP20]. Neural [BB12, SWCP20]. neuRosim [WDM^{+}11]. Neutral [Han07b]. Neuwirth [Fle11, Sco10b]. Newdistns [NR16]. News [The99]. Newton [MHJS16]. nhorton [Grö15a]. nhorton/r2/ [Grö15a]. NHPoisson [CAA15]. Nicholas [Grö15a, Liu17, Liu18, Pol11c]. Nick [Kha16]. Nicolas [Unw13b]. NIfTI [WST11]. Nina [Mil17a]. NJ [Lip21a]. nls2 [BH94]. No [Abr97, Ano16, Cox05, The99, Dal98, Har03, Hou07, Kel10, Lum08, Orm17, Pet02, Plu01, Put06, Rob17, Sta05, Vie11]. Noel [Peb21, Pet19]. noise [LF15]. Nolen [Oli17]. nomisweb [Smi17]. Nomogram [CG15]. Non [BH94, Dev86, Gas11, SP14a, BSG20, Cad13, Cox05, Fog17, KHJ^{+}23]. non-concurrent [KHJ^{+}23]. Non-Linear [Gas11, BH94, Cox05]. Non-Parametric [SP14a, Cad13]. non-target-species [BSG20]. Non-uniform [Dev86, Fog17]. Nonhomogeneous [CAA15]. Nonlinear [AC04, CLL17, Dal98, DMS14, Gan15a, Gra07, Hew05, PZK^{+}12, RS08, Rit09, VG09, Woo98, AGdSC20, H^{+}96, HBPJ04, Rao14, Ano09b, Car16, O’B09b, dL09a]. Nonnegative [FDGD16]. Nonparametric [AM14, Can04, CQZ^{+}10, Con03, DLN17, Far06, Far16, FDB12, Gal17, Har19, Hub11, JM15, JHQ^{+}11, KK15, KPSH15, MVS13, Mül16, NGBK12, Oja10, OCRC14, PP18, dREP12, Sav16, SVMMRP17, Sta08, Wan06, Wel18, WsT18, AKL^{+}21, HHJC14, Orm17, Ros09, Cho20]. NonParRolCor [PMLM23].

Nonstationary [Gra07]. nopp [CI17]. Normal [AD15, Phi10, PMM18, AD17, Kom09, MC18]. Normality [JJJ14, JV14]. Norman [IR12, Edd18]. Northcon [IEE93, IEE94]. Northcon/93 [IEE93]. Northcon/94 [IEE94]. Northorn [Ano06b]. Notes [R D11a, Mil92, VS02, Ven04]. November [Ree19]. np [Har08]. nparcomp [KPSH15]. nparLD [NGBK12]. npbr [DLN17]. NPCirc [OCRC14]. npregfast [SVMMRP17]. npRmpi [HHJC11]. nse [AB17b]. Nucleotide [SBMG06]. Null [ZFZ10]. Number [CGBN14, SIR^{+}11, Fog17, Kom09, Lu18]. Numbers [Kal24a, CHB14, GC23]. Numerical [AB17b, Blo14, Kha18, Kos15, Kuo03, Lud13, Rob19, SS18a, SS18b, Var14, How17, Mye12b, O’B12b]. NumPy [EW16]. NUOPT [Ano99, SM05]. Nutshell [Adl12, Adl10, Edd11, Leo10]. NW [Sha21c, Sha23a]. NY [Kim20b]. Nyhuis [Iac15, Sel17].

O [Car10, Cho20, Cur18, Sco11, NPP18]. obial [KP23]. Object [ATF10, Cal95, Cha14b, Lo 93, Lo 94, MC97, PR07]. Object-Oriented [ATF10, Cal95, Cha14b, Lo 93, MC97, PR07]. Objective [HWY18, VG09, Urb09]. Objects [CP11a, Wic11a]. Observation [OH11, SD18]. Observational [Ros10a]. Observations [TGJ17, VV16]. Observed [Alb16, KNI16, BI13]. Observed-Score [Alb16]. obtained [NPR13]. occasion [DO94]. Occupancy [MBGK18]. occupy [KP23]. Occurrence [FC11, GVM16]. October [IEE93, IEE94]. Odds [FH09, PCAS09]. Office [Pol09]. ofw [CC08a]. Ohri [Mat15]. Oja [Hub11]. Okan [Sab19]. OLAT [ZUL14]. Old [AB90]. Olga [Sha23b]. Olsen [Sis03]. Olson [Ano03c, Ano06d, Bar02, Bro07, Mac98, Zie98, Zie01c]. Olympia [HI97]. Omic [Sha21d, Wan22]. Oncology [WS11]. One [RV20]. One- [RV20]. OneArmPhaseTwoStudy [KWE^{+}17]. Online [Din06, XYC22]. onlineBcp [XYC22]. ontologies [RB11]. oop [ZMS21]. Open [The99, HBBP18, HBZ09, TRM16b, Aji17, Fri16, PL23, SA15]. Open-source [HBZ09, PL23, SA15]. OpenCPU [Kor18]. Opening [SGHY11]. OpenML [CBL^{+}19]. OpenMP [JPOJ12]. OpenMP-style [JPOJ12]. Operating [GG16]. operations [Tie09]. OptGS [Was15]. optim [Var14]. optima [dCCM^{+}22]. Optimal [CI17, GZP14, HWY18, LRRÁCSGS14, MdL10, RPVG11, SMWP20, Was15, You12, dLM09a, RK20, AA12, Grö11]. OptimalCutpoints [LRRÁCSGS14]. OptimaRegion [dCCM^{+}22]. Optimization [ACW12, Bar14, Bra14, Dre19, GT10, KM14, KR13, MS11, MAG^{+}11, Mul14, NV11, Nas14, PSM^{+}11, RGD12, Rui16, Sek11, Var14, Wol94, dLHM09, KKEM15, Pfa13, SM05]. Optimizing [SWHJ17, VG09, MUM16]. Optimum [PJSPC17]. optimx [NV11]. OptSig [Kim20a]. Order [HHB08, YWL02]. Ordering [BPSS09, Ber10, Ros10b]. Ordinal [AD15, Arc10, BP12, CGC11, GSD12, LX12, Beh04, PCAS09]. Ordinary [Sch14, Wu24]. OrdNor [AD15]. Oregon [IEE93]. oreo [LBL^{+}21]. organisation [Dir18]. Orientation [Mur03]. Oriented [ATF10, Cal95, Cha14b, Lo 93, Lo 94, MC97, PR07]. Orientlib [Mur03]. origami [CH18]. Orlando [Ano95a]. Ornstein [BI13]. Orthogonal [ADN15, BdMM15]. oscillatory [LBL^{+}21]. osDesign [HSL11]. Other [BML19, HM16, RG07]. Outcomes [TM15, LB12b, PFT^{+}12]. outliers [FBNRG21]. Outline [BPGC14]. Output [Han05, Lei13, Smi07, SVM^{+}17]. Outputs [MRC15, Pap16]. ová [Oli07]. overflow [ZGS^{+}18]. overhead [WPW15]. Overview [Cha95]. Owen [Ng11, O’B09a, Ree19]. Owen-Jones [O’B09a]. Ox [CN97]. Oxford [Ree19]. OXic [KP23].

P [AS24, Bra03, BL11, Cha14a, Eas03, Gle16a, Grö16, Har03, Hel15, IDE15, Joh20, Law02, Li11, Lip21b, Lum01, Mai10, Mat15, Mat16a, Mye12a, Neu11, Num20, O’B09c, Pet02, Rip11, Rus15, Sco10a, Zie99, Til96, ACG^{+}16, Cam12]. P-Splines [Cam12, ACG^{+}16]. P. [Dic12, Mig10, Mor09]. Package [ADH11, ALE10, ACW12, AR14, AMYR16, AVS20, Alb16, AGG13, ATF10, AT13, AL16, AD15, AWBM18, Arc10, AHvD09, ABC19, AMW14, AM14, BG18, BT05, BDMP15, BK11, Bar14, BFRP13, BY18, BS18, BPP17, BGH^{+}17, Bat15, Bea17, BBGL17, BCHY09, BKT14, BBG12, BSVT12, BK17, BdMM15, BVE^{+}15, BN10, Bon18, BDdM11, BPB09, BD18, BCAB07, Bra15, Bra14, Bra17, BS13, BPDD08, Bür17, Bur12, But08, CdM10, CSNF18, Cam12, CC08a, CB17, CFHBK11, CAA15, Cha12, CNA16, CQZ^{+}10, CGBN14, CKSLS12, CC11, CKY14, CGC11, CGC14, CFSR15, CO16, CF08, CGS09, CJM06, CI17, DMB18, DLN17, DBZ^{+}11, DC05, DMD15, Den16, DW17, DM18, DK18, DPSH18, DGP08, EE07, EBO^{+}13, FS10a, FBdlF12, FDB12, Fer11, FDGD16, FM19, Fie12, FBC07, FO15, FC11]. Package [FH09, FC12, FBB20, FM08, FKP17, FGEMI12, FS10b, GR16, GRD13, GSD12, Gan15a, Gas11, GGC^{+}15, GRK^{+}16, GP12, GG16, GKZ16, GK16, GKD14, Gio09, GGK10, GZ11, GLC^{+}15, GCA12, GFS14, GDMB08, GFC12, Gos11, Gra07, GT10, GvdMW09, GW18, Grö10, Grö14b, Grö18b, GvdL12, GH11, Gu14, GV12, Had10, HPWdL15, HHB08, HH14, HD18, HSL11, Han06b, Han07b, HBBP18, Har10, HD12, HXY12, HM18, HP09, HH19a, HH07, HPCS14, HM11, HMS16, Hoh18, Høj04, HHY05, HL07, Høj12, HMR^{+}13, HG14, HSG12, HGG08, HK08, HWY18, IP08, IDE15, IV05, ILS11, IRCA21, ICL16, Jac11, JPM19, Jal19, JM15, JJJ14, JKvT^{+}14, JGM18, JD15, Jon07, KMC^{+}12, KSHZ04, KL14, Kas16, Kav15, KK15, KHR21, Kie08, KWE^{+}17, KE14, KO06, Kim20a, KNI16]. Package [KF14, KN03, KG17, KR10, KY10, KK14, KSBZ16, KPSH15, KV13, KT16, KSP15, Kuh08, KR13, LAF^{+}17, LRGTA12, LJH08, LIL^{+}15, Lee13, LM18, Lee18, LRN18a, Lei10, Lei03, LHS08, LSPvdL17, Len16, LBW18, LM14b, LFF17, LZHC17, LL11, LBC^{+}16, LC10, LHA^{+}15, LRRÁCSGS14, dUJ13, LX12, LCSC14, MRC15, MR12, MF14, MHJS16, MH07, MdL10, Mar06, MH15, MBGK18, MF15, MP14, MPM14, ML13, MYK07, MZ08, MS11, MCM12, Mel16, MJGM10, Meu13, MW07, MVS13, ML11, MTvdM15, MBR11, MV14, MdUÁC10, MN14, MvSB^{+}07, MvS07, MAG^{+}11, Mur03, Mur09a, MG09, NR16, Nag18, ND12, NGBK12, NAA17, OH11, OHD17, Obe14, OCRC14, OL17, Oom13, OK14, PPGD15, PP17, Pan18, Pap16, PLRC10, PMW^{+}15, PSS^{+}17, PG15, PP18, Pet10, Pfa08b, PU13, PT07, PT09, PMM18]. Package [dREP12, RK20, RBB18, RLWP16, RC17, RJH14, Riz06, Riz10, Riz16, RR11, RMG12, Ros12, Rov17, SF18, SFS12, Sar16, SIR^{+}11, SZ11, SMM^{+}15, Sch16, SSH^{+}20, SlRC16, Sch17c, Scr13, SR18, Scu10, SVMMRP17, Sho13, SS19a, SKS15, SM21, Smi07, SYC08, SLS^{+}12, SP10, SPS10, SvdLN17, Sól09, SD18, Spe13, SM07, SLG05, SSV14, TP11, TMCE22, TM15, TMW18, TV11, TDRD13, TB17, TKM15, TMKD17, TMN16, Thi14, TFR16, TYH^{+}14, Tou15, TGJ17, TF12, Tyn16, ÜKD09, ÜS10, USHH18, VW13, VG09, Ver18, VKVC15, Vie10, VV16, VKM18, VdL09, VYD^{+}12, VS10, WGSL12, WMS17, WW11, Wan11, Wan13, WLH^{+}18, WHK21, Was15, Wee10, WF12, Wel18, WTB^{+}15, WDM^{+}11, WRD23, WM14, WCHB11, Wie04, WMR16, XMW10, XWHL15, Xie13a, YEL18, You10, YL17]. Package [Yua07, ZF15, ZPC^{+}16, ZQS16, ZP13, dSJdSF14, dLM09a, dLM09c, dWFP11, vdWG11, ATYK20, Alb15, Alb19, AA20, AD17, ABS23, Amo21, Ano13, ATCA20, ABEY18, AFM22, BSS^{+}24, BF17, BBG14, BMOF17, BI13, BdSvE23, BSG20, BBN10, CC23, Cal06, CBGGV17, Car17, CBL^{+}19, CPD^{+}20, CCJMR16, CKSLS18, CHB14, CLK21, CBHLG21, CPB^{+}23, DS20, DCMCPF20, Dir18, Dra23, DTDd19, FBNRG21, FFM09, FAM^{+}20, FVVL23, FM18b, GZ21a, GZ21b, Gas18, GSJC22, GSB19, GMF18, Gle16b, GEV18, GS19, Han20, HCSH15, HBA19, Hel22, HT11, HHJC11, HHJC14, HTWT23, HMRS14, Höh07, HF07, HRC20, HS18, IM15, IESdF18, JPH21, KF17, KOC21, Kom09, KHJ^{+}23, KP23, Lan18, Lan09b, LSM20, Lau18, LM03, LS16, LF15, LPR21, LLM^{+}20, LBL^{+}21, LHR16, MdN22, MM22, MS23, MUM16]. package [MT20, MCB18, MCSAGB20, MCSGBSA20, MN17, MZ18, MAY23, NPR13, NB13, NKHZ21, NWH21, NPP17, NPP18, OSH^{+}21, PL23, PCAS09, PFT^{+}12, PDH16, PSM^{+}11, PLGHG^{+}23, PMLM23, R D04, RRSPTR12, RRSPTR14, SVM^{+}17, SGPSzC22, Sek11, SAR11, SWCP20, SKD22, SMWP20, Smi18, SO23, SADU23, SR17, Tag23, Tan18, Tei22, TN22, TKM16, TLH11, Tho18, Tur18, TvEK20, VEF^{+}23, VRP23, VV19, Vid22, VWDB16, War23, Wil18, WD18, Wu24, Xie22, XLX^{+}19, XYC22, Yan17, Yan95, YS13, YÖYB23, Zag18, ZLHK02, ZY19, dSBSvE23, dCCM^{+}22]. Packages [Alm10, Boe17, DHF15, GASA15, HL09, KMTS14, RGD12, TD07, De 16, DG22, GIM^{+}24, Mil92, VC23, WML14, YCLL20, Cur18]. Packaging [MBM18]. page [KKEM15]. Pages [Ver12, Abr97, Ano03b, Cap19, Cox05, Cur18, Dal98, Har03, Hou07, Kel10, Laz24a, Liu18, Liu23, Lu18, Lum08, Num20, Orm17, Pet02, Plu01, Pod18, Pod21, Put06, Rao14, Rob17, Rui17, Sta05, Vie11, Mai11]. Paired [HD12]. Pairwise [SBMG06]. Panel [BL14a, BK11, Jac11, MP12]. Panel-Corrected [BK11]. paper [Tol23]. Paperback [Pol09, Mai21, Sha21c]. papers [DO94]. Paradigm [Lo 93, Lo 94]. Paradis [Mai09, Ano12a, Suc07]. Parallel [CQ95, Den16, Edd21, HM16, Hof11, Mat16b, Röh00, RTL07, dVSWAL17, SME^{+}09, TRL09, DP13, EKP^{+}11, HHJC11, KF17, PSM^{+}11, Edd18]. parallelism [JPOJ12]. parallelization [LMY^{+}11]. parallelize [ZMV^{+}18]. Parallelizing [PLZ^{+}15, Tie09]. Parameter [Asq14, BI13, CLL17, HWY18, PLGHG^{+}23]. Parameterization [Mug10]. Parameters [RC17, Sar16]. Parametric [Cul11, HH14, Jac16, KL14, KSP15, MRL12, SP14a, TP11, Cad13]. Parametrical [RMG12]. Parental [TV11]. parfm [MRL12]. Parkway [Sha21c, Sha23a]. Parser [Siu17]. Parsimonious [PMM18]. parsing [LPR21]. Part [HFY22, Lei02, Lei03]. Partial [MW07, Sch14, Sch17a, AY22, LB12b, SS18a, SS18b, Tu21]. Partially [KNI16, MT20]. Participatory [VEF^{+}23]. Partition [MH15, Sho13]. Partitioned [GKZ12]. Partitioning [GGC^{+}15]. Partitions [Han06a, HW07]. Parts [ZC10]. Party [CI17]. passage [RRSPTR12, RRSPTR14]. Passing [Hof11]. Past [Iha98]. patch [KOC21]. patches [MS23]. path [Shi16, VV19]. paths [Alb19, VV19]. patient [LB12b]. patient-reported [LB12b]. Patrizia [O’B10]. Pattern [BMGT15, Lab12, RD92, RD93, Val09]. Pattern-Mixture [BMGT15]. Patterns [BT05, GRD13, MR12, MTPL15, PG15]. Paul [Gun06, Lee21, Mai10, Ree19, Unw12b]. PAVA [dLHM09]. Pawel [Bas18]. pbk [Liu23]. pbkrtest [HH14]. pcalg [KMC^{+}12]. pcIRT [Hoh18]. pcnetmeta [LZHC17]. PCovR [VKVC15]. PCs [Mil92]. pcse [BK11]. pdc [Bra15]. PDEs [SS18b]. pdfCluster [AM14]. Peace [Par15a, Num13]. Peak [RG07]. Pearson [CM14]. Pebesma [Hoe09]. Penalized [RMG12]. Penalty [CS12]. Peng [Buc09]. People [Lip22, Piñ24]. PerFit [TMN16]. Perform [GGC^{+}15]. Performance [CQ95, MCM12, ES14, Gav10, Lan18, SADU23]. performance-ranking [SADU23]. Performing [EE07, LZHC17, Spe13, ELS17, LHR16]. Periodograms [TFR16]. Perl [Lab12, The99, Val09]. PerMallows [ICL16]. Permutation [BPSS09, PSM^{+}11, Ber10, Ros10b]. permutations [Han20]. persistent [WWDS18]. Person [TMN16]. Person-Fit [TMN16]. Personal [Gen98]. Perualila [Oli17]. Pesarin [Ber10, Ros10b]. pesticide [BSG20]. Petchey [Ree19]. Peter [Ano09a, Che11, Iac15, Ric11, San03, Sel17, PG20]. Petersen [MN03]. Petris [Mig10, O’B10]. Petrone [Mig10, O’B10]. Pfaff [Lüt11, Ano10]. Pharmaceutical [Bra03, Eas03, Har03, KM01]. Phase [HSL11, KWE^{+}17, PFT^{+}12]. Phil [Bat08a, Ko95b, Sau10]. Philentropy [Dro18]. Phillip [Pol09, Pol09]. Phonetic [How18]. photosearcher [FAM^{+}20]. Photovoltaic [Lam12]. PHP [MP06]. phtt [BL14a]. Phylogenetic [O’B14, Swe13, GL15, WD18]. Phylogenetics [Ano12a, Par12, Suc07, Par06a]. phylogram [WD18]. physiologically [HRC20]. phytoplankton [OSH^{+}21]. Picek [Oli07]. Pierre [Mye12b, O’B12b, Unw13b]. Pierre-André [Unw13b]. pigmentation [OSH^{+}21]. Pilz [AA12, Krä20]. Pinheiro [Woo01, Zie01a]. pipeline [Lan18, LWC^{+}09, WWDS18]. PIT [HBBP18]. PITR [HBBP18]. pksensi [HRC20]. PlackettLuce [TvEK20]. Planning [KWE^{+}17]. Plans [Kie08, MMB15]. Plant [For20]. Platform [AB12, Jac16, Rac12, CBL^{+}19, KHJ^{+}23, SA15, SCK^{+}23, SADU23, SKW17, YCLL20]. Platform-Independent [Rac12]. Plaza [All86]. plink [Wee10]. PLMIX [MT20]. Plot [MF14, GGS23, MÇSD14]. plotly [LB21, Sha21a]. plots [Tur18]. Plotting [GRD13]. pls [MW07]. Plug [TFH12]. Plug-In [TFH12]. PLUS [AC04, Cow03, Hew05, Pol09, SCD07, Sta99, Süs93, Utl05, Woo98, Zie04, dL05a, Ano99, Ano06c, Cox05, Dal98, Den98, Fox05a, Kra97, Ros00, Ano95b, Ano97, Ano98, Beh04, Beh05, BH94, But05, Can04, Car04, Cha99, Con03, Eve94, ERH01, Eve02, Eve05, FGG^{+}94, Gen98, Ger94, GZ11, Goo05, Gru95, Hal93, Har06, HC05, H^{+}96, HBPJ04, Kon04, Kuo03, Lan95, MJR93, MC97, Mat85, Sta92a, Sta92b, Sta93a, Sta93c, Sta93g, Sta93e, Sta93h, Sta93j, Mat94, Mil98, Mil02, PB00, SM05, SP05, Sta93b, Sta93d, Sta93f, Sta93i, Sta93k, Sta95c, Sta95d, Sta95e, Sta95g, Sta95f, SHR97, VR94, VR99, VR00a, Ano96, BA97, BG96, BHS00, Cra02, CN97, Dia05, Fox02, KO97, KO00, KM01, Kra05, Las97, Lo 93, Lo 94, Man03, Mat98, Mil00]. Plus [MN01, OS95, Pri05b, Rom07, RD92, RD93, SCK95, Sel98, Spe94, Sta95a, VR97, Wie04, ZW03, Ziv05, HH04, SA01, Yan95, YWL02, Abr97, Ano03c, Ano06d, Arg98, Bar02, Bro07, Daw03, Eas03, Fer02, Fis06, Fot07, Han98, Jam96, Kum07, Kus03, Lum01, Mac98, Sis03, Zie98, Zie01c, Zie01b, Zie01a, Zie02a, Zie02b, Zie05, Ano03b, Bra03, Bur07, Dav95, Dia06, Har03, Kim95, Ko95b, Ko95a, Law02, McN04, MN03, Mor03, Pet02, Til96, Tol03, Woo01, Zie99, Zie00, KK99, Lun02]. Plus(R) [Lun06, Ano03a]. Pnuts [KSC^{+}00]. Poggi [Mai21]. Point [BT05, EE07, FBC07, GRD13, Har10, JM15, KSS^{+}07, MTPL15, PG15, Pen03, RD92, ACG^{+}16, MC18, RD93, XYC22]. Point-referenced [FBC07]. Pointer [KSC^{+}00]. Points [LPLPD14, Cad13, Las97]. PoisNor [AD17].

Poisson [AD17, Cam12, CAA15, KN05, MC18]. poLCA [LL11]. Policies [Oom13]. Policy [HGG08]. Political [Lip21c, UC21, How16b]. Pollution [Buc09, PD08, YCLL20]. Polygons [LRN18b]. Polyhedral [LPLPD14]. Polymorphisms [SBMG06]. polynomials [SMHBR06]. Polytomous [Hoh18, LL11]. pomp [KNI16]. Pool [dLHM09]. Pool-Adjacent-Violators [dLHM09]. popbio [SM07]. Popes [SVCB18]. Population [Fie12, Fou09, SL09, Zha11]. Population-Based [SL09, Fou09, Zha11]. Populations [TV11, OSH^{+}21, RK20]. portfolio [BMOF17, Pfa13, SM05]. Portfolios [ABGF17]. Portland [IEE93]. Portnoy [Sta05]. Positions [CI17]. Positive [LA19]. post [GSB19]. post-processing [GSB19]. Posterior [Smi07]. Potentially [RG07]. £67.99 [Rui17]. Poursat [AC04, Cox05]. Poverty [Esp15, Lon15]. Power [LT16, LR15, ATCA20, JV14, LT16, Asq14]. powerful [Ger94, Mat94, YÖYB23]. PowTrAn [ATCA20]. pp [Aji17, Asz23, Ben21, Che23, Cho22a, Das21, Die21, Esk21, Fus22, GC23, Gle16a, Grö15a, Grö16, Hac17, Har23, Has18, Hel15, Hof15, How16b, Iac15, Joh20, Kal24a, Kal24b, Kha16, Kle17, Lee21, Len20, Lip20a, Lip20b, Lip21a, Lip21b, Lip21c, Lip22, Mai21, Mal21, Mat15, Mat16a, Peb21, Piñ24, Rus15, Sab19, Sha21c, Sha23a, Sha23b, Sta21, Tol23, Tu21, Wan22, Zei16]. PPGISr [VEF^{+}23]. PPtreeViz [Lee18]. Practical [AC04, CHT98, Dal98, Hew05, Iac15, IDE15, KF14, Lip21c, Mun14a, Otn17, Sel17, UC21, Woo98, ZM14, Cox05, Dév09, GL16a, H^{+}96, HBPJ04, RN17, SA20, Vin08, Wün13, Grö16, Kin20, Mil17a]. Practice [Nas14, KB11]. Practitioners [GL14]. Precision [dVSWAL17]. Predict [MBGK18]. Predicting [Arc10, Yua07, Xie22].

Prediction [CQ95, Iri19, Lip20a, SMM^{+}15, Che22, Asz23]. Predictive [BG18, Cap19, Kuh08, Mil15a, MZ18, YÖYB23]. Preferences [HD12]. preferred [HBA19]. prefmod [HD12]. PReMiuM [LHA^{+}15]. Preparation [Kra07]. Preparing [AL16]. prepdat [AL16]. preprocessing [NKHZ21]. Presence [FM08]. PresenceAbsence [FM08]. Presmoothed [dUJ13]. prespecified [Ros08]. Press [AS24, Ano16, Asz23, Ben21, Cun12, Cur18, Das21, Die21, For20, GC23, Grö15a, Har23, Hu21, Joh20, Kal24b, Kle17, Laz24a, Laz24b, Lee21, Len20, Lip22, Lu18, Pet02, Pet19, Piñ24, Pod18, Rao14, Ree19, Rui17, Sab19, Sha21c, Sha23a, Sha23b, Sta21, Tol23, Agr16, Cho22a, Cho20, Edd18, Gal17, Liu17, Par15a, Sán19, Tat18]. Press/Elsevier [Tol23]. Price [Abr97, Cox05, Dal98, Har03, Hou07, Kel10, Lum08, Orm17, Pet02, Plu01, Put06, Rob17, Sta05, Vie11]. Prices [Mil15b, BL14b]. Primal [LM14b]. Primer [GvdMW09, Hel18, Hor12c, Kim20a, Soe10, Sól10, Ste09, Aji17, Eks12, Eks16, Fri16, Kee12, Hac17, Häg12, Neu12, Grö13]. primitive [MGHR16]. Principal [Lei10, MW07, VKVC15]. Principles [CFZ09, SR16]. print [Kle17]. prioritizing [MS23]. Priors [Lee13, LRN18a, ZF15]. Prison [FM18b]. prisonbrief [FM18b]. Probabilistic [GR16, GVM16, Meu13, SAR11]. Probabilities [VV16]. Probability [BNB96, BNB07, BNB08, Hor09, KR10, Kur19, Laz11a, LQC^{+}12, Lu18, Mal09, Pic09, UMA08, vdWG11, Bac08, ST10, Bat08b, Ng09, Pie09, Sha22a]. Probit [IV05, Zag18]. Problem [Pap16, Bor16b, Krä20, VV19]. Problems [AVS20, EE07, VSV09, Wol99, LKD08, RRSPTR14]. PROC R [Wei12]. procedure [HBNS14]. Procedures [BGH^{+}17, TMW18]. Proceedings [All86, HI97, MN69]. Process [EN11, Gra07, GT10, MRC15, PLZ^{+}15, Pen03, SFS12, SKS15, BI13, CBHLG21, ELS17, MC18, VRP23]. Processes [CAA15, Gra16, Har10, KNI16, Kor22, LHA^{+}15, RLWP16, TDRD13, TDRD15, ACG^{+}16, RRSPTR12, RRSPTR14, Sha23b]. Processing [LBC^{+}16, Sól09, AFGH22, DP13, Fir03, GSB19, HCSH15, Röh00, Wün13]. processor [MZ18, SP14b]. prodest [Rov17]. producing [HBA19]. Product [Sho13, OOSM18]. Production [Rov17]. Products [Ano99]. profdpm [Sho13]. Professional [Ano99]. Profile [LHA^{+}15, Wie04]. Profiling [SP14b]. Program [BMGT15, But05, LHR16, Yan95]. Programmer [Sta92a, Sta93h, Sta93i, Sta97b]. Programmierung [Süs93]. Programming [Ano09d, BW88, BCW88, BCW18, Bel19, Ber09, Cal95, Cha86, Cha98, Cha08a, Cha14b, DN17, Hil10, Hly09, Hor12a, IR12, Lub91, Mat11, Ng11, Nor09, O’B09a, Oli10b, R D11a, Sco11, Ste00, Tho18, VR00a, Wei12, Wic08a, dL09b, Bia94, BM07, BM16, BM21, CN97, CNZ99, Dav15, Gen09, GL16a, Gro14a, JMR14, OJMR09, VR00b, VS02, Ven04, Lan17b, Plu01]. Programs [JS05, NK06, GDBK^{+}21, KKL^{+}15, SMHBR06]. Progression [BVE^{+}15]. Project [BCHR15, McN22]. Project-Based [McN22]. Projection [Lee18, FBNRG21, SA20]. projection-based [SA20]. Projections [LM14b, SAR11, WCHB11]. projects [SNR18]. Promises [GJK^{+}21]. Propensity [HP09, Sek11]. proportion [SR17, SR17]. Proportional [FH09, Gan15a, TKM06, ZQS16, PCAS09]. Proportional-Odds [FH09]. proposal [Lan09a]. Prospects [TLH11, MCSGBSA20]. Proven [LT19]. Providing [Den16, GMF18, NPP18]. Proximities [IP08]. PSAgraphics [HP09]. Pseudo [Fog17]. PST [GR16]. Psychoco [WSZ12]. Psychology [Veh13, RKY11]. Psychometric [WSZ12]. Psychometrics [DB18, Sab19]. PTAk [Lei10]. PtProcess [Har10]. Public [HGG08, VEF^{+}23]. publications [Pol11b]. Published [Lu18, Nun20]. Publishers [Aji17]. Publishing [Che23]. PubMed [RL15]. purchasing [BMOF17]. Pure [DC09]. Purposes [MP06]. Pursuit [Lee18, FBNRG21]. put [KB11]. Putler [Sco13a]. PY-SUMMA [AVS20]. PypeR [XMW10]. Python [AVS20, BP17a, Dan18, GH18, GHN19, Mil15a, Mül13, Smi17, Smi18, SMM^{+}22, XMW10, Lip22, Piñ24].

Q [LLS15, Tur18]. Q-Learning [LLS15]. Q-Q [Tur18]. QC [CPD^{+}20]. Qian [Fin11, Woo11]. qicharts2 [Anh18]. Qingzhao [Laz24b]. qqman [Tur18]. QtAC [SGPSzC22]. QTL [Doe10, BS09, HSG12, TV11, Bro14, Doe10]. QTS [OS95]. Qualark [OS95]. qualities [PSZ17]. Quality [Anh18, MG09, CC23]. Quandle [FW18]. Quantification [Dro18, SS15]. Quantile [SS19a, AA20]. Quantitative [Hof15, Sha22c, WS11, Day15]. Questionnaire [Fri12, O’B12a, Fal12]. Questionnaires [BBG16, Lal17]. Quizzes [ZUL14].

R

[AA12, Agr16, Aji17, AC04, Ano06c, Ano09c, Ano12b, Asz23, Ber09, Bos09, Bow10, Buc09, Bur09, Bur10, Cam09, Car10, Car16, Che23, Cho22b, Cow03, Cox05, Cur18, Dav07, DN17, Dia06, Doe06, Dur14, Edd09a, Edd18, Eme08, Esp15, Fus22, Gal17, GC23, Gil14, GL14, GR18, Gou05, Gre22a, Grö13, Grö15a, Gro18c, Hac17, Häg12, Han13a, Har19, Har23, Hel18, Hew05, Hew16, Hoe09, Hof15, Hor12b, Hou07, Hu21, Hub11, IR12, Iac15, IDE15, Kal24b, Kau13, Kha17, Kin20, Kos15, Lal17, Lan17b, Laz11b, Laz24b, Lig09, Lip20b, Lip22, Liu17, Liu23, Lor18, Lun06, Lun07a, Luo09, Mai08, Mal21, Mar12, Mat13, McG20, McN14, McN22, MCSAGB20, MCSGBSA20, Mor18, Mor03, Mül16, Mur14, Neu12, O’B13, Orm17, Otn17, Par15b, Piñ24]. R [Pol09, Pol11c, Pol13, Rah14, Rao14, Ree09, Ree19, Ros09, Rui17, SL09, San10a, Sar06, Sch08, Sco11, Sel17, Sha21a, Sha23b, Smi06, Sta07a, Ste11, Sto11, Tur12, Utl05, VM09b, VM09a, Veh13, Vey14, Wan06, Wan16, Wan22, Zha11, Zie04, dL05a, ADH11, AFGH22, ALE10, Adl10, Adl12, ACW12, AR14, AMYR16, AVS20, AFHD09, Aiz12, Alb16, ATYK20, Alb07, Alb09, AR12, Alb15, Alb19, Ald20, AGG13, ATF10, AT13, AA20, All11a, AL16, Alm10, ACG^{+}16, AD15, AD17, ABS23, Amo21, AB17a, AWBM18, Anh18, Swe13, Ano13, ADN15, AKL^{+}21, Arc10, AHvD09, AB17b, ABGF17, ABC19, ATCA20, ABEY18, AY22, Arm14, Arn17, AE21, Asq14, AMW14, AB12, AGdSC20, AFM22, AM14, BVFB19, BG18, Bac08, BL14a, BT05, BSS^{+}24, BDMP15, BK11]. R [BR07, BS15, BRF^{+}18, Bak13, BF17, Bar14, BFRP13, BP17a, BPL09, BY18, BS18, BBG14, BBG16, BPP17, BPSS09, BP12, BGH^{+}17, Bas18, Bat15, BU15, BKH17, BLY18, Bea17, BFA14, BBGL17, BML19, BP17b, BB18, Bee13, Beh12, Bel19, BCHY09, BKT14, BBG12, BB12, BC11, BAS12, BSVT12, BK17, BL15, BdMM15, BCG18, BLM^{+}15, BPGR08, BVE^{+}15, Blo14, BDT18, BMOF17, BS16, BLW12, Boe17, BN10, Bol08, Bon18, BPGC14, BDdM11, Bor16a, BPB09, Bos12, BD18, BCMR19, BCAB07, Bra15, BM07, Bra14, BM16, Bra17, BM21, BS13, BHW11, BPDD08, BS09, Bro14, Bro12, BI13, BdSvE23, BSG20, BBN10, BMGT15, Bür17, Bur12, But05, But08, CC23, Cad13, CF14, CPA20, CdM10, Cal06, CSNF18, Cam12, CM14, CBGGV17, CC08a, CG15, Car17, Car13]. R [Car14, CB17, CP11a, CBL^{+}19, CPD^{+}20, CFHBK11, CAA15, CCJMR16, Cha12, CH17, Cha08a, Cha14b, Cha16, Cha20, CNA16, Cha08b, CQZ^{+}10, Cha18, Cha15a, Cha15b, CGBN14, Cha10, CKSLS12, CKSLS18, CP11b, CP13, Che22, CP12, CHB14, CL11, CH11a, CC11, CKY14, CGC11, CA18, CLK21, Cic15, Cla08, CMS^{+}11, CS12, Coe18, CC08b, CGC14, CLL17, CFSR15, CRP13, CS07, CGH^{+}08, CGH^{+}10, CGH^{+}12a, CGH^{+}12b, CBHLG21, CPB^{+}23, Cot13, CM09, CH18, Cra05, Cra07, Cra12, Cra15, CNZ99, CNZ10, CO16, CF08, CsC08, CCP^{+}11, CGS09, CJM06, Cul11, CI17, DMB18, DC09, Dal02, Dal08, Dan18, DS20, DLN17, DHM11, DB13, Dav15, Day15, De 16, DBZ^{+}11, DC05, DP13, DMD15, Dem13, Dem20, DG22, Den13, Den16, DB18, DCB^{+}17, DLC06, Dév09]. R [DG23, DCMCPF20, Dir18, Dis23, DJS^{+}18, DW17, DM18, DMS14, Dra23, Dră12, DK18, Dro18, DTDd19, DPSH18, Duo07, DHF15, DGP08, EN11, EF11, ES14, EW16, ESO16, EB18, Edd21, EK23, Eks12, Eks16, EJ13, ELS17, Er12, EE07, EBO^{+}13, EK12, EL09, EKP^{+}11, Eve05, EH06, EH11, Eyl22, Fal12, Far05, Far06, Far16, FS10a, FBdlF12, FHM08, Fel12, FT08, FW18, FDB12, Fer11, FDGD16, FM19, FHH17, Fie12, FMF12, Fie15, Fil08, FGZ14, Fin14, Fin10, FBC07, Fir03, Fir05, FO15, FBNRG21, FC11, FM18a, Fou09, Fox02, Fox03, Fox05b, FH09, FFM09, FC12, FL16, FAM^{+}20, FVVL23, FBB20, FM08, FM18b, FBF14, FK17, Fri16, FKP17, FM16, FGEMI12, Fuj17, FS10b, GRMS11, GR16, GRD13, GH18, GHN19, GZ21a, GZ21b]. R [GB13, GSD12, Gan15b, Gan15a, Gas18, Gas11, GGC^{+}15, GSJC22, GSB19, GASA15, GIH^{+}05, Gen09, GRK^{+}16, GP12, GG16, GMF18, Ger20, GKZ16, GK16, GL16a, GKD14, Gio09, GL16b, GGK10, GHG19, Gle16b, GV19, GJK^{+}21, GDBK^{+}21, GLC^{+}15, GK14, GGS23, GCA12, GEV18, GFS14, GDMB08, Goo05, Goo13, GFC12, GS19, Gos11, GIM^{+}24, Gra07, GT10, Gra16, GvdMW09, GVM16, GW18, Gro14a, Grö10, Grö14b, Grö18b, GvdL12, GZP14, GL07, GZ09, GH11, GKZ12, Gu14, Gua13, GL15, Guh07, GAE23, GV12, Had10, HPWdL15, HHB08, HD10, HH14, HD18, HSL11, Han05, Han06a, Han06b, HW07, Han07a, Han07b, Han20, HJM08, Har06, HBBP18, Har08, HCSH15, Har10, Has18, HD12, HBA19, HXY12, HM18, Hec15, HH04, HN09, HH15, HP09, HH19a, Hel22]. R [HH07, HT11, Hic16, HCM^{+}18, HT19, Hla16, HHJC11, HHJC14, HTWT23, HPCS14, HM11, HM16, HL09, Hof11, HMRS14, HMS16, Hoh18, Höh07, HF07, Høj04, HHY05, HL07, Høj12, HEL12, Hor09, Hor12a, HFY22, HL04, HL05, HBZ09, HMR^{+}13, HG14, HBQ04, HK11, HK15, Hos22a, Hos22c, How17, How18, HRC20, HS18, HSG12, HBPJ04, HGG08, HLP11, HK08, HWY18, IP08, Iac08, IESdF18, IG96, Iha98, IV05, ILS11, IRCA21, Iri19, ICL16, IK20, Jac11, Jac16, JPM19, Jal19, JWHT13, JM15, JJJ14, JKvT^{+}14, JHQ^{+}11, JGM18, JPOJ12, JPH21, JD15, Joc14, JV14, Joh07, JS05, Jon07, JMR14, JP06, Kab11, KHLF^{+}10, KMMV14, KMC^{+}12, Kan17, KSHZ04, KMH06, KF10, KD11, KN05, KL14, Kas16, Kav15, KK15, Kee12, Kee10, Kee18, KHR21, KS14, KF17, Kie08]. R [KWE^{+}17, KE14, KO06, Kim20a, KOC21, KNI16, KSS^{+}07, KF14, KK22, KZ08, Kle17, KH10, KM08, KN03, KM14, KG17, KR10, KY10, KC14a, Kom09, KK14, KSBZ16, KPSH15, Kor22, Kor18, KKEM15, KKL^{+}15, KV13, KMTS14, KT16, KK21, Kra07, KSP15, KHJ^{+}23, KP23, Kub23, Kuh08, KR13, Kuo03, KB11, LPLPD14, LT16, LAF^{+}17, Lam12, Lan18, Lan14a, Lan17a, Lan09a, Lan09b, Lan14b, LBS17, LH12, LA19, Lan19, LSM20, LRGTA12, Lau18, LWC^{+}09, LL10, LM14a, LW16, LJH08, LeB18, LIL^{+}15, Lee13, LM18, Lee18, LRN18a, Lei10, Lei13, Lei02, Lei04, LHS08, LSPvdL17, Len09, Len16, LŠ20, LMY^{+}11, LB12a, LQC^{+}12, LB12b, LBW18, LB21, LM14b, LFF17, LM03, LS16, LF15, LZHC17, LLS15, LL11, LBC^{+}16, LC10, LHA^{+}15, LR15, Lof21, LPR21, LRN18b, LT19, Lon15, LRRÁCSGS14]. R [dUJ13, LH14, LLM^{+}20, LBL^{+}21, Luh13, LX12, LCSC14, Lup09, Lyu21, MGHR16, MRC15, Mäc07, MR12, MF14, MHJS16, MB03, MH07, MdL10, MMB15, MÇSD14, Mar06, MA14, MTPL15, MH15, MB15, MR14, Mar07, MBGK18, MQP11, MBM18, Mat11, Mat16b, MF15, MCAP19, MP14, MPM14, McE16, ML13, McL17, MYK07, MZ08, MH18, MdN22, MS11, MCM12, Mel16, MM22, MJGM10, MS23, Meu13, MW07, MH09, MUM16, MVS13, Mil15a, MP12, Mil15b, MP06, Mir14, ML11, MTvdM15, MT20, MCB18, MBR11, MV14, MHOV12, MdUÁC10, MN14, MN17, MZ18, MAY23, MBT^{+}20, Mue09, MH10, Mug10, Mul15, MvSB^{+}07, MvS07, MAG^{+}11, Mul14, Mül13, MRL12, Mun14a, Mur03, Mur06, Mur09a, MG09, Mur09b, Mur11, Mur18, Mur05, Mwi11, Mwi13a, Mwi13b, NK06, NPR13, NB13]. R [NR16, Nag18, NKHZ21, Nar05, NV11, Nas14, Nas08, ND12, NWH21, NGBK12, NL12, NL14, NPP17, NPP18, NdSL16, NRD16, NAA17, OH11, OHD17, Obe14, Ohr14, Oja10, OCRC14, OOSM18, OL17, OSH^{+}21, Oom13, OK14, OJMR09, PLZ^{+}15, PS22, PL23, PH16, PPGD15, PP17, Pan18, Pap16, Par06a, Par12, Par15b, Par15a, PCAS09, PFT^{+}12, PLRC10, PSS^{+}17, Peb12, Peb21, PG15, Pen03, Pen08, PD08, PDH16, PJSPC17, PZK^{+}12, PP18, Per14, PPC09, Pet10, PP11, PSM^{+}11, PR07, PNR13, Pfa06, Pfa08a, Pfa08b, Pfa13, Phi10, PU13, PK08, Pir10, PSZ17, Pla12, Pla19, PLGHG^{+}23, PC11, PMLM23, Pol11b, PT07, PT09, PLLC11, PLR^{+}16, Pri05a, Pri05b, PMM18, PK12, Qia10b, Qia16, QZLP21, dREP12, RH02, Rac12, Rah17, RHW21, RHG09, RN17, RPVG11, RKY11, RV20]. R [Rec10, RK20, Ree19, RFGD08, RBB18, RLWP16, RC17, RG07, RJH14, RS05, RS08, Rit09, RBHB15, Riz06, Riz10, Riz12, Riz16, Riz08, RC10, Rob18, Rob23, Rob08, RFKM12, RR11, RRSPTR12, RRSPTR14, Rom07, RMG12, Ros07, Ros12, RTL07, RL15, RGD12, Rov17, RVHG23, dVSWAL17, Ruf09, Run13, Sah22, SOD^{+}16, SF18, SS18a, SVM^{+}17, SFS12, Sar08, Sar16, SMHBR06, SH17, SP14a, Sav16, Saw09, SE18, SA15, SIR^{+}11, SZ11, Sch11, SK17, Sch14, Sch17a, SME^{+}09, Sch17b, Sch16, SSH^{+}20, SlRC16, SGPSzC22, Sch17c, Scr13, SR18, Scu10, SD15, SL05, Sek11, SVMMRP17, SAR11, Sha22b, Sha22c, Sha12, Sha21d, SWCP20, SKD22, She09, SNR18, SBMG06, Shi16, Sho13, SS19a, SC07, Sho18, SS06, SS11, SS19b, SA20, SKS15, SR16, SMWP20]. R [Siu17, SCK^{+}23, SM21, SCS13, Smi07, SYC08, Smi17, Smi18, SMM^{+}22, SLS^{+}12, SP10, SPS10, SCM12, SvdLN17, SO23, Sól09, SADU23, SD18, SPPP17, SHN17, Spa17, Spe08, Spe13, SP14b, SWHJ17, Sta08, SR07, SKW17, Ste23, Ste24, Ste09, SKZ05, SM07, SLG05, Su07, SGHY11, SR17, SSV14, ST10, TP11, TW11, Tae14, Tag23, TMCE22, Tak12, TM15, Tan18, TKM06, TMW18, TRM16a, Tat18, Tav17, TV11, TDRD13, TDRD15, TB17, Tee11, Tei22, TN22, TKM16, TKM15, TMKD17, TMN16, Ten18, TM05, TLH11, TFH12, Thi14, TFR16, Thi18, TD07, TR14, TT21, Tho18, Tie09, TRL09, TPAM07, TYH^{+}14, TNM17, Tou15, TGJ17, Tro09, Tsa13, Tsa14b, TGKV20, TF12, Tur18, TvEK20, Tus11, Tyn16, UMA08, UAK^{+}15, ÜKD09, ÜS10, Unw15, Urb09, UC21, USHH18]. R [VML12, Val09, VW13, VEF^{+}23, VSV09, VP16, Van18, VG09, Var14, VFV13, VRP23, VS02, Ven04, Ver18, VKVC15, Ver05, Ver12, Ver14, VV19, Vid22, VC23, Vie10, VV16, VKM18, Vin08, VdL09, Vin10, VYD^{+}12, VS10, VWDB16, WWDS18, WGSL12, WDT^{+}12, WMS17, WLK08, WW11, Wan11, Wan13, WPW15, WLH^{+}18, WHK21, War23, Was15, Wee10, WF12, Wei12, WML14, Wei02, Wel18, WTB^{+}15, WDM^{+}11, WRD23, WS11, WST11, WM14, WSZ12, WC07, Wic11a, WCHB11, WG17, Wic19, WRV23, WZMC19, Wil18, WD18, Wil14b, Wil11b, Wil12, WMR16, Wol94, Wol99, Woo06, WGE17, WsT18, Wu24, Wün13, WM18, XMW10, XWHL15, Xie13a, Xie13b, XAG19, XDR21, Xie22, XLX^{+}19, XYC22, Yan17, YS13, YCLL20, YEL18, YH14, YÖYB23, You10, You12, YL17, Yua07, ZGS^{+}18, Zag18, ZFZ10]. R [ZLHK02, ZKJ08, ZC10, ZUL14, ZF15, ZPC^{+}16, ZQS16, ZY19, ZMV^{+}18, ZP13, ZGZvB19, ZHL11, ZM09, ZML16, ZMS21, ZM14, ZS17, ZIM09, ZIW^{+}09, dSJdSF14, dLM09a, dLHM09, dLM09b, dLM09c, dSBSvE23, dVM12, dWFP11, dCCM^{+}22, lGlRGL09, vBGO11, vdA07, vdA12, vdWG11, AS24, AA12, All11b, Ano06b, Ano06a, Ano08, Ano09a, Ano09b, Ano09d, Ano10, Ano12a, Ano16, BN07, Bai11, Bat08a, Bat08b, Bau14, Ben21, Ber10, Boo10, BL11, Cap19, Cha14a, Cha22, Che11, Chi07, Cho09, Cho20, Das21, Dem17, Dic12, Doe10, Esk21, Eva14, Few09, Fin11, For20, Fri11, Gle16a, Grö15b, Gut11, Han13b, Han13a, Hec07, Hel15, Hly09, Hoe09, Hor12c, How16b, Hu09, Hüs18, Joh09, Kal24a, Kel10, Kha16, Kim20b, Krä20, Kuh10, Kum10, Lab12, Laz11a, Laz24a, Lee21]. R [Len20, Lep14, Li11, Lip20a, Lip21a, Lip21b, Liu16b, Liu18, Lu18, Lum08, Lüt11, Mai09, Mai10, Mai11, Mai14a, Mai14b, Mai21, Mar11, Mar12, Mat16a, Mig10, Mil12, Mil17a, Mor09, Mye09, Mye12b, Mye12a, Neu11, Ng06, Ng09, Ng11, Nor08, Nor09, Nor14, Num13, Num20, Nun20, O’B08, O’B09c, O’B09a, O’B09b, O’B10, O’B12b, O’B12a, O’B13, O’B14, Oli07, Oli10b, Oli10a, Oli17, PG20, Pet19, Pfa12, Pic09, Pod18, Pod21, Pol11a, Put06, Qia10a, Ree18, Ric11, Rie19, Rip11, Rob12, Rob05, Rob07, Rob17, Ros10b, Sab19, Sán19, San03, Sau10, Sau12, Sch09, Sco09, Sco10a, Sco11, Sco12, Sco13b, Sco13a, Sen14, Sha21c, Sha23a, Sól10, Sta21, Sta07b, Suc07, Tol23, Tsa14a, Tu21, Unw11, Unw12a, Unw13a, Ven10, Vie11]. R [Wil14a, Woo11, Yu12, Zei16, AC07, Ano09c, Arm19, Bar18, Bat04, Beb13, Ber09, Bos10, Bow09, Bro10, Bul06, Cap22, Cha21, Cho22a, Cho15, Chr09, CH11b, Dem18, Dia18, Die21, Dow17, Dre19, Dur15, Edd09c, Edd09b, Edd11, Edd12b, Edd12a, Eva11, Fle11, Fri12, Gep21, Gol11, Gou10, Grö11, Grö18a, Gro08, Gun06, Har07, Hel16, Hil06, Hil10, Hos22b, How11, How16a, Joh20, Jon13, Kha18, Kos16, Leo10, Lew16, Lha14, Lip21c, Liu15, Liu16a, Lud13, Lun07b, Mai06, Mai12, Mal13, Mal09, Mat15, Mil10, Mil17b, Mun14b, Nie11, Nie14, Nor15, Oom10, Pal15b, Pal15a, Pan15, Pav16, Pie09, Pod15, Ree09, Rob13, Rob19, Rob22, Rob09, Rui16, Rus15, San10b, San11, Sau11, Sau21, Sco10b, Sha16, Sha21b, Sha22a, Sha11, Sha19]. R [She11, Soe10, Str10, Sun15, Sun16, Tod22, Tus05, Tyl07, Unw12b, Unw13b, VSS^{+}17, VL21, Voe09, Wic08b, Wic08a, WG10, Yal10, dL05b, dL06, dL09a, dL09c, dL09b, Gre22b]. R-Based [DCB^{+}17, AFGH22, HLT09, SA15, WRV23]. R-INLA [BRF^{+}18]. R-Menu [BP12]. R-mode [MÇSD14]. R-package [CCJMR16, KHJ^{+}23, SGPSzC22]. R-packages [De 16]. R-php [MP06]. R-Software [ZS17]. R-Squared [Rec10]. R/MATLAB [VFV13]. R/PY [AVS20]. R/PY-SUMMA [AVS20]. R/Python [AVS20]. R/qtl [BS09, Bro14, Doe10]. R/S [Pol09, Goo05, Har06, Rom07]. R/S-PLUS [Pol09, Goo05, Har06, Rom07]. R/Stan [Dis23]. r2 [Grö15a]. R2GUESS [LBC^{+}16]. R2MLwiN [ZPC^{+}16]. R2WinBUGS [SLG05]. Rabe [Daw03, Zie02b]. Rabe-Hesketh [Daw03, Zie02b]. Radiation [Lam12]. Radisson [All86]. Rafael [Das21, Lip20a]. Rahlf [Hüs18]. Ramble [Siu17]. ramps [SYC08]. Ramsay [Bur10, Car10]. Randal [Car16]. Randall [Sco13b, Rao14]. Random [AE21, Bea17, CF14, Car17, DB13, FS10b, JJJ14, MR12, PB15, Sar16, SMM^{+}15, TKM16, Yan95, CHB14, Dev86, Fog17, SA20, Mai21]. Random-effects [Yan95]. RandomFields [SMM^{+}15]. Randomization [Han13a, USHH18, JPH21, ZHL11]. Randomized [HM18, TPAM07]. randomizeR [USHH18]. randomLCA [Bea17]. Randomly [MdUÁC10]. RandPro [SA20]. Rank [TM05]. Rank-Based [TM05]. ranked [KK21, MT20]. ranking [SADU23, dSBSvE23]. Rankings [HD12, TvEK20]. Ranks [Oja10, Hub11]. RankspeQ [SADU23]. RAppArmor [Oom13]. Raritan [Ano99]. Rasch [AA12, Krä20, Veh13, Hoh18, MH07]. Rate [SAR11]. Rates [MP14]. Ratings [HD12]. Ratio [MVS13]. Raton [Agr16, Asz23, Ben21, Cho22a, Cho20, Cur18, Das21, Die21, Edd18, Esk21, For20, Fus22, Gal17, GC23, Gle16a, Grö15a, Grö16, Hou07, Joh20, Kel10, Kha16, Laz24a, Lee21, Len20, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Liu17, Lum08, Mal21, Nun20, Orm17, Par15b, Pet19, Put06, Rob17, Sab19, Sán19, Sha21c, Sha23a, Sha23b, Sta21, Sta05, Tat18, Tu21, Vie11, Wan22]. Rattle [Sau12, Mwi13a, Wil11b]. rBeta2009 [CHB14]. Rcall [EK23]. Rcapture [BR07]. Rchoice [Sar16]. RcmdrPlugin.survival [FC12]. Rcpp [EF11, EB18]. RcppArmadillo [ES14]. RcppCNPy [EW16]. RCrawler [KF17]. Reach [Ano99]. reaching [Lof21]. Read [EW16, Sta95b]. Read-Write [EW16]. Ready [VYD^{+}12]. Ready-to-Use [VYD^{+}12]. REBayes [KG17]. Recapture [BR07]. Recipe [SH17]. Recipes [LT19]. recmap [Pan18]. record [IEE93, IEE94]. Recording [BK17]. Records [SD18]. Recovering [PLR^{+}16]. recruitment [Cad13]. Rectangular [Pan18]. reduce [WPW15]. reduced [BSS^{+}24]. reduced-impact [BSS^{+}24].

Reduction [AR14, Wei02]. Reference [R D11b, Cad13, R D04, WML14]. referenced [FBC07]. References [Cur18, McL17]. RefManageR [McL17]. Regimen [WMS17]. regimes [SMWP20]. Region [Bra14, HBNS14]. region-level [HBNS14]. Regional [LQC^{+}12]. Regions [WHK21, dCCM^{+}22]. Registration [BLY18]. Regression [AB17a, AC04, Ano09b, Bar18, BFRP13, BdMM15, Bon18, BKL05, Car13, Cha22, CGC11, CNZ10, Dal98, DLN17, DLC06, DM18, DPSH18, Eyl22, Far06, Far16, FM19, Gal17, GKZ16, Gou10, Gra07, GZP14, GKZ12, Har01, HM18, HB92, Hew05, Hos22b, KN05, KL14, KV13, Lee21, Lei04, LM14b, Lip22, LBC^{+}16, LHA^{+}15, Mar11, MW07, Mül16, Num20, O’B09b, Piñ24, RBHB15, She09, SS19a, SM21, Sta08, TFR16, TR14, TGJ17, UAK^{+}15, VKVC15, WMS17, WLK08, Wan06, Wei02, WMR16, Woo98, XWHL15, ZLHK02, ZKJ08, ZQS16, dL09a, AA20, AFM22, BH94, CCJMR16, Cox05, De 16, Fox02, Hos22c, H^{+}96, HBPJ04, MPV12, Orm17, Par06b, PCAS09, RS08, Rit09, Ros09, SMHBR06, TKM16, VP16, VRP23, Wei14, Grö16, Zie04, Cow03, Kal24b]. Regression-Based [WMS17]. Regression/Item [CGC11]. Regressions [Har23, Liu23, GL07]. regRSM [TKM16]. regular [PMLM23]. Regularization [Sch11]. Regularized [FHH17]. Reisensburg [DO94]. related [DG22]. Relational [But08, MC97]. relationships [PSZ17]. Relative [DHM11, DM18, PP18]. Relevant [CGBN14]. relevent [MB15]. Reliable [AGM07]. relsurv [PP18]. remBoot [Car17]. REML [Yan95]. rEMM [HD10]. Remote [Gos11, ADN15]. René [Zie05]. Repeated [PCAS09]. Replacement [Han07a]. Replication [Mil15b]. reported [LB12b]. reporttools [Ruf09]. REPPlab [FBNRG21]. Representation [Hic16]. Representations [Lau18]. Reproducibility [BBGL17, BML19, Lan18]. Reproducible [Gan15b, LT16, SvdLN17, CPB^{+}23, FAM^{+}20, MR09, SNR18, SMM^{+}22, McG20]. Reproducibly [MBM18]. Resampling [Har06, Pol09, Sco12, CH11a, Goo05, Goo13, HBNS14, Rom07, Tsa14a]. Research [Ano99, Gan15b, LT16, Mai14a, McG20, Rus15, Str10, Bee13, Cha15a, FVVL23, Hel15, LB12a, MR09, MS23, Vin10]. researchers [HBA19]. Resistant [Can04, Con03]. Resonance [CMS^{+}11, TW11, WS11]. Resource [Din06, RG96]. Response [Arc10, Bon18, Cha12, CGC11, CO16, DBZ^{+}11, Had10, Har23, HM18, IDE15, Joh07, KF14, Len09, Liu23, MR12, MPM14, MTvdM15, Riz06, BBG14, HFY22, dCCM^{+}22]. Response-Surface [Len09]. Responses [GSD12, KR13, LA19, LBC^{+}16, MTvdM15, Tou15, ZC10]. Restoring [BK17]. Results [BK17, Alb15, Tan18, Tur18]. Retention [KSS^{+}07]. Rethinking [Gre22a, Hew16, Mal21, Rui17, McE16]. rethnicity [Xie22]. retrieval [LPR21]. returns [FM18b]. reversible [GSB19]. Review [AS24, Abr97, AA12, Agr16, Aji17, All11b, AC04, Ano03a, Ano03c, Ano03b, Ano06b, Ano06a, Ano08, Ano09a, Ano09b, Ano09c, Ano09d, Ano10, Ano12a, Ano12b, Ano16, Arg98, Arm19, Asz23, Bai11, Bar02, Bar18, Bas18, Bat04, Bat08a, Bat08b, Bau14, Beb13, Ben21, Ber10, Ber09, Boo10, Bos09, Bos10, Bow09, Bow10, Bra03, Bro07, Bro10, Buc09, Bul06, Bur07, Bur09, Bur10, BL11, Cam09, Cap19, Cap22, Car10, Car16, Cha14a, Cha21, Cha22, Che23, Che11, Chi07, Cho22a, Cho15, Cho09, Cho20, Chr09, CH11b, Cow03, Cox05, Cur18, Dal98, Das21, Dav95, Daw03, Dem17, DN17, Dem18, Den98, Dia06, Dia18, Dic12, Die21, Doe06, Doe10, Dow17, Dre19, Dur14, Dur15, Eas03, Edd09c, Edd09a, Edd09b, Edd11]. Review [Edd12b, Edd12a, Edd18, Eme08, Esk21, Esp15, Eva11, Eva14, Fer02, Few09, Fin10, Fin11, Fis06, Fle11, For20, Fot07, Fox05a, Fri12, Fri11, Fus22, Gal17, GC23, Gen98, Gep21, Gil14, Gle16a, GL14, GR18, Gou05, Gou10, Gre22a, Gre22b, Grö11, Grö13, Grö15b, Grö15a, Grö16, Grö18a, Gro18c, Gro08, Gun06, Gut11, Hac17, Häg12, Hal93, Han13b, Han13a, Han98, Har03, Har19, Har23, Has18, Hec07, Hel15, Hel16, Hel18, Hew05, Hil06, Hil10, Hly09, Hoe09, Hof15, Hor12c, Hos22b, Hou07, How11, How16a, How16b, Hu09, Hu21, Hub11, Hüs18, IR12, Iac15, Jam96, Joh09, Joh20, Jon13, KK99, Kal24a, Kal24b, Kau13, Kha16, Kha17, Kha18, Kim20b, Kim95, Kin20, Kle17, Ko95b, Ko95a, Kos16, Krä20, Kuh10, Kum07, Kum10, Kus03]. Review [Lab12, Lal17, Lan17b, Law02, Laz11a, Laz11b, Laz24a, Laz24b, Lee21, Len20, Leo10, Lep14, Li11, Lig09, Lip20a, Lip20b, Lip21a, Lip21b, Lip21c, Lip22, Liu15, Liu16a, Liu16b, Liu17, Liu18, Liu23, Lor18, Lu18, Lud13, Lum01, Lum08, Lun02, Lun06, Lun07b, Luo09, Lüt11, Mac98, Mai06, Mai08, Mai09, Mai10, Mai11, Mai12, Mai14a, Mai14b, Mai21, Mal13, Mal21, Mal09, Man03, Mar11, Mar12, Mat13, Mat15, Mat16a, McG20, McN14, McN04, McN22, Mig10, Mil12, Mil10, MN03, Mil17a, Mil17b, Mor18, Mor03, Mor09, Mül16, Mun14b, Mur14, Mye09, Mye12b, Mye12a, Neu11, Neu12, Ng06, Ng09, Ng11, Nie11, Nie14, Nor08, Nor09, Nor14, Nor15, Num13, Num20, Nun20, O’B08, O’B09c, O’B09a, O’B09b, O’B10, O’B12b, O’B12a, O’B13, O’B14]. Review [Oli07, Oli10b, Oli10a, Oli17, Oom10, Orm17, Otn17, Pan15, Par15b, Par15a, Pav16, PG20, Peb21, Pet02, Pet19, Pfa12, Pic09, Pie09, Piñ24, Plu01, Pod15, Pod18, Pod21, Pol09, Pol11a, Pol11c, Pol13, Put06, Qia10a, Rao14, Ree09, Ree18, Ric11, Rie19, Rip11, Rob12, Rob13, Rob19, Rob05, Rob07, Rob09, Rob17, Ros09, Ros10b, Ros00, Rui16, Rui17, Rus15, SL09, Sab19, San10b, San10a, San11, Sán19, San03, SCD07, Sau10, Sau11, Sau12, Sau21, Sch08, Sch09, Sco09, Sco10b, Sco10a, Sco11, Sco12, Sco13b, Sco13a, Sel17, Sen14, Sha21a, Sha21b, Sha22a, Sha11, Sha19, Sha21c, Sha23a, Sha23b, She11, Sis03, Smi06, Soe10, Sól10, Sta21, Sta05, Sta07a, Sta07b, Ste11, Sto11, Str10, Suc07, Sun15, Sun16, Tat18, Til96]. Review [Tod22, Tol23, Tol03, Tsa14a, Tu21, Tur12, Tus05, Unw11, Unw12b, Unw12a, Unw13b, Unw13a, Utl05, VM09b, VM09a, Veh13, Ven10, Vie11, VSS^{+}17, VL21, Vis10, Voe09, Wan06, Wan16, Wan22, Wic08b, Wic08a, WG10, Wil14a, Wil11a, Woo01, Woo11, Woo98, Yal10, Yu12, Zei16, Zha11, Zie98, Zie99, Zie00, Zie01c, Zie01b, Zie01a, Zie02a, Zie02b, Zie04, Zie05, dL05b, dL05a, dL06, dL09a, dL09c, dL09b, AC07, BRF^{+}18, Cho22b, Dav07, Edd21, GIM^{+}24, Har08, Har07, Lun07a, Tyl07, Vey14]. Reviews [Ano06c, Kel10, FVVL23, Hew16, Kos15, Lew16, Lha14, Pal15a, Sha16]. Reweighted [HB92]. rfishbase [BLW12]. RGCCTranslationUnit [Lan09b]. rgenoud [MS11]. RgoogleMaps [LR15]. RGtk2 [LL10]. Rich [AB12]. Richard [Fle11, For20, Mal21, Rui17, Sco10b, Utl05]. Richness [SP05, Wan11]. right [Wil12]. RinRuby [DC09]. Ripley [Abr97, Arg98, Fer02, Jam96, Zie01b]. RISE [Wil18]. Risk [ABGF17, CNA16, DHM11, DM18, SZ11, Pfa13]. Risk-Based [ABGF17]. RiskPortfolios [ABGF17]. Risks [Mal13, dWFP11, BVFB19, BAS12, MN17]. Ritz [Ano09b, O’B09b]. River [WTB^{+}15]. Rizzo [Len20]. rjmcmc [GSB19]. RKWard [RFKM12]. RLadyBug [HF07]. RMatlab [VFV13]. RMatlab-app2web [VFV13]. rmcfs [DK18]. RMCriteria [MM22]. Rmixmod [LIL^{+}15]. RNetLogo [Thi14]. Rob [Krä20]. Roback [Lee21]. Robert [Gol11, Mai12, Neu11, Ng11, Nor14, Nun20, O’B09a, O’B13, Pol13, Rip11, Sau11, Sco10a, Sco13a, Voe09, Cha14a, Sch08]. ROBETH [MJR93]. Robin [Gle16a, Mai21]. Robinson [Mye12a, Ng11, O’B09a, Sco11]. RobPer [TFR16]. Robust [AMYR16, BGH^{+}17, HB92, H^{+}09, JP06, KSS^{+}07, KSBZ16, LH12, MJR93, TFR16, VW13, Wie04, ZFZ10, KK21, NKHZ21, SHR97, Oli07, Tyl07, dL06]. Robustbase [Fin10]. robustloggamma [AMYR16]. ROC [DCMCPF20, GZ21a]. Roger [HH14]. Roland [Lu18]. Roll [PLLC11]. rolling [PMLM23]. Root [Lup09]. rOpenSci [BCHR15]. Routine [CKY14, RL15]. Routines [LM14b, MJR93, Mül13, ELS17, Lan95]. Rouvière [Unw13b]. Row [YH14]. Row-column [YH14]. Rowman [Aji17]. Roxygen [Vid22, VC23]. rpanel [BCAB07]. rpartOrdinal [Arc10]. rpartScore [GSD12]. RProtoBuf [ESO16]. RRreg [HM18]. RSKC [KSBZ16]. rsm [Len09]. RSNNS [BB12]. rsppfp [VV19]. RStudio [Grö15a, HK15, Rac12]. rTensor [LBW18]. RTI [SVM^{+}17]. rtimicropem [SVM^{+}17]. Rubba [Iac15, Sel17]. Rubber [Kur19]. Rubio [Hoe09]. Ruby [DC09]. rucrdtw [BS16]. Ruey [Lep14]. Rule [RBHB15]. Rule-Based [RBHB15]. Run [ZPC^{+}16]. runjags [Den16]. Running [FK17, SLG05]. Runtime [KKL^{+}15, LMY^{+}11]. Ruppert [Num20].

S [AC04, Ano99, Ano06a, Ano06c, Bar02, Bra03, Bur10, Cam09, Car16, Cow03, Cox05, Dal98, Dav95, Den98, Fin11, Fox05a, Han13a, Hew05, Hoe09, How16b, Kim95, Ko95b, Ko95a, Kra97, Law02, Lep14, Li11, Liu18, Lun06, Mac98, Mai10, Mar11, Mig10, Nun20, Pfa12, Plu01, Pol13, Rao14, Ros00, SCD07, Sch08, Sco13a, Sta05, Sta07b, Sta99, Sto11, Süs93, Til96, Utl05, Vis10, Woo01, Woo11, Woo98, Zie98, Zie01c, Zie01a, Zie02a, Zie02b, Zie04, dL05a, Cun12, Ano95b, Ano96, Ano97, Ano98, BC81, BC84a, BC84b, BC85a, BC85b, BW88, BCW88, Bec94, BCW18, Beh04, Beh05, BH94, BA97, BG96, BHS00, But05, Cal95, Can04, CHT98, Car04, Cha86, CH92, CH93, Cha95, Cha98, Cha20, Cha99, Con03, Cra02, CN97, The99, Dia05]. S [Eve94, ERH01, Eve02, Eve05, FGG^{+}94, Fox02, Gen98, Ger94, GZ11, Gru95, Hal93, Här91, HB92, HH04, HC05, H^{+}96, HBPJ04, KVCS98, KSC^{+}00, Kon04, KO97, KO00, KM01, Kra05, Kuo03, Lan95, Las97, Lo 93, Lo 94, Lub91, Man03, MJR93, MC97, Mat85, Sta92a, Sta92b, Sta93a, Sta93c, Sta93g, Sta93e, Sta93h, Sta93j, Mat94, Mat98, Mil98, Mil00, MN01, Mil02, NS94, OS95, PB00, Pri05b, Röh00, RD92, RD93, SM05, SCK95, SA01, Sel98, SS92, SP05, Spe94, Sta93b, Sta93d, Sta93f, Sta93i, Sta93k, Sta95c, Sta95d, Sta95e, Sta95g, Sta95a, Sta95f, Ste00, SHR97, Süs93, VR94, VR97, VR99, VR00a, VR00b, VR02, WLK08, Wie04, Yan95, YWL02, ZW03, Ziv05, Abr97, Ano03a, Ano03c, Ano03b, Ano06d, Arg98, Bar02, Bra03]. S [Bro07, Bur07, Dav95, Daw03, Dia06, Eas03, Fer02, Fis06, Fot07, Han98, Har03, Jam96, Kim95, Ko95b, Ko95a, Kum07, Kus03, Law02, Lum01, Mac98, McN04, MN03, Mor03, Pet02, Sis03, Til96, Tol03, Woo01, Zie98, Zie99, Zie00, Zie01c, Zie01b, Zie01a, Zie02a, Zie02b, Zie05, KK99, Lun02]. S-language [Röh00]. S-PLUS [AC04, Ano99, Cow03, Hew05, Pol09, SCD07, Sta99, Süs93, Utl05, Woo98, Zie04, dL05a, Ano06c, Cox05, Dal98, Den98, Fox05a, Kra97, Ros00, Ano95b, Ano97, Ano98, Beh04, Beh05, BH94, But05, Can04, Car04, Cha99, Con03, Eve94, ERH01, Eve02, Eve05, FGG^{+}94, Gen98, Ger94, GZ11, Goo05, Gru95, Hal93, Har06, HC05, H^{+}96, HBPJ04, Kon04, Kuo03, Lan95, MJR93, MC97, Mat85, Sta92a, Sta92b, Sta93a, Sta93c, Sta93g, Sta93e, Sta93h, Sta93j, Mat94, Mil98, Mil02, PB00, SM05, SP05, Sta93b, Sta93d, Sta93f, Sta93i, Sta93k, Sta95c, Sta95d, Sta95e, Sta95g, Sta95f, SHR97, VR94, VR99, VR00a, Ano96, BA97, BG96, BHS00, Cra02, CN97, Dia05, Fox02, KO97, KO00, KM01, Kra05, Las97, Lo 93, Lo 94, Man03, Mat98, Mil00]. S-Plus [MN01, OS95, Pri05b, Rom07, RD92, RD93, SCK95, Sel98, Spe94, Sta95a, VR97, Wie04, ZW03, Ziv05, HH04, SA01, Yan95, YWL02, Abr97, Ano03c, Ano06d, Arg98, Bar02, Bro07, Daw03, Eas03, Fer02, Fis06, Fot07, Han98, Jam96, Kum07, Kus03, Lum01, Mac98, Sis03, Zie98, Zie01c, Zie01b, Zie01a, Zie02a, Zie02b, Zie05, Ano03b, Bra03, Bur07, Dav95, Dia06, Har03, Kim95, Ko95b, Ko95a, Law02, McN04, MN03, Mor03, Pet02, Til96, Tol03, Woo01, Zie99, Zie00, KK99, Lun02]. S-Plus(R) [Lun06, Ano03a]. S. [Car10, Doe06, Doe10, Fin11, Joh09, Zha11]. S4 [KSHZ04, KR10]. SAEM [CLL17]. saemix [CLL17]. Sahu [Sha23a]. Salmaso [Ber10, Ros10b]. Sample [LQC^{+}12, MVS13, PLRC10, Sta08, WMR16, WM18, CLK21, RK20, Tag23]. Sampler [MHJS16]. samples [CBHLG21]. Sampling [Bar14, CC11, Cho22a, GKZ16, Han07a, Kie08, Laz11a, PG15, VSV09, KK22, KK21, ST10]. SamplingStrata [Bar14]. Sandboxing [Oom13]. sanon [KK15]. Sarkar [Kuh10, Nor08]. SAS [Ano08, Ano09c, GL14, Hu09, Kel10, Lum08, Mai12, O’B08, Oli17, Ree09, Rob09, Sau11, Sta21, Voe09, BVFB19, BMGT15, Cha08b, GGS23, HH04, KH10, Liu17, Mue09, Pri05b, Rec10, SMHBR06, SC07, Sho18, Tae14, TR14, TPAM07, Wei12, Laz24b, Utl05, Oom10, Ste11, Tur12]. Satellite [MCAP19]. Saunders [BPL09]. SAVE [PPGD15]. Saveliev [Oli10a, Sch09]. Sawitzki [Mil12, Boo10]. Say [The99]. Scalable [CQ95, LM14a, NL12]. Scale [CB17, CFHBK11, Gra16, SR07, vdA07, vdA12, GDBK^{+}21, SNR18, ZMV^{+}18]. Scales [PLR^{+}16, BSG20]. Scaling [KM08, MdL10, PLLC11, ÜKD09, dLM09a, dLM09b]. Scatterplot3d [LM03]. scbursts [DTDd19]. Schemes [TDRD15]. Schloss [DO94]. Schumacker [Mor03]. Science [BB18, Ben21, Blo14, Cap19, Cha15b, Das21, Don17, Dow17, Dur15, DGP08, Edd18, Iri19, IK20, Kos15, Lip20a, Lip20b, Lip21c, Liu18, Mat16b, Mil17a, Mil17b, Pen17, Rao14, Rie19, SO13, Str10, UC21, All86, BKH17, BCMR19, Cha20, Che22, Fus22, Hor09, Mil15a, Pie15, Sán19, Sha22b, Sha22c, Vin10, WG17, ZM14, HBA19, HJ17, RL15, Gre22b]. Sciences [All86, DK09, Mun14b, SVMMRP17, Gua13, NL14, Kal24b]. Scientific [DN17, HI97, Hil10, Ng11, O’B09a, Sco11, FVVL23, JMR14, Kut13, OJMR09]. Scientist [Grö18a]. scientists [Hos22a, Ohr14]. Scope [GI00]. Score [Alb16, ABC19, HP09, Sek11, dSJdSF14, PCAS09]. scores [Gas18]. Scraping [Iac15, Mun14a, Sel17, KF17]. Screening [SF18]. Script [GHG19, KK22]. Scriptics [Ano99]. scripting [LMY^{+}11]. Scutari [Tat18]. sdcMicro [TKM15]. SDD [BDMP15]. Seamless [EF11, PFT^{+}12]. Search [ALE10, ARC04, BS16]. Searching [GFC12]. Seasonal [KMTS14, SE18]. SEATS [SE18]. Seattle [IEE94]. Sébastien [Mai11]. SEC [LPR21]. Second [Cha22, Dur14, Hac17, Has18, Mai12, Unw12b, VM09b, AC04, Gil14, Orm17, Tsa14a]. Security [KSC^{+}00, Oom13]. Segmentation [BLY18, TNM17]. Segmented [Mug10]. Select [CC08a]. Selecting [LRRÁCSGS14, Wil12]. Selection [CdM10, DK18, HPWdL15, HMS16, LAF^{+}17, LBC^{+}16, ND12, Sch11, SR18, TMW18, WF12, WMR16, Kom09, MUM16, PFT^{+}12, WM18, XLX^{+}19]. Selectiongain [MUM16]. selective [BSS^{+}24]. Selvin [Zie00]. semantics [GJK^{+}21]. Semi [Cul11, JHQ^{+}11, KL14, KSP15, LIL^{+}15, OH11, SP14a, Ste24, BBN10]. Semi- [JHQ^{+}11, SP14a]. Semi-automated [Ste24].

Semi-Markov [KSP15, ABS23, BBN10]. Semi-Parametric [Cul11, KL14]. Semi-Supervised [Cul11, LIL^{+}15]. SemiMarkov [KSP15]. Semiparametric [Bur12, Gra07, Num20]. SemiParSampleSel [WMR16]. Sen [Doe10]. Sensing [Gos11]. Sensitivity [GT10, SP10, HRC20]. September [Bia94]. seqHMM [HH19a]. seqtest [Yan17]. Sequence [HH19a, MB15, VV16, TNM17]. Sequences [GRMS11, GR16, Mel16, OH11]. Sequential [BVFB19, CC11, GG16, IDE15, KF14, Was15, Yan17]. Sergio [Tu21]. Serial [BDMP15]. Serialization [ESO16]. seriation [HHB08]. Series [All11b, Ano03a, Ano10, Bos10, Bra15, Bur07, Car16, CM09, CsC08, Dis23, DMS14, Edd09a, Fis06, Fot07, Gut11, Hac17, Har07, Has18, HK08, Joh09, Kas16, Kha17, Kum07, Lau18, Li11, LFF17, Lor18, Lu18, Lun07a, Lüt11, Mai10, MCAP19, MYK07, McN04, MV14, Nun20, Oli17, Pod15, Pol13, Qia10a, Rao14, San10b, San11, SCD07, SS19b, Tol03, VdL09, WGE17, ZML16, Amo21, BS16, Bos09, CCP^{+}11, ELS17, NKHZ21, Nie11, Per14, Pfa06, Pfa08a, PMLM23, Sch08, SS06, SS11, Tsa10, Tsa14b, ZW03, Ziv05, ZM09, Bos09]. Series-Theory [Rao14]. Server [Ano99, MC97, Kor18]. Service [RVHG23]. Services [Kra95, Aji17, Fri16]. Set [CGBN14, HW07, IRCA21, LPLPD14, dLHM09, KK21]. Sets [BDdM11, MH09]. Seventeenth [All86]. several [TPAM07]. shadows [GDBK^{+}21]. Shaken [GKZ12]. Shape [LPLPD14, SR07]. Shaped [PG15]. shapes [Ros08]. sharing [KKEM15]. Sheather [Mar11]. Shelter [Kim20b]. Shi [CCP^{+}11]. shifted [TN22]. Shiny [RVHG23, SCK^{+}23, LB21, Sha21a]. shinyHugePlot [Tag23]. Shkedy [Oli17]. Short [Bat08a, Bat08b, The99, Häg12, Han13b, Han13a, Kau13, Luo09, Mai08, Mai09, Mai10, Mai11, Mai12, Nor08, Nor09, Num13, O’B08, O’B09c, O’B09a, O’B09b, O’B10, O’B12b, O’B12a, O’B13, Pol11c, Pol13, Ric11, Rob12, Ros10b, Sco09, Sco10b, Sco10a, Sco12, Sco13b, Sco13a, Sta07a, Unw11, Unw12b, Unw12a, Unw13b, Unw13a, Veh13, Voe09]. Shortage [KSC^{+}00]. shortest [VV19]. Shortfall [CNA16]. Shoukri [Ano08, Lum08, Sta21]. Should [Cun12]. Shrinking [CQZ^{+}10]. Shumway [Nun20, Pol13, Sch08]. Sievert [LB21]. Sightability [Fie12]. SightabilityModel [Fie12]. sigma.js [Coe18]. sigmajs [Coe18]. Signatures [MBT^{+}20]. Significance [CFHBK11]. Significant [GFC12]. Signs [Hub11, Oja10]. Sik [Qia10a]. SILC [Lon15, Esp15]. simcausal [SvdLN17]. simecol [PR07]. SIMEX [HXY12]. simFrame [ATF10]. Simian [TNM17]. Simon [Chi07, Iac15, Rob07, Sel17]. simPH [Gan15a]. Simple [BCAB07, MN14, RTL07, SOD^{+}16, ZFZ10, dLM09c, Beh05]. simpleCache [SNR18]. SimpleITK [BLY18]. Simplex [ZQS16]. simplexreg [ZQS16]. simPop [TMKD17]. Simulate [DC05, MvSB^{+}07, BSS^{+}24, BSG20]. Simulating [AE21, BS18, BVE^{+}15, GRD13, Han07b, MCM12, Day15, KP23]. Simulation [ATF10, AHvD09, BGH^{+}17, DN17, Hil10, Iac08, LT16, Laz11a, MN14, Ng11, O’B09a, PK08, Rip87, SMM^{+}15, Sco11, SvdLN17, TMKD17, BF17, HT19, JMR14, KHJ^{+}23, MN17, OJMR09, ST10, XLX^{+}19, Bur09, Cam09, Luo09, Sto11]. Simulations [CGC11, HM16, PFT^{+}12]. Simulator [MRC15, BB12]. SimultAn [GZ11]. Simultaneous [GZ11, HH07, KPSH15, KR13]. Simultaneously [TMCE22]. Single [IDE15, KWE^{+}17, KF14, SBMG06, Su07, DTDd19, HBA19, SR17]. Single-Arm [IDE15, KWE^{+}17, KF14]. single-channel [DTDd19]. Singular [GK14, BFA14]. SIS [SF18]. Size [LQC^{+}12, RV20, CLK21, Tag23]. Sizes [MBGK18]. Skew [LM18, ML13]. Skills [Ald20]. Slave [Tyn16]. slope [Alb19]. slope-dependent [Alb19]. SMACOF [dLM09b]. Smale [GP12]. smarter [GL16a]. Smith [Sch09, Oli10a]. SmoothHazard [TGJ17]. Smoothing [Bar18, Cam12, Can04, Con03, Den98, Gu14, Här91, Kle09, MPM14, PT07, PT09, Ros00, BA97]. smoothROCtime [DCMCPF20]. sms [Kav15]. Snippet [Aiz12, AD15, BDMP15, BK11, BS18, CG15, Car13, CO16, DMB18, FKP17, GVM16, Han06a, HW07, Han07a, HXY12, Hof11, Hoh18, Jal19, LW16, LQC^{+}12, LRN18b, MF14, MHJS16, MTPL15, MBGK18, MF15, MP14, NK06, NAA17, Pan18, Pap16, PZK^{+}12, Phi10, Rec10, RBB18, Ruf09, SBMG06, TR14, VSV09, Wei12, XMW10, dSJdSF14]. Snow [TRL09]. sns [MHJS16]. Social [AT13, Kal24b, Str10, Hos22a, Vin10]. sociales [Gua13]. SOCR [Din06]. Soetaert [Lud13]. Soft [Sta99]. Software [AR14, BCHR15, But05, Cha08a, GL14, Har08, Hil96, IR12, KPSH15, Kra07, LZ10, Mat11, NV11, NGBK12, OS95, PN97, Sek11, SKZ05, TRM16b, Wie04, ZS17, AFGH22, Aji17, Bor16a, Cal06, FGZ14, Fin10, Fri16, GGS23, GS19, GIM^{+}24, LS05, PCAS09, PSZ17, SA15, Vis10, Ano09d, FL16, Oli10b]. soil [CPB^{+}23]. soiltestcorr [CPB^{+}23]. Solar [Lam12, Lam12]. Solari [Ber10, Ros10b]. Solutions [Dav07, Doe06, Lud13, GIH^{+}05, Krä20]. Solved [LKD08]. Solver [Tou15]. Solving [SPS10, SCM12, VG09]. Some [Mil92, AB90]. something [Wil12]. Song [Woo11]. Sonia [O’B10]. Sons [Ano03b, Hel15, Iac15, Lip21a]. Sophia [Daw03, Zie02b]. Søren [Han13b]. Sorin [Ano12b]. sorting [DTDd19]. Sound [Sha21c, Sha23a, MZ18]. Source [HBBP18, TRM16b, Aji17, Fri16, HBZ09, PL23, SA15]. Soviet [Bia94]. Soviet-Italian [Bia94]. Sozialwissenschaften [Luh13]. spa [Cul11].

Space

[CPA20, DLC06, Dis23, Jal19, PP11, Pod15, PLR^{+}16, ÜS10, Cal06, Per14, SS18a, SS18b]. Space-Time [Pod15, Per14]. Spaces [AE21]. spacetime [Peb12]. spam [FS10b]. sparr [DHM11]. Sparse [AMW14, Bra14, Bra17, FS10b, KN03, KSBZ16]. sparseHessianFD [Bra17]. SparseM [KN03]. spate [SKS15]. Spatial [Alm10, BT05, BRF^{+}18, BPGR08, BDdM11, DHM11, Eva14, FBC07, GGC^{+}15, GLC^{+}15, Gra16, Hoe09, HPCS14, Hub11, LRGTA12, Lee13, LRN18b, MP12, Oja10, PG15, Pir10, Pla12, Pla19, Pod15, RD92, SYC08, TDRD13, TDRD15, Arm14, CKSLS18, Kub23, LSM20, Per14, RD93, For20, Gle16a, VL21]. Spatial-Temporal [LRN18b]. Spatially [RC17, BSG20]. Spatially-Varying [RC17]. SpatialStats [Ano96, KVCS98]. Spatio [BS15, GRD13, LRN18a, Lei10, Mil15b, Peb12, RHW21, SKS15, TDRD13, WZMC19, ACG^{+}16, BL14b, LSM20, Sah22, Sha23a, Peb21, Pet19]. Spatio-Temporal [BS15, GRD13, LRN18a, Lei10, Peb12, SKS15, TDRD13, RHW21, WZMC19, ACG^{+}16, LSM20, Sah22, Sha23a, Peb21, Pet19]. Spatiotemporal [TDRD15, YCLL20]. spatstat [BT05]. spBayes [FBC07]. SPC [Cho22a]. Special [FGZ14, TW11]. Species [GVM16, MBGK18, SP05, Wan11, BSG20, Wan11]. Spector [Bat08a, Ko95b, Sau10]. Spectra [Ano99, Kra07]. Spectral [AE21]. spectrino [Kra07]. Spectrometry [CS12]. Spectroscopic [MvS07]. Spectrum [GK14]. speculation [SWHJ17]. speedR [VYD^{+}12]. Spherical [FDGD16]. sphet [Pir10]. Spider [EL09]. Spider-Man [EL09]. spikeSlabGAM [Sch11]. Splancs [RD92, RD93]. Spline [Gu14, Sch17a, WMR16, HHJC14]. Splines [Cam12, WLK08, ACG^{+}16]. split [OS95]. split-beam [OS95]. splm [MP12]. SPlus [Sta97a, Sta97b, Sta97c, Sta98a, Sta98b]. SPlus-4 [Sta97a, Sta97b, Sta97c]. SPlus-5 [Sta98a, Sta98b]. SPODT [GGC^{+}15]. spot [HT19]. sppmix [MC18]. spray [BSG20]. spray-drift [BSG20]. Spread [GW18]. Spreadsheet [Fle11, Sco10b, HN09]. Springer [Abr97, Bos09, Cap19, Che23, Cox05, Dal98, Har03, Hof15, How16b, Mai21, Mat15, Mat16a, Num20, Plu01, Rus15, Zei16]. Springer-Verlag [Abr97, Dal98, Hof15, How16b, Mat15, Mat16a, Rus15, Zei16]. SPRINT [PSM^{+}11]. SPSS [Ano09c, GL14, Mai12, Ree09, Sau11, Voe09, BP12, Hos22a, Mar07, Mue09, RKY11, Veh13]. spTest [Wel18]. spTimer [BS15]. square [CM14]. Squared [Rec10, QZLP21]. Squares [FHH17, HB92, Len16, MW07, Tu21]. SSDs [ZMV^{+}18]. ssMousetrack [CPA20]. SSN [HPCS14]. Stability [TMW18]. Stable [dVSWAL17]. stack [ZGS^{+}18]. Stage [JPM19, MMB15, KK22, MUM16]. stampr [LRN18b]. Stan [Bür17, Dis23, McE16, Har23, Liu23, Mal21, Rui17, Gre22a, Hew16]. Standard [AB17b, BK11]. Standardization [GHG19]. Standards [WST11]. Star [Kur19]. Starch [Ano16]. Start [SH17]. Started [Ree19, BP17b, Mat98]. Stat [Lub91, Nar05]. Stata [BBG16, Gol11, Tu21, SMHBR06, FHH17, MH10, SMM^{+}22, Cho22b, Lal17]. State [CPA20, CP11a, DLC06, Dis23, GRMS11, GR16, IEE94, Jac11, KSP15, PP11, SME^{+}09, dWFP11]. State-Space [CPA20, Dis23]. states [KP23, BL14b]. Stationary [EN11, VV16]. STATISTICA [Mar07]. Statistical [AFHD09, ATF10, AL16, Ano06b, Ano06a, AGM07, BBG16, Car04, Car14, Cha69, CH92, CH93, CRP13, Cra02, Cur18, Dav95, Dem18, Dia18, Dia05, DO94, EJ13, EK12, Eve94, Eve02, EH06, FBdlF12, FDGD16, Fou09, FL16, GI00, GL14, Grö15a, HI97, HKS08, HH04, HH15, Hil06, Hly09, HPCS14, HK15, Hos22c, H^{+}96, HBPJ04, IR12, IK20, Kim95, KNI16, KH10, Ko95a, KC14a, KC14b, Lei13, Liu17, Lon15, Lun02, Lun07b, Man03, Sta93e, Mat11, McE16, MN69, Mor03, MG09, Mwi13b, Nor09, Nor14, OHD17, Oli07, PPGD15, Pan18, Pen08, PD08, Pfa12, Pol11c, Pri05b, R D11b, Rao14, RFGD08, Riz08, Rob12, RFKM12, RTL07, SL09, SF18, Sch17b, Sta93f, Ste11, TP11, Tae14, TKM15, Tem97, Tro09, Tyl07, Ven10, Wic08a, Wil12]. Statistical [Xie13a, Zha11, Zie02a, dL06, Ano13, BNB96, BNB07, BNB08, Bos12, BM07, BM16, BM21, DS20, ELS17, FGZ14, Ger94, HLT09, HK11, JWHT13, JP06, KD11, LS05, Mat94, Mil92, NWH21, PCAS09, Röh00, SA01, Sta09, Tan18, Til96, Tur12, WML14, AC04, Ano03b, Buc09, Cap22, Cox05, Dal98, Dia06, Esp15, Fus22, Gre22a, Hew05, Hew16, Hu21, Kus03, Lal17, Laz24b, Len20, Lip20b, Mal21, Nor15, Rui17, Sco13b, Sha16, Utl05, Woo98, Zie05, dL09c]. Statistically [LAF^{+}17, MZ18]. Statistician [IDE15]. Statistics [Abr97, Ano09a, Arg98, Bat04, Bat08b, Boo10, BHH05, Bra03, BL11, CC08b, Cra05, Cra15, Dal02, Dal08, Dem20, Din06, DO94, DK09, Dră12, Eas03, Edd09c, Fer02, Fie15, Fin11, Fle11, Fox05b, Fri16, Fri11, Gep21, Har06, HH18, HH19b, HBQ04, Jam96, KK99, Kee18, Kha16, Kra95, Kur19, Law02, Lew16, Lip21a, Liu15, Liu16b, LT19, Lu18, Lub91, Lum01, Mäc07, MJR93, MB15, Mar07, MN03, Mor09, Mul15, Nas08, Ng09, O’B09c, Peb21, Pet02, Pol09, Qia10b, RKY11, Ree18, Ruf09, Rup04, Rup11, San03, Sco10b, Sco12, Sha22a, Sta97a, Sta98a, Unw11, Unw13a, VM09b, VR94, VR02, WF12, WG10, WZMC19, Woo11, Zei16, Zie01b, ATYK20, All86, Bak13, Bos09, Che23, CH11a, CGH^{+}12a, Dav15, Dév09, FMF12, Goo05, Goo13, Har03, Hec15, HN09]. statistics [Hel15, Hos22a, Kee10, Krä20, KM01, Lof21, Mil12, MN01, Pan15, Pet19, PNR13, Pie09, Qia16, Ree19, RFGD08, Rom07, RAM05, Saw09, Sha22c, Sha12, TRM16a, Tol23, UMA08, VR97, VR99, VR00a, Ver05, Ver14, Vie11, MN03, Ano06c, Smi06, Unw13b, Hou07, Aji17, Ano12b, Gou05, Hel15, McN22, Mil10, Ng06, Tsa14a, Veh13, Wan16]. Statistik [DK09]. Statistiksoftware [Luh13]. Statistiques [CGH^{+}08, CGH^{+}10, CGH^{+}12b, Gua13]. Stats [Zie99]. Stayton [Mai09]. Stefano [Bur09, Luo09]. Steffen [Han13b]. Step [Bel19, Cot13, GB13]. Step-by-Step [Bel19, Cot13, GB13]. Stephen [Sta05, Tol23]. Steve [Zie00]. Steven [Eas03, Har03, Lum01, Pet02, Zie99]. Stevens [Sól10, Hor12c]. Stirred [GKZ12]. StMoMo [VKM18]. Stochastic [BPSS09, Ber10, Bur09, Cam09, CC08a, CJM06, Iac08, Kas16, Kor22, Luo09, MHJS16, PK08, Rip87, Ros10b, SKS15, SS15, Sto11, VKM18, HF07, Sha23b]. stochvol [Kas16]. stock [Cad13]. stocking [KK22]. Stoffer [Nun20, Rao14, Sch08, Car16, Pol13]. storage [Dir18]. stpp [GRD13]. strangers [Urb09]. strategically [BMOF17]. Strategies [Har01]. strategy [KK22]. stratification [RK20]. Stratified [Bar14, KK15]. stratifyR [RK20]. Stream [HD10, HPCS14]. StreaMD [DJS^{+}18]. Streibig [Ano09b, O’B09b]. strict [GJK^{+}21]. string [NPP17]. strucchange [ZLHK02]. Structural [BKL05, Obe14, PT09, QZLP21, Ros12, ZLHK02, Shi16, Tu21]. Structured [UAK^{+}15, Jon07]. Stuart [O’B09a]. Student [Kau13, Mat13, Mur14, Den13, AHvD09, HJM08]. Student- [AHvD09]. Students [Ald20, GL14, Joc14]. Studies [ADH11, BPB09, Bra03, Eas03, Esp15, Fou09, HSL11, IDE15, KWE^{+}17, KF14, LT16, MF15, NAA17, PC11, Ros10a, SvdLN17, ZFZ10, AY22, FBB20, Har03, HBNS14, HT19, KM01, Lon15, SOD^{+}16, Sha21d, Wan22, Zha11, SL09]. Studio [Gan15b, McG20]. Study [Buc09, HM16, MCM12, PD08, BVFB19, GDBK^{+}21, JV14, YS13, ZGS^{+}18]. Studying [PG15]. Stuttgart [BB12]. Style [GHG19, JPOJ12]. Styling [Ste23]. Subroutines [BC76, BC77, Las97]. subsequence [BS16]. Subset [MZ08]. Subsets [LAF^{+}17]. Subspace [TKM16]. Suess [Laz11a]. Sufficient [AR14]. Suffix [GR16]. Suite [Alm10, LH14, Sha21c, Sha23a]. Sujit [Sha23a]. SUMMA [AVS20]. Summary [SHN17]. Sums [FDGD16]. Sung [Sta05]. Superheat [BY18]. Supervised [BD18, Cul11, LIL^{+}15, Tod22]. Supplement [Sta92b, Sta95g]. Supply [The99]. Support [EW16, HP09, KMH06, MM22, RHW21, SO23]. Supporting [Aiz12, GHG19, GMF18, SVM^{+}17]. Sure [SF18]. Surface [Len09, dCCM^{+}22]. surfaces [Alb19]. surrogate [Oli17, TPAM07]. surveillance [Höh07]. Survey [CG15, Kuo03, Obe14, GGS23, Gle16b]. Survey-Weighted [CG15]. Surveys [AT13]. Survival [CSNF18, CKY14, FC12, Har01, Jac16, dUJ13, MN14, OL17, PP18, RMG12, SVCB18, TM15, TR14, Dee99, Hu21, NB13, SMWP20, Tei22, Sta05]. survPresmooth [dUJ13]. survsim [MN14, MN17]. Sustainable [Bro14]. SV40 [TNM17]. SVAR [Pfa08b]. SVEC [Pfa08b]. Swayne [Eme08, Mai08]. Sweave [Fil08, Lei02, Lei03, Rac12, Sar06, MR09]. Swenson [O’B14]. Switching [Pap16]. Switzerland [Rus15, Zei16]. Sylvie [Cox05, Dal98]. Symbol [TKM06]. Symbolically [Phi10]. symmetric [VP16]. Symmetries [HL07]. Symposium [All86]. Synergism [BN10]. Synergism/Antagonism [BN10]. syntax [KMMV14]. Synth [ADH11]. synthesis [dSBSvE23]. Synthetic [ADH11, NRD16, TMKD17]. synthpop [NRD16]. System [BC76, BC77, BC81, BC84a, BC85b, NV11, TRL09, VG09, AGdSC20, BC85a, Kor18, Lan09a, MM22]. systematic [GIM^{+}24]. systemfit [HH07]. Systems [Ano99, HH07, Lam12, RBHB15, Cha99, Kut13, LBS17, NWH21, SGPSzC22, Wu24].

T [Bat08b, Dic12, Liu18, Ng09, Yu12, AWBM18, LM18]. t-Distributions [LM18]. T. [Ano06a]. Table [WGSL12]. Tableman [Sta05]. Tables [ILS11, Lei13, OK14, Ruf09, Sta08, SSV14]. Tai [Asz23]. takos [LLM^{+}20]. Takuya [Veh13]. talk [Urb09]. target [BSG20]. Targeted [GvdL12, LSPvdL17]. taxonomy [VC23]. Taylor [GC23, Har23, Lip20a, Lip20b, Lip21b, Lip21c, Lip22, Mal21, Nun20]. Taylor [Asz23]. TclPro [Ano99]. tcltk [BCAB07]. tclust [FGEMI12]. TDAstats [WWDS18]. teach [RH02]. Teaching [MP06]. Technical [Mor18]. Techniques [Agr16, Den98, FM16, Här91, MP14, Ros00, WMS17, Bar95, BA97, Mil15a, SHR97]. Technologies [Ano99, Mun14b, NL14]. technology [Ano96]. teigen [AWBM18]. Teleconnections [ADN15]. Telemetry [HBBP18]. Temperature [Mug10]. Tempering [GT10]. Templ [PG20]. Templates [Den16, Vin08]. Temporal [BS15, BDdM11, GRD13, LRN18a, Lei10, LRN18b, Mil15b, Peb12, SKS15, TDRD13, ACG^{+}16, BL14b, LSM20, RHW21, Sah22, Sha23a, WZMC19, Peb21, Pet19]. Tensor [Lei10, LBW18, PT09]. Term [PC11]. Terry [Fir05, TF12]. Test [Bat15, NAA17, Ver18, CPB^{+}23]. Testing [CFHBK11, GFC12, JJJ14, Kim20a, Lup09, MR12, Rob18, ZLHK02, ATYK20, GEV18, HBNS14, PSM^{+}11, Tae14, Tei22, War23, Gro18c]. Tests [BPSS09, Ber10, FS10a, FO15, HH14, JKvT^{+}14, LT16, LRRÁCSGS14, LX12, MVS13, PJSPC17, Ros10b, Wee10, Wel18, ZFZ10, dSJdSF14, AFM22, FFM09, GH18, JV14, Yan17]. texreg [Lei13]. Text [FHM08, HMR^{+}13, Iac15, Joc14, Mun14a, Sel17, SR16, TFH12, Tod22, KF10]. textcat [HMR^{+}13]. Texts [Rao14, Bos09]. Textual [BB18, Rie19, Ben21, Sán19]. tgcd [PDH16]. tgp [Gra07, GT10]. their [DH97, PSZ17]. Thematic [Ten18]. Theophile [Oli17]. Theoretic [Kim20a]. Theory [BPSS09, Ber10, Car16, Cha12, Cha08b, CGC11, CFZ09, DMS14, Dro18, Han07b, Hu09, KB11, MPM14, MTvdM15, O’B08, Rao14, Ros10b, ÜS10, Dem13, HFY22, Kel10, Krä20, Rob09]. thermal [LLM^{+}20]. thermoluminescence [PDH16]. Things [HHB08]. Third [Pol13, Ree19]. Thomas [Par15b, Hüs18]. Thorn [Hac17, Häg12, Neu12]. Threaded [Tem97]. Three [GKD14, LX12, RV20]. Three-Way [GKD14, RV20]. ThreeWay [GKD14]. threg [XWHL15]. Threshold [HS18, MMB15, PJSPC17, XWHL15]. Thresholding [JS05, KO06]. ThresholdROC [PJSPC17]. Throughput [PSS^{+}17]. Tian [Cho20]. Tibshirani [Nor14]. Tidy [Mil17b, SR16, FM18b, WG17]. tidytext [SR16]. Tidyverse [Fus22, IK20, Lip20b]. TIFF [NPP18]. TikZ [Ste24]. Tiles [LR15]. Tilman [Ano16]. Tim [Sco12]. Time [All11b, Ano03a, Ano10, BL14a, Bos10, Bra15, BHS00, Bur07, CHT98, Car16, CKY14, CS12, CM09, CsC08, Dis23, DMS14, Edd09a, Fis06, Fot07, Gio09, Gut11, Har10, Har07, Has18, HXY12, HK08, Jal19, Jon13, Kas16, Kha17, KSS^{+}07, Kum07, Lau18, Li11, LFF17, Lor18, Lu18, Lüt11, Mai10, MCAP19, MYK07, McN04, MV14, Pod15, Rao14, RG07, Riz10, Riz16, San10b, San11, SCD07, Sav16, SS06, SS11, SS19b, TR14, Tol03, VdL09, Wan13, WGE17, ZML16, Amo21, BS16, DCMCPF20, ELS17, Hel22, LS16, NKHZ21, Nie11, Per14, Pfa06, Pfa08a, PMLM23, Riz12, RRSPTR12, Sta05, Tsa10, Tsa14b, ZW03, Ziv05, ZM09, CCP^{+}11, Sch08, Bos09, Joh09, Lun07a, Nun20, Pol13, Qia10a]. time-dependent [DCMCPF20]. Time-frequency [CHT98]. Time-Indexed [Sav16]. Time-Intensity [RG07]. Time-Series [Dis23]. Time-to-Event [Jon13, Riz10, Riz16, Riz12]. Time-Varying [TR14, Hel22]. Time-Weighted [MCAP19]. timereg [SZ11]. Times [MTvdM15, RRSPTR14]. Timing [The99]. TIMP [MvS07, SLS^{+}12]. Tiny [Sen14, All11a]. Tissue [SSH^{+}20]. tm [TFH12]. tmap [Ten18]. tmle [GvdL12]. Tolerance [You10, FBB20, You10]. tolerate [FBB20]. Tomasz [Mai14b, Oli17]. Tone [LA19]. Tool [LT16, Cal06, MS23, Wil12]. Toolbox [HBQ04]. Toolkit [LL10, Lan18]. Tools [AC04, AGM07, Dal98, Fuj17, Hew05, Lan14b, LBS17, MTPL15, OS95, PJSPC17, RG07, Woo98, Bor16a, Cox05, DG22, H^{+}96, HBPJ04, Sha22c]. Topic [GH11]. topicmodels [GH11]. topological [WWDS18]. Torsten [Che11, Pfa12, Ric11, Unw12a]. Torster [Ano06b]. Total [Lu18, SAR11]. Tour [IR12, Mat11]. tourr [WCHB11]. Toxicokinetics [PSS^{+}17]. trace [ATCA20]. trackeR [FK17]. Tracking [CPA20, FK17]. trackr [BML19]. Tractography [CMS^{+}11]. TractoR [CMS^{+}11]. Trade [IEE94, Tyn16]. training [SWCP20]. TraMineR [GRMS11]. Trans [Tyn16]. Trans-Atlantic [Tyn16]. Transarc [Ano99]. Transform [Mil17b, NS94, KOC21, WG17]. Transforming [SD18]. Transforms [CHT98]. transition [WRV23]. Transparent [LMY^{+}11, SvdLN17]. Treatment [LQC^{+}12, TM15, WMS17, WLH^{+}18, PFT^{+}12, SMWP20]. trec [SO23]. Tree [Arc10, Ano99, KMMV14]. Treed [Gra07, GT10]. Trees [CSNF18, GR16, GSD12, GZP14, Lee18, GZ21a, GZ21b]. Trellis [BCS96, Gru95, Sta95e]. trend [SO23]. Trends [BL14a]. Trevor [Nor14]. Trial [Dre19, Num13, She11, CP11b, Hu21].

Trials [JPM19, LQC^{+}12, USHH18, BVFB19, Hu21, KHJ^{+}23, PFT^{+}12, TPAM07]. triangular [Yan17]. Trigonometric [FDGD16]. Trimming [FGEMI12]. Tristan [Mai09]. tropical [BSS^{+}24]. Trosset [Ven10]. Trumbo [Laz11a]. Truncated [KL14, MdUÁC10, NK06]. truncSP [KL14]. Trust [Bra14]. trustOptim [Bra14]. trustworthy [BDT18]. Tsai [Asz23]. Tsay [Lep14]. TSclust [MV14]. tscount [LFF17]. TSrepr [Lau18]. tsrobprep [NKHZ21]. TuneR [Bor16a]. tuning [Bor16a]. Tunnel [The99]. Tuples [Hic16]. Turf [CH17]. Tutorial [QZLP21, TR14, AY22, EKP^{+}11, FBB20, HT19, HMRS14, SMM^{+}22]. Twin [Sch17c]. Two [DHF15, HSL11, LA19, Lub91, MMB15, Meu13, MVS13, MBR11, RV20, Rec10, RGD12, KK22, PMLM23]. Two- [RV20]. Two-Level [Rec10]. Two-Mode [Meu13]. Two-Phase [HSL11]. Two-Sample [MVS13]. Two-Stage [MMB15, KK22]. Two-Way [Meu13]. Two-Zone [MBR11]. Type [FO15, FM18a, MBT^{+}20]. TYPEical [MBT^{+}20]. Types [RG07, TGKV20].

U [Cur18]. U.K [Plu01]. U.S. [LPR21].

uFTIR [CBHLG21]. Ugarte

[Bat08b, Ng09]. Uhlenbeck [BI13]. UK [Rao14, Smi17]. UKCensusAPI [Smi17]. UK£ [Cur18]. Ultrahigh [SF18]. Ultrahigh-Dimensional [SF18]. Unbiased [PG15, KK22]. Uncertain [BSVT12]. Uncertainty [SS15, Cad13]. Unconstrained [LC10]. Understanding [Kur19, SA01, Vid22, Hos22a, Lun02, Mor03]. Unfolding [LBW18]. Unified [BBG16, Han07b, Lal17, SYC08]. uniform [Dev86, Fog17, OOSM18]. Unifying [NV11]. Unit [LRN18a, LBC^{+}16, Lup09]. Unit-Based [LBC^{+}16]. United [BL14b]. Univariate [FBC07, Sha22c, GS19, Hos22a, LF15, RK20]. Universal [MC97]. University [Pod18, Ree19, MN69]. UNIX [KVCS98, Mil98, Mil00]. Unknown [VV16]. Unleashing [LR15]. unmarked [FC11]. Unsupervised [AVS20, LIL^{+}15]. untb [Han07b]. unweighted [GH18, GHN19]. UpSet [GGS23]. Urdinez [Lip21c]. Urn [Han07a]. USA [Har23, Tu21, Mil15b]. UScensus2000 [Alm10]. USD [Gle16a, Grö15a, Grö16, Hel15, How16b, Kha16, Mat15, Mat16a, Rus15, Zei16]. Use [HBQ04, SA15, Spa17, VYD^{+}12, Cal06, Fil08, GV19]. Useful [RG07, Cad13]. User [AB12, Fox05b, HL09, KVCS98, LL10, Mat85, Sta92b, Sta93j, Mat16a, Mil00, RFKM12, SLS^{+}12, Sta93k, Sta95e, Sta97c, Sta98b, VML12, WLH^{+}18, Mil98, Mil02, OS95, Shi16, Sta95f]. Users [LA19, NV11, WDT^{+}12, Mue09, MH10, NPP18, Tav17, Ano09c, Gol11, Mai12, Ree09, Sau11, Voe09]. Using [All11b, Ano06b, Ano06a, Ano12b, Ano13, Arm19, BS15, Bar18, BFRP13, BB12, BLM^{+}15, Blo14, BPGC14, Bos12, BCAB07, BHW11, Bür17, Cap19, CG15, Car13, CP11a, Cha08b, CP12, Che11, CGC11, Cic15, CLL17, Dav95, DHM11, Dav07, Daw03, DN17, DB18, Dic12, Doe06, DM18, Dre19, DPSH18, Eas03, EJ13, Esp15, Eva11, EH06, FDGD16, Fie12, GKD14, GLC^{+}15, Grö10, Grö15b, Gut11, HD18, Han13a, HB92, Hel15, Hil06, Hil10, HK11, HK15, How16b, Hu09, IDE15, KMC^{+}12, Kas16, Kha17, Kim95, KF14, Ko95a, KY10, Kuh08, Kus03, LPLPD14, LW16, Laz24a, Laz24b, LRN18a, Lei10, Len09, LB12b, LHA^{+}15, Lu18, Lun02, Lun07b, Mai11, Mai14b, MMB15, MB15, Mar07, MBM18, MS11, Mil15b, MP06, MTvdM15, Mor03, Ng06, Ng11, Num13, Nun20]. Using [O’B08, O’B09a, Oom13, Otn17, PB15, Pfa12, Pla12, Pla19, QZLP21, RH02, RG96, Rec10, Ric11, RS05, Riz16, RMG12, Sab19, San10a, San11, SIR^{+}11, SZ11, SCK95, Sco11, Sco13a, Sha22a, She11, SS19b, SS92, SR16, SP10, Str10, SM07, TV11, TKM15, TM05, TR14, Tyn16, Unw13a, Veh13, Ver05, Ver14, Wee10, Wie04, Wil14a, Wol94, Wol99, XMW10, Yal10, ZQS16, Zie00, Zie02a, Zie02b, ZML16, dLM09b, dLM09c, Aji17, Ano03b, AY22, Asq14, BSS^{+}24, Bas18, Beh12, BC11, CP11b, Cra02, Cra05, Cra15, Dan18, Dév09, Dia18, DJS^{+}18, Dră12, Er12, Eve94, ERH01, Eve02, FMF12, Fin14, Fir03, FBNRG21, For20, Fri16, GZ21a, GB13, GSB19, GIH^{+}05, Gle16b, GGS23, GL15, Har03, HCSH15, HT19, HMRS14, Hu21, HLP11, JMR14, Kel10, KK22, KM01]. using [Lab12, Lon15, Mäc07, MB03, Man03, MC18, Mir14, NWH21, OOSM18, OSH^{+}21, OJMR09, PH16, PFT^{+}12, PK12, RN17, RKY11, Rob09, RVHG23, SMHBR06, SA01, Sel98, Sha22c, Sha21c, SWCP20, SMM^{+}22, Sta05, Tak12, TPAM07, Til96, TNM17, Tu21, Tur18, Val09, Vin08, Vin10, WTB^{+}15, XYC22, Yan95, ZMV^{+}18, ZHL11, ZM09, Ano06c, Gou05, Grö15a, Hou07, Kos15, Pol11c, Smi06, Wan16]. Usual [PZK^{+}12]. Utilities [Den16, OS95].

V [Doe06, Mai10, Spa17]. V. [Hoe09, Li11]. V6.8A [Ano99]. Vacuolating [TNM17]. Validating [CAA15]. Validation [BPDD08, CH18, JKvT^{+}14, KO06, Yan95]. Valiente [Lab12]. Value [CNA16, GK16, HS18, MMB15, BFA14]. Values [MF15]. VAR [Pfa08b]. Variable [CS12, HMS16, LL11, LBC^{+}16, ND12, RJH14, Riz06, Sch11, SR18, TMW18, WF12]. Variables [Bon18, CC08a, CKSLS12, Liu23, BF17, Har23, Zag18]. variance [GS19]. variate [Dev86]. variation [Bor16b]. vars [Pfa08b]. Varying [RC17, TR14, Hel22]. VAT [Cox05]. vdg [SlRC16]. vdmR [Fuj17]. Vector [AE21, KMH06, Mur09a, NL12, Tie09]. Vectorization [WPW15]. vectors [CHB14, GMF18]. Venables [Abr97, Arg98, Fer02, Jam96, Zie01b]. Venturini [Tu21]. Verdooren [AA12, Krä20]. verification [AY22]. Verlag [Abr97, Dal98, Har03, Hof15, How16b, Mat15, Mat16a, Rus15, Zei16]. versatile [MS23]. Version [Ano99, Cha95, GT10, HL04, HL05, Mat85, R D11a, R D11c, R D11d, R D11e, R D11f, R D11g, Sta93d, Sta93f, Sta93i, Sta93k, Sta95d, Sta95e, Sta95g, Sta95b, Mil02, RRSPTR14, VS02, Ven04]. Versions [Mil00]. Vertex [HL07]. vertical [Gav10]. Verzani [Ano06c, Hou07, Smi06]. vGameReviews [FVVL23]. Via [IDE15, AWBM18, AM14, BBGL17, CPA20, Fil08, Fus22, Gra16, IK20, KNI16, KF14, Lip20b, ML13, MP06, PW18, SKD22, SCS13, SWHJ17, Su07, WMS17, Wan13]. video [FVVL23]. View [Kal24b, Ger94, Mat94]. Views [The99]. Vignettes [Lei03]. Vikram [Hof15]. VIM [KT16]. Vine [BS13]. Violators [dLHM09]. virus [TNM17]. Visas [The99]. Visual [BCS96, Fuj17]. Visualisation [Hüs18, Rah17]. Visualisierungsbeispiele [Rah14]. Visualization [Agr16, Fer11, FM16, GGK10, Gro08, Kra07, Kuh10, Mül16, Nor08, OHD17, PU13, Sha21a, Wil14a, Coe18, Fri16, Kle09, LB21, MÇSD14, Mir14, Sar08, Aji17]. Visualizations [BML19, Die21, Ger20, Tan18]. Visualize [Mil17b, WG17]. Visualizing [BY18, BKT14, Cle93, FFM09, GRMS11, Gio09, Kie08, Lee18, LM03, MBT^{+}20, Tyn16, BLW12, Tur18]. VNM [HWY18]. Volatility [Kas16]. vollstandigen [Kra97]. Volume [TW11, VR99, VR00a, SS18a, SS18b]. Votes [PLLC11]. vrmlgen [GGK10]. vtreat [MZ18].

W [Abr97, All11b, Arg98, Bai11, Den98, Die21, Doe06, Doe10, Fer02, Gut11, Hly09, Jam96, Nor09, Ros00, Sha21c, Zie01b, Lip21b, Lip21c]. W. [Ven10]. Walker [Oli10a, Sch09, Hel22]. Walter [Lu18]. Wand [Num20]. Wang [Fis06, Ano03a, Bur07, Fot07, Kum07, McN04, Tol03]. Warping [CS12, Gio09, MCAP19, SCS13]. Wars [Kur19]. Washington [HI97, IEE94]. water [CC23]. Wavelet [CHT98, EN11, FM19, Han98, NS94, Nas08, BG96, BL11, Mor09, O’B09c]. Wavemulcor [FM19]. Way [GKD14, Kur19, Meu13, MvS07, RV20]. Wayne [Has18]. ways [Urb09]. Weather [RVHG23, SHN17]. Web [Iac15, Mun14b, Sel17, Sha21a, KF17, LB21, Fir03, Fuj17, GGK10, HBA19, Mun14a, NL14, RL15, SVC^{+}19, VFV13, Ver12]. Web-Based [Sha21a, LB21, Fuj17]. website [FM18b]. Wedding [MC97]. Weighted [CG15, DPSH18, FS10a, GLC^{+}15, MCAP19, WS11, GH18, GHN19, LAJJ18, SOD^{+}16]. Weighting [vdWG11]. weights [GGS23]. Weihs [Cur18]. Weka [HBZ09]. Well [SH17]. Westfall [Dic12, Che11, Ric11]. wgaim [TV11]. WGEEs [SOD^{+}16]. WhatIf [SKZ05]. Whitney [FO15]. Wickham [Mil17b, Wil11a]. Wikle [Peb21, Pet19]. Wildlife [FC11]. Wiley [Ano03b, Hel15, Iac15, Lip21a, Bas18, Sel17]. Williams [Sau12]. Wim [Oli17]. WinBUGS [LB12b, SLG05]. Windows [PG15, SGHY11, Tak12, Hil96, KVCS98, Mat85, Mil98, Mil00]. Wingz [Ano99]. Wisconsin [MN69]. Wissenschaften [DK09]. Within [Mug10, Pfa08b, Ver12, ZPC^{+}16]. Without [Han07a, BG18]. Witten [Nor14]. wnominate [PLLC11]. Wood [Chi07, Rob07, Sta07b]. Woodward [Has18]. word [DS20]. word.alignment [DS20]. WordPerfect [Ano99]. WordPress [BP17a]. Work [MBM18, LBS17]. worked [PG20, Sha19]. Worker [KSC^{+}00]. Workflow [GL16b]. Working [Gle16b, Lan09b, WST11]. World [FM18b]. Wrangling [MH18]. Wrapper [CC08a].

Wrappers [WLK08]. Write [EW16]. Writing [R D11g]. Wu [Cho20, Hu21]. www.amherst.edu [Grö15a]. www.crcpress.com [Gle16a, Grö16, Kha16]. www.imperial.ac.uk [Hel15]. www.r-datacollection.com [Iac15]. www.springer.com [Hof15, How16b, Mat15, Mat16a, Rus15, Zei16].

x [Mai21, Put06, Sha23b, SE18, Til96]. X-13ARIMA-SEATS [SE18]. x12 [KMTS14]. x12GUI [KMTS14]. XDrift [BSG20]. xi [CCP^{+}11, Num20]. Xie [Joh20, Vey14]. xii [Kal24b]. xiii [Abr97, Asz23, Sha21c]. Xin [Cho20]. xiv [Cox05, Len20]. xix [Lip21c]. XML [NL14, Mun14b]. xsample [VSV09]. Xu [Har23, Liu23, CCP^{+}11]. xv [Har23, Liu23, Plu01, Sta21, Sta05]. xvi [Hou07, Lip22, Mal21]. xvii [Ben21, Cho22a, Hac17, Lee21, Lip21a]. xviii [Har03, Kha16]. xxi [Kal24a, Oli17]. xxii [Kel10, Pod21, Sha23a]. xxiii [Sab19]. xxx [AS24, Lip20a, Lip20b]. xxxiv [Cap19, Joh20].

yaImpute [CF08]. yan [CCP^{+}11]. Yanagida [Veh13]. Ye [Han98]. Years [Bro14, Don17, HJ17, Pen17]. yield [CPB^{+}23]. Yihui [Joh20, Vey14]. ying [CCP^{+}11]. yong [CCP^{+}11]. York [Abr97, Ano03b, Cox05, Dal98, How16b, Mat15, Mat16a, Nun20]. Yu [Laz24b, CCP^{+}11]. Yuelin [Mai14a]. Yuima [BI13, IM15].

Z [BC76, BC77]. Zammit [Peb21, Pet19]. Zammit-Mangion [Peb21, Pet19]. Zaven [Rob12]. Zeileis [Mye09, Sco09]. Zelterman [Zei16]. Zero [LC10, MAY23]. zeroaltered [MAY23]. Zero-Inflated [LC10, MAY23]. Zhao [Vis10]. Zieffler [Han13a]. Ziv [Oli17]. Zivot [Ano03a, Bur07, Fot07, Kum07, McN04, Tol03, Fis06]. Zobitz [Laz24a]. Zoë [Unw13a]. Zone [MBR11]. Zucchini [All11b, Gut11, Lu18]. zum [Ano97]. Zumel [Mil17a]. Zurbenko [ZS17]. Zuur [Bai11, Oli10a, Sch09].

References

Adams:2012:BRB

[AA12] Wayne Adams and Mark Anderson. Book review: Optimal Experimental Design with R by Dieter Rasch; Jurgen Pilz; L. R. Verdooren; Albrecht Gebhardt. Technometrics, 54 (2):203, May 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                                      //www.jstor.org/stable/
                                                      41714877.

Alhamzawi:2020:BRP

[AA20] Rahim Alhamzawi and Haithem Taha Mohammad Ali. Brq: an R package for Bayesian quantile regression. Metron, 78(3):313– 328, December 2020. CO-DEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic). URL

                                                      https://link.springer.
                                                      com/article/10.1007/s40300-
                                                      020-00190-6.

Azzalini:1990:LSD

[AB90] A. Azzalini and A. W. Bowman. A look at some data on the Old Faithful Geyser. Applied Statistics, 39(3):357–365, 1990.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Austenfeld:2012:GUI

[AB12] Marcel Austenfeld and Wolfram Beyschlag. A graphical user interface for R in a rich client platform for ecological modeling. Journal of Statistical Software, 49(4):??, June 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v49/i04.

Anderson-Bergman:2017:IRM

[AB17a] Clifford Anderson-Bergman.

icenReg: Regression models for interval censored data in R. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081i12;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081i12/
v81i12.pdf.

Ardia:2017:PNC

[AB17b] David Ardia and Keven Bluteau. nse: Computation of numerical standard errors in R. Journal of Open Source Software, 2 (10):172:1–172:2, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00172.

Ardia:2019:GAS

[ABC19] David Ardia, Kris Boudt, and Leopoldo Catania. Generalized autoregressive score models in R: The GAS package. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v088i06;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v088i06/
v88i06.pdf.

Arel-Bundock:2018:PCR

[ABEY18] Vincent Arel-Bundock, Nils Enevoldsen, and C. J.

Yetman.  countrycode:

An R package to convert country names and country codes. Journal of Open Source Software, 3(28): 848:1–848:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00848.

Ardia:2017:PRC

[ABGF17] David Ardia, Kris Boudt, and Jean-Philippe Gagnon-

Fleury. RiskPortfolios:

Computation of risk-based portfolios in R. Journal of Open Source Software, 2(10):171:1, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00171.

Abrams:1997:BRM

[Abr97] Keith R. Abrams. Book review: Modern Applied Statistics with S-PLUS. W. N. Venables and B. D. Ripley, Springer-Verlag, New York, 1994. No. of pages: xiii + 462 + 3.5 inch diskette. Price: £24.50. ISBN: 0-387-94350-1. Statistics in Medicine, 16(13):1548–1550, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Amini:2023:HRP

[ABS23] Morteza Amini, Afarin Bayat, and Reza Salehian. hhsmm: an R package for hidden hybrid Markov/semi-Markov models. Computational Statistics, 38(3): 1283–1335, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
022-01248-x.

Anderson-Cook:2004:BRBc

[AC04] Christine M. Anderson-Cook. Book review: Statistical Tools for Nonlinear Regression: A Practical Guide with S-PLUS and R Examples, second edition, by S. Huet; A. Bouvier; M.-A. Poursat; E. Jolivet. Journal of the American Statistical Association, 99(467):902– 903, September 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

                                        /miranda.asa.catchword.
                                        org/vl=1515603/cl=66/nw=
                                        1/rpsv/cw/asa/01621459/
                                        v99n467/s52/p904;   http:
                                        //www.jstor.org/stable/
                                        27590464.

Anderson-Cook:2007:BRBb

[AC07] Christine M. Anderson-Cook. Book review: Generalized Additive Models: An Introduction With R. Journal of the American Statistical Association, 102(478): 760–761, June 2007. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Altieri:2016:BPS

[ACG^{+}16] L. Altieri, D. Cocchi, F. Greco, J. B. Illian, and E. M. Scott. Bayesian Psplines and advanced computing in R for a changepoint analysis on spatio-temporal point processes. Journal of Statistical Computation and Simulation, 86(13): 2531–2545, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Adragni:2012:GRP

[ACW12] Kofi Placid Adragni, R. Dennis Cook, and Seongho Wu. GrassmannOptim: An R package for Grassmann manifold optimization. Journal of Statistical Software, 50(5):??, July 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i05.

Amatya:2015:CSO

[AD15] Anup Amatya and Hakan Demirtas. Code snippet: OrdNor: An R package for concurrent generation of correlated ordinal and normal data. Journal of Statistical Software, 68(CS-2):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068c02;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v068c02/v68c02.pdf.

Amatya:2017:PRP

[AD17] Anup Amatya and Hakan Demirtas. PoisNor: An R package for generation of multivariate data with Poisson and normal marginals. Communications in Statistics: Simulation and Computation, 46(3):2241–2253, 2017. CODEN CSSCDB. ISSN 0361-0918.

Abadie:2011:SRP

[ADH11] Alberto Abadie, Alexis Diamond, and Jens Hainmueller. Synth: An R package for synthetic control methods in comparative case studies. Journal of Statistical Software, 42 (13):??, June 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v42/i13.

Adler:2010:RN

[Adl10]  Joseph Adler.  R in a

nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2010. ISBN 0-596-80170-X. xx + 611 pp. LCCN QA276.45.R3eb; QA276.45.R3; Internet.

Adler:2012:RN

[Adl12] Joseph Adler. R in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2012. ISBN 1-4493-1208-X (paperback). xix + 699 pp. LCCN ???? EUR 41.00.

Appelhans:2015:REO

[ADN15] Tim Appelhans, Florian Detsch, and Thomas Nauss. remote: Empirical orthogonal teleconnections in R. Journal of Statistical Software, 65(10):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v65/
                                        i10.

Arroyo:2021:ARI

[AE21] Daisy Arroyo and Xavier Emery. Algorithm 1013: an R implementation of a continuous spectral algorithm for simulating vector Gaussian random fields in Euclidean spaces. ACM Transactions on Mathematical Software, 47(1):8:1–8:25, January 2021. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3421316.

Abdollahi:2022:CRB

[AFGH22] Masoud Abdollahi, Babak Farjad, Anil Gupta, and Quazi K. Hassan. CMIP6-D&A: an R-based software with GUI for processing climate data available in network common data format. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022000395.

Aitkin:2009:SMR

[AFHD09] Murray A. Aitkin, Brian Francis, John Hinde, and Ross Darnell. Statistical modelling in R, volume 35 of Oxford statistical science series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-921913-3 (paperback), 0-19-921914-1 (hardcover). xii + 576 pp. LCCN QA276.45.R3 S73 2009. URL http://

www.loc.gov/catdir/toc/
ecip0824/2008031898.html.

Azais:2022:CRP

[AFM22] R. Azaïs, S. Ferrigno, and M.-J. Martinez. cvmgof: an R package for Cramér–von Mises goodness-of-fit tests in regression models. Journal of Statistical Computation and Simulation, 92(6): 1246–1266, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ayala:2020:RLN

[AGdSC20] Helon Vicente Hultmann Ayala, Marcos Cesar Gritti, and Leandro dos Santos Coelho. An R library for nonlinear blackbox system identification. SoftwareX, 11(??):Article 100495, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711019302663.

Alfaro:2013:ARP

[AGG13] Esteban Alfaro, Matias Gamez, and Noelia García. adabag: An R package for classification with boosting and bagging. Journal of Statistical Software, 54(2): ??, September 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v54/i02.

Atman:2007:ATA

[AGM07] Micah Atman, Jeff Gill, and Michael P. McDonald. accuracy: Tools for accurate and reliable statistical computing. Journal of Statistical Software, 21 (1):1–30, July 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v21/i01.

Agresti:2013:CDA

[Agr13] Alan Agresti. Categorical data analysis, volume 792 of Wiley series in probability and statistics. Wiley, New York, NY, USA, third edition, 2013. ISBN 0-470-46363-5 (hardcover). xvi + 714 pp. LCCN QA278 .A353 2013.

Agresti:2016:BRF

[Agr16] Alan Agresti. Book review: Friendly, Michael and Meyer, David. Discrete Data Analysis with R: Visualization and Modeling Techniques for Categorical and Count Data. Boca Raton, FL: CRC Press. Biometrics, 72(2):659–660, June 2016. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ardia:2009:AMS

[AHvD09] David Ardia, Lennart F.

Hoogerheide, and Herman K. van Dijk. Adaptive mixture of Student-t distributions as a flexible candidate distribution for efficient simulation: The R package AdMit. Journal of Statistical Software, 29(3): ??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/i03.

Aizaki:2012:CSB

[Aiz12] Hideo Aizaki. Code snippet: Basic functions for supporting an implementation of choice experiments in R. Journal of Statistical Software, 50(CS-2): ??, September 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v50/c02.

Ajiferuke:2017:BRS

[Aji17] Isola Ajiferuke. Book review: Statistics for library and information services: a primer for using open source R software for accessibility and visualization. Alon Friedman. Lanham, MD: Rowman & Littlefield Publishers, 2015. 376 pp. $95.00 (hardback). (ISBN 978-1-4422-4992-9). Journal of the Association for Information Science and Technology, 68(6):1594–1595, June 2017. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Arbel:2021:BBN

[AKL^{+}21] J. Arbel, G. Kon Kam King, A. Lijoi, L. Nieto-Barajas, and I. Prünster. BNPdensity: Bayesian nonparametric mixture modelling in R. Australian & New Zealand Journal of Statistics, 63(3):542–564, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Allon:2016:PPR

[AL16] Ayala Allon and Roy Luria.

prepdat — an R package for preparing experimental data for statistical analysis. Journal of Open Research Software, 4 (1):e43–??, November 25, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.134/.

Albert:2007:BCR

[Alb07] Jim Albert. Bayesian computation with R. Use R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-71384-0 (paperback). x + 267 pp. LCCN QA279.5 .A53 2007.

URL http://www.loc.gov/
catdir/enhancements/fy0824/
2007929182-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0824/2007929182-t.html.

Albert:2009:BCR

[Alb09] Jim Albert. Bayesian computation with R. Use R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-92297-0 (paperback). xii + 298 pp. LCCN QA279.5 .A53 2009.

Alberti:2015:CRP

[Alb15] Gianmarco Alberti. CAinterprTools: an R package to help interpreting Correspondence Analysis’ results. SoftwareX, 1–2(??):19–25, September 2015. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711015000060.

Albano:2016:ERP

[Alb16] Anthony D. Albano. equate:

An R package for observedscore linking and equating. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v074i08;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v074i08/
                                        v74i08.pdf.

Alberti:2019:MRP

[Alb19] Gianmarco Alberti. movecost: an R package for calculating accumulated slopedependent anisotropic costsurfaces and least-cost paths. SoftwareX, 10(??):Article 100331, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711019302341.

Aldridge:2020:CSB

[Ald20] Caleb A. Aldridge. Computing skills beyond R for fisheries students. Fisheries, 45(3):116–120, March 2020. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:/

/afspubs.onlinelibrary.
wiley.com/doi/10.1002/
fsh.10392.

Abreu:2010:HDG

[ALE10] Gabriel C. G. Abreu, Rodrigo Labouriau, and David Edwards. High-dimensional graphical model search with the gRapHD R package. Journal of Statistical Software, 37 (1):??, November 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v37/i01.

Allen:1986:CSS

[All86] David M. Allen, editor.

Computer science and statistics: Proceedings of the Seventeenth Symposium on the Interface of Computer Sciences and Statistics, Radisson Plaza Hotel, Lexington, Kentucky, March 17– 19, 1985. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL

http://catalog.hathitrust.
org/api/volumes/oclc/13455449.
html;   http://www.dtic.
mil/dtic/tr/fulltext/u2/
a166019.pdf;   http://
zbmath.org/?q=an:0647.
62008.

Allerhand:2011:THR

[All11a] Mike Allerhand. A tiny handbook on R. Springer-Briefs in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-17979-7, 3-642-17980-0. ix + 83 pp. LCCN QA276.45.R3 A45

                                        2011.  URL http://d-
                                        nb.info/1008556300/04;
                                         http://link.springer.
                                        com/openurl?genre=book&
                                        isbn=978-3-642-17979-2.

Alliot:2011:BRH

[All11b] Pierre Alliot. Book review: Hidden Markov Models for Time Series: an Introduction Using R, by W. Zucchini and I. L. Macdonald. Australian & New Zealand Journal of Statistics, 53(2): 265–266, June 2011. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Almquist:2010:UCS

[Alm10] Zack W. Almquist. US census spatial and demographic data in R: The

                                        UScensus2000 suite of pack-

ages. Journal of Statistical Software, 37(6):??, November 2010. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v37/i06.

Azzalini:2014:CND

[AM14] Adelchi Azzalini and Giovanna Menardi. Clustering via nonparametric density estimation: The R package pdfCluster. Journal of Statistical Software, 57 (11):??, May 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v57/i11.

Amornbunchornvej:2021:MRP

[Amo21] Chainarong Amornbunchornvej. mFLICA: an R package for inferring leadership of coordination from time series. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000935.

Augugliaro:2014:PRP

[AMW14] Luigi Augugliaro, Angelo Mineo, and Ernst C. Wit. dglars: An R package to estimate sparse generalized linear models. Journal of Statistical Software, 59(8): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v59/i08.

Agostinelli:2016:REG

[AMYR16] Claudio Agostinelli, Alfio Marazzi, Víctor J. Yohai, and Alex Randriamiharisoa. Robust estimation of the generalized loggamma model: The R package

robustloggamma. Journal

of Statistical Software, 70 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v070i07;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v070i07/v70i07.pdf.

Anhoj:2018:PQQ

[Anh18] Jacob Anhøj. qicharts2: Quality improvement charts for R. Journal of Open Source Software, 3(25): 699:1–699:2, May 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00699.

Anonymous:1995:ACA

[Ano95a] Anonymous, editor. 20th Annual conference: — April 1995, Orlando, FL, SUGI 1995; 20th. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-211-0. LCCN ????

Anonymous:1995:S

[Ano95b] Anonymous. S-PLUS. In Bartley [Bar95], page 299. ISSN 0265-1491.

Anonymous:1996:SBP

[Ano96] Anonymous. S+SpatialStats builds on S-Plus technology. Silicon Graphics World, 6 (6):15, June 1996. ISSN 1057-7041.

Anonymous:1997:PLD

[Ano97] Anonymous. S-PLUS: learning by doing; eine Anleitung zum Arbeiten mit S-PLUS. Lucius und Lucius, Stuttgart, Germany, 1997. ISBN 3-8282-0049-4. xii + 126 pp. LCCN ????

Anonymous:1998:S

[Ano98] Anonymous. S-PLUS 4.5.

Duxbury Press, North Scituate, MA, USA, student edition, 1998. ISBN 0-534-35873-X. 440 pp. LCCN ???? Includes CD-ROM.

Anonymous:1999:NPE

[Ano99] Anonymous. New products: ENVI Version 3.1, Research Systems, Inc.; FTLinuxCourse, Future Technologies; Alexandria 4.50, Spectra Logic Corporation; Applixware for Linux, Applix, Inc.; c-tree Plus V6.8A, FairCom Corporation; AFS Server and AS Client, Transarc; Hamilton, Microstate Corporation; Cat5 Reach, Raritan Computer; TclPro 1.1, Scriptics Corporation; Wingz Professional 3.0 for Linux, IISC; Word-Perfect 8 for Linux, Corel Computer; NUOPT for S-PLUS, MathSoft. Linux Journal, 59:??, March 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2003:BRBam

[Ano03a] Anonymous. Book review:

Modeling Financial Time Series with S-PLUS(R) by Eric Zivot; Jiahui Wang. Technometrics, 45(4):373– 374, November 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25047114.

Anonymous:2003:BRSa

[Ano03b] Anonymous. Book review: Statistical computing: an introduction to data analysis using S-Plus: by Michael J. Crawley. John Wiley & Sons, New York. (2002). 761 pages. $85. Computers and Mathematics with Applications, 45(10–11):1775, May/June 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0898122103801311.

Anonymous:2003:BRBu

[Ano03c] Anonymous. Book review: The Basics of S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 45 (2):183, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25047039.

Anonymous:2006:BRHa

[Ano06a] Anonymous. Book review: A Handbook of Statistical Analyses Using R, Edited by Everitt, B. S. and Hothorn, T. Biometrics, 62(4):1279, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRBam

[Ano06b]  Anonymous.  Book re-

view: A Handbook of Statistical Analysis Using R by Brian Everitt; Torster Northorn. Technometrics, 48(4):577, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471266.

Anonymous:2006:BRU

[Ano06c] Anonymous. Book reviews:

Using R for Introductory Statistics, by J. Verzani and An R and S-Plus Companion to Multivariate Analysis by B. S. Everitt. Biometrics, 62(1):311, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BBP

[Ano06d] Anonymous. The Basics of S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 48(2):314, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471184.

Anonymous:2008:BRAb

[Ano08] Anonymous. Book review:

Analysis of Correlated Data with SAS and R, 3rd edition. by Shoukri, M. M. and Chaudhary, M. A. Biometrics, 64(2):660–661, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRBf

[Ano09a] Anonymous. Book review: Introductory Statistics With R (2nd ed.) by Peter Dalgaard. Technometrics, 51 (2):225, May 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        40586615.

Anonymous:2009:BRN

[Ano09b] Anonymous. Book review: Nonlinear Regression with R, by Ritz, C. and Streibig, J. C. Biometrics, 65(3):1001, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRRb

[Ano09c] Anonymous. Book review: R for SAS and SPSS Users, by Muenchen, R. A. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRSc

[Ano09d] Anonymous. Book review: Software for Data Analysis: Programming with R, by Chambers, J. M. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2010:BRBb

[Ano10] Anonymous. Book review:

Analysis of Integrated and Cointegrated Time Series With R (2nd ed.) by Bernhard Pfaff. Technometrics, 52(1):145, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40586696.

Anonymous:2012:BRBj

[Ano12a]  Anonymous.  Book re-

view: Analysis of Phylogenetics and Evolution With R (2nd ed.) by Emmanuel Paradis. Technometrics, 54 (3):330–331, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714914.

Anonymous:2012:BRBm

[Ano12b] Anonymous. Book review:

Statistics and Data Analysis for Microarrays Using R and Bioconductor, 2nd Edition by Sorin Dräghici. Technometrics, 54(4):445– 446, November 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        41714939.

Anonymous:2013:URA

[Ano13]  Anonymous.  Using the

R advanced statistical package. Linux Journal, 2013 (232):1:1–1:??, August 2013. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2016:BRBb

[Ano16]  Anonymous.  Book re-

view: The Book of R, by Tilman Davies. No Starch Press. ISBN 978-1-59327-651-5. Network Security, 2016(9):4, September 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1353485816300848.

Albert:2012:RE

[AR12] Jim Albert and Maria L. Rizzo. R by example. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-1364-8, 1-4614-1365-6 (e-book). xiii + 359 pp. LCCN QA276.45.R3 A53 2012.

Adragni:2014:PRS

[AR14] Kofi Placid Adragni and Andrew M. Raim. ldr: An R software package for likelihood-based sufficient dimension reduction. Journal of Statistical Software, 61 (3):??, November 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i03.

Atkinson:2004:EMD

[ARC04] Anthony C. Atkinson, Marco Riani, and Andrea Cerioli. Exploring Multivariate Data with the Forward Search, volume ?? of Springer Series in Statistics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN ???? ISBN 0-387-21840-8. ISSN 0172-7397. xxi + 621 pp. LCCN QA278 .A85 2004. URL http://

link.springer.com/book/
10.1007/978-0-387-21840-
3.

Archer:2010:RRP

[Arc10] Kellie J. Archer. rpartOrdinal:

An R package for deriving a classification tree for predicting an ordinal response. Journal of Statistical Software, 34(7):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v34/
i07.

Argyle:1998:BRBb

[Arg98]  Jennifer Argyle.  Book

review: Modern Applied Statistics with S-PLUS, by W. N. Venables; B. D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 47(4):

730, ???? 1998. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        2988400.

Armstrong:2014:ASM

[Arm14] David A. Armstrong II. Analyzing spatial models of choice and judgment with R. Chapman and Hall/CRC statistics in the social and behavioral sciences series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-1715-8 (hardcover). xx + 336 pp. LCCN QA278.2 .A76 2014. URL http://

                                        www.crcnetbase.com/doi/
                                        book/10.1201/b16486.

Armero:2019:BRD

[Arm19] Carmen Armero. Book review: Data Analysis Using Hierarchical Generalized Linear Models with R. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v088b01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v088b01/
                                        v88b01.pdf.

Arnold:2017:PKR

[Arn17] Taylor B. Arnold. kerasR: R interface to the Keras Deep Learning Library. Journal of Open Source Software, 2(14):296:1, June 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00296.

Aalabaf-Sabaghi:2024:BRL

[AS24] Morteza Aalabaf-Sabaghi.

Book review: Learning Microeconometrics with R, by Christopher P Adams , CRC Press , Chapman & Hall. 2021. xxx + 368pp. ISBN: 978-0-367-25538-1, £77.99 (hbk). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187 (2):545–546, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

academic.oup.com/jrsssa/
article/187/2/545/7263080.

Asquith:2014:PEP

[Asq14] William H. Asquith. Parameter estimation for the 4parameter Asymmetric Exponential Power distribution by the method of L-moments using R. Computational Statistics & Data Analysis, 71(??):955–970, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947312004471.

Aszani:2023:BRM

[Asz23] Aszani Aszani. Book review: Machine Learning for Knowledge Discovery with R: Methodologies for Modeling, Inference, and Prediction, Kao-Tai Tsai, Boca Raton, FL: CRC Press, Taylor Francis Group, LLC, 2022, xiii + 260 pp., $88.00, ISBN: 978-1-032-06536-6 (H). Technometrics, 65(4):607–609, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Alfons:2013:ESE

[AT13] Andreas Alfons and Matthias Templ. Estimation of social exclusion indicators from complex surveys: The R package laeken. Journal of Statistical Software, 54(15): ??, September 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v54/i15.

Ardito:2020:PRP

[ATCA20] Luca Ardito, Marco Torchiano, Riccardo Coppola, and Giulio Antoniol. PowTrAn: an R package for power trace analysis. SoftwareX, 12(??):Article 100512, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S235271102030025X.

Alfons:2010:OOF

[ATF10] Andreas Alfons, Matthias Templ, and Peter Filzmoser. An object-oriented framework for statistical simulation: The R package simFrame. Journal of Statistical Software, 37(3): ??, November 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v37/i03.

Albayrak:2020:KRP

[ATYK20] Muammer Albayrak, Kemal Turhan, Yasemin Yavuz, and Zeliha Aydin Kasap. kaphom: an R package for testing the homogeneity of intra-class kappa statistics. Communications in Statistics: Simulation and Computation, 49 (12):3283–3298, 2020. CO-DEN CSSCDB. ISSN 0361-0918.

Ahsen:2020:PRP

[AVS20] Mehmet Eren Ahsen, Robert Vogel, and Gustavo A. Stolovitzky. R/PY-SUMMA: an R/Python package for unsupervised ensemble learning for binary classification problems in bioinformatics. Journal of Computational Biology, 27(9):1337–1340, September 2020. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.
liebertpub.com/doi/abs/
10.1089/cmb.2019.0348;
 https://www.liebertpub.
com/doi/pdf/10.1089/cmb.
2019.0348.

Andrews:2018:TRP

[AWBM18] Jeffrey L. Andrews, Jaymeson R. Wickins, Nicholas M. Boers, and Paul D. McNicholas. teigen: An R package for model-based clustering and classification via the multivariate t distribution. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v083i07;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v083i07/
                                        v83i07.pdf.

Arifin:2022:CPV

[AY22] Wan Nor Arifin and Umi Kalsom Yusof. Correcting for partial verification bias in diagnostic accuracy studies: a tutorial using R. Statistics in Medicine, 41(9):1709– 1727, April 30, 2022. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bowman:1997:AST

[BA97] Adrian W. Bowman and Adelchi Azzalini. Applied smoothing techniques for data analysis: the kernel approach with S-Plus illustrations, volume 18 of Oxford statistical science series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-852396-3. xi + 193 pp. LCCN QA278.B68 1997.

Baclawski:2008:IPR

[Bac08] Kenneth Baclawski. Introduction to probability with R, volume 75 of Chapman and Hall/CRC texts in statistical science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-4200-6521-1 (hardcover). xvi + 363 pp. LCCN QA273 .B2535 2008.

URL http://www.loc.gov/
catdir/enhancements/fy0803/
2007041288-d.html;   http:
//www.loc.gov/catdir/toc/
ecip082/2007041288.html.

Bailey:2011:BRB

[Bai11] Mark Bailey. Book review:

A Beginner’s Guide to R by Alain F. Zuur; Elena N. Ieno; Erik H. W. G. Meesters. Technometrics, 53(1):105, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40997308.

Bakouch:2013:RS

[Bak13]  Hassan S. Bakouch.  R

for statistics. Journal of Applied Statistics, 40(4): 924, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Bartley:1995:FET

[Bar95] M. R. Bartley, editor. Field experiment techniques: 11– 13 December 1995, Churchill College, Cambridge, volume 43 of Aspects of Applied Biology. The Association of Applied Biology, Wellesbourne, Warwick, UK, 1995. ISSN 0265-1491.

Barber:2002:BRBb

[Bar02] Stuart Barber. Book review: The Basics of S and S-PLUS, by A. Krause; M. Olson. Journal of the Royal Statistical Society. Series D (The Statistician), 51(2):296–297, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
3650339.

Barcaroli:2014:PRP

[Bar14] Giulio Barcaroli. SamplingStrata: An R package for the optimization of stratified sampling. Journal of Statistical Software, 61(4):??, November 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i04.

Barber:2018:BRF

[Bar18] Xavier Barber. Book review: Flexible Regression and Smoothing: Using GAMLSS in R. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085b02;   https:/

/www.jstatsoft.org/index.
php/jss/article/view/v085b02/
v85b02.pdf.

Beyersmann:2012:CRM

[BAS12] Jan Beyersmann, Arthur Allignol, and Martin Schumacher. Competing risks and multistate models with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-2034-2 (paperback), 1-4614-2035-0 (e-book). xi + 245 pp. LCCN QA276.4 .B49 2012.

Basu:2018:BRC

[Bas18] Sumanta Basu. Book review:

Cichosz, Pawel. Data Mining Algorithms explained using R. Hoboken, Wiley. Biometrics, 74(4):1519–1520, December 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bates:2004:BRB

[Bat04] Samantha C. Bates. Book review: Introductory Statistics with R. The American Statistician, 58(1):88– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//oberon.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/00031305/
v58n1/s20/p88.

Bates:2008:SBRa

[Bat08a] Douglas M. Bates. Short book review: Data Manipulation with R by Phil Spector. International Statistical Review = Revue Internationale de Statistique, 76 (2):324–325, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919643.

Bates:2008:SBRb

[Bat08b] Douglas M. Bates. Short book review: Probability and Statistics with R by Maria Dolores Ugarte, Ana F. Militino, Alan T. Arnholt. International Statistical Review = Revue Internationale de Statistique, 76 (3):447–448, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919667.

Battauz:2015:ERP

[Bat15] Michela Battauz. equateIRT: An R package for IRT test equating. Journal of Statistical Software, 68(7): ??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068i07;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v068i07/v68i07.pdf.

Baumer:2014:BRA

[Bau14]  Ben S. Baumer.  Book

review: Analyzing Baseball Data with R by Max Marchi, Jim Albert. International Statistical Review = Revue Internationale de Statistique, 82(2):313– 315, August 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Bergmeir:2012:NNR

[BB12] Christoph Bergmeir and José M. Benítez. Neural networks in R using the Stuttgart Neural Network Simulator: RSNNS. Journal of Statistical Software, 46 (7):??, January 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/i07.

Becue-Bertaut:2018:TDS

[BB18] Monica Bécue-Bertaut. Textual Data Science with R. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2018. ISBN 1-138-62691-0 (hardcover), 1-315-21266-8 (e-book). xvii + 194 pp. LCCN HF5548.2 .B36116 2018.

Berge:2012:HRP

[BBG12] Laurent Bergé, Charles Bouveyron, and Stéphane Girard. HDclassif: An R package for model-based clustering and discriminant analysis of high-dimensional data. Journal of Statistical Software, 46(6):??, January 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/i06.

Bartolucci:2014:MRP

[BBG14] Francesco Bartolucci, Silvia Bacci, and Michela Gnaldi. MultiLCIRT: an R package for multidimensional latent class item response models. Computational Statistics & Data Analysis, 71 (??):971–985, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947313002053.

Bartolucci:2016:SAQ

[BBG16] Francesco Bartolucci, Silvia Bacci, and Michela Gnaldi. Statistical Analysis of Questionnaires: a Unified Approach Based on R and Stata. Chapman and Hall/CRC interdisciplinary statistics series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4665-6849-6 (hardcover). xxi + 308 pp. LCCN HA32 .B36 2016.

Becker:2017:ERC

[BBGL17] Gabriel Becker, Cory Barr, Robert Gentleman, and Michael Lawrence. Enhancing reproducibility and collaboration via management of R package Cohorts. Journal of Statistical Software, 82(??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v082i01;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v082i01/v82i01.pdf.

Bulla:2010:HRP

[BBN10] Jan Bulla, Ingo Bulla, and Oleg Nenadić. hsmm — an R package for analyzing hidden semi-Markov models. Computational Statistics & Data Analysis, 54 (3):611–619, March 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016794730800426X.

Becker:1976:GZS

[BC76] Richard A. Becker and John M. Chambers. GR-Z: a system of graphical subroutines for data analysis. In David C. (David Caster) Hoaglin and Roy E. Welsch, editors, Proceedings of the 9th Interface Symposium on Computer Science and Statistics: Harvard University, Massachusetts Institute of Technology, April 1–2, 1976, page ?? Prindle, Weber, Schmidt, Boston, MA, USA, 1976. ISBN 0-87150-237-2. LCCN QA276.A1 I53

Becker:1977:GZS

[BC77] Richard A. Becker and John M. Chambers. GR-Z: a system of graphical subroutines for data analysis. In David C. (David Caster) Hoaglin and Dennis W. Fife, editors, Computer Science and Statistics: Tenth Annual Symposium on the Interface: Proceedings of the 10th Annual Symposium held at the National Bureau of Standards, Gaithersburg, Maryland, April 14– 15, 1977, pages 409–415. United States Government Printing Office, Washington, DC, USA, 1977. ISBN 1-390-52421-3. LCCN QA276.A1 I53 1976. NBS Special publication 503.

Becker:1981:LSD

[BC81] Richard A. Becker and John M. Chambers. S: a Language and System for Data Analysis. AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981.

Becker:1984:DSD

[BC84a] R. A. Becker and J. M. Chambers. Design of the S system for data analysis. Communications of the ACM, 27(5):486–495, May 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Becker:1984:IED

[BC84b] Richard A. Becker and John M. Chambers. S: An Interactive Environment for Data Analysis and Graphics. Wadsworth, Pacific Grove, CA, USA, 1984. ISBN 0-534-03313-X. xii + 550 pp. LCCN QA76.73.S15 B44 1984.

Becker:1985:DSD

[BC85a] R. A. Becker and J. M.

Chambers. Design of the S system for data analysis. AT&T Technical Journal, 64(9):2131–2151, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Becker:1985:ESS

[BC85b] Richard A. Becker and John M. Chambers. Extending the S System. Wadsworth, Pacific Grove, CA, USA, 1985. ISBN 0-534-05016-6. viii + 166 pp. LCCN QA276.4 .B4241 1985.

Berridge:2011:MGL

[BC11] Damon Berridge and Robert Crouchley. Multivariate generalized linear mixed models using R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1326-4 (hardcover). xxiii + 280 pp. LCCN HA31.35 .B47 2011.

Bowman:2007:RSI

[BCAB07] Adrian Bowman, Ewan Crawford, Gavin Alexander, and Richard W. Bowman. rpanel: Simple interactive controls for R functions using the tcltk package. Journal of Statistical Software, 17 (9):1–18, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v17/i09.

Bion:2018:HRH

[BCG18] Ricardo Bion, Robert Chang, and Jason Goodman. How R helps Airbnb make the most of its data. The American Statistician, 72(1):46– 52, 2018. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

amstat.tandfonline.com/
doi/full/10.1080/00031305.
2017.1392362.

Boettiger:2015:BSB

[BCHR15] Carl Boettiger, Scott Chamberlain, Edmund Hart, and Karthik Ram. Building software, building community: Lessons from the rOpenSci project. Journal of Open Research Software, 3(1):e8–??, November 16, 2015. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.bu/.

Benaglia:2009:MRP

[BCHY09] Tatiana Benaglia, Didier Chauveau, David R. Hunter, and Derek S. Young. mixtools: An R package for analyzing mixture models. Journal of Statistical Software, 32 (6):??, October 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/i06.

Bouveyron:2019:MBCh

[BCMR19] Charles Bouveyron, Gilles Celeux, T. Brendan Murphy, and Adrian E. Raftery. Model-based clustering and classification for data science: with applications in R, volume 50 of Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-49420-X (hardcover), 1-108-64059-1, 1-108-64418-X (e-book). xvii + 427 pp. LCCN QA278.55 .M63 2019.

Becker:1996:VDC

[BCS96] R. A. Becker, W. S. Cleveland, and M.-J. Shyu. The visual design and control of trellis display. Journal of Computational and Graphical Statistics, 5(2):123–155, 1996. ISSN 1061-8600 (print), 1537-2715 (electronic).

Becker:1988:NPLa

[BCW88] Richard A. Becker, John M.

Chambers, and Allan R. Wilks. The New S Programming Language: a Programming Environment for Data Analysis and Graphics. Wadsworth and Brooks/ Cole Publishing Co., Pacific Grove, CA, USA and Pacific Grove, CA, USA, 1988. ISBN 0-534-09192-X, 0-534-09193-8. xvii + 702 pp. LCCN QA76.73.S15 B43 1988.

Becker:2018:NLP

[BCW18] Richard A. Becker, John M. Chambers, and Allan R. Wilks. The New S Language: a Programming Environment for Data Analysis and Graphics. CRC revivals. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2018. ISBN 1-315-89588-9, 1-351-07498-9 (e-book). ???? pp. LCCN QA76.73.S15.

Bougeard:2018:SMA

[BD18] Stéphanie Bougeard and Stéphane Dray. Supervised multiblock analysis in R with the ade4 package. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v086i01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v086i01/
                                        v86i01.pdf.

Bordier:2011:TSI

[BDdM11] Cécile Bordier, Michel Dojat, and Pierre Lafaye de Micheaux. Temporal and spatial independent component analysis for fMRI data sets embedded in the AnalyzeFMRI R package. Journal of Statistical Software, 44 (9):??, October 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v44/i09.

Bilodeau:2015:RPG

[BdMM15] Martin Bilodeau, Pierre Lafaye de Micheaux, and Smail Mahdi. The R package groc for generalized regression on orthogonal components. Journal of Statistical Software, 65(1):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/
i01.

Bagnato:2015:CSP

[BDMP15] Luca Bagnato, Lucio De Capitani, Angelo Mazza, and Antonio Punzo. Code snippet: SDD: An R package for serial dependence diagrams. Journal of Statistical Software, 64(CS-2): ??, March 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/c02.

Brown:2023:ARP

[BdSvE23] David Brown, Kauê de Sousa, and Jacob van Etten. ag5Tools: an R package for downloading and extracting agrometeorological data from the AgERA5 database. SoftwareX, 21 (??):??, February 2023.

CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022001856.

Bodin:2018:TMF

[BDT18] Martin Bodin, Tomás Diaz, and Éric Tanter. A trustworthy mechanized formalization of R. ACM SIGPLAN Notices, 53(8):13–24, October 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:/

                                        /dl.acm.org/doi/abs/10.
                                        1145/3393673.3276946.

Beath:2017:RRP

[Bea17] Ken J. Beath. randomLCA: An R package for latent class with random effects analysis. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v081i13;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v081i13/
                                        v81i13.pdf.

Bebbington:2013:BRE

[Beb13] Mark Bebbington. Book review: Event History Analysis with R. Australian & New Zealand Journal of Statistics, 55(3):344–346, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Becker:1994:BH

[Bec94] R. A. Becker. A brief history of S. In Dirschedl and Ostermannn [DO94], pages 81– 110. ISBN 3-7908-0813-X. LCCN QA276.4 .C569 1994.

Beeley:2013:BRD

[Bee13] Chris Beeley. Behavioural research data analysis with R. Journal of Applied Statistics, 40(4):922, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Beh:2004:PCO

[Beh04] Eric J. Beh. S-PLUS code for ordinal correspondence analysis. Computational Statistics, 19(4):593–612, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02753914.

Beh:2005:PCS

[Beh05] Eric J. Beh. S-PLUS code for simple and multiple correspondence analysis. Computational Statistics, 20(3): 415–438, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/BF02741306.

Beh:2012:EMA

[Beh12] Eric J. Beh. Exploratory multivariate analysis by example using R. Journal of Applied Statistics, 39 (6):1381–1382, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bell:2019:RPS

[Bel19] Daniel Bell. R Programming: A Step-by-Step Guide for Absolute Beginners. Guzzler Media LLC, ????, 2019. ISBN 1-69676-964-7. 170 (est.) pp. LCCN ????

Benoit:2021:BRT

[Ben21] Kenneth R. Benoit. Book review: Textual Data Science with R, Mónica Bécue-Bertaut. Boca Raton, FL: Chapman & Hall/CRC Press, 2019, xvii + 194 pp., $79.95 (H), $54.95 (e-book), ISBN: 978-1-138-62691-1 (H), 978-1-315-21266-1 (ebook). The American Statistician, 75(4):453–454, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/00031305.
2021.1985864.

Berry:2009:BRB

[Ber09] Diane Berry. Book review: R Programming for Bioinformatics, by R. Gentleman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):696, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
20622535.

Berger:2010:BRP

[Ber10] Vance W. Berger. Book review: Permutation Tests for Stochastic Ordering and ANOVA: Theory and Applications with R, by Basso, D., Pesarin, F., Salmaso, L., and Solari, A. Biometrics, 66(1):319–321, March 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barbiero:2017:RPS

[BF17] Alessandro Barbiero and Pier Alda Ferrari. An R package for the simulation of correlated discrete variables. Communications in Statistics: Simulation and Computation, 46(7):5123–5140, 2017. CODEN CSSCDB. ISSN 0361-0918.

Beaton:2014:EMA

[BFA14] Derek Beaton, Cherise R. Chin Fatt, and Hervé Abdi. An ExPosition of multivariate analysis with the singular value decomposition in R. Computational Statistics & Data Analysis, 72(??):176–189, April 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947313004441.

Barragan:2013:IRP

[BFRP13] Sandra Barragán, Miguel Fernández, Cristina Rueda, and Shyamal Peddada. isocir: An R package for constrained inference using isotonic regression for circular data, with an application to cell biology. Journal of Statistical Software, 54 (4):??, September 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v54/i04.

Bruce:1996:AWA

[BG96] Andrew Bruce and Hong-Ye Gao. Applied wavelet analysis with S-Plus. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94714-0 (softcover). xxi + 338 pp. LCCN TK5102.9.B78 1996. URL

http://www.springer-ny.
com/catalog/np/jul96np/
DATA/0-387-94714-0.html.

Backlin:2018:DFC

[BG18] Christofer L. Bäcklin and Mats G. Gustafsson. Developerfriendly and computationally efficient predictive modeling without information leakage: The emil package for R. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i13;   https:/

/www.jstatsoft.org/index.
php/jss/article/view/v085i13/
v85i13.pdf.

Basturk:2017:RPM

[BGH^{+}17] Nalan Bastürk, Stefano Grassi, Lennart Hoogerheide, Anne Opschoor, and Herman K. van Dijk. The R package MitISEM: Efficient and robust simulation procedures for Bayesian inference. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v079i01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v079i01/
v79i01.pdf.

Bouvier:1994:NNL

[BH94] Annie Bouvier and Sylvie Huet. nls2: Non-linear regression by S-PLUS — functions. Computational Statistics & Data Analysis, 18 (1):187–190, August 1994. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0167947394901384.

Box:2005:SED

[BHH05] George E. P. Box, J. Stuart Hunter, and William Gordon Hunter. Statistics for experimenters: design, innovation, and discovery. Wiley series in probability and statistics. Wiley-Intersci-ence, New York, NY, USA, second edition, 2005. ISBN 0-471-71813-0. xvii + 633 pp. LCCN QA279 .B69 2005.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0619/
                                        2004063780-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0619/2004063780-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley051/2004063780.
                                        html;   http://www.wiley.
                                        com/WileyCDA/WileyTitle/
                                        productCd-0471718130.html.

Bureau:2000:PIH

[BHS00] Alexandre Bureau, James P. Hughes, and Stephen C. Shiboski. An S-Plus implementation of hidden Markov models in continuous time. Journal of Computational and Graphical Statistics, 9 (4):621–??, December 2000. ISSN 1061-8600 (print), 1537-2715 (electronic). URL

                                        http://www.amstat.org/
                                        publications/jcgs/abstracts00/
                                        Bureau.htm.

Bretz:2011:MCU

[BHW11] Frank Bretz, Torsten Hothorn, and Peter H. Westfall. Multiple Comparisons Using R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-58488-574-2 (hardcover). xvii + 187 pp. LCCN QA278.4 .B74 2011.

Brouste:2013:PED

[BI13] Alexandre Brouste and Stefano M. Iacus. Parameter estimation for the discretely observed fractional Ornstein–Uhlenbeck process and the Yuima R package. Computational Statistics, 28(4):1529–1547, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-012-0365-6.

Bianco:1994:MAM

[Bia94] L. Bianco, editor. Methods and applications of mathematical programming: Biennial Soviet-Italian conference; 1st — September 1992, Cosenza, Italy, volume 3(1) of Optimization Methods and Software. Gordon and Breach, Langhorne, PA, USA, 1994. CODEN OM-SOE2. ISSN 1055-6788.

Bailey:2011:CSI

[BK11] Delia Bailey and Jonathan N.

Katz. Code snippet: Implementing panel-corrected standard errors in R: The pcse package. Journal of Statistical Software, 42(CS-1):??, June 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v42/c01.

Biecek:2017:ARP

[BK17] Przemyslaw Biecek and Marcin Kosiński. archivist: An R package for managing, recording and restoring data analysis results. Journal of Statistical Software, 82(??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v082i11;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v082i11/v82i11.pdf.

Baumer:2017:MDS

[BKH17] Benjamin Baumer, Daniel Kaplan, and Nicholas J. Horton. Modern data science with R. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2017. ISBN 1-4987-2448-5. xvi + 551 pp. LCCN QA76.9.D343 B38 2017.

Brezger:2005:BAB

[BKL05] Andreas Brezger, Thomas Kneib, and Stefan Lang. BayesX: Analyzing Bayesian structural additive regression models. Journal of Statistical Software, 14(11):1– 22, September 9, 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=136&url=v14/i11&
                                        ct=2;   http://www.jstatsoft.
                                        org/counter.php?id=136&
                                        url=v14/i11/v14i11.pdf&
                                        ct=1.

Berentsen:2014:ILR

[BKT14] Geir Drage Berentsen, Tore Selland Kleppe, and Dag Bjarne Tjøstheim. Introducing

localgauss, an R pack-

age for estimating and visualizing local Gaussian correlation. Journal of Statistical Software, 56(12): ??, February 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v56/i12.

Burr:2011:BRBa

[BL11] Tom Burr and Claire Longo.

Book review: Wavelet Methods in Statistics With R by G. P. Nason. Technometrics, 53(2):212–213, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
23209736.

Bada:2014:PPD

[BL14a] Oualid Bada and Dominik Liebl. phtt: Panel data analysis with heterogeneous time trends in R. Journal of Statistical Software, 59(6): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v59/i06.

Baltagi:2014:FES

[BL14b] Badi H. Baltagi and Jing Li. Further evidence on the spatio–temporal model of house prices in the United States. Journal of Applied Econometrics, 29(3): 515–522, April–May 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See replication report [Mil15b].

Bilder:2015:ACD

[BL15] Christopher R. Bilder and Thomas M. Loughin. Analysis of Categorical Data with R. Chapman and Hall/CRC texts in statistical science. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4398-5567-6 (hardcover), 1-4987-0676-2 (e-book). xiii + 533 pp. LCCN QA169 .B55 2015.

Bischl:2015:PPA

[BLM^{+}15] Bernd Bischl, Michel Lang, Olaf Mersmann, Jörg Rahnenführer, and Claus Weihs.

                                        BatchJobs and BatchExperiments:

Abstraction mechanisms for using R in batch environments. Journal of Statistical Software, 64(11): ??, March 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v64/i11.

Bloomfield:2014:URN

[Blo14] Victor A. Bloomfield. Using R for Numerical Analysis in Science and Engineering. Chapman and Hall/CRC the R series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2014. ISBN 1-4398-8448-X (hardcover). xxii + 335 pp. LCCN Q183.9 .B56 2014. URL

http://www.crcnetbase.
com/isbn/9781439884492.

Boettiger:2012:PRE

[BLW12] C. Boettiger, D. T. Lang, and P. C. Wainwright. rfishbase: exploring, manipulating and visualizing FishBase data from R. Journal of Fish Biology, 81(6): 2030–2039, November 2012. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beare:2018:ISR

[BLY18] Richard Beare, Bradley Lowekamp, and Ziv Yaniv. Image segmentation, registration and characterization in R with SimpleITK. Journal of Statistical Software, 86(??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v086i08;
https://www.jstatsoft.
org/index.php/jss/article/
view/v086i08/v86i08.pdf.

Braun:2007:FCS

[BM07] John Braun and Duncan James Murdoch. A first course in statistical programming with R. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-87265-0 (hardcover), 0-521-69424-8 (paperback). x + 163 pp. LCCN QA276.45.R3 B73 2007.

URL http://www.loc.gov/

catdir/enhancements/fy0827/
2008295814-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0827/2008295814-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0827/2008295814-
t.html.

Braun:2016:FCS

[BM16] John Braun and Duncan James Murdoch. A first course in statistical programming with R. Cambridge University Press, Cambridge, UK, second edition, 2016. ISBN 1-107-57646-6 (paperback), 1-316-45109-7 (e-book). xiv + 215 pp. LCCN QA276.45.R3 B73 2016.

Braun:2021:FCS

[BM21] John Braun and Duncan James Murdoch. A first course in statistical programming with R. Cambridge University Press, Cambridge, UK, third edition, 2021. ISBN 1-108-99514-4 (paperback), 1-108-99941-7 (PDF e-book). xx + 260 pp. LCCN QA276.45.R3 B73 2021. URL https:

//www.cambridge.org/us/
academic/subjects/statistics-
probability/computational-
statistics-machine-learning-
and-information-sc/first-
course-statistical-programming-
r-3rd-edition.

Bunouf:2015:SPC

[BMGT15] Pierre Bunouf, Geert Molenberghs, Jean-Marie Grouin, and Herbert Thijs. A SAS program combining R functionalities to implement pattern-mixture models. Journal of Statistical Software, 68(8):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068i08;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v068i08/v68i08.pdf.

Becker:2019:TFE

[BML19] Gabriel Becker, Sara E.

Moore, and Michael Lawrence. trackr: A framework for enhancing discoverability and reproducibility of data visualizations and other artifacts in R. Journal of Computational and Graphical Statistics, 28(3):644–658, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Boehmke:2017:PKR

[BMOF17] Bradley C. Boehmke, Robert T.

Montgomery, Jeffrey A. Ogden, and Jason K. Freels.

KraljicMatrix: An R pack-

age for implementing the Kraljic Matrix to strategically analyze a firm’s purchasing portfolio. Journal of Open Source Software, 2 (10):170:1–170:2, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00170.

Baier:2007:ECR

[BN07] Thomas Baier and Erich Neuwirth. Excel :: COM :: R. Computational Statistics, 22(1):91–108, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

                                        link.springer.com/article/
                                        10.1007/s00180-007-0023-
                                        6.

Boik:2010:RPA

[BN10] John C. Boik and Balasubramanian Narasimhan. An R package for assessing drug synergism/antagonism. Journal of Statistical Software, 34(6):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v34/
                                        i06.

Bartoszynski:1996:PSI

[BNB96] Robert Bartoszyński and Magdalena Niewiadomska-Bugaj. Probability and statistical inference. Wiley series in probability and statistics. Wiley, New York, NY, USA, 1996. ISBN 0-471-31073-5 (cloth). xvi + 826 pp. LCCN QA273 .B2584 1996.

                                        URL http://www.loc.gov/
                                        catdir/description/wiley031/
                                        96004105.html;   http:/
                                        /www.loc.gov/catdir/enhancements/
                                        fy0806/96004105-b.html;

 http://www.loc.gov/catdir/
toc/onix01/96004105.html.

Bartoszynski:2007:PSI

[BNB07] Robert Bartoszyński and Magdalena Niewiadomska-Bugaj. Probability and statistical inference. Wiley-Interscience, New York, NY, USA, second edition, 2007. ISBN 0-470-19158-9 (e-book). xii + 647 pp. LCCN QA273 .B2584 2007. URL http://

www.wiley.com/WileyCDA/
WileyTitle/productCd-0470191589.
html.

Bartoszynski:2008:PSI

[BNB08] Robert Bartoszyński and Magdalena Niewiadomska-Bugaj. Probability and statistical inference. Wiley-In-terscience, New York, NY, USA, second edition, 2008. ISBN 0-471-69693-5 (hardcover). xii + 647 pp. LCCN QA273 .B2584 2008.

URL http://www.loc.gov/
catdir/enhancements/fy0826/
2008271954-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0826/2008271954-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0826/2008271954-
t.html;   http://www.wiley.
com/WileyCDA/WileyTitle/
productCd-0471696935.html.

Boettiger:2017:GPC

[Boe17]  Carl Boettiger.  Gener-

ating CodeMeta metadata for R packages. Journal of Open Source Software, 2(19):454:1, November 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00454.

Bolker:2008:EMD

[Bol08] Benjamin M. Bolker. Ecological Models and Data in R. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-12522-

8. vii + 396 pp. LCCN QH541.15.S72 B65 2008.

Bonat:2018:MRV

[Bon18] Wagner Hugo Bonat. Multiple response variables regression models in R: The mcglm package. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v084i04;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v084i04/
                                        v84i04.pdf.

Booth:2010:BRBa

[Boo10] David Booth. Book review: Computational Statistics: An Introduction to R by Gunther Sawitzki. Technometrics, 52(3):364, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27867261.

Borg:2016:TFT

[Bor16a] Markus Borg. TuneR: a framework for tuning software engineering tools with hands-on instructions in R. Journal of Software: Evolution and Process, 28(6): 427–459, June 2016. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Borja:2016:VBP

[Bor16b] Mario Cortina Borja. A variation of the birthday problem. Significance (Oxford, England), 13(6):12– 13, December 2016. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Boshnakov:2009:BRT

[Bos09] Georgi N. Boshnakov. Book review: Time series analysis with applications in R series: Springer texts in statistics, 2nd edition. Journal of Time Series Analysis, 30 (6):708–709, November 2009. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2010:BRI

[Bos10] Georgi N. Boshnakov. Book review: Introductory Time Series with R. Journal of Time Series Analysis, 31(5):406, September 2010. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Boshnakov:2012:URD

[Bos12] Georgi N. Boshnakov. Using R for data management, statistical analysis, and graphics. Journal of Applied Statistics, 39(6):1382– 1383, June 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Bowman:2009:BRB

[Bow09] Adrian Bowman. Book review: Morphometrics with R. Journal of Statistical Software, 31(BR-1): ??, September 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v31/b01.

Bowman:2010:BRB

[Bow10]  Adrian Bowman.  Book

review: Functional Data Analysis with R and MAT-LAB. Journal of Statistical Software, 34(BR-3):??, April 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v34/b03.

Basto:2012:SRM

[BP12] Mário Basto and José Manuel Pereira. An SPSS Rmenu for ordinal factor analysis. Journal of Statistical Software, 46(4):??, January 2012. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v46/i04.

Barrios:2017:DIG

[BP17a] David Barrios and Carlos Prieto. D3GB: An interactive genome browser for R, Python, and Word-Press. Journal of Computational Biology, 24(5):447– 449, May 2017. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.
liebertpub.com/doi/abs/
10.1089/cmb.2016.0213;
 https://www.liebertpub.
com/doi/pdf/10.1089/cmb.
2016.0213.

Beckerman:2017:GSR

[BP17b] Andrew P. Beckerman and Owen L. Petchey. Getting started with R: an introduction for biologists. Oxford biology. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2017. ISBN 0-19-878784-7 (paperback). xviii + 231 pp. LCCN QA276.45.R3 B43 2017.

Bornkamp:2009:MRP

[BPB09] Björn Bornkamp, José Pinheiro, and Frank Bretz. MCPMod: An R package for the design and analysis of dose-finding studies. Journal of Statistical Software, 29 (7):??, February 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/i07.

Brock:2008:CRP

[BPDD08] Guy Brock, Vasyl Pihur, Susmita Datta, and Somnath Datta. clValid: An R package for cluster validation. Journal of Statistical Software, 25(4):1– 22, March 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v25/i04.

Bonhomme:2014:MOA

[BPGC14] Vincent Bonhomme, Sandrine Picq, Cédric Gaucherel, and Julien Claude. Momocs: Outline analysis using R. Journal of Statistical Software, 56(13):??, February 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:

                                        //www.jstatsoft.org/v56/
                                        i13.

Bivand:2008:ASD

[BPGR08] Roger S. Bivand, Edzer J. Pebesma, and Virgilio Gómez-Rubio. Applied Spatial Data Analysis with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-78170-6. xiv + 374 pp. LCCN ???? EUR 48.10; SFR 75.00.

                                        URL http://deposit.d-
                                        nb.de/cgi-bin/dokserv?
                                        id=3087668&prov=M&dok_
                                        var=1&dok_ext=htm.

Barros:2009:RIG

[BPL09] Michelli Barros, Gilberto A. Paula, and Víctor Leiva. An R implementation for generalized Birnbaum–Saunders distributions. Computational Statistics & Data Analysis, 53(4):1511–1528, February 15, 2009. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947308005434.

Bartolucci:2017:LRP

[BPP17] Francesco Bartolucci, Silvia Pandolfi, and Fulvia Pennoni. LMest: An R package for latent Markov models for longitudinal categorical data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081i04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081i04/
v81i04.pdf.

Basso:2009:PTS

[BPSS09] Dario Basso, Fortunato Pesarin, Luigi Salmaso, and Aldo Solari. Permutation Tests for Stochastic Ordering and ANOVA: Theory and Applications with R, volume 194 of Lecture notes in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-85955-1 (paperback), 0-387-85956-X (e-book). xiii + 217 pp. LCCN QA165 .P437 2009.

Baillargeon:2007:RLM

[BR07] Sophie Baillargeon and Louis-Paul Rivest. Rcapture: Loglinear models for capturerecapture in R. Journal of Statistical Software, 19 (5):1–31, April 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v19/i05.

Brayshaw:2003:BRB

[Bra03] Nigel Brayshaw. Book review: Applied Statistics in the Pharmaceutical Industry, with Case Studies in S-Plus, by S. P. Millard; A. Krause. Journal of the Royal Statistical Society. Series D (The Statistician), 52(2):243–244, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4128260.

Braun:2014:PRP

[Bra14] Michael Braun. trustOptim: An R package for trust region optimization with sparse Hessians. Journal of Statistical Software, 60(4): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v60/i04.

Brandmaier:2015:PPR

[Bra15] Andreas M. Brandmaier. pdc: An R package for complexity-based clustering of time series. Journal of Statistical Software, 67 (5):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v067i05;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v067i05/v067i05.pdf.

Braun:2017:SRP

[Bra17] Michael Braun. sparseHessianFD: An R package for estimating sparse Hessian matrices. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i10;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i10/
v82i10.pdf.

Bakka:2018:SMR

[BRF^{+}18] Haakon Bakka, Håvard Rue, Geir-Arne Fuglstad, Andrea Riebler, David Bolin, Janine Illian, Elias Krainski, Daniel Simpson, and Finn Lindgren. Spatial modeling with R-INLA: a review. WIREs Computational Statistics, 10(6): e1443:1–e1443:??, November/December 2018. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Brooms:2007:BRB

[Bro07] A. C. Brooms. Book review: The Basics of S-PLUS, by A. Krause; M. Olson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):256– 257, January 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4623151.

Brooms:2010:BRB

[Bro10] A. C. Brooms. Book review: Data Manipulation with R. Journal of Applied Statistics, 37(12):2121, December 2010. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [Spe08].

Brostrom:2012:EHA

[Bro12] Göran Broström. Event history analysis with R. The R series; Chapman and Hall book. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-3164-5 (hardcover), 1-4398-3167-X (e-book). xx + 216 pp. LCCN H61 .B338 2012.

Broman:2014:FYP

[Bro14]  Karl Broman.  Fourteen

years of R/qtl: Just barely sustainable. Journal of Open Research Software, 2 (1):e11–??, July 09, 2014. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.at/.

Broman:2009:GQM

[BS09] Karl W. Broman and Saunak Sen. A Guide to QTL Mapping with R/qtl. Statistics for biology and health. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-92124-9 (hardcover), 0-387-92125-7 (e-book). ISSN 1431-8776. xv + 396 pp. LCCN QH447 .B755 2009.

Brechmann:2013:MDC

[BS13] Eike Christian Brechmann and Ulf Schepsmeier. Modeling dependence with Cand D-Vine copulas: The R package CDVine. Journal of Statistical Software, 52 (3):??, February 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v52/i03.

Bakar:2015:SST

[BS15] Khandoker Shuvo Bakar and Sujit K. Sahu. spTimer: Spatio-temporal Bayesian modeling using R. Journal of Statistical Software, 63 (15):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i15.

Boersch-Supan:2016:PRF

[BS16] Philipp H. Boersch-Supan. rucrdtw: Fast time series subsequence search in R. Journal of Open Source Software, 1(7):100:1–100:2, November 2016. CO-DEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/
10.21105/joss.00100.

Barthelemy:2018:CSM

[BS18] Johan Barthélemy and Thomas Suesse. Code snippet: mipfp: An R package for multidimensional array fitting and simulating multivariate Bernoulli distributions. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086c02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086c02/
v86c02.pdf.

Bub:2020:XRP

[BSG20] Sascha Bub, Thorsten Schad, and Zhenglei Gao. XDrift

— an R package to simulate spatially explicit pesticide spray-drift exposure of nontarget-species habitats at landscape scales. SoftwareX, 12(??):Article 100610, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S235271102030323X.

Badouard:2024:LPS

[BSS^{+}24] Vincyane Badouard, Sylvain Schmitt, Guillaume Salzet, Thomas Gaquiere, Margaux Rojat, Caroline Bedeau, Olivier Brunaux, and Géraldine Derroire. LoggingLab: an R package to simulate reduced-impact selective logging in tropical forests using forest inventory data. Ecological Modelling, 487(??):??, January 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0304380023002697.

Biecek:2012:RPB

[BSVT12] Przemyslaw Biecek, Ewa Szczurek, Martin Vingron, and Jerzy Tiuryn. The R package bgmm: Mixture modeling with uncertain knowledge. Journal of Statistical Software, 47 (3):??, April 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v47/i03.

Baddeley:2005:SRP

[BT05] Adrian Baddeley and Rolf Turner. spatstat: An R package for analyzing spatial point patterns. Journal of Statistical Software, 12 (6):1–42, ???? 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=113&url=v12/i06&
                                        ct=2;   http://www.jstatsoft.
                                        org/counter.php?id=113&
                                        url=v12/i06/v12i06.pdf&
                                        ct=1.

Baumer:2015:RM

[BU15] Benjamin Baumer and Dana Udwin. R Markdown. WIREs Computational Statistics, 7(3):167–177, May/ June 2015. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Buckeridge:2009:BRS

[Buc09] David Buckeridge. Book review: Statistical Methods for Environmental Epidemiology with R: a Case Study in Air Pollution and Health, by Peng, R. D. and Dominici, F. Biometrics, 65(3): 996–997, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bullard:2006:BRB

[Bul06] James Bullard. Book review: R Graphics. Journal of Statistical Software, 16(BR-7):1–3, October 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/b07.

Burr:2007:BRB

[Bur07] Thomas L. Burr. Book review: Modeling Financial Time Series with S-Plus by Eric Zivot; Jiahui Wang. Technometrics, 49 (1):105–106, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471290.

Burr:2009:BRBa

[Bur09] Thomas L. Burr. Book review: Simulation and Inference for Stochastic Differential Equations With R Examples by Stefano Iacus. Technometrics, 51 (3):340–341, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40586630.

Burr:2010:BRB

[Bur10] Tom Burr. Book review:

Functional Data Analysis With R and MATLAB by J. Ramsay; G. Hooker; S. Graves. Technometrics, 52 (4):466, November 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40997262.

Burr:2012:BRP

[Bur12]  Deborah Burr.  bspmma:

An R package for Bayesian semiparametric models for meta-analysis. Journal of Statistical Software, 50(4): ??, July 2012. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v50/i04.

Burkner:2017:BRP

[Bür17] Paul-Christian Bürkner. brms: An R package for Bayesian multilevel models using Stan. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v080i01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v080i01/
                                        v80i01.pdf.

Buttrey:2005:CLL

[But05] Samuel E. Buttrey. Calling the lp solve linear program software from R, S-PLUS and Excel. Journal of Statistical Software, 14(4): 1–13, May 1, 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=129&url=v14/i04&
                                        ct=2;   http://www.jstatsoft.
                                        org/counter.php?id=129&
                                        url=v14/i04/v14i04.pdf&
                                        ct=1.

Butts:2008:NPM

[But08] Carter T. Butts. network: a package for managing relational data in R. Journal of Statistical Software, 24 (2):1–36, May 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v24/i02.

Blaser:2015:PRP

[BVE^{+}15] Nello Blaser, Luisa Salazar Vizcaya, Janne Estill, Cindy Zahnd, Bindu Kalesan, Matthias Egger, Thomas Gsponer, and Olivia Keiser. gems: An R package for simulating from disease progression models. Journal of Statistical Software, 64 (10):??, March 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/i10.

Baayen:2019:STC

[BVFB19] C. Baayen, C. Volteau, C. Flamant, and P. Blanche. Sequential trials in the context of competing risks: Concepts and case study, with R and SAS code. Statistics in Medicine, 38(19): 3682–3702, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1988:NLPb

[BW88] Richard A. Becker and Allan Reeve Wilks. The New S Language: a Programming Environment for Data Analysis and Graphics. Chapman and Hall, Ltd., London, UK, 1988. ISBN 0-412-74150-4. xvii + 702 pp. LCCN ????

Barter:2018:SRP

[BY18] Rebecca L. Barter and Bin

Yu. Superheat: An R

package for creating beautiful and extendable heatmaps for visualizing complex data. Journal of Computational and Graphical Statistics, 27 (4):910–922, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chollet:2018:DLR

[CA18] François Chollet and Joseph J. Allaire. Deep learning with R. Manning Publications, Greenwich, CT, USA, 2018. ISBN 1-61729-554-X (paperback). xxi + 335 pp. LCCN QA276.45.R3 C46 2018.

Cebrian:2015:PRP

[CAA15] Ana C. Cebrián, Jesús Abaurrea, and Jesús Asín.

NHPoisson: An R pack-

age for fitting and validating nonhomogeneous Poisson processes. Journal of Statistical Software, 64 (6):??, March 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/i06.

Cadigan:2013:FNP

[Cad13]  Noel G. Cadigan.  Fit-

ting a non-parametric stock– recruitment model in R that is useful for deriving MSY reference points and accounting for model uncertainty. ICES Journal of Marine Science, 70 (1):56–67, January 2013. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/
article/70/1/56/663017.

Calaway:1995:OOP

[Cal95] Richard Calaway. Objectoriented programming in S. Dr. Dobb’s Journal of Software Tools, 20(10):44, 46, 48–49, 51, 113–115, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Calenge:2006:PPA

[Cal06] Clément Calenge. The package “adehabitat” for the R software: a tool for the analysis of space and habitat use by animals. Ecological Modelling, 197(3–4): 516–519, August 25, 2006. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0304380006001414.

Campbell:2009:BRS

[Cam09] Dave Campbell. Book review: Simulation and Inference for Stochastic Differential Equations with R Examples, by Iacus, S. M. Biometrics, 65(1):335–336, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Camarda:2012:MRP

[Cam12] Carlo G. Camarda. MortalitySmooth: An R package for smoothing Poisson counts with P-splines. Journal of Statistical Software, 50(1):??, July 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i01.

Cantoni:2004:RNS

[Can04]  Eva Cantoni.  Resistant

nonparametric smoothing with S-PLUS. Journal of Statistical Software, 10(2): 1–6, ???? 2004. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=93&url=v10/i02/
Mspline4.pdf&ct=1.

Capaldi:2019:BRD

[Cap19] Mindy Capaldi. Book review: Data Science and Predictive Analytics: Biomedical and Health Applications Using R, Ivo D. Dinov, Springer, 2018, xxxiv + 832 pages, £89.99, hardcover ISBN: 978-3-319-72346-4. International Statistical Review = Revue Internationale de Statistique, 87(1):181– 182, April 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Caporin:2022:BRS

[Cap22] Massimiliano Caporin. Book review: Statistical Analysis of Financial Data: with Examples In R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):432–433, January 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Carmona:2004:SAF

[Car04] R. (René) Carmona. Statistical analysis of financial data in S-PLUS. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-20286-2. xvi + 451 pp. LCCN HG106 .C37 2004.

URL http://www.loc.gov/
catdir/enhancements/fy0818/
2003066218-d.html;   http:
//www.loc.gov/catdir/toc/
fy044/2003066218.html.

Cardot:2010:BRF

[Car10] Hervé Cardot. Book review:

Functional Data Analysis with R and Matlab, by Ramsay, J. O., Hooker, G., and Graves, S. Biometrics, 66(4): 1312–1313, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carbon:2013:CSB

[Car13]  Claus-Christian  Carbon.

Code snippet: BiDimRegression: Bidimensional regression modeling using R. Journal of Statistical Software, 52 (CS-1):??, March 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v52/c01.

Carmona:2014:SAF

[Car14] R. (René) Carmona. Statistical analysis of financial data in R. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2014. ISBN 1-4614-8787-0 (hardcover). ISSN 1431-875X. xvii + 588 pp. LCCN HG106 .C37 2014.

URL http://www.loc.gov/
catdir/enhancements/fy1404/
2013951152-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1404/2013951152-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1404/2013951152-
t.html.

Caron:2016:BRN

[Car16] François Caron. Book review: Nonlinear Time Series: Theory, methods, and applications with R examples, by Randal Douc, Eric Moulines, David S. Stoffer. Statistics and Computing, 26(1–2):571–572, January 2016. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s11222-
015-9575-4;   http://link.
springer.com/article/10.
1007/s11222-015-9575-4.

Caravaggi:2017:PRR

[Car17] Anthony Caravaggi. remBoot: An R package for random encounter modelling. Journal of Open Source Software, 2(10):176:1–176:3, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00176.

Caro:2017:IRP

[CB17] Daniel H. Caro and Przemyslaw Biecek. intsvy: An R package for analyzing international large-scale assessment data. Journal of Statistical Software, 81 (??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v081i07;
https://www.jstatsoft.
org/index.php/jss/article/
view/v081i07/v81i07.pdf.

Cano-Berlanga:2017:ECG

[CBGGV17] Sebastián Cano-Berlanga, José-Manuel Giménez-Gómez, and Cori Vilella. Enjoying cooperative games: the R package GameTheory. Applied Mathematics and Computation, 305(??):381– 393, July 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300317301145.

Corradini:2021:URP

[CBHLG21] Fabio Corradini, Nicolas Beriot, Esperanza Huerta-Lwanga, and Violette Geissen. uFTIR: an R package to process hyperspectral images of environmental samples captured with μ FTIR microscopes. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001321.

Casalicchio:2019:ORP

[CBL^{+}19] Giuseppe Casalicchio, Jakob Bossek, Michel Lang, Dominik Kirchhoff, Pascal Kerschke, Benjamin Hofner, Heidi Seibold, Joaquin Vanschoren, and Bernd Bischl. OpenML: An R package to connect to the machine learning platform OpenML. Computational Statistics, 34(3):977–991, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-017-0742-2.

Cao:2008:ORP

[CC08a] Kim-Anh Lê Cao and Patrick Chabrier. ofw: An R package to select continuous variables for multiclass classification with a stochastic wrapper method. Journal of Statistical Software, 28 (9):??, November 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v28/i09.

Cohen:2008:SDR

[CC08b] Yosef Cohen and Jeremiah Y.

Cohen. Statistics and data with R: an applied approach through examples. Wiley, New York, NY, USA, 2008. ISBN 0-470-75805-8 (cloth), 0-470-72188-X (e-book). xviii + 599 pp. LCCN QA276.45.R3 C64 2008. URL http://

www.loc.gov/catdir/toc/
ecip0824/2008032153.html;
 http://www.wiley.com/
WileyCDA/WileyTitle/productCd-
047072188X.html;   http:
//www.wiley.com/WileyCDA/
WileyTitle/productCd-0470758058.
html.

Childs:2011:MFS

[CC11] Aaron Childs and Yalin Chen. Multilevel fixed and sequential acceptance sampling: The R package MFSAS. Journal of Statistical Software, 43(6): ??, July 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v43/i06.

Cabrera:2023:PMR

[CC23] Juan M. Cabrera and Julieta Capeletti. metrix: an R package for assessing the biological quality of water. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S030438002300203X.

Cepeda-Cuervo:2016:BBR

[CCJMR16] Edilberto Cepeda-Cuervo, Daniel Jaimes, Margarita Marín, and Javier Rojas. Bayesian beta regression

                                        with Bayesianbetareg R-

package. Computational Statistics, 31(1):165–187, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-015-0591-9.

Chambers:1983:GMD

[CCKT83] John M. Chambers, W. S. Cleveland, B. Kleiner, and P. A. Tukey, editors. Graphical methods for data analysis. The Wadsworth statistics/probability series. Wadsworth, Pacific Grove, CA, USA, 1983. ISBN 0-534-98052-X. xiv + 395 pp. LCCN QA276.3 .G73 1983.

Cryer:2011:SJX

[CCP^{+}11] Jonathan D. Cryer, Kung-Sik Chan, Hongyu Pan, et al. Shi jian xu lie fen xi ji ying yong: R yu yan = Time series analysis with applications in R. Hua zhang shu xue yi cong. Ji xie gong ye chu ban she, Beijing, China, 2011. ISBN 7-111-32572-9. 350 pp. LCCN ????

Calcagno:2010:GRP

[CdM10] Vincent Calcagno and Claire de Mazancourt. glmulti: An R package for easy automated model selection with (generalized) linear models. Journal of Statistical Software, 34(12):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v34/
i12.

Crookston:2008:YRP

[CF08] Nicholas L. Crookston and Andrew O. Finley. yaImpute: An R package for kNN imputation. Journal of Statistical Software, 23(10):1– 16, January 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i10.

Caimo:2014:PBE

[CF14] Alberto Caimo and Nial Friel. Bergm: Bayesian exponential random graphs in R. Journal of Statistical Software, 61(2):??, October 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i02.

Causeur:2011:FAM

[CFHBK11] David Causeur, Chloe Friguet, Magalie Houee-Bigot, and Maela Kloareg. Factor analysis for multiple testing (FAMT): An R package for large-scale significance testing under dependence. Journal of Statistical Software, 40(14):??, May 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/i14.

Conde:2015:DRP

[CFSR15] David Conde, Miguel Fernández, Bonifacio Salvador, and Cristina Rueda. dawai: An R package for discriminant analysis with additional information. Journal of Statistical Software, 66(10):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v066i10;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v066i10/v66i10.pdf.

Clarke:2009:PTD

[CFZ09] Bertrand Salem Clarke, Ernest Fokoué, and Hao Helen Zhang. Principles and Theory for Data Mining and Machine Learning, volume ?? of Springer Series in Statistics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN ???? ISBN 0-387-98134-9, 0-387-98135-7 (e-book). ISSN 0172-7397. xv + 781 pp. LCCN QA76.9.D343 C53 2009. URL http://

link.springer.com/book/
10.1007/978-0-387-98135-
2.

Capanu:2015:CSB

[CG15] Marinela Capanu and Mithat Gönen. Code snippet: Building a nomogram for surveyweighted Cox models using R. Journal of Statistical Software, 64(CS-1): ??, March 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/c01.

Charrad:2014:PRP

[CGBN14] Malika Charrad, Nadia Ghazzali, Véronique Boiteau, and Azam Niknafs. NbClust: An R package for determining the relevant number of clusters in a data set. Journal of Statistical Software, 61 (6):??, November 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i06.

Choi:2011:LRP

[CGC11] Seung W. Choi, Laura E.

Gibbons, and Paul K. Crane. lordif: An R package for detecting differential item functioning using iterative hybrid ordinal logistic regression/item response theory and Monte Carlo simulations. Journal of Statistical Software, 39 (8):??, March 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v39/i08.

Colombi:2014:PRP

[CGC14] Roberto Colombi, Sabrina Giordano, and Manuela Cazzaro. hmmm: An R package for hierarchical multinomial marginal models. Journal of Statistical Software, 59(11): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v59/i11.

Cornillon:2008:SAR

[CGH^{+}08] Pierre-André Cornillon, Arnaud Guyader, François Husson, Nicolas Jégou, Julie Josse, Maela Kloareg, Eric Matzner-Løber, and Laurent Rouvière. Statistiques avec R. Didact statistique. Presses universitaires de Rennes, Rennes, France, 2008. ISBN 2-7535-0722-8. v + 257 pp. LCCN QA276.45 R3 S797 2008.

Cornillon:2010:SAR

[CGH^{+}10] Pierre-André Cornillon, Arnaud Guyader, François Husson, Nicolas Jégou, Julie Josse, Maela Kloareg, Eric Matzner-Løber, and Laurent Rouvière. Statistiques avec R. Pratique de la statistique / [collection] dirigée par Éric Matzner-Løber. Presses universitaires de Rennes, Rennes, France, second edition, 2010. ISBN 2-7535-1087-3 (paperback). v + 274 pp. LCCN QA276.45R3 C67.2010. URL http:/

/catalogue.bnf.fr/ark:/
12148/cb42210870p.

Cornillon:2012:RS

[CGH^{+}12a] Pierre-André Cornillon, Arnaud Guyader, François Husson, Nicolas Jégou, Julie Josse, Maela Kloareg, Eric Matzner-Løber, and Laurent Rouvière. R for statistics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-8145-6 (paperback). xiv + 306 pp. LCCN QA276.45.R3 S7413

2012. URL http://www.
crcpress.com/product/isbn/
9781439881453.

Cornillon:2012:SAR

[CGH^{+}12b] Pierre-André Cornillon, Arnaud Guyader, François Husson, Nicolas Jégou, Julie Josse, Maela Kloareg, Eric Matzner-Løber, and Laurent Rouvière. Statistiques avec R. Pratique de la statistique. Presses universitaires de Rennes, Rennes, France, third edition, 2012. ISBN 2-7535-1992-7 (paperback). ISSN 1295-1765. v + 296 pp. LCCN QA276.45R3 C67.2012. URL http:/

/catalogue.bnf.fr/ark:/
12148/cb42674494g.

Cule:2009:LRP

[CGS09] Madeleine Cule, Robert B.

Gramacy, and Richard Sam-

worth. LogConcDEAD: An R

package for maximum likelihood estimation of a multivariate log-concave density. Journal of Statistical Software, 29(2):??, January 2009. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v29/i02.

Chambers:1992:SM

[CH92] John M. Chambers and Trevor Hastie, editors. Statistical models in S. Wadsworth and Brooks/Cole computer science series. Wadsworth and Brooks/Cole Advanced Books and Software, Pacific Grove, CA, USA, 1992. ISBN 0-534-16764-0 (casebound), 0-534-16765-9 (paperback). xv + 608 pp. LCCN QA276.4 .S65 1991.

Chambers:1993:SMS

[CH93] John M. Chambers and Trevor J. Hastie, editors. Statistical Models in S. Chapman and Hall, Ltd. and Wadsworth, London, UK and Pacific Grove, CA, USA, 1993. ISBN 0-412-05301-2 (Chapman and Hall), 0-534-16765-9 (Wadsworth paperback), 0-534-16764-0 (Wadsworth hardcover). xv + 608 pp. LCCN QA276.4 .S65 1992. US$40.95.

Chihara:2011:MSR

[CH11a] Laura Chihara and Tim Hesterberg. Mathematical statistics with resampling and R. Wiley, New York, NY, USA, 2011. ISBN 1-118-02985-2 (hardcover). xiv + 418 pp. LCCN QA278.8 .C45 2011.

Cook:2011:BRB

[CH11b] Dianne Cook and Heike Hofmann. Book review: R Graphics (2nd Edition). Journal of Statistical Software, 43(BR-3): ??, September 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v43/b03.

Chamberlain:2017:PLR

[CH17] Scott Chamberlain and Jeffrey W. Hollister. lawn: An R client for the Turf Javascript Library for Geospatial Analysis. Journal of Open Source Software, 2(11):194:1, March 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00194.

Coyle:2018:POG

[CH18] Jeremy R. Coyle and Nima S.

Hejazi. origami: A generalized framework for crossvalidation in R. Journal of Open Source Software, 3(21):512:1–512:3, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00512.

Chambers:1969:SDL

[Cha69] J. M. Chambers. A statistical data language. In Milton and Nelder [MN69], pages 179–199. ISBN 0-12-498150-X. LCCN QA276.4 .C6 1969.

Chambers:1986:PED

[Cha86]  John M. Chambers.  S

as a programming environment for data analysis and graphics. In Allen [All86], pages 206–209 (physical). ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http:

                                        //www.dtic.mil/dtic/tr/
                                        fulltext/u2/a166019.pdf.

Chambers:1995:OV

[Cha95] J. M. Chambers. Overview of version 4 of S. Technical report, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 23, 1995.

Chambers:1998:PDG

[Cha98] John M. Chambers. Programming with Data: a Guide to the S Language. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-91577-X. xv + 469 pp. LCCN QA76.73.S15C483 1998.

Chao:1999:IDA

[Cha99] Edward C. Chao. Interactive data analysis based on dialog systems in S-PLUS. Computational Statistics, 14(1): 23–38, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.

springer.com/article/10.
1007/PL00022703.

Chambers:2008:SDA

[Cha08a] John M. Chambers. Software for Data Analysis: Programming with R. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-75935-2, 0-387-75936-0. xiv + 498 pp. LCCN QA276.45.R3 C43 2008.

Chang:2008:ADT

[Cha08b] Mark Chang. Adaptive Design Theory and Implementation Using SAS and R. Chapman and Hall/CRC biostatistics series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-962-4. xxii + 418 pp. LCCN R853.C55 C42 2008.

URL http://www.loc.gov/
catdir/enhancements/fy0801/
2007011412-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0713/2007011412.html.

Chausse:2010:CGM

[Cha10] Pierre Chaussé. Computing generalized method of moments and generalized empirical likelihood with R. Journal of Statistical Software, 34(11):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v34/
i11.

Chalmers:2012:MMI

[Cha12] R. Philip Chalmers. mirt: a multidimensional item response theory package for the R environment. Journal of Statistical Software, 48(6):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v48/i06.

Chakraborty:2014:BRB

[Cha14a] Sounak Chakraborty. Book review: Bayesian Essentials with R by Jean-Michel Marin and Christian P. Robert. International Statistical Review = Revue Internationale de Statistique, 82 (3):488–490, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Chambers:2014:OOP

[Cha14b] John M. Chambers. Objectoriented programming, functional programming and R. Statistical Science, 29(2): 167–180, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

                                        projecteuclid.org/euclid.
                                        ss/1408368569.

Chapman:2015:RMR

[Cha15a] Chris Chapman. R for marketing research and analytics. Springer Science+Business Media, LLC, New York, NY, 2015.

ISBN 3-319-14435-9 (paperback), 3-319-14436-7 (ebook). xviii + 454 pp. LCCN QA276.45.R3.

URL http://www.loc.gov/
catdir/enhancements/fy1504/
2014960277-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1504/2014960277-d.html;
http://www.loc.gov/catdir/
enhancements/fy1504/2014960277-
t.html.

Charpentier:2015:CAS

[Cha15b] Arthur Charpentier, editor. Computational Actuarial Science with R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-9259-1 (hardcover). xxxi + 618 pp. LCCN HG8781 .C637 2014.

Chambers:2016:ER

[Cha16] John M. Chambers. Extending R. Chapman and Hall/CRC the R series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2016. ISBN 1-4987-7571-3 (paperback), 1-4987-7572-1 (ebook). xvii + 364 pp. LCCN QA276.45.R3 C436 2016.

Chang:2018:RGC

[Cha18] Winston Chang. R Graphics Cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2018. ISBN 1-4919-7860-0 (paperback), 1-4919-7855-4 (e-book), 1-4919-7857-0 (ebook), 1-4919-7859-7. xiii + 425 pp. LCCN QA276.45.R3 C46 2018.

Chambers:2020:RDS

[Cha20] John M. Chambers. S, R, and data science. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):84:1–84:17, June

2020.  URL https://
dl.acm.org/doi/abs/10.
1145/3386334.

Chaturvedi:2021:BRA

[Cha21] Anoop Chaturvedi. Book review: Analyzing highdimensional gene expression and DNA methylation data with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184 (3):1154, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chaturvedi:2022:BRH

[Cha22] Anoop Chaturvedi. Book review: Handbook of Regression Analysis with Applications in R (Second Edition). Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S777–S778, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/185/Supplement_
2/S777/7069518.

Cheng:2014:GBR

[CHB14] Ching-Wei Cheng, Ying-Chao Hung, and Narayanaswamy Balakrishnan. Generating beta random numbers and Dirichlet random vectors in R: the package rBeta2009. Computational Statistics & Data Analysis, 71(??):1011– 1020, March 2014. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947313000753.

Chernick:2011:BRBc

[Che11] Michael R. Chernick. Book review: Multiple Comparisons Using R by Frank Bretz; Torsten Hothorn; Peter Westfall. Technometrics, 53(3):327–328, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
23210417.

Chen:2022:IDS

[Che22] Li-Pang Chen. Introduction to data science: Data analysis and prediction algorithms with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):733– 734, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Chen:2023:BRF

[Che23] Li-Pang Chen. Book review: Fundamentals of highdimensional statistics: with exercises and R Labs. By Johannes Lederer, Springer International Publishing, 2021. pp. 355. ISBN: 978-3-030-73791-7. Biometrics, 79(3): 2772–2773, September 2023. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiang:2007:BRB

[Chi07]  Alan Y. Chiang.  Book

review: Generalized Additive Models: An Introduction with R by Simon N. Wood. Technometrics, 49 (3):360–361, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        25471358.

Chopin:2009:BRJ

[Cho09] Nicolas Chopin. Book review: Jim Albert: Bayesian computation with R. Statistics and Computing, 19 (1):111–112, March 2009. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://

                                        link.springer.com/accesspage/
                                        article/10.1007/s11222-
                                        008-9069-8.

Choi:2015:BRB

[Cho15] Taeryon Choi. Book review: Bayesian networks with examples in R. Biometrics, 71(3):864–865, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chowdhury:2020:BRN

[Cho20]  Mohammed  Chowdhury.

Book review: Nonparametric models for longitudinal data: With implementation in R, Wu, Colin O., Tian, Xin, Boca Raton, FL: CRC Press. Biometrics, 76(1):363–364, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choe:2022:BRI

[Cho22a] Youngjun Choe. Book review: An Introduction to Acceptance Sampling and SPC with R John Lawson, boca raton, FL, Chapman & Hall/CRC Press, 2021, xvii + 280 pp., $150.00(H), $59.95(e-book), ISBN 978-0-367-56995-2(H), 978-1-003-10027-0(e-book). Journal of the American Statistical Association, 117(537): 528, 2022. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chowdhry:2022:BRA

[Cho22b] Amit K. Chowdhry. Book review of Applied Meta-Analysis with R and Stata. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):728– 729, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Christ:2009:BRB

[Chr09] Aaron Christ. Book review: Mixed Effects Models and Extensions in Ecology with R. Journal of Statistical Software, 32(BR-1): ??, November 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/b01.

Carmona:1998:PTF

[CHT98] René Carmona, Wen-Liang Hwang, and Bruno Torrésani. Practical Time-frequency Analysis. Gabor and Wavelet Transforms with an Implementation in S. Academic Press, New York, NY, USA, 1998. ISBN 0-12-160170-6. 490 pp. With a preface by Ingrid Daubechies.

Curini:2017:NOP

[CI17] Luigi Curini and Stefano M. Iacus. Nash optimal party positions: The nopp R package. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v081i11;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v081i11/
                                        v81i11.pdf.

Cichosz:2015:DMA

[Cic15] Pawel Cichosz. Data Mining Algorithms: Explained Using R. Wiley, New York, NY, USA, 2015. ISBN 1-118-33258-X (hardcover), 1-118-95080-1 (e-book), 1-118-95084-4 (e-book), 1-118-95095-X. xxxi + 683 pp. LCCN QA76.9.D343 C472 2015. URL http://

www.wiley.com/WileyCDA/
WileyTitle/productCd-111833258X.
html.

Culp:2006:ARP

[CJM06] Mark Culp, Kjell Johnson, and George Michailides. ada: An R package for stochastic boosting. Journal of Statistical Software, 17(2):1– 27, September 2006. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v17/i02.

Chavent:2012:CRP

[CKSLS12] Marie Chavent, Vanessa Kuentz-Simonet, Benoît Liquet, and Jérôme Saracco. ClustOfVar: An R package for the clustering of variables. Journal of Statistical Software, 50(13): ??, September 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i13.

Chavent:2018:CRP

[CKSLS18] Marie Chavent, Vanessa Kuentz-Simonet, Amaury Labenne, and Jérôme Saracco.

ClustGeo: an R package for hierarchical clustering with spatial constraints. Computational Statistics, 33(4):1799–1822, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-018-0791-1.

Chiou:2014:FAF

[CKY14] Sy Han Chiou, Sangwook Kang, and Jun Yan. Fitting accelerated failure time models in routine survival analysis with R package aftgee. Journal of Statistical Software, 61(11): ??, November 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i11.

Chernick:2011:IBM

[CL11] Michael R. Chernick and Robert A. LaBudde. An introduction to bootstrap methods with applications to R. Wiley, New York, NY, USA, 2011. ISBN 0-470-46704-5 (hardcover). xvii + 216 pp. LCCN QA276.8 .C478 2011.

Claude:2008:MR

[Cla08] Julien Claude. Morphometrics with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-77789-X (paperback). xviii + 316 pp. LCCN QH351 .C53 2008.

Cleveland:1985:EGD

[Cle85] William S. Cleveland. The elements of graphing data. Wadsworth, Pacific Grove, CA, USA, 1985. ISBN 0-534-03730-5 (paperback). xii + 323 pp. LCCN QA90 .C54 1985. US$18.95.

Cleveland:1993:VD

[Cle93] William S. Cleveland. Visualizing Data. Hobart Press, Summit, NJ, USA, 1993. ISBN 0-9634884-0-6. 360 pp. LCCN QA276.3.C57 1993. US$50.00.

Cleveland:1994:EGD

[Cle94] William S. Cleveland. The elements of graphing data. AT&T Bell Laboratories, Murray Hill, NJ, USA, revised edition, 1994. ISBN 0-9634884-1-4 (cloth). 297 pp. LCCN QA90 .C54 1994.

Chung:2021:BRP

[CLK21] Jong-Hee Chung, Yong-Bin Lim, and Donghoh Kim. BDEsize: an R package for efficient determination of sample size in a balanced design of experiments. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001229.

Comets:2017:PEN

[CLL17] Emmanuelle Comets, Audrey Lavenu, and Marc Lavielle. Parameter estimation in nonlinear mixed effect models using saemix, an R implementation of the SAEM algorithm. Journal of Statistical Software, 80 (??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v080i03;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v080i03/v80i03.pdf.

Cowpertwait:2009:ITS

[CM09] Paul S. P. Cowpertwait and Andrew V. Metcalfe. Introductory Time Series with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-88697-4. xv + 254 pp. LCCN QA280 .C69 2009.

URL http://www.loc.gov/
catdir/enhancements/fy1002/
2009928496-d.htm;   http:
//www.loc.gov/catdir/enhancements/
fy1002/2009928496-t.htm.

Canal:2014:CSC

[CM14] Luisa Canal and Rocco Micciolo. The chi-square controversy: what if Pearson had R? Journal of Statistical Computation and Simulation, 84(5):1015–1021, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Clayden:2011:TMR

[CMS^{+}11] Jonathan D. Clayden, Susana Muñoz Maniega, Amos J. Storkey, Martin D. King, Mark E. Bastin, and Chris A. Clark. TractoR: Magnetic resonance imaging and tractography with R. Journal of Statistical Software, 44 (8):??, October 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v44/i08.

Cribari-Neto:1997:EPE

[CN97] Francisco Cribari-Neto. Econometric programming environments: GAUSS, Ox and S-Plus. Journal of Applied Econometrics, 12(1):77–89, January 1997. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Chan:2016:RPV

[CNA16] Stephen Chan, Saralees Nadarajah, and Emmanuel Afuecheta. An R package for value at risk and expected shortfall. Communications in Statistics: Simulation and Computation, 45(9):3416–??, 2016.

Cribari-Neto:1999:RYA

[CNZ99] Francisco Cribari-Neto and Spyros G. Zarkos. R: yet another econometric programming environment. Journal of Applied Econometrics, 14(3):319–329, May– June 1999. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Cribari-Neto:2010:BRR

[CNZ10] Francisco Cribari-Neto and Achim Zeileis. Beta regression in R. Journal of Statistical Software, 34 (2):??, April 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v34/i02.

Crippa:2016:CSM

[CO16] Alessio Crippa and Nicola Orsini. Code snippet: Multivariate dose-response metaanalysis: The dosresmeta R package. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v072c01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v072c01/
                                        v72c01.pdf.

Coene:2018:PSR

[Coe18] Jean-Philippe Coene. sigmajs:

                                        An R htmlwidget inter-

face to the sigma.js visualization library. Journal of Open Source Software, 3(28):814:1–814:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00814.

Contoni:2003:RNS

[Con03]  Eva Contoni.  Resistant

nonparametric smoothing with S-PLUS. Journal of Statistical Software, 8(3): ??, ???? 2003. CO-DEN JSSOBK. ISSN

????  URL http://www.
jstatsoft.org/v08/i03;
 http://www.jstatsoft.
org/v08/i03/; http://www.
jstatsoft.org/v08/i03/
updates;   http://www.
jstatsoft.org/v08/i03/
v8.txt.

Cotton:2013:LRS

[Cot13]  Richard Cotton.  Learn-

ing R: [a step-by-step function guide to data analysis]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2013. ISBN 1-4493-5710-5 (paperback), 1-4493-5718-0, 1-4493-5719-9. xix + 377 pp. LCCN QA276.45.R3 C68 2013.

Cowles:2003:BRB

[Cow03] Mary Kathryn Cowles. Book review: An R and S-PLUS Companion to Applied Regression. The American Statistician, 57(4):316– ??, November 2003. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//oberon.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/00031305/

                                        v57n4/s16/p316. Review of

[Fox02].

Cox:2005:BRS

[Cox05] Christopher Cox. Book review: 1. Statistical tools for non-linear regression. A practical guide with S-Plus and R examples (2nd edn). Sylvie Huet, Annie Bouvier, Marie-Anne Poursat and Emmanuel Jolivet, Springer, New York, 2004. No. of pages: xiv + 232. Price: $74.94, £57.50, EUR74.95 (+ VAT). ISBN 0-387-40081-8. Statistics in Medicine, 24(13):2105–2106, July 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carstensen:2011:ULO

[CP11a] Bendix Carstensen and Martyn Plummer. Using Lexis objects for multi-state models in R. Journal of Statistical Software, 38(6): ??, January 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v38/i06.

Chen:2011:CTD

[CP11b] Ding-Geng Chen and Karl E. Peace. Clinical trial data analysis using R. Chapman and Hall/CRC biostatistics series. CRC Press, Boca Raton, Florida, 2011. ISBN 1-4398-4020-2 (hardcover), 1-4398-4021-0 (e-book). xxiii + 363 pp. LCCN R853.C55 C44 2011.

Cheng:2012:IDM

[CP12] Ya-Shan Cheng and Chien-Yu Peng. Integrated degradation models in R using iDEMO. Journal of Statistical Software, 49(2): ??, June 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v49/i02.

Chen:2013:AMA

[CP13] Ding-Geng Chen and Karl E.

Peace. Applied metaanalysis with R. Chapman and Hall/CRC biostatistics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4665-0599-0 (hardcover). xxiv + 321 pp. LCCN R853.M48 C44 2013. URL

http://images.tandf.co.
uk/common/jackets/websmall/
978146650/9781466505995.
jpg.

Calcagni:2020:SAC

[CPA20] Antonio Calcagnì, Massimiliano Pastore, and Gianmarco Altoé. ssMousetrack — analysing computerized tracking data via Bayesian state-space models in R. Mathematical and Computational Applications, 25(3): ??, September 2020. CO-DEN ???? ISSN 2297-8747.

URL https://www.mdpi.
com/2297-8747/25/3/41.

Correndo:2023:SRP

[CPB^{+}23] Adrian A. Correndo, Austin Pearce, Carl H. Bolster, John T. Spargo, Deanna Osmond, and Ignacio A. Ciampitti. The soiltestcorr R package: an accessible framework for reproducible correlation analysis of crop yield and soil test data. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022001935.

Castelluzzo:2020:MQD

[CPD^{+}20] Michele Castelluzzo, Alessio Perinelli, Simone Detassis, Michela Alessandra Denti, and Leonardo Ricci.

                                        MiRNA-QC-and-Diagnosis:

an R package for diagnosis based on MiRNA expression. SoftwareX, 12 (??):Article 100569, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711020301977.

Clement:1995:DPP

[CQ95] Mark J. Clement and Michael J. Quinn. Dynamic performance prediction for scalable parallel computing. Technical Report 95-80-04, Oregon State University, Corvallis, OR, USA, 1995.

Chang:2010:CRP

[CQZ^{+}10] Fang Chang, Weiliang Qiu, Ruben H. Zamar, Ross Lazarus, and Xiaogang Wang. clues: An R package for nonparametric clustering based on local shrinking. Journal of Statistical Software, 33(4): ??, February 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v33/i04.

Crawley:2002:SCI

[Cra02] Michael J. Crawley. Statistical computing: an introduction to data analysis using S-Plus. Wiley, New York, NY, USA, 2002. ISBN 0-471-56040-5. ix + 761 pp. LCCN QA276.4 .C73 2002.

Crawley:2005:SIU

[Cra05] Michael J. Crawley. Statistics: an introduction using R. Wiley, New York, NY, USA, 2005. ISBN 0-470-02297-3 (hardback), 0-470-02298-1 (paperback). xiii + 327 pp. LCCN QA276.12 .C73 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.imperial.ac.
uk/bio/research/crawley/
statistics;   http://www.
loc.gov/catdir/toc/ecip053/
2004026793.html.

Crawley:2007:RB

[Cra07] Michael J. Crawley. The R book. Wiley, New York, NY, USA, 2007. ISBN 0-470-51024-2, 0-470-51506-6 (ebook). viii + 942 pp. LCCN QA276.45.R3 C73 2007.

URL http://www.loc.gov/
catdir/enhancements/fy0741/
2007296951-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2007296951-d.html;
 http://www.loc.gov/catdir/
toc/fy0713/2007296951.
html;   http://www.wiley.
com/WileyCDA/WileyTitle/
productCd-0470515066.html.

Crawley:2012:RB

[Cra12] Michael J. Crawley. The R book. Wiley, New York, NY, USA, second edition, 2012. ISBN 0-470-97392-7 (hardback). xxiv + 1051 pp. LCCN QA276.45.R3 C73 2013.

Crawley:2015:SIU

[Cra15] Michael J. Crawley. Statistics: an introduction using R. Wiley, New York, NY, USA, second edition, 2015. ISBN 1-118-94110-1, 1-118-94111-X (e-book), 1-118-94109-8 (e-book). xiii + 339 pp. LCCN QA276.12; QA276.12 .C73 2015eb.

Contreras-Reyes:2013:SAA

[CRP13] Javier E. Contreras-Reyes and Wilfredo Palma. Statistical analysis of autoregressive fractionally integrated moving average models in R. Computational Statistics, 28(5):2309–2331, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-013-0408-7.

Cook:2007:IDG

[CS07] Dianne Cook and Deborah F. Swayne. Interactive and Dynamic Graphics for Data Analysis: with R and GGobi. Use R! Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-71761-7. vii + 188 pp. LCCN QA276.3 .C66 2007.

URL http://www.loc.gov/
catdir/enhancements/fy0825/
2007925720-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0825/2007925720-t.html.

Clifford:2012:VPD

[CS12] David Clifford and Glenn Stone. Variable penalty dynamic time warping code for aligning mass spectrometry chromatograms in R. Journal of Statistical Software, 47(8):??, April 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v47/i08.

Cryer:2008:TSA

[CsC08] Jonathan D. Cryer and Kung sik Chan. Time Series Analysis: with Applications in R. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 1-4419-2613-5 (paperback), 0-387-75958-1, 0-387-75959-X (e-book). xiii + 491 pp. LCCN QA280 .C78 2008. URL http://

www.loc.gov/catdir/toc/
fy0804/2008923058.html.

Calhoun:2018:CMS

[CSNF18] Peter Calhoun, Xiaogang Su, Martha Nunn, and Juanjuan Fan. Constructing multivariate survival trees: The MST package for R. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v083i12;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v083i12/
v83i12.pdf.

Culp:2011:SSS

[Cul11] Mark Culp. spa: Semisupervised semi-parametric graph-based estimation in R. Journal of Statistical Software, 40(10):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v40/
i10.

Cunningham:2012:SGM

[Cun12] J. Kelly Cunningham. Should S get more press? The American Statistician, 66(4): 237, 2012. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Curran:2018:BRF

[Cur18] J. M. Curran. Book review: Foundations of Statistical Algorithms With References to R Packages, By C. Weihs, O. Mersmann and U. Ligges. Boca Raton, Florida: CRC Press. 2015. 500 pages. UK£44.99 (hardback). ISBN 978-1-4398-7885-9. Australian & New Zealand Journal of Statistics, 60(2):276– 277, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Conway:2012:MLH

[CW12] Drew Conway and John Myles White. Machine learning for hackers. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2012. ISBN 1-4493-0371-

4. xiii + 303 pp. LCCN QA76.9.A43 C674 2012.

URL http://www.loc.gov/
catdir/enhancements/fy1307/
2012277057-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1307/2012277057-d.html;
http://www.loc.gov/catdir/
enhancements/fy1307/2012277057-
t.html.

Dalgaard:1998:BRSb

[Dal98] Peter Dalgaard. Book review: Statistical Tools for Nonlinear Regression. A Practical Guide with S-Plus examples. Sylvie Huet, Annie Bouvier, Marie-Anne Gruet and Emmanuel Jolivet, Springer-Verlag, New York, 1996. No. of pages: ix + 154. Price: £28. ISBN 0-387-94727-2. Statistics in Medicine, 17(23):2807, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dalgaard:2002:ISR

[Dal02] Peter Dalgaard. Introductory Statistics with R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95475-9. xv + 267 pp. LCCN QA276.4 .D33 2002.

Dalgaard:2008:ISR

[Dal08] Peter Dalgaard. Introductory Statistics with R. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 0-387-79053-5 (paperback). ISSN 1431-8784. xvi + 363 pp. LCCN QA276.45.R3 D35 2008.

Dan:2018:LJE

[Dan18]  Toomey Dan.  Learning

Jupyter 5: explore interactive computing using Python, Java, JavaScript, R, Julia, and JupyterLab. Packt Publishing, Birmingham, 2018. ISBN 1-78913-740-3, 1-78913-744-6. 282 pp. LCCN Q183.9; QA76.9.I52 .T666 2018.

Dasgupta:2021:BRI

[Das21] Nairanjana Dasgupta. Book review: Introduction to Data Science: Data Analysis and Algorithms with R, By Rafael Irrizarry Boca Raton, FL: CRC Press, 2020. Hard cover. pp. 711. Biometrics, 77(3):1131–1132, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davies:1995:BRBd

[Dav95] Neville Davies. Book review: A Handbook of Statistical Analysis Using S-Plus, by B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):633– 634, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
2983455.

Davis:2007:BRBa

[Dav07] J. Wade Davis. Book review:

Bioinformatics and Computational Biology Solutions Using R and Bioconductor. Journal of the American Statistical Association, 102 (477):388–389, March 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Davies:2015:BRF

[Dav15] Tilman M. Davies. The book of R: a first course in programming and statistics. No Starch Press, San Francisco, CA, USA, 2015. ISBN 1-59327-651-6 (paperback), 1-59327-779-2 (e-book). xxxii + 792 pp. LCCN QA76.73.R3 D38 2016.

Dawson:2003:BRB

[Daw03] Jeffrey D. Dawson. Book review: Analyzing Medical Data Using S-PLUS, by Brian Everitt; Sophia Rabe-Hesketh. Journal of the American Statistical Association, 98(463):773–774, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        30045317.

Dayal:2015:IRQb

[Day15] Vikram Dayal. Introduction to R for quantitative economics: graphing, simulating and computing. SpringerBriefs in economics. Springer, New Delhi, India, 2015. ISBN 81-322-2339-X, 81-322-2340-3. xv + 109 pp. LCCN ????

Davies:2013:CEG

[DB13] Tilman M. Davies and David Bryant. On circulant embedding for Gaussian random fields in R. Journal of Statistical Software, 55(9): ??, November 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v55/i09.

Desjardins:2018:HEM

[DB18] Christopher D. Desjardins and Okan Bulut. Handbook of Educational Measurement and Psychometrics Using R. Chapman and Hall/CRC = The R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2018. ISBN 1-315-15426-9 (e-book), 1-351-64077-1 (e-book), 1-351-65030-0 (ebook), 1-4987-7013-4 (hardcover), 1-4987-7014-2. xxiii + 302 pp. LCCN LB3051 .D4465 2018.

DeBoeck:2011:EIR

[DBZ^{+}11] Paul De Boeck, Marjan Bakker, Robert Zwitser, Michel Nivard, Abe Hofman, Francis Tuerlinckx, and Ivailo Partchev. The estimation of item response models with the lmer function from the lme4 package in R. Journal of Statistical Software, 39 (12):??, March 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v39/i12.

Dejean:2005:FRP

[DC05] Sébastien Déjean and Serge Cohen. FracSim: An R package to simulate multifractional Lévy motions. Journal of Statistical Software, 14(18):1–19, December 2005. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v14/
i18.

Dahl:2009:RAR

[DC09] David B. Dahl and Scott Crawford. RinRuby: Accessing the R interpreter from pure Ruby. Journal of Statistical Software, 29 (4):??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/i04.

Deslauriers:2017:PFB

[DCB^{+}17] David Deslauriers, Steven R. Chipps, James E. Breck, James A. Rice, and Charles P.

Madenjian. Fish Bioenergetics

4.0: an R-based modeling application. Fisheries, 42(11):586–596, 2017. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/03632415.
2017.1377558.

delCastillo:2022:CRL

[dCCM^{+}22] Enrique del Castillo, Peng Chen, Adam Meyers, John Hunt, and James Rapkin. Confidence regions for the location of response surface optima: the R package

OptimaRegion. Communi-

cations in Statistics: Simulation and Computation, 51 (12):7074–7094, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Diaz-Coto:2020:SRP

[DCMCPF20] Susana Díaz-Coto, Pablo Martínez-Camblor, and Sonia Pérez-Fernández. smoothROCtime: an R package for timedependent ROC curve estimation. Computational Statistics, 35(3):1231–1251, September 2020. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
020-00955-7.

DeBin:2016:BCR

[De 16] Riccardo De Bin. Boosting in Cox regression: a comparison between the likelihoodbased and the model-based approaches with focus on the R-packages CoxBoost and mboost. Computational Statistics, 31(2):513– 531, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-015-0642-2.

Deegan:1999:ISM

[Dee99] Leo Joseph Deegan. Improper survival models. Thesis (Ph.D. in Biostatistics), University of California, Berkeley, Berkeley, CA, USA, 1999. vii + 114 pp. Includes S-PLUS code: isurvreg, Improper Survival Regression.

Demidenko:2013:MMT

[Dem13] Eugene Demidenko. Mixed models: theory and applications with R, volume 893 of Wiley series in probability and statistics. Wiley, New York, NY, USA, second edition, 2013. ISBN 1-118-09157-4 (hardcover), 1-118-59299-9, 1-118-59301-4, 1-118-59306-5 (e-book), 1-118-65153-7 (e-book). xxvii + 717 pp. LCCN QA279.D457 2013.

Demirtas:2017:BRB

[Dem17]  Hakan Demirtas.  Book

review: Bayesian Computation with R (2nd Edition). Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v079b01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v079b01/
                                        v79b01.pdf.

Demirtas:2018:BRH

[Dem18] Hakan Demirtas. Book review: Handbook of Fitting Statistical Distributions with R. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/

view/v086b02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086b02/
v86b02.pdf.

Demidenko:2020:ASA

[Dem20] Eugene Demidenko. Advanced Statistics with Applications in R. Wiley, New York, NY, USA, 2020. ISBN 1-118-38798-8 (hardcover), 1-118-59413-4 (PDF), 1-118-59461-4 (ePub), 1-119-44919-7 (e-book),. xvii + 856 pp. LCCN QA276.45.R3.

Denison:1998:BRB

[Den98] David G. T. Denison. Book review: Applied Smoothing Techniques for Data Analysis — The Kernel Approach with S-Plus Illustrations, by A. W. Bowman; A. Azzalini. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):539–540, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
2988634.

Dennis:2013:RSC

[Den13] Brian Dennis. The R student companion. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2013. ISBN 1-4398-7540-5 (paperback). xvii + 339 pp. LCCN QA276.45.R3 D46 2013. URL http:/

/jacketsearch.tandf.co.

                                        uk/common/jackets/covers/
                                        websmall/978143987/9781439875407.
                                        jpg.

Denwood:2016:RRP

[Den16] Matthew J. Denwood. runjags: An R package providing interface utilities, model templates, parallel computing methods and additional distributions for MCMC models in JAGS. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v071i09;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v071i09/
                                        v71i09.pdf.

Devroye:1986:NUR

[Dev86] Luc Devroye. Non-uniform random variate generation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-96305-7. xvi + 843 pp. LCCN QA274 .D48 1986. US$54.00.

Deviere:2009:ALD

[Dév09] Elisabeth Dévière. Analyzing linguistic data: a practical introduction to statistics using R. Journal of Applied Statistics, 36(4):471– 472, April 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Demirtas:2022:MDG

[DG22] H. Demirtas and R. Gao.

Mixed data generation packages and related computational tools in R. Communications in Statistics: Simulation and Computation, 51 (8):4520–4563, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

DiNardo:2023:GNB

[DG23] Elvira Di Nardo and Giuseppe Guarino. On the generation of necklaces and bracelets in R. Communications in Statistics: Simulation and Computation, 52(11):5727– 5737, 2023. CODEN CSS-CDB. ISSN 0361-0918.

Dutang:2008:ARP

[DGP08] Christophe Dutang, Vincent Goulet, and Mathieu Pigeon. actuar: An R package for actuarial science. Journal of Statistical Software, 25(7): 1–37, March 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i07.

Davison:1997:BMT

[DH97] A. C. (Anthony Christopher) Davison and D. V. Hinkley. Bootstrap methods and their application. Cambridge series on statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-57391-2 (hardcover), 0-521-57471-4 (paperback). x + 582 pp. LCCN QA276.8 .D38 1997.

Dupuy:2015:DDT

[DHF15] Delphine Dupuy, Céline Helbert, and Jessica Franco.

                                        DiceDesign and DiceEval:

Two R packages for design and analysis of computer experiments. Journal of Statistical Software, 65(11):??, June 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v65/i11.

Davies:2011:SAS

[DHM11] Tilman M. Davies, Martin L. Hazelton, and Jonathan C. Marshall. sparr: Analyzing spatial relative risk using fixed and adaptive kernel density estimation in R. Journal of Statistical Software, 39(1):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v39/
                                        i01.

Dias:2005:SAF

[Dia05] Alexandra Dias. Statistical analysis of financial data in S-Plus. The American Statistician, 59(4):351, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dialsingh:2006:BRBc

[Dia06] Isaac Dialsingh. Book review: Statistical Analysis of Financial Data in S-Plus, by R. A. Carmona. Journal of the Royal Statistical Society. Series A (Statistics in Society), 169(4):1004– 1005, October 2006. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
3877409.

Dialsingh:2018:BRH

[Dia18] Isaac Dialsingh. Book review: Hurricane Climatology: a Modern Statistical Guide using R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1260–1261, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dickhaus:2012:BRM

[Dic12] Thorsten Dickhaus. Book review: Multiple Comparisons Using R, by Bretz, F., Hothorn, T., and Westfall, P. Biometrics, 67(3):995, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dietz:2021:BRR

[Die21]  Sebastian Dietz.  Book

review: R Visualizations — Derive Meaning from Data, Gerbing, D. W., 2020. Chapman and Hall/CRC Press, Boca Raton. 237 pp.

68,99 GBP. ISBN 978-1-138-59963-5. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(1):401–402, January 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Dinov:2006:SSO

[Din06] Ivo D. Dinov. SOCR: Statistics online computational resource. Journal of Statistical Software, 16(11):1– 16, October 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/i11.

Dirmeier:2018:PDR

[Dir18] Simon Dirmeier. datastructures: An R package for organisation and storage of data. Journal of Open Source Software, 3(28): 910:1–910:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00910.

Dissanayake:2023:TSA

[Dis23] Gnanadarsha Sanjaya Dissanayake. Time-series analysis for the state-space model with R/Stan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 186(1):164–165, January 2023. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/186/1/164/7017743.

Dombrowsky:2018:SAA

[DJS^{+}18] Maximilian J. Dombrowsky, Sven Jager, Benjamin Schiller, Benjamin E. Mayer, Sebastian Stammler, and Kay Hamacher. StreaMD: Advanced analysis of molecular dynamics using R. Journal of Computational Chemistry, 39(21):1666–1674, August 5, 2018. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Dormann:2009:ASB

[DK09] Carsten F. Dormann and Ingolf Kühn. Angewandte Statistik für die biologischen Wissenschaften. (German) [Applied Statistics for the Biological Sciences], 2009. x + 245 pp. URL http://

cran.r-project.org/doc/
contrib/Dormann+Kuehn_
AngewandteStatistik.pdf.

Draminski:2018:RRP

[DK18] Michal Dramiński and Jacek Koronacki. rmcfs: An R package for Monte Carlo feature selection and interdependency discovery. Journal of Statistical Software, 85(??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v085i12;

                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v085i12/v85i12.pdf.

deLeeuw:2005:BRBb

[dL05a] Jan de Leeuw. Book review: An R and S-PLUS Companion to Multivariate Analysis. Journal of Statistical Software, 14(BR-6): 1–3, September 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v14/b06.

deLeeuw:2005:BRBa

[dL05b] Jan de Leeuw. Book review: Correspondence Analysis and Data Coding with Java and R. Journal of Statistical Software, 14(BR-5): 1–2, September 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v14/b05.

deLeeuw:2006:BRB

[dL06] Jan de Leeuw. Book review: Robust Statistical Methods with R. Journal of Statistical Software, 16(BR-2):1–2, July 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/b02.

deLeeuw:2009:BRBa

[dL09a] Jan de Leeuw. Book review: Nonlinear Regression with R. Journal of Statistical Software, 29(BR-6): ??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/b06.

deLeeuw:2009:BRBc

[dL09b] Jan de Leeuw. Book review: R Programming for Bioinformatics. Journal of Statistical Software, 29(BR-8):??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/b08.

deLeeuw:2009:BRBb

[dL09c] Jan de Leeuw. Book review: Statistical Methods for Environmental Epidemiology with R. Journal of Statistical Software, 29(BR-7): ??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/b07.

Dethlefsen:2006:FSS

[DLC06] Claus Dethlefsen and Søren Lundbye-Christensen. Formulating state space models in R with focus on longitudinal regression models. Journal of Statistical Software, 16 (1):1–15, April 2006. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v16/i01.

deLeeuw:2009:IOR

[dLHM09] Jan de Leeuw, Kurt Hornik, and Patrick Mair. Isotone optimization in R: Pool-adjacent-violators algorithm (PAVA) and active set methods. Journal of Statistical Software, 32(5): ??, October 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/i05.

deLeeuw:2009:GMO

[dLM09a] Jan de Leeuw and Patrick Mair. Gifi methods for optimal scaling in R: The package homals. Journal of Statistical Software, 31 (4):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v31/i04.

deLeeuw:2009:MSU

[dLM09b] Jan de Leeuw and Patrick Mair. Multidimensional scaling using majorization: SMACOF in R. Journal of Statistical Software, 31 (3):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v31/i03.

deLeeuw:2009:SCC

[dLM09c] Jan de Leeuw and Patrick Mair. Simple and canonical correspondence analysis using the R package anacor. Journal of Statistical Software, 31(5):??, August 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v31/
                                        i05.

Daouia:2017:NPN

[DLN17] Abdelaati Daouia, Thibault Laurent, and Hohsuk Noh. npbr: A package for nonparametric boundary regression in R. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v079i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v079i09/
v79i09.pdf.

Donoghoe:2018:LRP

[DM18] Mark W. Donoghoe and Ian C. Marschner. logbin: An R package for relative risk regression using the logbinomial model. Journal of Statistical Software, 86 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v086i09;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v086i09/v86i09.pdf.

DEnza:2018:CSI

[DMB18] Alfonso Iodice D’Enza, Angelos Markos, and Davide Buttarazzi. Code snippet: The idm package: Incremental decomposition methods in R. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086c04;   https:/
/www.jstatsoft.org/index.

                                        php/jss/article/view/v086c04/
                                        v86c04.pdf.

Delignette-Muller:2015:PRP

[DMD15] Marie Laure Delignette-Muller and Christophe Du-

                                        tang. fitdistrplus: An

R package for fitting distributions. Journal of Statistical Software, 64(4): ??, March 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v64/i04.

Douc:2014:NTS

[DMS14] Randal Douc, Eric Moulines, and David S. Stoffer. Nonlinear Time Series: Theory, Methods and Applications with R Examples. Chapman and Hall/CRC texts in statistical science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-0225-8 (hardcover). xx + 531 pp. LCCN QA280 .D68 2014. URL

                                        http://images.tandf.co.
                                        uk/common/jackets/websmall/
                                        978146650/9781466502253.
                                        jpg.

Demirtas:2017:BRI

[DN17] Hakan Demirtas and Rachel Nordgren. Book review: Introduction to Scientific Programming and Simulation Using R (2nd Edition). Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK.

ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v078b04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v078b04/
v78b04.pdf.

Dirschedl:1994:CSP

[DO94] Peter Dirschedl and Rudiger Ostermannn, editors. Computational Statistics: papers collected on the occasion of the 25th Conference on Statistical Computing at Schloss Reisensburg, Contributions to statistics. Physica-Verlag, Heidelberg, Germany, 1994. ISBN 3-7908-0813-X. LCCN QA276.4 .C569 1994.

Doerge:2006:BRB

[Doe06] R. W. Doerge. Book review: Bioinformatics and Computational Biology Solutions Using R and Bioconductor, edited by Gentleman, R., Carey, V., Huber, W., Irizarry, R., and Dudoit, S. Biometrics, 62(4): 1270–1271, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Doerge:2010:BRG

[Doe10] R. W. Doerge. Book review:

A Guide to QTL Mapping with R/qtl, by Broman, K. W. and Sen, S. Biometrics, 66(2):658–659, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donoho:2017:YDS

[Don17] David Donoho. 50 years of data science. Journal of Computational and Graphical Statistics, 26(4):745–766, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See comments [Pen17, HJ17].

Downie:2017:BRM

[Dow17] Tim Downie. Book review: Modern Data Science with R. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v080b02;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v080b02/
                                        v80b02.pdf.

Delgado:2013:EEE

[DP13] Michael S. Delgado and Christopher F. Parmeter. Embarrassingly easy embarrassingly parallel processing in R. Journal of Applied Econometrics, 28 (7):1224–1230, November– December 2013. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Dunkler:2018:WCR

[DPSH18] Daniela Dunkler, Meinhard Ploner, Michael Schemper, and Georg Heinze. Weighted Cox regression using the R package coxphw. Journal of Statistical Software, 84 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v084i02;
https://www.jstatsoft.
org/index.php/jss/article/
view/v084i02/v84i02.pdf.

Draghici:2012:SDA

[Dră12] Sorin Drăghici. Statistics and data analysis for microarrays using R and Bioconductor. Chapman and Hall/CRC mathematical and computational biology. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-0975-5 (hardback). xlviii + 1042 pp. LCCN QP624.5.D726 D74 2012.

Drachal:2023:DRP

[Dra23] Krzysztof Drachal. “dynmix”:

an R package for the estimation of dynamic finite mixtures. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711023000845.

Dressler:2019:BRC

[Dre19] Emily Dressler. Book review: Clinical Trial Optimization Using R. The American Statistician, 73(2): 210–211, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

amstat.tandfonline.com/
doi/full/10.1080/00031305.
2019.1603479.

Quintela-del-Rio:2012:NKD

[dREP12] Alejandro Quintela del Río and Graciela Estévez-Pérez. Nonparametric kernel distribution function estimation with kerdiest: An R package for bandwidth choice and applications. Journal of Statistical Software, 50 (8):??, August 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i08.

Drost:2018:PPI

[Dro18] Hajk-Georg Drost. Philentropy: Information theory and distance quantification with R. Journal of Open Source Software, 3(26): 765:1–765:4, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00765.

Daneshgar:2020:WAR

[DS20] Neda Daneshgar and Majid

Sarmad. word.alignment:

an R package for computing statistical word alignment and its evaluation. Computational Statistics, 35(4): 1597–1619, December 2020.

CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
020-00979-z.

deSousa:2023:GRP

[dSBSvE23] Kauê de Sousa, David Brown, Jonathan Steinke, and Jacob van Etten. gosset: an R package for analysis and synthesis of ranking data in agricultural experimentation. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000985.

daSilva-Junior:2014:CSP

[dSJdSF14] Antonio Hermes M. da Silva-Júnior, Damião Nóbrega da Silva, and Silvia L. P. Ferrari. Code snippet: mdscore: An R package to compute improved score tests in generalized linear models. Journal of Statistical Software, 61(CS-2):??, October 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v61/
c02.

Drummond:2019:SRP

[DTDd19] Blair R. Drummond, Christian J. G. Tessier, Mathieu F. Dextraze, and Corrie J. B. daCosta. scbursts: an R package for analysis and sorting of singlechannel bursts. SoftwareX, 10(??):Article 100285, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019300810.

Lopez-de-Ullibarri:2013:SRP

[dUJ13] Ignacio López de Ullibarri and M. Amalia Jácome.

                                        survPresmooth: An R pack-

age for presmoothed estimation in survival analysis. Journal of Statistical Software, 54(11):??, September 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v54/i11.

Duong:2007:KKD

[Duo07] Tarn Duong. ks: Kernel density estimation and kernel discriminant analysis for multivariate data in R. Journal of Statistical Software, 21 (7):1–16, October 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v21/i07.

Durante:2014:BRBa

[Dur14] Fabrizio Durante. Book review: The R Book, Second Edition by Michael J. Crawley. International Statistical Review = Revue Internationale de Statistique, 82(1):145–146, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Durante:2015:BRC

[Dur15] Fabrizio Durante. Book review: Computational Actuarial Science with R. International Statistical Review = Revue Internationale de Statistique, 83(3):511, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

deVries:2012:RDb

[dVM12] Andrie de Vries and Joris Meys. R for dummies. For dummies. Wiley, New York, NY, USA, 2012. ISBN 1-119-96284-6 (paperback), 1-119-96312-5 (e-book), 1-119-96313-3 (e-book), 1-119-96314-1 (e-book). xviii + 387 pp. LCCN QA276.45.R3 V75 2012.

Royuela-del-Val:2017:LFP

[dVSWAL17] Javier Royuela del Val, Federico Simmross-Wattenberg, and Carlos Alberola-López. libstable: Fast, parallel, and high-precision computation of α-stable distributions in R, C/C++, and MAT-LAB. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v078i01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v078i01/
v78i01.pdf.

Dong:2017:BRP

[DW17] Chen Dong and Michel Wedel. BANOVA: An R package for hierarchical Bayesian ANOVA. Journal of Statistical Software, 81(??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v081i09;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v081i09/v81i09.pdf.

deWreede:2011:MRP

[dWFP11] Liesbeth C. de Wreede, Marta Fiocco, and Hein Putter. mstate: An R package for the analysis of competing risks and multistate models. Journal of Statistical Software, 38(7): ??, January 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v38/i07.

Eastwood:2003:BRB

[Eas03] Brian J. Eastwood. Book review: Applied Statistics in the Pharmaceutical Industry with Case Studies Using S-PLUS, by Steven P. Millard; Andreas Krause. Journal of the American Statistical Association, 98(461): 250, March 2003. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        30045215.

Eddelbuettel:2018:ERC

[EB18] Dirk Eddelbuettel and James Joseph Balamuta. Extending R with C++: A brief introduction to Rcpp. The American Statistician, 72(1):28– 36, 2018. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

amstat.tandfonline.com/
doi/full/10.1080/00031305.
2017.1375990.

Erni:2013:MRP

[EBO^{+}13] Birgit Erni, Bo T. Bonnevie, Hans-Dieter Oschadleus, Res Altwegg, and Les G. Underhill. moult: An R package to analyze moult in birds. Journal of Statistical Software, 52 (8):??, February 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v52/i08.

Eddelbuettel:2009:BRBb

[Edd09a] Dirk Eddelbuettel. Book review: Analysis of Integrated and Cointegrated Time Series with R (2nd Edition). Journal of Statistical Software, 30(BR-5):??, April 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v30/b05.

Eddelbuettel:2009:BRBc

[Edd09b] Dirk Eddelbuettel. Book review: Applied Econometrics with R. Journal of Statistical Software, 29(BR-14):

??, February 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v29/b14.

Eddelbuettel:2009:BRBa

[Edd09c] Dirk Eddelbuettel. Book review: Computational Statistics: An Introduction to R. Journal of Statistical Software, 32(BR-2): ??, December 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/b02.

Eddelbuettel:2011:BRB

[Edd11] Dirk Eddelbuettel. Book review: R in a Nutshell. Journal of Statistical Software, 40 (BR-1):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v40/b01.

Eddelbuettel:2012:BRBb

[Edd12a] Dirk Eddelbuettel. Book review: R for Dummies. Journal of Statistical Software, 50(BR-1):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v50/
                                        b01.

Eddelbuettel:2012:BRBa

[Edd12b] Dirk Eddelbuettel. Book review: R in Action. Journal of Statistical Software, 46(BR-2):??, February 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v46/
                                        b02.

Eddelbuettel:2018:BRN

[Edd18] Dirk Eddelbuettel. Book review: Norman Matloff. Parallel Computing for Data Science: With Examples in R, C++, and CUDA. Boca Raton: CRC Press. Biometrics, 74(2):770, June 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eddelbuettel:2021:PCR

[Edd21] Dirk Eddelbuettel. Parallel computing with R: A brief review. WIREs Computational Statistics, 13(2): e1515:1–e1515:??, March– April 2021. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Erdman:2007:BRP

[EE07] Chandra Erdman and John W.

Emerson. bcp: An R package for performing a Bayesian analysis of change point problems. Journal of Statistical Software, 23(3): 1–13, December 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i03.

Eddelbuettel:2011:RSR

[EF11] Dirk Eddelbuettel and Romain François. Rcpp: Seamless R and C++ integration. Journal of Statistical Software, 40(8):1–18, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v40/
i08.

Everitt:2006:HSA

[EH06] Brian Everitt and Torsten Hothorn. A Handbook of Statistical Analyses Using R. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-539-

4. xvi + 275 pp. LCCN QA276.45.R3 E94 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0648/
2005056944-d.html;   http:
//www.loc.gov/catdir/toc/
fy0608/2005056944.html.

Everitt:2011:IAM

[EH11] Brian Everitt and Torsten Hothorn. An introduction to applied multivariate analysis with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-9649-4, 1-4419-9650-8. xiv + 273 pp. LCCN QA278 .E87 2011.

URL http://www.loc.gov/
catdir/enhancements/fy1304/
2011926793-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1304/2011926793-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1304/2011926793-
t.html.

Elsner:2013:HCM

[EJ13] James B. Elsner and Thomas H. Jagger. Hurricane Climatology: a Modern Statistical Guide Using R. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-982763-X (hardcover). xiv + 373 pp. LCCN QC944 .E47 2013.

Eubank:2012:SCC

[EK12] Randall L. Eubank and Ana Kupresanin. Statistical computing in C++ and R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4200-6650-1 (hardcover). xv + 540 pp. LCCN QA276.4 .E87 2012.

Egert:2023:RRI

[EK23] Janine Egert and Clemens Kreutz. Rcall: an R interface for MATLAB. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001947.

Eugster:2011:HTP

[EKP^{+}11] Manuel J. A. Eugster, Jochen Knaus, Christine Porzelius, Markus Schmidberger, and Esmeralda Vicedo. Hands-on tutorial for parallel computing with R. Computational Statistics, 26(2):219– 239, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-010-0206-4.

Ekstrom:2012:RP

[Eks12] Claus Thorn Ekstrøm. The R primer. CRC Press/Taylor and Francis, Boca Raton, FL, USA, 2012. ISBN 1-4398-6206-0 (hardcover). xii + 287 pp. LCCN QA276.45.R3 E37 2012.

Ekstrom:2016:RP

[Eks16] Claus Thorn Ekstrøm. The R primer. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2016. ISBN 1-4987-7255-2 (hardcover), 1-138-63197-3 (paperback), 1-138-05856-4 (paperback). xvii + 408 pp. LCCN QA276.45.R3 E37 2016.

Eugster:2009:SMH

[EL09] Manuel J. A. Eugster and Friedrich Leisch. From Spider-Man to hero — archetypal analysis in R. Journal of Statistical Software, 30(8):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v30/
                                        i08.

Emura:2017:RRP

[ELS17] Takeshi Emura, Ting-Hsuan Long, and Li-Hsien Sun. R routines for performing estimation and statistical process control under copulabased time series models. Communications in Statistics: Simulation and Computation, 46(4):3067–3087, 2017. CODEN CSSCDB. ISSN 0361-0918.

Emerson:2008:BRI

[Eme08] John W. Emerson. Book review: Interactive and Dynamic Graphics for Data Analysis: With R and GGobi, by Cook, D. and Swayne, D. Biometrics, 64(4):1301–1303, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eckley:2011:LIL

[EN11] Idris A. Eckley and Guy P.

Nason. LS2W: Implementing the locally stationary 2D wavelet process approach in R. Journal of Statistical Software, 43 (3):??, July 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v43/i03.

Er:2012:MGL

[Er12] Sebnem Er. Multivariate generalized linear models using R. Journal of Applied Statistics, 39(8):1851– ??, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Everitt:2001:AMD

[ERH01] Brian Everitt and Sophia Rabe-Hesketh. Analyzing medical data using S-PLUS. Statistics for biology and health. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-98862-9. ???? pp. LCCN RA409.5 E94 2001.

Eddelbuettel:2014:RAR

[ES14] Dirk Eddelbuettel and Conrad Sanderson. RcppArmadillo: Accelerating R with high-performance C++ linear algebra. Computational Statistics & Data Analysis, 71(??):1054–1063, March 2014. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0167947313000492.

Eskelson:2021:BRB

[Esk21] Bianca N. I. Eskelson. Book review: Biometry for forestry and environmental data with examples in R. Mehtätalo, Lauri and Lappi, Juha. Boca Raton, FL : Chapman and Hall/CRC, 2020. pp. 426. Biometrics, 77(4):1500–1502, December 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eddelbuettel:2016:REC

[ESO16] Dirk Eddelbuettel, Murray Stokely, and Jeroen Ooms. RProtoBuf: Efficient crosslanguage data serialization in R. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK.

ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v071i02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v071i02/
v71i02.pdf.

Espejo:2015:BRSb

[Esp15] Mariano Ruiz Espejo. Book review: Statistical Studies of Income, Poverty and Inequality in Europe: Computing and Graphics in R Using EU-SILC. International Statistical Review = Revue Internationale de Statistique, 83(1):173–174, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Evans:2011:BRB

[Eva11]  Gary Evans.  Book re-

view: Exploratory Multivariate Analysis by Example Using R. Journal of Statistical Software, 40(BR-2):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/b02.

Evans:2014:BRA

[Eva14]  Gary Evans.  Book re-

view: Analyzing Spatial Models of Choice and Judgment with R. Journal of Statistical Software, 58(BR-1):??, July 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v58/b01.

Everitt:1994:HSA

[Eve94] Brian S. Everitt. A Handbook of Statistical Analyses using S-PLUS. Chapman and Hall, Ltd., London, UK, 1994. ISBN 0-412-56310-X. viii + 143 pp. LCCN QA276.4.E93 1994.

Everitt:2002:HSA

[Eve02] Brian S. Everitt. A Handbook of Statistical Analyses using S-PLUS. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2002. ISBN 1-58488-280-8. xiii + 240 pp. LCCN QA276.4.E93 1994.

Everitt:2005:RPC

[Eve05] Brian S. Everitt. An R and S-PLUS companion to multivariate analysis. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-882-2. xiii + 221 pp. LCCN QA278 .E926 2005.

Eddelbuettel:2016:PRR

[EW16] Dirk Eddelbuettel and Wush Wu. RcppCNPy: Readwrite support for NumPy files in R. Journal of Open Source Software, 1 (5):55:1, September 2016. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00055.

Eyles:2022:BML

[Eyl22] Emily Eyles. Beyond multiple linear regression: Applied generalized linear models and multilevel models in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (2):726–727, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Falissard:2012:AQD

[Fal12] Bruno Falissard. Analysis of questionnaire data with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-1766-9 (hardcover), 1-4398-1767-7. ix + 269 pp. LCCN QA276.45.R3 F35 2012.

Fox:2020:PPR

[FAM^{+}20] Nathan Fox, Tom August, Francesca Mancini, Katherine E. Parks, Felix Eigenbrod, James M. Bullock, Louis Sutter, and Laura J.

Graham. “photosearcher”

package in R: an accessible and reproducible method for harvesting large datasets from Flickr. SoftwareX, 12 (??):Article 100624, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S235271102030337X.

Faraway:2005:LMR

[Far05] Julian James Faraway. Linear models with R, volume 63 of Texts in statistical science. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-425-8. x + 229 pp. LCCN QA279 .F37 2005.

Faraway:2006:ELM

[Far06] Julian James Faraway. Extending the Linear Model with R: Generalized Linear, Mixed Effects and Nonparametric Regression Models. Chapman and Hall/CRC texts in statistical science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-424-X. ix + 301 pp. LCCN QA279 .F368 2006.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0646/
                                        2005054822-d.html.

Faraway:2016:ELM

[Far16] Julian James Faraway. Extending the Linear Model with R: Generalized Linear, Mixed Effects and Nonparametric Regression Models. Chapman and Hall/CRC texts in statistical science series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, second edition, 2016. ISBN 1-4987-2096-X. xiii + 399 pp. LCCN QA279 .F368 2016.

Francq:2020:TAT

[FBB20] Bernard G. Francq, Marion Berger, and Charles Boachie. To tolerate or to agree: a tutorial on tolerance intervals in method comparison studies with BivRegBLS R package. Statistics in Medicine, 39(28):4334–4349, December 10, 2020. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finley:2007:SRP

[FBC07] Andrew O. Finley, Sudipto Banerjee, and Bradley P. Carlin. spBayes: An R package for univariate and multivariate hierarchical pointreferenced spatial models. Journal of Statistical Software, 19(4):1–24, April 2007. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v19/
i04.

Febrero-Bande:2012:SCF

[FBdlF12] Manuel Febrero-Bande and Manuel Oviedo de la Fuente. Statistical computing in functional data analysis: The R package fda.usc. Journal of Statistical Software, 51(4):??, October 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v51/
i04.

Fresno:2014:LLM

[FBF14] Cristóbal Fresno, Mónica Balzarini, and Elmer Fernández. lmdme: Linear models on designed multivariate experiments in R. Journal of Statistical Software, 56(7): ??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/i07.

Fischer:2021:PRR

[FBNRG21] Daniel Fischer, Alain Berro, Klaus Nordhausen, and Anne Ruiz-Gazen. REPPlab: an R package for detecting clusters and outliers using exploratory projection pursuit. Communications in Statistics: Simulation and Computation, 50(11):3397– 3419, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Fiske:2011:URP

[FC11] Ian Fiske and Richard Chandler. unmarked: An R package for fitting hierarchical models of wildlife occurrence and abundance. Journal of Statistical Software, 43 (10):??, August 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/i10.

Fox:2012:RSP

[FC12] John Fox and Marilia S. Carvalho. The RcmdrPlugin.survival package: Extending the R Commander interface to survival analysis. Journal of Statistical Software, 49(7):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v49/
i07.

Ferguson:2012:MRP

[FDB12] Nicole Ferguson, Somnath Datta, and Guy Brock. msSurv: An R package for nonparametric estimation of multistate models. Journal of Statistical Software, 50(14):??, September 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v50/
i14.

Fernandez-Duran:2016:CRP

[FDGD16] Juan José Fernández-Durán and María Mercedes Gregorio-Domínguez. CircNNTSR: An R package for the statistical analysis of circular, multivariate circular, and spherical data using nonnegative trigonometric sums. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v070i06;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v070i06/
v70i06.pdf.

Fellows:2012:DDA

[Fel12] Ian Fellows. Deducer: a data analysis GUI for R. Journal of Statistical Software, 49(8):??, June 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v49/i08.

Fernandez:2002:BRB

[Fer02] Carmen Fernández. Book review: Modern Applied Statistics with S-PLUS, by W. N. Venables; B. D. Ripley. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (2):302–303, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        3650346.

Ferl:2011:DRP

[Fer11] Gregory Z. Ferl. DATforDCEMRI: An R package for deconvolution analysis and visualization of DCE-MRI data. Journal of Statistical Software, 44(3): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v44/i03.

Fewster:2009:BRE

[Few09] Rachel M. Fewster. Book review: Ecological Models and Data in R, by Bolker, B. M. Biometrics, 65(2):669– 670, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fox:2009:VHT

[FFM09] John Fox, Michael Friendly, and Georges Monette. Visualizing hypothesis tests in multivariate linear models: the heplots package for R. Computational Statistics, 24(2):233– 246, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-008-0120-1.

Fritz:2012:TRP

[FGEMI12] Heinrich Fritz, Luis A.

García-Escudero, and Agustín Mayo-Iscar. tclust: An R package for a trimming approach to cluster analysis. Journal of Statistical Software, 47(12):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v47/
i12.

Farrell:1994:HDB

[FGG^{+}94] W. E. Farrell, J. Gaffney, J. Given, R. D. Jenkins, and N. Hall. A hydrographic database built on Montage and S-PLUS. Technical Report S2K-94-47, University of California, Berkeley, Berkeley, CA, USA, March 1994. 31 pp.

Filzmoser:2014:SIS

[FGZ14] Peter Filzmoser, Cristian Gatu, and Achim Zeileis. Special issue on statistical algorithms and software in R. Computational Statistics & Data Analysis, 71 (??):887–888, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167947313003654.

Fox:2009:EDR

[FH09] John Fox and Jangman Hong. Effect displays in R for multinomial and proportional-odds logit models: Extensions to the effects package. Journal of Statistical Software, 32 (1):??, October 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/i01.

Ferwerda:2017:KBR

[FHH17] Jeremy Ferwerda, Jens Hainmueller, and Chad J. Hazlett. Kernel-based regularized least squares in R (KRLS) and Stata (krls). Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v079i03;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v079i03/
                                        v79i03.pdf.

Feinerer:2008:TMI

[FHM08] Ingo Feinerer, Kurt Hornik, and David Meyer. Text mining infrastructure in R. Journal of Statistical Software, 25 (5):1–54, March 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v25/i05.

Fieberg:2012:EPA

[Fie12] John R. Fieberg. Estimating population abundance using sightability models: R SightabilityModel package. Journal of Statistical Software, 51(9):??, November 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i09.

Fieller:2015:BMA

[Fie15]  N. R. J. Fieller.  Ba-

sics of Matrix Algebra for Statistics with R. Chapman and Hall/CRC the R series. Taylor and Francis, Boca Raton, FL, USA, 2015. ISBN 1-4987-1236-3 (paperback). xviii + 226 pp. LCCN QA272 .F54 2015.

Filzmoser:2008:LRS

[Fil08] Peter Filzmoser. L^{A}TEX and R via Sweave: An example document how to use Sweave. Technical report, Vienna University of Technology, Wiedner Haupstraße 8-10, A-1040 Wien, Austria, March 12, 2008. 6

pp.  URL http://www.
statistik.tuwien.ac.at/
public/filz/students/SweaveExa.
pdf.

Finger:2010:RRS

[Fin10] Robert Finger. Review of ‘robustbase’ software for R. Journal of Applied Econometrics, 25(7):1205–1210, November–December 2010. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Finley:2011:BRE

[Fin11] Andrew Finley. Book review: Environmental and Ecological Statistics with R, by Qian, S. S. Biometrics, 67(2):674–675, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finch:2014:MMU

[Fin14] W. Holmes (William Holmes) Finch. Multilevel modeling using R, volume 16 of Chapman and Hall/CRC statistics in the social and behavioral sciences. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-1585-6 (paperback). xiii + 216 pp. LCCN HA31.35 .F56 2014.

Firth:2003:CFP

[Fir03] David Firth. CGIwithR: Facilities for processing Web forms using R. Journal of Statistical Software, 8(10): 1–8, ???? 2003. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v08/i10;
                                         http://www.jstatsoft.
                                        org/v08/i10/CGIwithR-overview.
                                        pdf.

Firth:2005:BTM

[Fir05] David Firth. Bradley–Terry models in R. Journal of Statistical Software, 12(1): 1–12, ???? 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=107&url=v12/i01&
ct=2;   http://www.jstatsoft.
org/counter.php?id=107&
url=v12/i01/v12i01.pdf&
ct=1.

Fischer:2006:BRZ

[Fis06] Matthias Fischer. Book review: Zivot Eric and Jiahui Wang, Modeling Financial Time Series with S-PLUS. AStA. Advances in Statistical Analysis, 90(4):631– 632, December 2006. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10182-
006-0013-y.

Frick:2017:TIR

[FK17] Hannah Frick and Ioannis Kosmidis. trackeR: Infrastructure for running and cycling data from GPSenabled tracking devices in R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i07;   https:/
/www.jstatsoft.org/index.

                                        php/jss/article/view/v082i07/
                                        v82i07.pdf.

Friedrich:2017:CSG

[FKP17] Sarah Friedrich, Frank Konietschke, and Markus Pauly. Code snippet: GFD: An R package for the analysis of general factorial designs. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v079c01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v079c01/
                                        v79c01.pdf.

Fox:2016:RJS

[FL16] John Fox and Allison Leanage. R and the Journal of Statistical Software. Journal of Statistical Software, 73(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v073i02;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v073i02/
                                        v73i02.pdf.

Fletcher:2011:BRR

[Fle11] David Fletcher. Book review: R Through Excel: a Spreadsheet Interface for Statistics, Data Analysis, and Graphics, by Richard M. Heiberger and Erich Neuwirth. Australian & New Zealand Journal of Statistics, 53(4):491–492, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Freeman:2008:PRP

[FM08] Elizabeth A. Freeman and Gretchen Moisen. PresenceAbsence: An R package for presence absence analysis. Journal of Statistical Software, 23(11): 1–31, January 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i11.

Friendly:2016:DDA

[FM16] Michael Friendly and David Meyer. Discrete Data Analysis with R: Visualization and Modeling Techniques for Categorical and Count Data, volume 120 of Chapman and Hall/CRC texts in statistical science series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4987-2583-X (pack book and ebook). xvii + 544 pp. LCCN QA300 .F744 2016.

Foss:2018:KCM

[FM18a] Alexander H. Foss and Marianthi Markatou. kamila: Clustering mixed-type data in R and Hadoop. Journal of Statistical Software, 83(??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/

                                        jss/article/view/v083i13;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v083i13/v83i13.pdf.

Freire:2018:PPR

[FM18b] Danilo Freire and Robert Myles

                                        McDonnell. prisonbrief:

An R package that returns tidy data from the World Prison Brief website. Journal of Open Source Software, 3(22):361:1, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00361.

Fernandez-Macho:2019:PPW

[FM19] J. Fernández-Macho. Package Wavemulcor: Wavelet multiple regression and correlation in R. Computing in Science and Engineering, 21 (6):63–73, November 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Field:2012:DSU

[FMF12] Andy P. Field, Jeremy Miles, and Zoë Field. Discovering statistics using R. Sage, London, UK, 2012. ISBN 1-4462-0045-0 (hardcover), 1-4462-0046-9. xxxiv + 957 pp. LCCN HA32 .F537 2012.

Fischer:2015:MWT

[FO15] Daniel Fischer and Hannu Oja. Mann–Whitney type tests for microarray experiments: The R package gMWT.

Journal of Statistical Software, 65(9):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/
i09.

Fog:2017:PRN

[Fog17] Agner Fog. Pseudo random number generators: uniform and non-uniform distributions. Web site, 2017.

URL http://www.agner.
org/random/.

Fortin:2020:BRS

[For20] Marie-Josée Fortin. Book review: Spatial data analysis in ecology and agriculture using R, Plant, Richard E., Boca Raton, FL: CRC Press, 2019. Biometrics, 76(1):362, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fotopoulos:2007:BRBb

[Fot07] Stergios B. Fotopoulos. Book review: Modeling Financial Time Series with S-PLUS by Eric Zivot; Jiahui Wang. Technometrics, 49(3):360, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471357.

Foulkes:2009:ASG

[Fou09] Andrea S. Foulkes. Applied Statistical Genetics with R: for Population-based Association Studies. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-89553-1. xxiii + 252 pp. LCCN QH438.4.S73 F68 2009. URL http://

springerlink.com/openurl.
asp?genre=book&isbn=978-
0-387-89553-6;   http:/
/www.loc.gov/catdir/enhancements/
fy0913/2009284624-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0913/2009284624-
t.html.

Fox:2002:RPC

[Fox02] John Fox. An R and S-Plus companion to applied regression. Sage Publications, Thousand Oaks, CA, USA, 2002. ISBN 0-7619-2279-2, 0-7619-2280-6 (paperback). xvi + 312 pp. LCCN QA278.2 .F628 2002.

Fox:2003:EDR

[Fox03] John Fox. Effect displays in R for generalised linear models. Journal of Statistical Software, 8(15): 1–27, ???? 2003. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v08/i15;
 http://www.jstatsoft.
org/v08/i15/effect-displays-
revised.pdf.

Fox:2005:BRB

[Fox05a] John Fox. Book review: The Basics of S-Plus, Fourth Edition. Journal of Statistical Software, 14(BR-8): 1–3, December 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v14/b08.

Fox:2005:RCB

[Fox05b] John Fox. The R Commander: a basic-statistics graphical user interface to R. Journal of Statistical Software, 14(9):1–42, August 19, 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=134&url=v14/i09&
ct=2;   http://www.jstatsoft.
org/counter.php?id=134&
url=v14/i09/v14i09.pdf&
ct=1.

Friendly:2011:BRG

[Fri11] Michael Friendly. Book review: Graphics for Statistics and Data Analysis with R, by Keen, K. J. Biometrics, 68(3):1177, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fricker:2012:BRB

[Fri12] Ronald D. Fricker. Book review: Analysis of Questionnaire Data with R. Journal of Statistical Software, 46(BR-1):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v46/
b01.

Friedman:2016:SLI

[Fri16] Alon Friedman. Statistics for library and information services: a primer for using open source R software for accessibility and visualization. Rowman and Littlefield, Lanham, UK, 2016. ISBN 1-4422-4992-7 (hardcover), 1-4422-4993-5 (ebook). xvi + 355 pp. LCCN Z669.8 .F75 2016.

Fay:2010:EAW

[FS10a] Michael P. Fay and Pamela A. Shaw. Exact and asymptotic weighted logrank tests for interval censored data: The interval R package. Journal of Statistical Software, 36 (2):??, August 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v36/i02.

Furrer:2010:SSM

[FS10b] Reinhard Furrer and Stephan R. Sain. spam: a sparse matrix R package with emphasis on MCMC methods for Gaussian Markov random fields. Journal of Statistical Software, 36(10): ??, September 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v36/i10.

Feng:2008:CDI

[FT08] Dai Feng and Luke Tierney. Computing and displaying isosurfaces in R. Journal of Statistical Software, 28 (1):??, September 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v28/i01.

Fujino:2017:VGW

[Fuj17] Tomokazu Fujino. vdmR:

Generating Web-based visual data mining tools with R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i06;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i06/
v82i06.pdf.

Fusfeld:2022:BRS

[Fus22] Zachary Fusfeld. Book review: Statistical inference via data science: a modern dive into R and the Tidyverse, Ismay, Chester and Kim, Albert Y. Boca Raton, FL: Chapman and Hall/ CRC, 2019. pp. 460. Biometrics, 78(1):410–412, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fox:2023:VRP

[FVVL23] Nathan Fox, Derek Van Berkel, Ramiro Serrano Verge, and Mark Lindquist.

vGameReviews: an R pack-

age for harnessing video game reviews for scientific research. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S235271102300119X.

Fenn:2018:QBH

[FW18] Roger Fenn and Ansgar Wenzel. Quandle and biquandle homology calculation in R. Journal of Open Research Software, 6 (1):3–??, January 16, 2018. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.53/.

Guillen:2023:BBA

[GAE23] Maria D. Guillen, Juan Aparicio, and Victor J.

                                        España.  boostingDEA:

a boosting approach to Data Envelopment Analysis in R. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023002455.

Galecki:2017:BRF

[Gal17] Andrzej Galecki. Book review: Faraway, Julian J., Extending the Linear Model with R: Generalized Linear, Mixed Effects, and Nonparametric Regression Models, 2nd edition. Boca Raton: CRC Press. Biometrics, 73(4):1465–1466, December 2017. CODEN BIOMB6.

ISSN 0006-341X (print), 1541-0420 (electronic).

Gandrud:2015:SRP

[Gan15a] Christopher Gandrud. simPH:

An R package for illustrating estimates from Cox proportional hazard models including for interactive and nonlinear effects. Journal of Statistical Software, 65 (3):??, June 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v65/i03.

Gandrud:2015:RRR

[Gan15b] Christopher Gandrud. Reproducible Research with R and R Studio. Chapman and Hall / CRC The R Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2015. ISBN 1-4987-1537-0 (paperback), 1-4987-1538-9 (e-book). xxvii + 295 pp. LCCN Q180.55.S7G36

2015. URL http://public.
eblib.com/choice/publicfullrecord.
aspx?p=4710300;   http:
//public.ebookcentral.
proquest.com/choice/publicfullrecord.
aspx?p=4710300.

Gasparrini:2011:DLL

[Gas11] Antonio Gasparrini. Distributed lag linear and nonlinear models in R: The package dlnm. Journal of Statistical Software, 43 (8):??, July 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/i08.

Gasparini:2018:PCR

[Gas18] Alessandro Gasparini. comorbidity: An R package for computing comorbidity scores. Journal of Open Source Software, 3(23):648:1–648:3, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00648.

Genolini:2015:KKR

[GASA15] Christophe Genolini, Xavier Alacoque, Mariane Sentenac, and Catherine Arnaud. kml and kml3d: R packages to cluster longitudinal data. Journal of Statistical Software, 65 (4):??, June 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v65/i04.

Gavrilovska:2010:AHP

[Gav10] Ada Gavrilovska, editor. Attaining high performance communications: a vertical approach. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-9308-8 (hardcover). xxxii + 411 pp. LCCN QA76.88 .A77 2010.

Galecki:2013:LME

[GB13] Andrzej T. Gałecki and Tomasz Burzykowski. Linear mixed-effects models using R: a step-by-step approach.

Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-3899-3. ISSN 1431-875X. 542 pp. LCCN QA279 .G35 2013.

Garcia-Ceja:2023:BRC

[GC23] Enrique Garcia-Ceja. Book review: A Criminologist’s Guide to R: Crime by the Numbers, Jacob Kaplan, Boca Raton, FL: Chapman and Hall/CRC Press, Taylor & Francis Group, 2022, 432 pp., ISBN 978-1-032-24407-5. Technometrics, 65(4):610–611, 2023. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Goncalves:2012:RPB

[GCA12] M. Helena Gonçalves, M. Salomé Cabral, and Adelchi Azzalini. The R package bild for the analysis of binary longitudinal data. Journal of Statistical Software, 46 (9):??, February 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/i09.

Gelman:2014:BDA

[GCS^{+}14] Andrew Gelman, John B.

Carlin, Hal S. Stern, David B. Dunson, Aki Vehtari, and Donald B. Rubin. Bayesian data analysis. Chapman and Hall/CRC texts in statistical science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2014. ISBN 1-4398-4095-4 (hardcover). xiv + 661 pp. LCCN QA279.5 .G45 2014. URL

                                        http://images.tandf.co.
                                        uk/common/jackets/websmall/
                                        978143984/9781439840955.
                                        jpg.

Goel:2021:WWE

[GDBK^{+}21] Aviral Goel, Pierre Donat-Bouillud, Filip Krikava, Christoph M. Kirsch, and Jan Vitek. What we eval in the shadows: a large-scale study of eval in R programs. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA): 125:1–125:23, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3485502.

Gonzalez:2008:CRP

[GDMB08] Ignacio González, Sébastien Déjean, Pascal G. P. Martin, and Alain Baccini. CCA: An R package to extend canonical correlation analysis. Journal of Statistical Software, 23(12):1– 14, January 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v23/i12.

Gentleman:1998:PCRb

[Gen98] Robert Gentleman. Personal crunching: Review of S-PLUS 4.5. Chance, 11 (4):42–45, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:2009:RPB

[Gen09] Robert Gentleman. R programming for bioinformatics. Chapman and Hall/CRC computer science and data analysis series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6367-7. xii + 314 pp. LCCN QH324.2 .G46 2009.

URL http://www.loc.gov/
catdir/enhancements/fy0901/
2008011352-d.html;   http:
//www.loc.gov/catdir/toc/
fy0805/2008011352.html.

Gepp:2021:BRA

[Gep21] Adrian Gepp. Book review:

Advanced Statistics with Applications in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1610–1611, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Gersten:1994:PAI

[Ger94] R. Carey Gersten. S-PLUS for ARC/INFO: your link to a powerful statistical view, 1994. Credits: Producer, R. Carey Gersten. Demonstrates the advantages of combining the ARC/INFO computer program with S-PLUS, as a means of GIS analysis to provide more productive and insightful spatial analysis and statistical modeling. Applications are broad including natural resource planning, business geodemographic analysis and urban and rural planning.

Gerbing:2020:RVD

[Ger20] David Gerbing. R Visualizations: Derive Meaning from Data. CRC Press LLC, Milton, UK, 2020. ISBN 0-429-89491-0 (e-book), 0-429-89492-9 (e-book), 1-138-59963-8 (hardcover). 237 pp. LCCN QA76.9.I52.

Gonzalez-Estrada:2018:RPT

[GEV18] E. González-Estrada and J. A. Villaseñor. An R package for testing goodness of fit: goft. Journal of Statistical Computation and Simulation, 88(4):726–751, 2018. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Giudici:2009:ADM

[GF09] Paolo Giudici and Silvia Figini. Applied data mining for business and industry. Wiley, New York, NY, USA, second edition, 2009. ISBN 0-470-05886-2 (cloth), 0-470-05887-0 (paperback), 0-470-74582-7 (e-book). viii + 249 pp. LCCN QA76.9.D343 G75 2009. URL http://

www.wiley.com/WileyCDA/
WileyTitle/productCd-0470058862.
html.

Gopal:2012:BRP

[GFC12] Vikneswaran Gopal, Claudio Fuentes, and George Casella. bayesclust: An R package for testing and searching for significant clusters. Journal of Statistical Software, 47(14):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v47/i14.

Gonzalez-Fernandez:2014:PRP

[GFS14] Yasser Gonzalez-Fernandez and Marta Soto. copulaedas: An R package for estimation of distribution algorithms based on copulas. Journal of Statistical Software, 58(9):??, June 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v58/i09.

Gerber:2016:GRP

[GG16] Florian Gerber and Thomas Gsponer. gsbDesign: An R package for evaluating the operating characteristics of a group sequential Bayesian design. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069i11;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069i11/
v69i11.pdf.

Gaudart:2015:PRP

[GGC^{+}15] Jean Gaudart, Nathalie Graffeo, Drissa Coulibaly, Guillaume Barbet, Stanilas Rebaudet, Nadine Dessay, Ogobara K. Doumbo, and Roch Giorgi. SPODT: An R package to perform spatial partitioning. Journal of Statistical Software, 63(16): ??, February 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v63/i16.

Glaab:2010:VRP

[GGK10] Enrico Glaab, Jonathan M. Garibaldi, and Natalio Krasnogor. vrmlgen: An R package for 3D data visualization on the Web. Journal of Statistical Software, 36 (8):??, August 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v36/i08.

Gomez:2023:CUP

[GGS23] Camilo Gomez, Alexander V. Goponenko, and Julia N. Soulakova. Constructing UpSet plot for survey data with weights using SAS and R software. Communications in Statistics: Simulation and Computation, 52 (6):2320–2326, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Grun:2011:TRP

[GH11] Bettina Grün and Kurt

                                        Hornik. topicmodels: An

R package for fitting topic models. Journal of Statistical Software, 40(13): ??, May 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/i13.

Gagunashvili:2018:CCG

[GH18] Nikolay D. Gagunashvili and Helgi Halldorsson. CHI-WEI: A code of goodness of fit tests for weighted and unweighted histograms in Fortran-77, C++, R and Python. Computer Physics Communications, 231(??): 245, October 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465518301449.

Glassic:2019:CFR

[GHG19] Hayley C. Glassic, Kurt C.

Heim, and Christopher S. Guy. Creating figures in R that meet the AFS Style Guide: Standardization and supporting script. Fisheries, 44(11):539–544, November 2019. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:/

/afspubs.onlinelibrary.
wiley.com/doi/10.1002/
fsh.10272.

Gagunashvili:2019:CCC

[GHN19] Nikolay D. Gagunashvili, Helgi Halldorsson, and Helmut Neukirchen. CHICOM:

Code for comparing weighted or unweighted histograms in Fortran-77, C++, R and Python. Computer Physics Communications, 245(??):Article 106872, December 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/S0010465519302590.

Gentleman:2000:LSS

[GI00] Robert Gentleman and Ross Ihaka. Lexical scope and statistical computing. Journal of Computational and Graphical Statistics, 9(3): 491–508, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL

http://www.amstat.org/
publications/jcgs/abstracts00/
Ihaka.htm;   http://www.
tandfonline.com/doi/abs/
10.1080/10618600.2000.
10474895.

Gentleman:2005:BCB

[GIH^{+}05] Robert Gentleman, Rafael A. Irizarry, Wolfgang Huber, Sandrine Dudoit, and Vincent J. Carey, editors. Bioinformatics and computational biology solutions using R and Bioconductor. Statistics for biology and health. Springer Science+Business Media, New York, NY, USA, 2005. ISBN 0-387-25146-4. xix + 473 pp. LCCN QH324.2 .B5473

2005. URL http://www.
loc.gov/catdir/enhancements/
fy0662/2005923843-d.html;
http://www.loc.gov/catdir/
enhancements/fy0818/2005923843-
t.html.

Gillard:2014:BRB

[Gil14] Jonathan Gillard. Book review: The R book, second edition. Journal of Applied Statistics, 41(4): 909, 2014. ISSN 0266-4763 (print), 1360-0532 (electronic).

Govindasamy:2024:SRE

[GIM^{+}24] Priyalatha Govindasamy, Nor Junainah Mohd Isa, Nor Firdous Mohamed, Amelia Mohd Noor, Lin Ma, Antonio Olmos, and Kathy Green. A systematic review of exploratory factor analysis packages in R software. WIREs Computational Statistics, 16(1):e1630:1– e1630:??, January–February 2024. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https:

//wires.onlinelibrary.
wiley.com/doi/epdf/10.
1002/wics.1630.

Giorgino:2009:CVD

[Gio09] Toni Giorgino. Computing and visualizing dynamic time warping alignments in R: The dtw package. Journal of Statistical Software, 31 (7):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v31/i07.

Goel:2021:PMB

[GJK^{+}21] Aviral Goel, Jan Jecmen, Sebastián Krynski, Olivier Flückiger, and Jan Vitek. Promises are made to be broken: migrating R to strict semantics. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):101:1–101:20, October 2021. CODEN ???? ISSN 2475-1421 (electronic).

URL https://dl.acm.org/
doi/10.1145/3485478.

Golyandina:2014:BSS

[GK14] Nina Golyandina and Anton Korobeynikov. Basic singular spectrum analysis and forecasting with R. Computational Statistics & Data Analysis, 71 (??):934–954, March 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947313001394.

Gilleland:2016:EEV

[GK16] Eric Gilleland and Richard W. Katz. extRemes 2.0: An extreme value analysis package in R. Journal of Statistical Software, 72(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v072i08;

 https://www.jstatsoft.
org/index.php/jss/article/
view/v072i08/v72i08.pdf.

Giordani:2014:TWC

[GKD14] Paolo Giordani, Henk A. L.

Kiers, and Maria Antonietta Del Ferraro. Three-way component analysis using the R package ThreeWay. Journal of Statistical Software, 57(7):??, April 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v57/i07.

Grun:2012:EBR

[GKZ12] Bettina Grün, Ioannis Kosmidis, and Achim Zeileis. Extended beta regression in R: Shaken, stirred, mixed, and partitioned. Journal of Statistical Software, 48 (11):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v48/i11.

Ghysels:2016:MFD

[GKZ16] Eric Ghysels, Virmantas Kvedaras, and Vaidotas Zemlys. Mixed frequency data sampling regression models: The R package midasr. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v072i04;   https:/
/www.jstatsoft.org/index.

                                        php/jss/article/view/v072i04/
                                        v72i04.pdf.

Grun:2007:FFM

[GL07] Bettina Grün and Friedrich Leisch. Fitting finite mixtures of generalized linear regressions in R. Computational Statistics & Data Analysis, 51(11):5247–5252, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167947306002787.

Godfrey:2014:SRS

[GL14] A. Jonathan R. Godfrey and M. Theodor Loots. Software review: Statistical Software (R, SAS, SPSS, and Minitab) for Blind Students and Practitioners. Journal of Statistical Software, 58 (SR-1):??, July 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v58/s01.

Guerin:2015:MPE

[GL15] Greg R. Guerin and Andrew J. Lowe. Mapping phylogenetic endemism in R using georeferenced branch extents. SoftwareX, 3–4 (??):13–21, December 2015. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711015000126.

Gillespie:2016:ERP

[GL16a] Colin (Colin Stevenson)

Gillespie and Robin Lovelace. Efficient R programming: a practical guide to smarter programming. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2016. ISBN 1-4919-5078-1 (paperback), 1-4919-5072-2 (e-book), 1-4919-5075-7 (e-book), 1-4919-5073-0 (ebook). xiii + 203 pp. LCCN QA276.45.R3 G55 2016.

URL http://proquest.
safaribooksonline.com/
9781491950777.

Giraud:2016:PCC

[GL16b] Timothée Giraud and Nicolas Lambert. cartography: Create and integrate maps in your R workflow. Journal of Open Source Software, 1(4):54:1–54:2, August 2016. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00054.

Gollini:2015:PRP

[GLC^{+}15] Isabella Gollini, Binbin Lu, Martin Charlton, Christopher Brunsdon, and Paul Harris. GWmodel: An R package for exploring spatial heterogeneity using geographically weighted models. Journal of Statistical Software, 63 (17):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v63/i17.

Glenn:2016:BRS

[Gle16a] Ezra Haber Glenn. Book review: Spatial Microsimulation with R, Robin Lovelace and Morgane Dumont, Chapman & Hall/CRC, Boca Raton, 2016. ISBN 978-1-4987-1154-8. 259 pp. USD 89.95 (P), 62.97

                                        (eBook). https://www.crcpress.com/9781498711548.

Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

                                        www.jstatsoft.org/index.
                                        php/jss/article/view/v072b01;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v072b01/v72b01.pdf.

Glenn:2016:WAC

[Gle16b] Ezra Haber Glenn. Working with the American Community Survey in R: a guide to using the ACS package. Springer Science+Business Media, New York, NY, USA, 2016. ISBN 3-319-45771-

3. viii + 53 pp. LCCN HA201.122. URL http:

                                        //link.springer.com/10.
                                        1007/978-3-319-45772-7.

Gerber:2018:DRP

[GMF18] F. Gerber, K. Mösinger, and R. Furrer. dotCall64: an R package providing an efficient interface to compiled C, C++, and Fortran code supporting long vectors. SoftwareX, 7(??):211– 216, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018300785.

Goldstein:2011:BSU

[Gol11]  Richard Goldstein.  R

for Stata Users by Robert A. Muenchen; Joseph M. Hilbe. Technometrics, 53 (4):442, November 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714959.

Good:2005:IST

[Goo05]  Phillip I. Good.  Intro-

duction to statistics through resampling methods and R/S-PLUS. Wiley-Inter-science, New York, NY, USA, 2005. ISBN 0-471-71575-1 (paperback). xii + 229 pp. LCCN QA278.8 .G63 2005. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/wiley051/2004061922.
html.

Good:2013:IST

[Goo13] Phillip I. Good. Introduction to statistics through resampling methods and R. Wiley, New York, NY, USA, second edition, 2013. ISBN 1-118-42821-8 (paperback), 1-118-49756-2 (e-book). xii + 210 pp. LCCN QA278.8 .G63 2013. URL http:

                                        //onlinelibrary.wiley.
                                        com/book/10.1002/9781118497593.

Goslee:2011:ARS

[Gos11] Sarah C. Goslee. Analyzing remote sensing data in R: The landsat package. Journal of Statistical Software, 43(4):??, July 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/i04.

Gould:2005:BRB

[Gou05] Robert Gould. Book review: Using R for Introductory Statistics. Journal of Statistical Software, 14(BR-7):1–3, November 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v14/b07.

Gould:2010:BRB

[Gou10] Robert Gould. Book review: A Modern Approach to Regression with R. Journal of Statistical Software, 33 (BR-3):??, March 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v33/b03.

Gerber:2012:DAM

[GP12] Samuel Gerber and Kristin Potter. Data analysis with the Morse–Smale complex: The msr package for R. Journal of Statistical Software, 50(2):??, July 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i02.

Gabadinho:2016:ASS

[GR16] Alexis Gabadinho and Gilbert Ritschard. Analyzing state sequences with probabilistic suffix trees: The PST R package. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v072i03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v072i03/
v72i03.pdf.

Gomez-Rubio:2018:BRG

[GR18] Virgilio Gomez-Rubio. Book review: Generalized Additive Models: An Introduction with R (2nd Edition). Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086b01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086b01/
v86b01.pdf.

Gramacy:2007:TRP

[Gra07] Robert B. Gramacy. tgp:

An R package for Bayesian nonstationary, semiparametric nonlinear regression and design by treed Gaussian process models. Journal of Statistical Software, 19 (9):1–46, June 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v19/i09.

Gramacy:2016:LLS

[Gra16] Robert B. Gramacy. laGP: Large-scale spatial modeling via local approximate Gaussian processes in R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v072i01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v072i01/
                                        v72i01.pdf.

Gabriel:2013:SRP

[GRD13] Edith Gabriel, Barry S. Rowlingson, and Peter J. Diggle. stpp: An R package for plotting, simulating and analyzing spatio-temporal point patterns. Journal of Statistical Software, 53 (2):??, April 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v53/i02.

Green:2022:BRS

[Gre22a] Nathan Green. Book review: Statistical Rethinking: A Bayesian Course with Examples in R and Stan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(1):426– 427, January 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Greenacre:2022:BRR

[Gre22b] Michael Greenacre. Book review: R for Health Data Science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S765–S766, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/185/Supplement_
2/S765/7069548.

George:2016:RPC

[GRK^{+}16] Ann Cathrice George, Alexander Robitzsch, Thomas Kiefer, Jürgen Groß, and Ali Ünlü. The R package CDM for cognitive diagnosis models. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v074i02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v074i02/
v74i02.pdf.

Gabadinho:2011:AVS

[GRMS11] Alexis Gabadinho, Gilbert Ritschard, Nicolas S. Müller, and Matthias Studer. Analyzing and visualizing state sequences in R with TraMineR. Journal of Statistical Software, 40(4):??, April 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v40/
                                        i04.

Grothendieck:2008:BRB

[Gro08] Gabor Grothendieck. Book review: lattice: Multivariate Data Visualization with R. Journal of Statistical Software, 25(BR-2):1– 3, May 2008. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v25/b02.

Gromping:2010:ILE

[Grö10] Ulrike Grömping. Inference with linear equality and inequality constraints using R: The package ic.infer. Journal of Statistical Software, 33(10): ??, February 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v33/i10.

Gromping:2011:BRB

[Grö11] Ulrike Grömping. Book review: Optimal Experimental Design with R. Journal of Statistical Software, 43(BR-5):??, October 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/b05.

Gromping:2013:BRB

[Grö13] Ulrike Grömping. Book review: The R Primer. Journal of Statistical Software, 52(BR-3):??, February 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v52/
b03.

Grolemund:2014:HPR

[Gro14a] Garrett Grolemund. Handson programming with R. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2014. ISBN 1-4493-5901-9 (paperback). xiii + 230 pp. LCCN QA276.45.R3 G76 2014.

Gromping:2014:RPF

[Grö14b] Ulrike Grömping. R package FrF2 for creating and analyzing fractional factorial 2-level designs. Journal of Statistical Software, 56(1): ??, January 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v56/i01.

Gromping:2015:BRU

[Grö15a] Ulrike Grömping. Book review: Using R and RStudio for Data Management, Statistical Analysis and Graphics (2nd Edition), Nicholas J. Horton and Ken Kleinman, CRC Press, Boca Raton, 2015. ISBN 978-1-4822-3736-

8. 313 pp. USD 59.46 (H).

http://www.amherst.edu/ nhorton/r2/.

Journal of Statistical Software, 68(BR-4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/index.

                                        php/jss/article/view/v068b04;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v068b04/v68b04.pdf.

Gromping:2015:BRM

[Grö15b] Ulrike Grömping. Book review: Multilevel Modeling Using R. Journal of Statistical Software, 62(BR-1): ??, January 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v62/b01.

Gromping:2016:BRP

[Grö16] Ulrike Grömping. Book review: Practical Guide to Logistic Regression, Joseph M. Hilbe, Chapman & Hall/CRC, Boca Raton, 2016. ISBN 978-1-4987-0957-

6. 174 pp. USD 49.95 (P).

                                        https://www.crcpress.com/9781498709576.

Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

                                        www.jstatsoft.org/index.
                                        php/jss/article/view/v071b03;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v071b03/v71b03.pdf.

Gromping:2018:BRD

[Grö18a] Ulrike Grömping. Book review: A Data Scientist’s Guide to Acquiring, Cleaning and Managing Data in R. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086b03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086b03/
v86b03.pdf.

Gromping:2018:RPD

[Grö18b] Ulrike Grömping. R package DoE.base for factorial experiments. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085i05/
v85i05.pdf.

Grose:2018:BRT

[Gro18c] Daniel Grose. Book review:

Testing R Code. Journal of Statistical Software, 85 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v085b01;
https://www.jstatsoft.
org/index.php/jss/article/
view/v085b01/v85b01.pdf.

Grunsky:1995:PTD

[Gru95]  Eric Grunsky.  S-PLUS

trellis displays. Computers and Geosciences, 21(7): 917–??, 1995. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Gorecki:2019:FRS

[GS19] Tomasz Górecki and Lukasz Smaga. fdANOVA: an R software package for analysis of variance for univariate and multivariate functional data. Computational Statistics, 34(2):571– 597, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-018-0842-7;
 http://link.springer.
com/content/pdf/10.1007/
s00180-018-0842-7.pdf.

Gelling:2019:RPR

[GSB19] Nicholas Gelling, Matthew R. Schofield, and Richard J. Barker. R package rjmcmc: reversible jump MCMC using post-processing. Australian & New Zealand Journal of Statistics, 61(2):189– 212, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Galimberti:2012:CTO

[GSD12] Giuliano Galimberti, Gabriele Soffritti, and Matteo Di Maso. Classification trees for ordinal responses in R: The rpartScore package. Journal of Statistical Software, 47(10):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v47/i10.

Gelis:2022:DPC

[GSJC22] M. M. Nicolosi Gelis, M. B.

Sathicq, J. Jupke, and J. Cochero. DiaThor: R package for computing diatom metrics and biotic indices. Ecological Modelling, 465(??):Article 109859, March 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0304380021003975.

Gramacy:2010:CIS

[GT10] Robert B. Gramacy and Matthew Alan Taddy. Categorical inputs, sensitivity analysis, optimization and importance tempering with tgp version 2, an R package for treed Gaussian process models. Journal of Statistical Software, 33(6): ??, February 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v33/i06.

Gu:2014:SSA

[Gu14]  Chong Gu.  Smoothing

spline ANOVA models: R package gss. Journal of Statistical Software, 58(5): ??, June 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v58/i05.

Guay:2013:SSS

[Gua13] Jean H. Guay. Statistiques en sciences sociales avec R.

Collection Méthodes des sciences humaines. Presses de l’Université Laval, Laval, Québec, Canada, 2013. ISBN 2-7637-1773-X. xiii + 212 pp. LCCN QA276.45.R3 G918 2013.

Guha:2007:CIF

[Guh07]  Rajarshi Guha.  Chemi-

cal informatics functionality in R. Journal of Statistical Software, 18(5):1– 16, January 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v18/i05.

Gunter:2006:BRB

[Gun06] Berton H. Gunter. Book review: R Graphics by Paul Murrell. Technometrics, 48(2):308–309, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        25471176.

Guttorp:2011:BRH

[Gut11] Peter Guttorp. Book review: Hidden Markov Models for Time Series: an Introduction Using R, by Zucchini, W. and MacDonald, I. L. Biometrics, 68 (3):1178, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guolo:2012:RPM

[GV12] Annamaria Guolo and Cristiano Varin. The R package metaLik for likelihood inference in meta-analysis. Journal of Statistical Software, 50 (7):??, August 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i07.

Goel:2019:DIU

[GV19] Aviral Goel and Jan Vitek.

On the design, implementation, and use of laziness in R. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 153:1–153:27, October 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3360579.

Gruber:2012:TRP

[GvdL12] Susan Gruber and Mark van der Laan. tmle: An R package for targeted maximum likelihood estimation. Journal of Statistical Software, 51(13): ??, November 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i13.

Grasman:2009:FCC

[GvdMW09] Raoul Grasman, Han L. J.

van der Maas, and Eric-Jan Wagenmakers. Fitting the cusp catastrophe in R: a cusp package primer. Journal of Statistical Software, 32 (8):??, November 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/i08.

Griffith:2016:CSC

[GVM16] Daniel M. Griffith, Joseph A. Veech, and Charles J. Marsh. Code snippet: cooccur: Probabilistic species cooccurrence analysis in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

                                        www.jstatsoft.org/index.
                                        php/jss/article/view/v069c02;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v069c02/v69c02.pdf.

Groendyke:2018:ERP

[GW18] Chris Groendyke and David Welch. epinet: An R package to analyze epidemics spread across contact networks. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v083i11;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v083i11/
                                        v83i11.pdf.

Grun:2009:AGE

[GZ09] Bettina Grün and Achim Zeileis. Automatic generation of exams in R. Journal of Statistical Software, 29 (10):??, February 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v29/i10.

Goitisolo:2011:SAP

[GZ11] Beatriz Goitisolo and Amaya Zárraga. Simultaneous analysis in S-PLUS: The SimultAn package. Journal of Statistical Software, 40(11):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/i11.

Gajowniczek:2021:IRPa

[GZ21a] Krzysztof Gajowniczek and Tomasz Zabkowski. ImbTreeAUC: an R package for building classification trees using the area under the ROC curve (AUC) on imbalanced datasets. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000807.

Gajowniczek:2021:IRPb

[GZ21b] Krzysztof Gajowniczek and Tomasz Zabkowski. ImbTreeEntropy: an R package for building entropy-based classification trees on imbalanced datasets. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001242.

Grubinger:2014:PEL

[GZP14] Thomas Grubinger, Achim Zeileis, and Karl-Peter Pfeiffer. evtree: Evolutionary learning of globally optimal classification and regression trees in R. Journal of Statistical Software, 61 (1):??, October 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i01.

Huet:1996:STN

[H^{+}96] S. (Sylvie) Huet et al. Statistical tools for nonlinear regression: a practical guide with S-PLUS examples. Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94727-

2. ix + 154 pp. LCCN QA278.2.S73 1996.

Heritier:2009:RMB

[H^{+}09] Stephane Heritier et al., editors. Robust methods in biostatistics. Wiley series in probability and statistics. Wiley, New York, NY, USA, 2009. ISBN 0-470-02726-6 (cloth), 0-470-74054-X (e-book). xiv + 268 pp. LCCN QH323.5 .R615 2009. URL http://

www.wiley.com/WileyCDA/
WileyTitle/productCd-0470027266.
html.

Hackl:2017:BRC

[Hac17]  Peter Hackl.  Book re-

view: Claus Thorn Ekstrøm (2017): The R Primer, Second Edition, Chapman & Hall/CRC, The R Series, xvii + 408 pp., £39.99, ISBN 978-1-138-63197-7. Statistical Papers, 58(4):1281–1282, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00362-017-0954-0.

Hadfield:2010:MMM

[Had10] Jarrod D. Hadfield. MCMC methods for multi-response generalized linear mixed models: The MCMCglmm R package. Journal of Statistical Software, 33(2):??, February 2010. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v33/i02.

Haggstrom:2012:SBRa

[Häg12] Olle Häggström. Short book review: The R Primer by Claus Thorn Ekstrøm. International Statistical Review = Revue Internationale de Statistique, 80 (1):197, April 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
23257192.

Hallman:1993:RP

[Hal93] J. Hallman. Review of S-PLUS. Journal of Applied Econometrics, 8(2):213–219, April 1993. CODEN JAE-CET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hansen:1998:BRB

[Han98] Mark Hansen. Book review: Applied Wavelet Analysis with S-PLUS by Andrew Bruce; Hong-Ye Gao. Technometrics, 40 (2):159, May 1998. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1270657.

Hankin:2005:IBR

[Han05] Robin K. S. Hankin. Introducing BACCO, an R bundle for Bayesian analysis of computer code output. Journal of Statistical Software, 14(16):1–21, October 19, 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=141&url=v14/i16&
                                        ct=2;   http://www.jstatsoft.
                                        org/counter.php?id=141&
                                        url=v14/i16/v14i16.pdf&
                                        ct=1.

Hankin:2006:CSA

[Han06a] Robin K. S. Hankin. Code snippet: Additive integer partitions in R. Journal of Statistical Software, 16(CS-1):1–3, May 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/c01.

Hankin:2006:IER

[Han06b] Robin K. S. Hankin. Introducing elliptic, an R package for elliptic and modular functions. Journal of Statistical Software, 15(7): 1–22, January 2006. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i07.

Hankin:2007:CSU

[Han07a] Robin K. S. Hankin. Code snippet: Urn sampling without replacement: Enumerative combinatorics in R. Journal of Statistical Software, 17(CS-1):1– 6, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v17/c01.

Hankin:2007:IUR

[Han07b] Robin K. S. Hankin. Introducing untb, an R package for simulating ecological drift under the unified neutral theory of biodiversity. Journal of Statistical Software, 22(12):1– 15, September 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v22/i12.

Hand:2013:SBRk

[Han13a]  David J. Hand.  Short

book review: Comparing Groups: Randomization and Bootstrap Methods Using R by Andrew S. Zieffler, Jeffrey R. Harring, Jeffrey D. Long. International Statistical Review = Revue Internationale de Statistique, 81(2):326–328, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hand:2013:SBRi

[Han13b] David J. Hand. Short book review: Graphical Models with R by Søren Højsgaard, David Edwards, Steffen Lauritzen. International Statistical Review = Revue Internationale de Statistique, 81(2): 316, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Hankin:2020:IPR

[Han20] Robin K. S. Hankin. In-

                                        troducing the permutations

R package. SoftwareX, 11 (??):Article 100453, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711019302055.

Hardle:1991:STI

[Här91] Wolfgang Härdle. Smoothing Techniques: with Implementation in S. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97367-2 (New York), 3-540-97367-2 (Berlin). xi + 261 pp. LCCN QA278 .H348 1990.

Harrell:2001:RMS

[Har01] Frank E. Harrell, Jr. Regression Modeling Strategies:

With Applications to Linear Models, Logistic Regression, and Survival Analysis, volume ?? of Springer Series in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CO-DEN ???? ISBN 1-4419-2918-5, 1-4757-3462-X (ebook). ISSN 0172-7397. xxiii + 571 pp. LCCN QA276-280. URL http://

link.springer.com/book/
10.1007/978-1-4757-3462-
1.

Harbo:2003:BRA

[Har03] Ingrid Sofie Harbo. Book review: Applied statistics in the pharmaceutical industry: With case studies using S-Plus. Steven P. Millard and Andreas Krause (eds), Springer Verlag, Heidelberg, Germany, 2001. No. of pages: xviii + 513. Price: £62.50. ISBN 0-387-98814-9. Statistics in Medicine, 22(17):2819–2820, September 15, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardin:2006:IST

[Har06] Jo Hardin. Introduction to statistics through resampling methods and R/ S-PLUS. The American Statistician, 60(4):343–344, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Harvill:2007:BRB

[Har07] Jane L. Harvill. Book review: Analysis of Integrated and Co-Integrated Time Series With R. Journal of the American Statistical Association, 102(477):389–390, March 2007. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harrison:2008:SRR

[Har08] Teresa D. Harrison. Software review: Review of np software for R. Journal of Applied Econometrics, 23(6):861–865, September– October 2008. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Harte:2010:PRP

[Har10] David Harte. PtProcess: An R package for modelling marked point processes indexed by time. Journal of Statistical Software, 35 (8):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v35/i08.

Harel:2019:BRM

[Har19] Ofer Harel. Book review: Missing and Modified Data in Nonparametric Estimation: With R Examples. Journal of the American Statistical Association, 114 (527):1421–1423, 2019. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Harris:2023:BRM

[Har23] Mark N. Harris. Book review: Modern Applied Regressions: Bayesian and Frequentist Analysis of Categorical and Limited Response variables with R and Stan, Jun Xu, Florida, USA: CRC Press (Taylor & Francis Group), 2023, xv + 281 pp., ISBN 978-0-367-17387-6. Journal of the American Statistical Association, 118(543):2209–2211, 2023. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hassler:2018:BRW

[Has18] Uwe Hassler. Book review:

Wayne A. Woodward, Henry L. Gray and Alan C. Elliott (2017): Applied Time Series Analysis with R, Second Edition, Chapman & Hall/ CRC, 618 pp., $109.95, ISBN 978-1-4987-3422-6. Statistical Papers, 59(1):417– 418, March 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00362-017-0977-6.

Heiberger:1992:DFR

[HB92] Richard M. Heiberger and Richard A. Becker. Design of an S function for robust regression using iteratively reweighted least squares. Journal of Computational and Graphical Statistics, 1 (3):181–196, 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474580.

Haunschild:2019:RPP

[HBA19] Robin Haunschild, Lutz Bornmann, and Jonathan Adams. R package for producing beamplots as a preferred alternative to the h index when assessing single researchers (based on downloads from Web of Science). Scientometrics, 120(2):925– 927, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11192-019-03147-3.

Harding:2018:PPN

[HBBP18] Joel M. S. Harding, Douglas C. Braun, Nicholas J. Burnett, and Annika Putt. PITR: a new open source R package for PIT telemetry data. Fisheries, 43 (1):5, January 2018. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). URL https:/

/afspubs.onlinelibrary.

wiley.com/doi/10.1002/
fsh.10027.

Hieke:2014:MRB

[HBNS14] Stefanie Hieke, Harald Binder, Alexandra Nieters, and Martin Schumacher. minPtest: a resampling based gene region-level testing procedure for genetic case-control studies. Computational Statistics, 29(1–2):51–63, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-012-0391-4.

Huet:2004:STN

[HBPJ04] S. (Sylvie) Huet, Annie Bouvier, Marie-Anne Poursat, and Emmanuel Jolivet. Statistical tools for nonlinear regression: a practical guide with S-PLUS and R examples. Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 0-387-40081-8. xiv + 232 pp. LCCN QA278.2 .H84 2004.

Horton:2004:URT

[HBQ04] Nicholas J. Horton, Elizabeth R. Brown, and Linjuan Qian. Use of R as a toolbox for mathematical statistics exploration. The American Statistician, 58(4): 343–357, November 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27643595.

Hornik:2009:OSM

[HBZ09] Kurt Hornik, Christian Buchta, and Achim Zeileis. Open-source machine learning: R meets Weka. Computational Statistics, 24(2): 225–232, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-008-0119-7.

Hin:2005:ACD

[HC05] Lin Yee Hin and Vincent Carey. An algorithm for clustered data generalized additive modelling with S-PLUS. Journal of Statistical Software, 14(8):1– 24, August 31, 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=133&url=v14/i08&
                                        ct=2;   http://www.jstatsoft.
                                        org/counter.php?id=133&
                                        url=v14/i08/v14i08.pdf&
                                        ct=1.

Hiebert:2018:PCC

[HCM^{+}18] James Hiebert, Alex J. Cannon, Trevor Murdock, Stephen Sobie, and Arelia Werner. ClimDown: Climate downscaling in R. Journal of Open Source Software, 3(22):360:1, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00360.

Harrison:2015:RPP

[HCSH15] Lisa-Marie K. Harrison, Martin J. Cox, Georg Skaret, and Robert Harcourt. The R package EchoviewR for automated processing of active acoustic data using Echoview. Frontiers in Marine Science, 2, February 2015. ISSN 2296-7745.

Hahsler:2010:REM

[HD10] Michael Hahsler and Margaret H. Dunham. rEMM: Extensible Markov model for data stream clustering in R. Journal of Statistical Software, 35(5):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v35/
i05.

Hatzinger:2012:PRP

[HD12] Reinhold Hatzinger and Regina Dittrich. prefmod: An R package for modeling preferences based on paired comparisons, rankings, or ratings. Journal of Statistical Software, 48 (10):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v48/i10.

Han:2018:GRP

[HD18] Zifei Han and Victor De Oliveira. gcKrig: An R package for the analysis of geostatistical count data using Gaussian copulas. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v087i13;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v087i13/
v87i13.pdf.

Heckler:2007:BRB

[Hec07] Charles E. Heckler. Book review: Correspondence Analysis and Data Coding with Java and R by Fionn Murtagh. Technometrics, 49(1):100, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471279.

Hector:2015:NSR

[Hec15] Andy Hector. New statistics with R: an introduction for biologists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-872905-7, 0-19-872906-5 (paperback), 0-19-104526-8 (e-book), 1-322-57064-7 (ebook). xi + 199 pp. LCCN QH324.2 .H43 2015.

Hojsgaard:2012:GMR

[HEL12] Søren Højsgaard, David Edwards, and Steffen L. Lauritzen. Graphical models with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-2298-1 (paperback), 1-4614-2299-X (e-book). ix + 182 pp. LCCN QA279 .H64 2012.

Helmreich:2015:BRS

[Hel15] James E. Helmreich. Book review: Statistics: An Introduction Using R (2nd Edition), Michael J. Crawley, John Wiley & Sons, Chichester, 2015. ISBN 978-1-118-94109-6. 339 pp. USD 45.00

(P). http://www.imperial.ac.uk/bio/research/crawley/statistics/.

Journal of Statistical Software, 67(BR-5):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/index.
php/jss/article/view/v067b05;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v067b05/v67b05.pdf.

Helmreich:2016:BRL

[Hel16] James E. Helmreich. Book review: Learning Base R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069b04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069b04/
v69b04.pdf.

Helmreich:2018:BRR

[Hel18] James E. Helmreich. Book review: The R Primer (2nd ed.). Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v083b02;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v083b02/
                                        v83b02.pdf.

Helske:2022:EBG

[Hel22]  Jouni Helske.  Efficient

Bayesian generalized linear models with time-varying coefficients: the walker package in R. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S235271102200022X.

Hewson:2005:BRBa

[Hew05] Paul Hewson. Book review: Statistical Tools for Nonlinear Regression: a Practical Guide with S-PLUS and R Examples, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (1):254, January 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        3559726.

Hewson:2016:BRS

[Hew16] Paul Hewson. Book reviews: Statistical Rethinking: a Bayesian Course with Examples in R and Stan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1131–1132, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hohle:2007:RRP

[HF07] Michael Höhle and Ulrike Feldmann. RLadyBug — an R package for stochastic epidemic models. Computational Statistics & Data Analysis, 52(2):680–686, October 15, 2007. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947306004221.

Hori:2022:IRTa

[HFY22] Kazuki Hori, Hirotaka Fukuhara, and Tsuyoshi Yamada. Item response theory and its applications in educational measurement. Part I: Item response theory and its implementation in R. WIREs Computational Statistics, 14(2):e1531:1– e1531:??, March–April 2022. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hornik:2014:PRP

[HG14] Kurt Hornik and Bettina Grün. movMF: An R package for fitting mixtures of von Mises–Fisher distributions. Journal of Statistical Software, 58(10):??, July 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v58/
                                        i10.

Hui:2008:LRP

[HGG08] Wallace Hui, Yulia R. Gel, and Joseph L. Gastwirth. lawstat: An R package for law, public policy and biostatistics. Journal of Statistical Software, 28(3): ??, October 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v28/i03.

Heiberger:2004:SAD

[HH04] Richard M. Heiberger and Burt Holland. Statistical analysis and data display: an intermediate course with examples in S-plus, R, and SAS. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-40270-5. xxiv + 729 pp. LCCN QA276.4 .H445 2004.

Henningsen:2007:SPE

[HH07] Arne Henningsen and Jeff D.

                                        Hamann. systemfit: a

package for estimating systems of simultaneous equations in R. Journal of Statistical Software, 23(4):1– 40, December 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i04.

Halekoh:2014:KRA

[HH14] Ulrich Halekoh and Søren Højsgaard. A Kenward– Roger approximation and parametric bootstrap methods for tests in linear mixed models — the R package pbkrtest. Journal of Statistical Software, 59(9): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v59/i09.

Heiberger:2015:SAD

[HH15] Richard M. Heiberger and Burt Holland. Statistical Analysis and Data Display: an Intermediate Course with Examples in R. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2015. ISBN 1-4939-2121-5 (hardcover), 1-4939-2122-3 (e-book). xxxi + 898 pp. LCCN QA276.4 .H445 2015.

Holmes:2018:MSM

[HH18] Susan Holmes and Wolfgang Huber. Modern Statistics for Modern Biology. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-108-42702-2, 1-108-70529-

4. LCCN ???? URL https:

//www.cambridge.org/us/
academic/subjects/statistics-

                                        probability/statistics-
                                        life-sciences-medicine-
                                        and-health/modern-statistics-
                                        modern-biology.

Helske:2019:MHM

[HH19a] Satu Helske and Jouni Helske. Mixture hidden Markov models for sequence data: The seqHMM package in R. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v088i03;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v088i03/
                                        v88i03.pdf.

Holmes:2019:MSMe

[HH19b] Susan Holmes and Wolfgang Huber. Modern Statistics for Modern Biology. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-70529-4 paperback. xxiii + 382 pp. LCCN QH323.5 .H65 2019.

Hahsler:2008:GTO

[HHB08] Michael Hahsler, Kurt Hornik, and Christian Buchta. Getting things in order: An introduction to the R package seriation. Journal of Statistical Software, 25(3): 1–34, March 2008. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v25/i03.

Ho:2011:NPP

[HHJC11] Anson T. Y. Ho, Kim P.

Huynh, and David T. Jacho-Chávez. npRmpi: A package for parallel distributed kernel estimation in R. Journal of Applied Econometrics, 26(2):344–349, March 2011. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ho:2014:CPN

[HHJC14] Anson T. Y. Ho, Kim P.

Huynh, and David T. Jacho-Chávez. CRS: a package for nonparametric spline estimation in R. Journal of Applied Econometrics, 29(2):348–352, March 2014. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Hojsgaard:2005:RPG

[HHY05] Søren Højsgaard, Ulrich Halekoh, and Jun Yan. The R package geepack for generalized estimating equations. Journal of Statistical Software, 15(2):1– 11, December 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i02.

Hansen:1997:ICS

[HI97] David Hansen and Yannis Ioannidis, editors. 9th International Conference on Scientific and Statistical Database Management: proceedings: August 11–13, 1997, Olympia, Washington. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7952-2. LCCN QA76.9.D3 I6 1997.

Hickey:2016:RMG

[Hic16]  Peter F. Hickey.  Rep-

resentation and manipulation of genomic tuples in R. Journal of Open Source Software, 1(1):20:1, May 2016. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00020.

Hilbe:1996:WFC

[Hil96] Joseph Hilbe. Windows file conversion software. The American Statistician, 50 (3):268–??, ???? 1996. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hilbe:2006:BRB

[Hil06] Joseph Hilbe. Book review: A Handbook of Statistical Analyses Using R. Journal of Statistical Software, 16(BR-6):1–6, August 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/b06.

Hilbe:2010:BRB

[Hil10] Joseph Hilbe. Book review: Introduction to Scientific Programming and Scientific Simulation Using R. Journal of Statistical Software, 36 (BR-4):??, September 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v36/
b04.

Hilbe:2014:MCD

[Hil14] Joseph M. Hilbe. Modeling Count Data. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02833-7 (hardcover), 1-107-61125-3 (paperback). xv + 283 pp. LCCN QA278 .H56 2014.

Holmes:2017:DSY

[HJ17] Susan Holmes and Julie Josse. Discussion of ‘50 Years of Data Science. Journal of Computational and Graphical Statistics, 26(4): 768–769, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Don17].

Hanley:2008:SWI

[HJM08] James A. Hanley, Marilyse Julien, and Erica E. M. Moodie. Student’s z, t, and s: What if Gosset had R? The American Statistician, 62(1):64–69, February 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hyndman:2008:ATS

[HK08] Rob J. Hyndman and Yeasmin Khandakar. Automatic time series forecasting: The forecast package for R. Journal of Statistical Software, 27(3):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v27/
i03.

Horton:2011:URD

[HK11] Nicholas J. Horton and Ken Kleinman. Using R for data management, statistical analysis, and graphics. CRC Press, Boca Raton, FL, USA, 2011. ISBN 1-4398-2755-9 (paperback). xxii + 275 pp. LCCN QA76.73.R3 H67 2011.

Horton:2015:URR

[HK15] Nicholas J. Horton and Ken Kleinman. Using R and RStudio for Data Management, Statistical Analysis, and Graphics. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, second edition, 2015. ISBN 1-4822-3736-9, 1-4822-3737-7 (e-book). xxiv + 289 pp. LCCN QA276.45.R3 H67 2015.

Hartung:2008:SMA

[HKS08] Joachim Hartung, Guido Knapp, and Bimal K. Sinha. Statistical metaanalysis with applications. Wiley series in probability and statistics. Wiley, New York, NY, USA, 2008. ISBN 0-470-29089-7 (cloth), 0-470-38633-9 (ebook). xv + 247 pp. LCCN QA277 .H373 2008.

URL http://www.loc.gov/
catdir/enhancements/fy0811/
2008009435-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0814/2008009435-t.html;
http://www.wiley.com/
WileyCDA/WileyTitle/productCd-
0470386339.html.

Hornik:2004:RV

[HL04] Kurt Hornik and Friedrich Leisch. R Version 2.0.0. Computational Statistics, 19 (4):661–664, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/BF02753918.

Hornik:2005:RV

[HL05] Kurt Hornik and Friedrich Leisch. R Version 2.1.0. Computational Statistics, 20 (2):197–202, June 2005. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/BF02789699.

Hojsgaard:2007:IGG

[HL07] Søren Højsgaard and Steffen L. Lauritzen. Inference in graphical Gaussian models with edge and vertex symmetries with the gRc package for R. Journal of Statistical Software, 23(6):1– 26, December 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i06.

Hoffmann:2009:FMA

[HL09] Thomas J. Hoffmann and Nan M. Laird. fgui: a method for automatically creating graphical user interfaces for command-line R packages. Journal of Statistical Software, 30(2): ??, April 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v30/i02.

Hlavac:2016:EEB

[Hla16] Marek Hlavac. ExtremeBounds: Extreme bounds analysis in R. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v072i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v072i09/
v72i09.pdf.

Husson:2011:EMA

[HLP11] François Husson, Sébastien Lê, and Jérôme Pagès. Exploratory multivariate analysis by example using R. Chapman and Hall/CRC computer science and data analysis. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-3580-2 (hardcover). xii + 228 pp. LCCN QA278 .H87 2011.

Harner:2009:JJR

[HLT09] E. James Harner, Dajie Luo, and Jun Tan. JavaStat: a Java/R-based statistical computing environment. Computational Statistics, 24(2):295– 302, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-008-0131-y.

Hlynka:2009:BRBa

[Hly09] Myron Hlynka. Book review: A First Course in Statistical Programming With R by W. John Braun; Duncan J. Murdoch. Technometrics, 51(3):340, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40586629.

Hofert:2011:NAC

[HM11] Marius Hofert and Martin Maechler. Nested Archimedean copulas meet R: The nacopula package. Journal of Statistical Software, 39(9):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v39/
i09.

Hofert:2016:POS

[HM16] Marius Hofert and Martin Mächler. Parallel and other simulations in R made easy: An end-to-end study. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069i04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069i04/
v69i04.pdf.

Heck:2018:RRP

[HM18] Daniel W. Heck and Morten Moshagen. RRreg: An R package for correlation and regression analyses of randomized response data. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085i02/
v85i02.pdf.

Hornik:2013:TPG

[HMR^{+}13] Kurt Hornik, Patrick Mair, Johannes Rauch, Wilhelm Geiger, Christian Buchta, and Ingo Feinerer. The textcat package for ngram based text categorization in R. Journal of Statistical Software, 52(6): ??, February 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v52/i06.

Hofner:2014:MBB

[HMRS14] Benjamin Hofner, Andreas Mayr, Nikolay Robinzonov, and Matthias Schmid. Modelbased boosting in R: a hands-on tutorial using the R package mboost. Computational Statistics, 29(1–2): 3–35, February 2014. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/s00180-012-0382-
5.

Hofner:2016:GRP

[HMS16] Benjamin Hofner, Andreas Mayr, and Matthias Schmid.

gamboostLSS: An R pack-

age for model building and variable selection in the GAMLSS framework. Journal of Statistical Software, 74(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v074i01;
https://www.jstatsoft.
org/index.php/jss/article/
view/v074i01/v74i01.pdf.

Heiberger:2009:RTE

[HN09] Richard M. Heiberger and Erich Neuwirth. R through Excel: a spreadsheet interface for statistics, data analysis, and graphics. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4419-0051-9 (paperback), 1-4419-0052-7 (e-book). xxv + 342 pp. LCCN QA276.45.R3 H45 2009.

Hoef:2009:BRA

[Hoe09] Jay M. Ver Hoef. Book review: Applied Spatial Data Analysis with R. by Bivand, R. S., Pebesma, E. J., and Gomez-Rubio, V. Biometrics, 65(2):660–661, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoffmann:2011:CSP

[Hof11] Thomas J. Hoffmann. Code snippet: Passing in command line arguments and parallel cluster/multicore batching in R with batch. Journal of Statistical Software, 39(CS-1):??, March 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v39/
c01.

Hofmarcher:2015:BRI

[Hof15]  Paul Hofmarcher.  Book

review: An Introduction to R for Quantitative Economics, Vikram Dayal. Springer-Verlag, New Dehli, 2015. ISBN 978-81-322-2339-

9. 109 pp. EUR 51.99.

http://www.springer.com/book/9788132223399.

Journal of Statistical Software, 67(BR-4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/index.
php/jss/article/view/v067b04;

 http://www.jstatsoft.
org/index.php/jss/article/
view/v067b04/v67b04.pdf.

Hohle:2007:RPM

[Höh07] Michael Höhle. surveillance:

an R package for the monitoring of infectious diseases. Computational Statistics, 22(4):571–582, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-007-0074-8.

Hohensinn:2018:CSP

[Hoh18] Christine Hohensinn. Code snippet: pcIRT: An R package for polytomous and continuous Rasch models. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084c02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084c02/
v84c02.pdf.

Hojsgaard:2004:MPG

[Høj04] Søren Højsgaard. The mimR package for graphical modelling in R. Journal of Statistical Software, 11(6): 1–13, ???? 2004. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.

                                        php?id=102&url=v11/i06/
                                        v11i06.pdf&ct=1.

Hojsgaard:2012:GIN

[Høj12] Søren Højsgaard. Graphical independence networks with the gRain package for R. Journal of Statistical Software, 46(10): ??, February 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v46/i10.

Horgan:2009:PRI

[Hor09] Jane M. Horgan. Probability with R: an introduction with computer science applications. Wiley, New York, NY, USA, 2009. ISBN 0-470-28073-5 (cloth). xviii + 393 pp. LCCN QA76.9.M35 H863 2009. URL http://

                                        www.loc.gov/catdir/toc/
                                        ecip0819/2008022817.html;
                                         http://www.wiley.com/
                                        WileyCDA/WileyTitle/productCd-
                                        0470280735.html.

Horgan:2012:PR

[Hor12a] Jane M. Horgan. Programming in R. WIREs Computational Statistics, 4(1):75– 84, January/February 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hornik:2012:CRA

[Hor12b] Kurt Hornik. The Comprehensive R Archive Network. WIREs Computational Statistics, 4(4):394– 398, July/August 2012. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Hornikova:2012:BRBd

[Hor12c] Adriana Horníková. Book review: A Primer of Ecology With R by Martin Henry H. Stevens. Technometrics, 54 (4):439–440, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714927.

Hossain:2022:AUB

[Hos22a] Md. Moyazzem Hossain. Applied univariate, bivariate, and multivariate statistics: Understanding statistics for social and natural scientists, with applications in SPSS and R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185 (2):727–728, April 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Hossain:2022:BRL

[Hos22b] Md. Moyazzem Hossain.

Book review: Linear Regression Models: Applications in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S771, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

                                        academic.oup.com/jrsssa/
                                        article/185/Supplement_
                                        2/S771/7069536.

Hossain:2022:SRM

[Hos22c] Md. Moyazzem Hossain. Statistical regression modeling with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):743–744, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Houseman:2007:BRU

[Hou07] E. Andres Houseman. Book review: Using R for introductory statistics. John Verzani, Chapman and Hall/ CRC, Boca Raton, FL, 2005. No. of pages: xvi + 414. Price: $44.95. ISBN 1-58488-450-9. Statistics in Medicine, 26(11):2431–2432, May 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Howard:2011:BRB

[How11] James P. Howard. Book review: R Cookbook. Journal of Statistical Software, 40 (BR-3):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v40/b03.

Howard:2016:BRM

[How16a] James P. Howard II. Book review: Meta-Analysis with R. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v070b01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v070b01/
v70b01.pdf.

Howard:2016:BRP

[How16b] James P. Howard II. Book review: Political Analysis Using R, James E. Monogan III, Springer-Verlag, New York, 2015. ISBN 978-3-319-23446-5. 242 pp. USD 59.99 (S).

http://www.springer.com/9783319234458.

Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

www.jstatsoft.org/index.
php/jss/article/view/v074b01;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v074b01/v74b01.pdf.

Howard:2017:CMN

[How17] James Patrick Howard, II. Computational methods, for numerical analysis with R. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, 2017. ISBN 1-4987-2363-2 (hardcover), 1-4987-2364-0 (ebook), 1-315-12019-4 (ebook). xx + 257 pp. LCCN QA297 .H67 2017. URL

                                        http://www.crcnetbase.
                                        com/isbn/9781498723640.

Howard:2018:PAR

[How18] James P. Howard II. Phonetic algorithms in R. Journal of Open Source Software, 3(22):480:1–480:2, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00480.

Helmreich:2009:PRP

[HP09] James E. Helmreich and Robert M. Pruzek. PSAgraphics: An R package to support propensity score analysis. Journal of Statistical Software, 29(6):??, February 2009. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v29/i06.

Hoef:2014:SRP

[HPCS14] Jay Ver Hoef, Erin Peterson, David Clifford, and Rohan Shah. SSN: An R package for spatial statistical modeling on stream networks. Journal of Statistical Software, 56 (3):??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/i03.

Haggstrom:2015:CRP

[HPWdL15] Jenny Häggström, Emma Persson, Ingeborg Waernbaum, and Xavier de Luna. CovSel: An R package for covariate selection when estimating average causal effects. Journal of Statistical Software, 68(1):??, ???? 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068i01;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v068i01/v68i01.pdf.

Hsieh:2020:PRP

[HRC20] Nan-Hung Hsieh, Brad Reisfeld, and Weihsueh A. Chiu. pksensi: an R package to apply global sensitivity analysis in physiologically based kinetic modeling. SoftwareX, 12(??):Article 100609, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711020303228.

Hu:2018:ERP

[HS18] Yang Hu and Carl Scarrott. evmix: An R package for extreme value mixture modeling, threshold estimation and boundary corrected kernel density estimation. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084i05/
v84i05.pdf.

Huang:2012:DRP

[HSG12] B. Emma Huang, Rohan Shah, and Andrew W. George. dlmap: An R package for mixed model QTL and association analysis. Journal of Statistical Software, 50(6):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v50/
                                        i06.

Haneuse:2011:ORP

[HSL11] Sebastien Haneuse, Takumi Saegusa, and Thomas Lumley. osDesign: An R package for the analysis, evaluation, and design of two-phase and case-control studies. Journal of Statistical Software, 43 (11):??, August 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/i11.

Henningsen:2011:PMP

[HT11] Arne Henningsen and Ott Toomet. maxLik: A package for maximum likelihood estimation in R. Computational Statistics, 26(3):443–458, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-010-0217-1.

Hirschfeld:2019:CBS

[HT19] G. Hirschfeld and C. Thiele. Cloud-based simulation studies in R — a tutorial on using doRedis with Amazon spot fleets. Statistics in Medicine, 38(20):3947– 3959, September 10, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ho:2023:CFR

[HTWT23] George To Sum Ho, Yung Po Tsang, Qin Wu, and Valerie Tang. ck-FARM: an R package to discover big data associations for business intelligence. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000377.

Hu:2009:BRA

[Hu09] Feifang Hu. Book review:

Adaptive Design Theory and Implementation Using SAS and R, by Chang, M. Biometrics, 65(2):668–669, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2021:BRS

[Hu21] Chen Hu. Book review:

Statistical methods for survival trial design-With applications to cancer clinical trials using R by Jianrong Wu, CRC Press, 2018, ISBN 978-0-367-73432-9. Biometrics, 77(2):777–778, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hubert:2011:BRM

[Hub11]  Mia Hubert.  Book re-

view: Multivariate Nonparametric Methods with R. An Approach Based on Spatial Signs and Ranks, by Oja, H. Biometrics, 67(4): 1674–1675, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Husch:2018:BRT

[Hüs18] Marc Hüsch. Book review: Thomas Rahlf: Data Visualisation with R. Statistical Papers, 59(2):847– 848, June 2018. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00362-018-1003-3.

Hankin:2007:CSS

[HW07] Robin K. S. Hankin and Luke J. West. Code snippet: Set partitions in R. Journal of Statistical Software, 23(CS-2):1–12, December 2007. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v23/c02.

Hyun:2018:VRP

[HWY18] Seung Won Hyun, Weng Kee Wong, and Yarong Yang. VNM: An R package for finding multiple-objective optimal designs for the 4parameter logistic model.

Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://

www.jstatsoft.org/index.
php/jss/article/view/v083i05;
https://www.jstatsoft.
org/index.php/jss/article/
view/v083i05/v83i05.pdf.

He:2012:CSS

[HXY12] Wenqing He, Juan Xiong, and Grace Y. Yi. Code snippet: SIMEX R package for accelerated failure time models with covariate measurement error. Journal of Statistical Software, 46(CS-1):??, January 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/c01.

Iacus:2008:SIS

[Iac08] Stefano M. Iacus. Simulation and Inference for Stochastic Differential Equations: With R Examples, volume ?? of Springer Series in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN ???? ISBN 0-387-75839-9, 0-387-75838-

0. ISSN 0172-7397. xviii + 284 pp. LCCN QA274.23 .I23 2008. URL http://

link.springer.com/book/
10.1007/978-0-387-75839-
8.

Iacus:2015:BRA

[Iac15] Stefano M. Iacus. Book review: Automated Data Collection with R — A Practical Guide to Web Scraping and Text Mining, Simon Munzert, Christian Rubba, Peter Meißner, Dominic Nyhuis, John Wiley & Sons, Chichester, 2014. ISBN 978-1-118-83481-

7. 480 pp. GBP 50.00 (H).

http://www.r-datacollection.com/.

Journal of Statistical Software, 68(BR-3):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/index.
php/jss/article/view/v068b03;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v068b03/v68b03.pdf.

Irurozki:2016:PRP

[ICL16] Ekhine Irurozki, Borja Calvo, and Jose A. Lozano.

PerMallows: An R pack-

age for Mallows and generalized Mallows models. Journal of Statistical Software, 71(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v071i12;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v071i12/v71i12.pdf.

Ignatova:2015:KJF

[IDE15] Iliana Ignatova, Roland Deutsch, and Don Edwards. Kirk, J. L., and Fay, M. P. “An Introduction to Practical Sequential Inferences Via Single-Arm Binary Response Studies Using the Binseqtest R Package,” The American Statistician, 68, 230– 242. The American Statistician, 69(3):256–257, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/00031305.
2015.1053523. See [KF14].

IEEE:1993:NCR

[IEE93] IEEE, editor. Northcon/93:

conference record, Oregon Convention Center, Portland, Oregon, October 12– 14, 1993, NORTHCON 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-9972-2, 0-7803-9973-0. LCCN TK7868.D5N67 1993.

IEEE:1994:NCR

[IEE94] IEEE, editor. Northcon/ 94: conference record, Washington State Convention and Trade Center, Seattle, Washington: October 11–13, 1994, NORTHCON 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-9994-3, 0-7803-9995-1. LCCN TK 7801 N67 1994.

Ibarra-Espinosa:2018:PER

[IESdF18] Sergio Ibarra-Espinosa, Daniel Schuch, and Edmilson Dias de Freitas. eixport: An R package to export emissions to atmospheric models. Journal of Open Source Software, 3(24): 607:1–607:4, April 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00607.

Ihaka:1996:RLD

[IG96] Ross Ihaka and Robert Gentleman. R: a language for data analysis and graphics. Journal of Computational and Graphical Statistics, 5(3):299–314, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:

//www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474713. This is the

paper that introduced the open-source R system that grew out of S and S-Plus, independently developed without reference to the source code of those predecessors, but intended to be largely upward compatible with version of the S language. R version 1.0.0 was released on 29 February 2000.

Ihaka:1998:RPF

[Iha98] Ross Ihaka. R: Past and future history. Technical report, ????, 1998. URL

https://cran.r-project.
org/doc/html/interface98-
paper/paper.html. Draft

for 1998 Interface Symposium on Computer Science and Statistics.

Ismay:2020:SIDd

[IK20] Chester Ismay and Albert Young-Sun Kim. Statistical Inference via Data Science: a ModernDive, Into R and the Tidyverse. Chapman and Hall/CRC the r series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2020. ISBN 0-367-40982-8 (paperback), 0-367-40987-9 (hardcover), 0-367-40991-7 (e-book), 1-000-76346-3 (e-book), 1-000-76351-X (Mobipocket ebook), 1-000-76356-0 (EPUB e-book). xxx + 430 pp. LCCN QA276.45.R3 I86 2020.

Imai:2011:ERP

[ILS11] Kosuke Imai, Ying Lu, and Aaron Strauss. eco: R package for ecological inference in 2 × 2 tables. Journal of Statistical Software, 42(5):??, June 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v42/i05.

Iacus:2015:ILC

[IM15] Stefano M. Iacus and Lorenzo Mercuri. Implementation of Lévy CARMA model in Yuima package. Computational Statistics, 30(4):1111–1141, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-015-0569-7.

Iacus:2008:IMF

[IP08] Stefano Iacus and Giuseppe Porro. Invariant and metric free proximities for data matching: An R package. Journal of Statistical Software, 25(11):1– 22, April 2008. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v25/i11.

Iacobucci:2012:BRB

[IR12] Alessandra Iacobucci and Christian Robert. Book review: The Art of R Programming: A Tour of Statistical Software Design, by Norman Matloff. Chance, 25(2):55–56, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Innis:2021:GPC

[IRCA21] Sarah E. Innis, Kelsie Reinaltt, Mete Civelek, and Warren D. Anderson. GSEAplot: a package for customizing gene set enrichment analysis in R. Journal of Computational Biology, 28(6):629– 631, June 2021. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

                                        tronic). URL https://www.

liebertpub.com/doi/abs/
10.1089/cmb.2020.0426;
 https://www.liebertpub.
com/doi/reader/10.1089/
cmb.2020.0426.

Irizarry:2019:IDS

[Iri19] Rafael A. Irizarry. Introduction to Data Science: Data Analysis and Prediction Algorithms with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2019. ISBN 0-367-35798-4 (hardcover), 0-367-35799-2 (paperback), 0-429-34183-0 (ebook (PDF)). xxx + 713 pp. LCCN QA276.45.R3 I75 2019.

Imai:2005:MRP

[IV05] Kosuke Imai and David A.

Van Dyk. MNP: R package for fitting the Multinomial Probit Model. Journal of Statistical Software, 14(3): 1–32, May 8, 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=128&url=v14/i03&
ct=2;   http://www.jstatsoft.
org/counter.php?id=128&
url=v14/i03/v14i03.pdf&
ct=1.

Jackson:2011:MSM

[Jac11] Christopher Jackson. Multistate models for panel data: The msm package for R. Journal of Statistical Software, 38 (8):??, January 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v38/i08.

Jackson:2016:FPP

[Jac16] Christopher Jackson. flexsurv: A platform for parametric survival modeling in R. Journal of Statistical Software, 70(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v070i08;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v070i08/
                                        v70i08.pdf.

Jalilian:2019:CSE

[Jal19] Abdollah Jalilian. Code snippet: ETAS: An R package for fitting the space–time ETAS model to earthquake data. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v088c01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v088c01/
                                        v88c01.pdf.

James:1996:BRB

[Jam96] David A. James. Book review: Modern Applied Statistics with S-PLUS by W. N. Venables; B. D. Ripley. Technometrics, 38 (1):77–78, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1268908.

Jing:2015:PRP

[JD15] Liang Jing and Victor De Oliveira. geoCount: An R package for the analysis of geostatistical count data. Journal of Statistical Software, 63(11): ??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i11.

Jenness:2018:ERP

[JGM18] Samuel M. Jenness, Steven M.

Goodreau, and Martina Morris. EpiModel: An R package for mathematical modeling of infectious disease over networks. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084i08;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084i08/
v84i08.pdf.

Jara:2011:DBS

[JHQ^{+}11] Alejandro Jara, Timothy Hanson, Fernando A. Quintana, Peter Müller, and Gary L. Rosner. DPpackage: Bayesian semi- and nonparametric modeling in R. Journal of Statistical Software, 40(5):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v40/i05.

Jamshidian:2014:MRP

[JJJ14] Mortaza Jamshidian, Siavash Jalal, and Camden Jansen. MissMech: An R package for testing homoscedasticity, multivariate normality, and missing completely at random (MCAR). Journal of Statistical Software, 56 (6):??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/i06.

Jan:2014:PRP

[JKvT^{+}14] Budczies Jan, Daniel Kosztyla, Christian von Törne, Albrecht Stenzinger, Silvia Darb-Esfahani, Manfred Dietel, and Carsten Denkert.

                                        cancerclass: An R package

for development and validation of diagnostic tests from high-dimensional molecular data. Journal of Statistical Software, 59(1):??, August 2014. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v59/i01.

James:2015:PRP

[JM15] Nicholas A. James and David S. Matteson. ecp: An R package for nonparametric multiple change point analysis of multivariate data. Journal of Statistical Software, 62(7):

??, January 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v62/i07.

Jones:2014:ISP

[JMR14] Owen (Owen Dafydd) Jones, Robert Maillardet, and Andrew (Andrew P.) Robinson. Introduction to scientific programming and simulation using R. The R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2014. ISBN 1-4665-6999-9, 1-4665-7001-6. xxiv + 582 pp. LCCN Q183.9 .J65 2014.

Jockers:2014:TAR

[Joc14] Matthew Lee Jockers. Text Analysis with R for Students of Literature. Quantitative methods in the humanities and social sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-030-39643-6 (e-book), 3-319-03163-5 (hardcover), 3-319-03164-3 (e-book). xvi + 194 pp. LCCN QA76.9.T48 J63 2014.

Johnson:2007:MML

[Joh07] Matthew S. Johnson. Marginal maximum likelihood estimation of item response models in R. Journal of Statistical Software, 20(10):1– 24, February 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v20/i10.

Johnson:2009:BRT

[Joh09] Timothy D. Johnson. Book review: Time Series Analysis with Applications in R, 2nd edition by Cryer, J. D. and Chan, K.-S. Biometrics, 65(1):337, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2020:BRR

[Joh20] Paul Johnson. Book review: R Markdown: The Definitive Guide, Yihui Xie, J. J. Allaire, and Garrett Grolemund. Boca Raton, FL: Chapman & Hall/ CRC Press, 2018, xxxiv + 303 pp., $38.95(P), ISBN: 978-1-138-35933-8. The American Statistician, 74 (2):209–210, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/00031305.
2020.1745577.

Jones:2007:DPC

[Jon07] James Holland Jones. demogR: a package for the construction and analysis of age-structured demographic models in R. Journal of Statistical Software, 22(10):1– 28, September 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v22/i10.

Jones:2013:BRJ

[Jon13] Geoff Jones. Book review:

Joint Models for Longitudinal and Time-to-Event Data: with Applications in R. Australian & New Zealand Journal of Statistics, 55(2):198– 199, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jureckova:2006:RSM

[JP06] Jana Jurečková and Jan Picek. Robust statistical methods with R. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-454-1. xi + 193 pp. LCCN QA276 .J868 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0646/
2005053192-d.html.

Jin:2021:AMR

[JPH21] Man Jin, Adam Polis, and Jonathan Hartzel. Algorithms for minimization randomization and the implementation with an R package. Communications in Statistics: Simulation and Computation, 50(10):3077– 3087, 2021. CODEN CSS-CDB. ISSN 0361-0918.

Jaki:2019:RPM

[JPM19] Thomas Jaki, Philip Pallmann, and Dominic Magirr. The R package MAMS for designing multiarm multi-stage clinical trials. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v088i04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v088i04/
v88i04.pdf.

Jiang:2012:OSP

[JPOJ12] Lei Jiang, Pragneshkumar B. Patel, George Ostrouchov, and Ferdinand Jamitzky. OpenMP-style parallelism in data-centered multicore computing with R. ACM SIGPLAN Notices, 47(8): 335–336, August 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Johnstone:2005:ERP

[JS05] Iain M. Johnstone and Bernard W. Silverman. EbayesThresh: R programs for empirical Bayes thresholding. Journal of Statistical Software, 12(8):1– 38, ???? 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=125&url=v12/i08&
ct=2;   http://www.jstatsoft.
org/counter.php?id=125&
url=v12/i08/v12i08.pdf&
ct=1.

Joenssen:2014:PSG

[JV14] Dieter W. Joenssen and Jürgen Vogel. A power study of goodness-of-fit tests for multivariate normality implemented in R. Journal of Statistical Computation and Simulation, 84(5): 1055–1078, 2014. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

James:2013:ISL

[JWHT13] Gareth James, Daniela Witten, Trevor Hastie, and Robert Tibshirani. An introduction to statistical learning: with applications in R, volume 103 of Springer texts in statistics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7137-0, 1-4614-7138-9 (e-book). ISSN 1431-875X. xvi + 426 pp. LCCN QA276 .I585 2013.

Kabacoff:2011:RAD

[Kab11]  Robert Kabacoff.  R in

action: data analysis and graphics with R. Manning Publications, Greenwich, CT, USA, 2011. ISBN 1-935182-39-0 (paperback). xxiv + 447 pp. LCCN QA276.45.R3 K33 2011.

Kalyani:2024:BRC

[Kal24a] V. Kalyani. Book review:

Crime by the Numbers: a Criminologist’s Guide to R, by Jacob Kaplan, 1st Edition, 2023, Chapman and Hall/CRC, xxi + 432 pp, £59.99 (Hardback) and £59.99 (eBook), ISBN 978-1-003-27921-1. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):266, January 2024. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

academic.oup.com/jrsssa/
article/187/1/266/7235728.

Kalyani:2024:BRR

[Kal24b] V. Kalyani. Book review: Regression Analysis in R: a Comprehensive View for the Social Sciences, by Jocelyn H. Bolin, 1st ed., CRC Press, 2023. xii + 180 pp, ISBN 978-1-032-25775-4, £145.00 (Hardback), £51.29 (eBook). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):262– 263, January 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

academic.oup.com/jrsssa/
article/187/1/262/7219780.

Kane:2017:PBR

[Kan17] Michael Kane. bittrex: An R client for the Bittrex Crypto-Currency Exchange. Journal of Open Source Software, 2(17):402:1, September 2017. CO-DEN ???? ISSN 2475-9066. URL http://

joss.theoj.org/papers/
10.21105/joss.00402.

Kastner:2016:DSV

[Kas16] Gregor Kastner. Dealing with stochastic volatility in time series using the R package stochvol. Journal of Statistical Software, 69(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v069i05;
https://www.jstatsoft.
org/index.php/jss/article/
view/v069i05/v69i05.pdf.

Kauppinen:2013:SBR

[Kau13] Joonas Kauppinen. Short book review: The R Student Companion by Brian Dennis. International Statistical Review = Revue Internationale de Statistique, 81(2): 309, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kavroudakis:2015:SRP

[Kav15] Dimitris Kavroudakis. sms:

An R package for the construction of microdata for geographical analysis. Journal of Statistical Software, 68(2):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068i02;
http://www.jstatsoft.
org/index.php/jss/article/
view/v068i02/v68i02.pdf.

Kuparinen:2011:TPP

[KB11] Anna Kuparinen and Mats Björklund. Theory put into practice: an R implementation of the infinitedimensional model. Ecological Modelling, 222(12): 2027–2030, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0304380011001931.

Kolaczyk:2014:SAN

[KC14a] Eric D. Kolaczyk and Gábor Csárdi. Statistical analysis of network data with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4939-0982-7 (paperback). xiii + 207 pp. LCCN QA402 .K6483 2014.

Kroese:2014:SMC

[KC14b] Dirk P. Kroese and Joshua (Joshua C. C.) Chan. Statistical Modeling and Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4614-8774-9 (hardcover), 1-4614-8775-7 (e-book). xx + 400 pp. LCCN QA276 .K6986 2014.

Karian:2011:HFS

[KD11] Zaven A. Karian and Edward J. Dudewicz. Handbook of fitting statistical distributions with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-58488-711-7 (hardcover), 1-58488-712-5 (e-book: PDF). xlv + 1672 pp. LCCN QA273.6 .K374 2011. URL

http://marc.crcnetbase.
com/isbn/9781584887126.

Includes CD-ROM.

Killick:2014:PRP

[KE14] Rebecca Killick and Idris A.

Eckley. changepoint: An

R package for changepoint analysis. Journal of Statistical Software, 58 (3):??, June 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v58/i03.

Keen:2010:GSD

[Kee10] K. J. Keen. Graphics for statistics and data analysis with R. Chapman and Hall/CRC texts in statistical science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2010. ISBN 1-58488-087-2 (hardcover). xxxiv + 447 + 8 pp. LCCN QA276.45.R3 K44 2010.

URL http://www.loc.gov/
catdir/enhancements/fy1013/
2009053524-b.html.

Keeble:2012:RP

[Kee12]  Claire Keeble.  The R

primer. Journal of Applied Statistics, 39(10):2303– ??, October 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Keen:2018:TSS

[Kee18] K. J. Keen. Graphics for Statistics and Data Analysis with R. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2018. ISBN 1-4987-7983-2 (hardcover). xx + 590 pp. LCCN QA276.45.R3 K44 2018.

Kelly:2010:BRA

[Kel10] Patrick Kelly. Book reviews: Adaptive design theory and implementation using SAS and R. Mark Chang, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: xxii + 418. Price: $89.95. ISBN 10: 1-58488-962-4, ISBN 13: 978-1-58488-962-5. Statistics in Medicine, 29(9):1058, April 30, 2010. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karatzoglou:2010:KBM

[KF10] Alexandros Karatzoglou and Ingo Feinerer. Kernelbased machine learning for fast text mining in R. Computational Statistics & Data Analysis, 54(2):290– 297, February 1, 2010. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947309003533.

Kirk:2014:IPS

[KF14] Jennifer L. Kirk and Michael P.

Fay. An introduction to practical sequential inferences via single-arm binary response studies using the binseqtest R package. The American Statistician, 68(4):230–242, 2014. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Khalil:2017:RRP

[KF17] Salim Khalil and Mohamed Fakir. RCrawler: an R package for parallel web crawling and scraping. SoftwareX, 6(??):94–97, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300110.

Koenker:2017:RRP

[KG17] Roger Koenker and Jiaying Gu. REBayes: An R package for empirical Bayes mixture methods. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i08;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i08/
v82i08.pdf.

Kleinman:2010:SRD

[KH10] Ken Kleinman and Nicholas J. Horton. SAS and R: Data Management, Statistical Analysis, and Graphics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-7057-6 (hard back). xix + 323 pp. LCCN QA76.73.S27 K54 2010.

Khademi:2016:BRB

[Kha16] Abdolvahab Khademi. Book review: Basics of Matrix Algebra for Statistics with R, Nick Fieller, Chapman & Hall/CRC, Boca Raton, 2016. ISBN 978-1-4987-1236-1. xviii + 226 pp. USD 59.95 (H).

https://www.crcpress.com/9781498712361.

Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

www.jstatsoft.org/index.
php/jss/article/view/v071b01;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v071b01/v71b01.pdf.

Khademi:2017:BRH

[Kha17] Abdolvahab Khademi. Book review: Hidden Markov Models for Time Series: An Introduction Using R (2nd Edition). Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/

view/v080b01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v080b01/
v80b01.pdf.

Khademi:2018:BRCb

[Kha18] Abdolvahab Khademi. Book review: Computational Methods for Numerical Analysis with R. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v087b02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v087b02/
v87b02.pdf.

Krotka:2023:NRP

[KHJ^{+}23] Pavla Krotka, Katharina Hees, Peter Jacko, Dominic Magirr, Martin Posch, and Marta Bofill Roig. NCC: an R-package for analysis and simulation of platform trials with nonconcurrent controls. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023001334.

Kahm:2010:GFB

[KHLF^{+}10] Matthias Kahm, Guido Hasenbrink, Hella Lichtenberg-Fraté, Jost Ludwig, and Maik Kschischo. grofit: Fitting biological growth curves with R. Journal of Statistical Software, 33(7):

??, February 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v33/i07.

Kelley:2021:ARP

[KHR21] Dan E. Kelley, Jaimie Harbin, and Clark Richards. argoFloats: an R package for analyzing Argo data. Frontiers in Marine Science, 8, May 2021. ISSN 2296-7745.

Kiermeier:2008:VAA

[Kie08] Andreas Kiermeier. Visualizing and assessing acceptance sampling plans: The R pack-

                                        age AcceptanceSampling.

Journal of Statistical Software, 26(6):??, July 2008. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v26/
                                        i06.

Kimber:1995:BRB

[Kim95] Alan Kimber. Book review: A Handbook of Statistical Analyses Using S-Plus, by B. S. Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 44(3):408–409, ???? 1995. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        2348714.

Kim:2020:DTH

[Kim20a]  Jae H. Kim.  Decision-

theoretic hypothesis testing: A primer with R package OptSig. The American Statistician, 74 (4):370–379, 2020. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/00031305.
2020.1750484.

Kim:2020:BRD

[Kim20b] Sehee Kim. Book review:

Deep learning with R, Chollet, François, Allaire, Joseph J., Shelter Island, NY: Manning. Biometrics, 76(1):361– 362, March 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2020:BRP

[Kin20] Thomas King. Book review:

Practical R for Mass Communication and Journalism. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183(1):405, January 2020. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kafadar:1999:BRB

[KK99] Karen Kafadar and James R.

Koehler. Book review: Modern Applied Statistics with S-Plus, 2nd edition. The American Statistician, 53 (1):86–??, February 1999. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). Review of [VR97].

Komarek:2014:CRP

[KK14] Arnost Komárek and Lenka Komárková. Capabilities of R package mixAK for clustering based on multivariate continuous and discrete longitudinal data. Journal of Statistical Software, 59(12): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v59/i12.

Kawaguchi:2015:PSR

[KK15] Atsushi Kawaguchi and Gary G. Koch. sanon: An R package for stratified analysis with nonparametric covariable adjustment. Journal of Statistical Software, 67(9):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v067i09;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v067i09/v67i09.pdf.

Koyuncu:2021:DRM

[KK21] Nursel Koyuncu and Derya Karagöz. Designing robust modified R control charts for asymmetric distributions under ranked set and median ranked set sampling. Computational Statistics, 36 (2):1093–1121, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

                                        https://link.springer.

com/article/10.1007/s00180-
020-01051-6.

Kitada:2022:SUE

[KK22] Shuichi Kitada and Hirohisa Kishino. An R script for unbiased estimation of stocking effectiveness using a two-stage sampling strategy based on fishery landings. Fisheries Research, 250(??): Article 106266, June 2022. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0165783622000431.

Kotthaus:2015:DPS

[KKEM15] Helena Kotthaus, Ingo Korb, Michael Engel, and Peter Marwedel. Dynamic page sharing optimization for the R language. ACM SIG-PLAN Notices, 50(2):79–90, February 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kotthaus:2015:RMC

[KKL^{+}15] Helena Kotthaus, Ingo Korb, Michel Lang, Bernd Bischl, Jörg Rahnenführer, and Peter Marwedel. Runtime and memory consumption analyses for machine learning R programs. Journal of Statistical Computation and Simulation, 85(1):14–29, 2015. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Karlsson:2014:TRP

[KL14] Maria Karlsson and Anita Lindmark. truncSP: An R package for estimation of semi-parametric truncated linear regression models. Journal of Statistical Software, 57(14):??, May 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v57/
                                        i14.

Klemela:2009:SMD

[Kle09]  Jussi Klemelä.  Smooth-

ing of multivariate data: density estimation and visualization. Wiley series in probability and statistics. Wiley, New York, NY, USA, 2009. ISBN 0-470-29088-9. xxvi + 603 pp. LCCN QA278 .K584 2009.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy1012/
                                        2008032166-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy1012/2008032166-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0824/2008032166.
                                        html;   http://www.wiley.
                                        com/WileyCDA/WileyTitle/
                                        productCd-0470290889.html.

Kleiber:2017:BRJ

[Kle17] Christian Kleiber. Book review: John M. Chambers (2016): Extending R. Chapman and Hall/CRC Press, 364 pp., ISBN 978-1-4987-7571-7, GBP 44.99 (print), GBP 31.49 (eBook). Statistical Papers, 58(4):1279–1280, December 2017. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00362-017-0942-4.

Krause:2001:ASP

[KM01] Andreas Krause and Steven P.

Millard, editors. Applied statistics in the pharmaceutical industry: with case studies using S-Plus. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-98814-9. xviii + 513 pp. LCCN S122 .A665 2001.

Knoblauch:2008:MML

[KM08] Kenneth Knoblauch and Laurence T. Maloney. MLDS: Maximum likelihood difference scaling in R. Journal of Statistical Software, 25 (2):1–26, March 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i02.

Koenker:2014:COR

[KM14] Roger Koenker and Ivan Mizera. Convex optimization in R. Journal of Statistical Software, 60(5): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v60/i05.

Kalisch:2012:CIU

[KMC^{+}12] Markus Kalisch, Martin Mächler, Diego Colombo, Marloes H. Maathuis, and Peter Bühlmann. Causal inference using graphical models with the R package pcalg. Journal of Statistical Software, 47(11): ??, May 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v47/i11.

Karatzoglou:2006:SVM

[KMH06] Alexandros Karatzoglou, David Meyer, and Kurt Hornik. Support vector machines in R. Journal of Statistical Software, 15(9): 1–28, April 2006. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i09.

Kalibera:2014:FAS

[KMMV14] Tomas Kalibera, Petr Maj, Floreal Morandat, and Jan Vitek. A fast abstract syntax tree interpreter for R. ACM SIGPLAN Notices, 49(7): 89–102, July 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kowarik:2014:SAR

[KMTS14] Alexander Kowarik, Angelika Meraner, Matthias Templ, and Daniel Schopfhauser. Seasonal adjustment with the R packages x12 and x12GUI. Journal of Statistical Software, 62(2):??, December 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v62/i02.

Koenker:2003:SSM

[KN03] Roger Koenker and Pin Ng.

SparseM: a sparse matrix package for R. Journal of Statistical Software, 8 (6):1–9, ???? 2003. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v08/i06/
SparseM.pdf;   http://
www.jstatsoft.org/v08/
i6.

Karlis:2005:BPD

[KN05] Dimitris Karlis and Ioannis Ntzoufras. Bivariate Poisson and diagonal inflated bivariate Poisson regression models in R. Journal of Statistical Software, 14(10):1– 36, September 5, 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=135&url=v14/i10&
ct=2;   http://www.jstatsoft.
org/counter.php?id=135&
url=v14/i10/v14i10.pdf&
ct=1.

King:2016:SIP

[KNI16] Aaron A. King, Dao Nguyen, and Edward L. Ionides. Statistical inference for partially observed Markov processes via the R package pomp. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v069i12;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v069i12/
                                        v69i12.pdf.

Ko:1995:BRBb

[Ko95a] Daijin Ko. Book review: A Handbook of Statistical Analyses Using S-Plus by Brian S. Everitt. Technometrics, 37(2):232–233, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1269630.

Ko:1995:BRBa

[Ko95b] Daijin Ko. Book review: An Introduction to S and S-Plus by Phil Spector. Technometrics, 37(2):232, May 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1269629.

Krause:1997:BP

[KO97] Andreas Krause and Melvin Olson. The basics of S and S-Plus. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94985-2. xi + 242 pp. LCCN QA276.4 .K73 1997.

Krause:2000:BP

[KO00] Andreas Krause and Melvin Olson. The basics of S and S-Plus. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-98961-7. xxii + 382 pp. LCCN QA276.4 .K73 2000.

Kim:2006:CRP

[KO06] Donghoh Kim and Hee-Seok Oh. CVTresh: R package for level-dependent cross-validation thresholding. Journal of Statistical Software, 15(10):1– 13, April 2006. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i10.

Kim:2021:ERP

[KOC21] Donghoh Kim, Hee-Seok Oh, and Guebin Choi. EPT: an R package for ensemble patch transform. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000492.

Komarek:2009:NRP

[Kom09] Arnošt Komárek. A new R package for Bayesian estimation of multivariate normal mixtures allowing for selection of the number of components and interval-censored data. Computational Statistics & Data Analysis, 53(12): 3932–3947, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167947309001935.

Koning:2004:FAF

[Kon04]  Ruud H. Koning.  Fin-

Metrics: analysis of financial data in S-PLUS. Journal of Applied Econometrics, 19(2): 283–290, March–April 2004. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Kortschak:2018:PAG

[Kor18] R. Daniel Kortschak. arrgh: a Go interface to the OpenCPU R server system. Journal of Open Source Software, 3(21):517:1, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00517.

Korosteleva:2022:SPR

[Kor22] Olga Korosteleva. Stochastic Processes with R: an Introduction. Chapman and Hall/CRC texts in statistical science. Chapman and Hall/CRC, Boca Raton, FL, USA, 2022. ISBN 1-032-15373-3 (hardcover), 1-032-15473-X (paperback), 1-000-53733-1 (PDF), 1-000-53737-4 (EPUB), 1-003-24428-9 (ebook). ix + 190 pp. LCCN QA274 .K6745 2022.

Kostenko:2015:BRU

[Kos15] Andrey Kostenko. Book reviews: Using R for Numerical Analysis in Science and Engineering. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1098, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Kostenko:2016:BRG

[Kos16] Andrey Kostenko. Book review: Graphical Data Analysis with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(3):880, June 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Krug:2023:MRP

[KP23] Rainer M. Krug and Owen L.

Petchey. Microxanox: an R package for simulating an aquatic MICR obial ecosystem that can occupy OXic or ANOXic states. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002352.

Konietschke:2015:PRS

[KPSH15] Frank Konietschke, Marius Placzek, Frank Schaarschmidt, and Ludwig A. Hothorn. nparcomp: An R software package for nonparametric multiple comparisons and simultaneous confidence intervals. Journal of Statistical Software, 64 (9):??, March 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/i09.

Kohl:2010:RPD

[KR10] Matthias Kohl and Peter Ruckdeschel. R package distrMod: S4 classes and methods for probability models. Journal of Statistical Software, 35 (10):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/i10.

Kuhnt:2013:SOM

[KR13] Sonja Kuhnt and Nikolaus Rudak. Simultaneous optimization of multiple responses with the R package JOP. Journal of Statistical Software, 54(9): ??, September 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v54/i09.

Krause:1995:ESS

[Kra95] Andreas Krause. Electronic services in statistics. Statistical Software Newsletter in Computational Statistics and Data Analysis, 19(5): 595–604, 1995. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (elec-

tronic). URL http://www.
genedata.com/~andreas/
internet/internet.html.

Krause:1997:EPA

[Kra97] Andreas Krause. Einführung in S und S-Plus: mit Aufgaben und vollstandigen Lösungen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-60932-6 (paperback). xviii + 319 pp. LCCN ????

Krause:2005:BP

[Kra05] Andreas Krause. The basics of S-Plus. Statistics, computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2005. ISBN 0-387-26109-5. xxii + 442 pp. LCCN QA276.4 .K73 2005.

Krastev:2007:SSS

[Kra07] Theodor Krastev. spectrino software: Spectra visualization and preparation for R. Journal of Statistical Software, 18(10):1– 16, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v18/i10.

Kramer:2020:BRD

[Krä20]  Walter Krämer.  Book

review: Dieter Rasch, Rob Verdooren and Jürgen Pilz: Applied statistics: theory and problem solutions with R. Statistical Papers, 61(2):919– 920, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00362-020-01164-
6;   http://link.springer.
com/content/pdf/10.1007/
s00362-020-01164-6.pdf.

Kenkel:2014:EEF

[KS14] Brenton Kenkel and Curtis Signorino. Estimating extensive form games in R. Journal of Statistical Software, 56 (8):??, January 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v56/i08.

Kondo:2016:RRP

[KSBZ16] Yumi Kondo, Matias Salibian-Barrera, and Ruben Zamar. RSKC: An R package for a robust and sparse K-means clustering algorithm. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v072i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v072i05/
v72i05.pdf.

Kautz:2000:LLI

[KSC^{+}00] Frederick Kautz, Dimitrios Souflis, Robert Carbonari, DDJ, Al Gettier, Tina Femia, Andrew D. Todd, Anonymous, Alan Horn, Anonymous, Brad Clawsie, and Samuel Pearlman. Letters: The S language; IP security; Pnuts pointer; flyby; worker shortage?; design by interface. Dr. Dobb’s Journal of Software Tools, 25(4): 10, 12, April 2000. CODEN DDJOEB. ISSN 1044-789X. See [Ste00].

Karatzoglou:2004:KSP

[KSHZ04] Alexandros Karatzoglou, Alex Smola, Kurt Hornik, and Achim Zeileis. kernlab — an S4 package for kernel methods in R. Journal of Statistical Software, 11 (9):1–20, ???? 2004. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=105&url=v11/i09/
v11i09.pdf&ct=1.

Krl:2015:SRP

[KSP15] Agnieszka Król and Philippe Saint-Pierre. SemiMarkov: An R package for parametric estimation in multi-state semi-Markov models. Journal of Statistical Software, 66(6):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v066i06;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v066i06/v66i06.pdf.

Kirchner:2007:ARP

[KSS^{+}07] Marc Kirchner, Benjamin Saussen, Hanno Steen, Judith A. J. Steen, and Fred A. Hamprecht. amsrpm: Robust point matching for retention time alignment of LC/MS data with R. Journal of Statistical Software, 18 (4):1–12, January 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v18/i04.

Kowarik:2016:IRP

[KT16] Alexander Kowarik and Matthias Templ. Imputation with the R package VIM. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v074i07;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v074i07/
                                        v74i07.pdf.

Kubota:2023:DMS

[Kub23] Takafumi Kubota. Diseases maps of spatial epidemiological data by R. WIREs Computational Statistics, 15 (4):e1604:1–e1604:??, July– August 2023. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic). URL https:

                                        //wires.onlinelibrary.
                                        wiley.com/doi/epdf/10.
                                        1002/wics.1604.

Kuhn:2008:BPM

[Kuh08] Max Kuhn. Building predictive models in R using the caret package. Journal of Statistical Software, 28 (5):??, November 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v28/i05.

Kuhn:2010:BRB

[Kuh10]  Max Kuhn.  Book re-

view: Lattice: Multivariate Data Visualization With R by Deepayan Sarkar. Technometrics, 52(3):363– 364, August 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
27867260.

Kumar:2007:BRB

[Kum07] Kuldeep Kumar. Book review: Modelling Financial Time Series with S-PLUS, by E. Zivot; J. Wang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(4):1189– 1190, October 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
4623247.

Kume:2010:BRM

[Kum10] Alfred Kume. Book review: Morphometrics with R, by Claude, J. Biometrics, 66(2):662–663, June 2010.

CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kuonen:2003:NIP

[Kuo03] Diego Kuonen. Numerical integration in S-PLUS or R: a survey. Journal of Statistical Software, 8(13):1–14, 2003. CO-DEN JSSOBK. ISSN

                                        ????  URL http://www.
                                        jstatsoft.org/v08/i13;
                                         http://www.jstatsoft.
                                        org/v08/i13/article.pdf.

Kurt:2019:BSF

[Kur19] Will Kurt. Bayesian Statistics the Fun Way: Understanding Statistics and Probability with Star Wars, LEGO, and Rubber Ducks. No Starch Press, San Francisco, CA, USA, 2019. ISBN 1-59327-956-6. xxii + 234 pp. LCCN QA279.5.

Kushler:2003:BRB

[Kus03] Robert Kushler. Book review: Statistical Computing: An Introduction to Data Analysis Using S-PLUS by Michael J. Crawley. Technometrics, 45(4): 369, November 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        25047104.

Kutz:2013:DDM

[Kut13] Jose Nathan Kutz. Datadriven modeling and scientific computation: methods for complex systems and big data. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966033-6 (hardcover), 0-19-966034-4 (paperback). xvii + 638 pp. LCCN Q183.9 .K88 2013.

Kovalchik:2013:FAB

[KV13] Stephanie Kovalchik and Ravi Varadhan. Fitting additive binomial regression models with the R package blm. Journal of Statistical Software, 54(1): ??, September 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v54/i01.

Kaluzny:1998:SUM

[KVCS98] Stephen P. Kaluzny, Silvia C. Vega, Tamre P. Cardoso, and Alice A. Shelly. S+SpatialStats: User’s manual for Windows and UNIX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98226-4. xvi + 327 pp. LCCN QA278.2 .S18 1998.

Kieser:2017:ORP

[KWE^{+}17] Meinhard Kieser, Marius Wirths, Stefan Englert, Cornelia Ursula Kunz, and Geraldine Rauch. OneArmPhaseTwoStudy: An R package for planning, conducting, and analysing single-arm phase II studies. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v081i08;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v081i08/
                                        v81i08.pdf.

Kojadinovic:2010:MMD

[KY10] Ivan Kojadinovic and Jun Yan. Modeling multivariate distributions with continuous margins using the copula R package. Journal of Statistical Software, 34(9):??, May 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v34/i09.

Kleiber:2008:AER

[KZ08] Christian Kleiber and Achim Zeileis. Applied Econometrics with R. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-77316-9. x + 221 pp. LCCN QA276.45.R3 K58 2008.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0903/
                                        2008934356-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0903/2008934356-t.html.

Lanovaz:2019:CCT

[LA19] M. J. Lanovaz and B. Adams. Comparing the communication tone and responses of users and developers in two R mailing lists: Measuring positive and negative emails. IEEE Software, 36(5):46–50, September/October 2019. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Labarre:2012:RBP

[Lab12]  Anthony Labarre.  Re-

view of Combinatorial Pattern Matching Algorithms in Computational Biology using Perl and R, by Gabriel Valiente. ACM SIGACT News, 43(3):48–50, September 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Lagani:2017:FSR

[LAF^{+}17] Vincenzo Lagani, Giorgos Athineou, Alessio Farcomeni, Michail Tsagris, and Ioannis Tsamardinos. Feature selection with the R package MXM: Discovering statistically equivalent feature subsets. Journal of Statistical Software, 80 (??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v080i07;
https://www.jstatsoft.
org/index.php/jss/article/
view/v080i07/v80i07.pdf.

Lin:2018:FEA

[LAJJ18] Jie Lin, Donald A. Adjeroh, Bing-Hua Jiang, and

Yue Jiang. fastWKendall:

an efficient algorithm for weighted Kendall correlation. Computational Statistics, 33(4):1823–1845, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-017-0775-6.

Lalanne:2017:BRS

[Lal17] Christophe Lalanne. Book review: Statistical Analysis of Questionnaires: A Unified Approach Based on R and Stata. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082b01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082b01/
v82b01.pdf.

Lamigueiro:2012:SSR

[Lam12] Oscar Perpiñán Lamigueiro. solaR: Solar radiation and photovoltaic systems with R. Journal of Statistical Software, 50(9):??, August 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v50/
i09.

Langston:1995:CMR

[Lan95] R. Langston. A comparison of the MODULE routines with S-PLUS extensibility. In Anonymous [Ano95a], pages 1402–1405. ISBN 1-55544-211-0. LCCN ????

Lang:2009:MPA

[Lan09a] Duncan Temple Lang. A modest proposal: an approach to making the internal R system extensible. Computational Statistics, 24 (2):271–281, May 2009. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/s00180-008-0127-
7.

Lang:2009:WMD

[Lan09b] Duncan Temple Lang. Working with meta-data from C/C++ code in R: the

RGCCTranslationUnit pack-

age. Computational Statistics, 24(2):283–293, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

http://link.springer.com/
article/10.1007/s00180-
008-0128-6.

Lander:2014:REA

[Lan14a] Jared P. Lander. R for everyone: advanced analytics and graphics. Addison-Wesley, Reading, MA, USA, 2014. ISBN 0-321-88803-0 (paperback), 0-13-325715-0 (e-book), 0-13-325717-7 (ebook), 0-13-325718-5 (Web). 464 pp. LCCN QA76.73.R3.

URL http://proquest.

                                        safaribooksonline.com/
                                        book/programming/r/9780133257182.

Lang:2014:ERA

[Lan14b] Duncan Temple Lang. Enhancing R with advanced compilation tools and methods. Statistical Science, 29(2):181–200, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

                                        projecteuclid.org/euclid.
                                        ss/1408368570.

Lander:2017:REA

[Lan17a] Jared P. Lander. R for Everyone: Advanced Analytics and Graphics. Addison-Wes-ley, Reading, MA, USA, second edition, 2017. ISBN 0-13-454692-X (paperback), 0-13-454698-9, 0-13-454699-7 (e-book). xxiv + 525 pp. LCCN QA76.73.R3 L36 2017.

Lang:2017:BRE

[Lan17b] Michel Lang. Book review: Efficient R Programming. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v080b03;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v080b03/
                                        v80b03.pdf.

Landau:2018:PDR

[Lan18] William Michael Landau. The drake R package: a pipeline toolkit for reproducibility and highperformance computing. Journal of Open Source Software, 3(21):550:1, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00550.

Lantz:2019:MLR

[Lan19] Brett Lantz. Machine Learning with R. Packt Publishing, Birmingham, 2019. ISBN 1-78829-586-2 (paperback). xiii + 437 pp. LCCN Q325.5 .L367 2019.

Lasecki:1997:PFF

[Las97] Jennifer L. Lasecki. S-Plus functions and Fortran subroutines for multiple comparison critical points. Thesis (M.S.), University of South Carolina, Columbia, SC, USA, 1997. v + 68 pp.

Laurinec:2018:PTR

[Lau18] Peter Laurinec. TSrepr R package: Time series representations. Journal of Open Source Software, 3(23): 577:1–577:2, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00577.

Lawson:2002:BRB

[Law02] Andrew Lawson. Book review: Environmental Statistics with S-Plus, by S. P. Millard; N. K. Neerchal. Journal of the Royal Statistical Society. Series D (The Statistician), 51(3):425–426, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
3650306.

Lazar:2011:BRB

[Laz11a] Nicole Lazar. Book review: Introduction to Probability Simulation and Gibbs Sampling With R by Eric A. Suess; Bruce E. Trumbo. Technometrics, 53(3):327, August 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
23210415.

Lazic:2011:BRBb

[Laz11b] S. E. Lazic. Book review: The R Book, by M. J. Crawley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (3):843–844, July 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
23013530.

Lazic:2024:BRE

[Laz24a] Stanley E. Lazic. Book review: Exploring Modeling with Data and Differential Equations Using R, by John Zobitz, 2022, Chapman & Hall/CRC Press, Boca Raton, FL, 356 pages, £74.99 (Hardcover), ISBN: 978-1-032-25948-2. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):550– 551, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

academic.oup.com/jrsssa/
article/187/2/550/7420307.

Lazic:2024:BRS

[Laz24b] Stanley E. Lazic. Book review: Statistical Methods for Mediation, Confounding and Moderation Analysis Using R and SAS, by Qingzhao Yu and Bin Li, CRC Press, 2022. 294pp, ISBN: 978-0-367-36547-9, £125.00. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(1):265, January 2024. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

academic.oup.com/jrsssa/
article/187/1/265/7235727.

Li:2012:BRD

[LB12a] Yuelin Li and Jonathan Baron. Behavioral research data analysis with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-1237-4 (paperback), 1-4614-1238-2 (e-book). xii + 245 pp. LCCN QA276.45.R3 L5

Li:2012:URW

[LB12b] Yuelin Li and Ray Baser. Using R and WinBUGS to fit a generalized partial credit model for developing and evaluating patient-reported outcomes assessments. Statistics in Medicine, 31(18):2010–2026, August 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:2021:IWB

[LB21] Ran Li and Usama Bilal. Interactive web-based data visualization with R, plotly, and shiny (Carson Sievert). Biometrics, 77(2):776– 777, June 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liquet:2016:RGP

[LBC^{+}16] Benoît Liquet, Leonardo Bottolo, Gianluca Campanella, Sylvia Richardson, and Marc Chadeau-Hyam. R2GUESS: A graphics processing unit-based R package for Bayesian variable selection regression of multivariate responses. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v069i02;   https:/
                                        /www.jstatsoft.org/index.

php/jss/article/view/v069i02/
v69i02.pdf.

Luciano:2021:ORP

[LBL^{+}21] Giorgio Luciano, Serena Berretta, Kristian Hovde Liland, Gavin J. Donley, and Simon A. Rogers. oreo: an R package for large amplitude oscillatory analysis. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102100087X.

Lang:2017:PBT

[LBS17] Michel Lang, Bernd Bischl, and Dirk Surmann.

batchtools: Tools for R

to work on batch systems. Journal of Open Source Software, 2(10): 135:1–135:2, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00135.

Li:2018:RRP

[LBW18] James Li, Jacob Bien, and Martin T. Wells. rTensor: An R package for multidimensional array (tensor) unfolding, multiplication, and decomposition. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v087i10;   https:/

                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v087i10/
                                        v87i10.pdf.

Liu:2010:ICR

[LC10] Hai Liu and Kung-Sik Chan. Introducing COZIGAM: An R package for unconstrained and constrained zero-inflated generalized additive model analysis. Journal of Statistical Software, 35(11): ??, July 2010. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v35/i11.

Luo:2014:MRP

[LCSC14] Sheng Luo, Yong Chen, Xiao Su, and Haitao Chu. mmeta: An R package for multivariate meta-analysis. Journal of Statistical Software, 56 (11):??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/i11.

LeBeau:2018:PHC

[LeB18] Brandon LeBeau. highlightHTML: CSS formatting of R Markdown documents. Journal of Open Source Software, 3(21):185:1–185:2, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00185.

Lee:2013:CRP

[Lee13] Duncan Lee. CARBayes: An R package for Bayesian spatial modeling with conditional autoregressive priors. Journal of Statistical Software, 55(13):??, November 2013. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v55/i13.

Lee:2018:PRP

[Lee18] Eun-Kyung Lee. PPtreeViz:

An R package for visualizing projection pursuit classification trees. Journal of Statistical Software, 83 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v083i08;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v083i08/v83i08.pdf.

Lee:2021:BRB

[Lee21] Youjin Lee. Book review:

Beyond Multiple Linear Regression: Applied Generalized Linear Models and Multilevel Models in R, by Paul Roback and Julie Legler. Boca Raton, FL: Chapman & Hall/CRC Press, 2020 (H)/2021 (e-book), xvii + 418 pp., $99.95 (H/e-book), ISBN: 978-1-4398-8538-3 (H), 978-0-429-06666-5 (ebook). The American Statistician, 75(4):450–451, 2021. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/

                                        doi/full/10.1080/00031305.
                                        2021.1985862.

Leisch:2002:SPM

[Lei02] Friedrich Leisch. Sweave, Part I: Mixing R and L^{A}TEX. R News, 2(3):28–31, December 2002. URL http:

                                        //www.statistik.lmu.de/
                                        ~leisch/Sweave/Sweave-
                                        Rnews-2002-3.pdf.

Leisch:2003:SPI

[Lei03] Friedrich Leisch. Sweave, Part II: Package vignettes. R News, 3(2):21–24, October 2003. URL http:

                                        //www.statistik.lmu.de/
                                        ~leisch/Sweave/Sweave-
                                        Rnews-2003-2.pdf.

Leisch:2004:FGF

[Lei04] Friedrich Leisch. FlexMix: a general framework for finite mixture models and latent class regression in R. Journal of Statistical Software, 11 (8):1–18, ???? 2004. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=104&url=v11/i08/
                                        v11i08.pdf&ct=1.

Leibovici:2010:STM

[Lei10] Didier G. Leibovici. Spatiotemporal multiway data decomposition using principal tensor analysis on k-modes: The R package PTAk. Journal of Statistical Software, 34(10):??, May 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v34/i10.

Leifeld:2013:TCS

[Lei13] Philip Leifeld. texreg: Conversion of statistical model output in R to L^{A}TEX and HTML tables. Journal of Statistical Software, 55(8): ??, November 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v55/i08.

Lenth:2009:RSM

[Len09] Russell V. Lenth. Responsesurface methods in R, using rsm. Journal of Statistical Software, 32(7):??, October 2009. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/i07.

Lenth:2016:LSM

[Len16] Russell V. Lenth. Leastsquares means: The R package lsmeans. Journal of Statistical Software, 69(??): ??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v069i01;
https://www.jstatsoft.
org/index.php/jss/article/
view/v069i01/v69i01.pdf.

Leng:2020:BRS

[Len20]  Ling Leng.  Book re-

view: Statistical Computing With R, (2nd ed.), by Maria L. Rizzo. Boca Raton, FL: Chapman & Hall/ CRC Press, 2019, xiv + 474 pp., $99.95(H), ISBN: 978-1-4665-5332-3. Journal of the American Statistical Association, 115(531): 1557–1558, 2020. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Leong:2010:BRB

[Leo10] Chee Kian Leong. Book review: R in a Nutshell. Journal of Statistical Software, 36(BR-2):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v36/
b02.

Leppisaari:2014:BRB

[Lep14] Matias Leppisaari. Book review: An Introduction to Analysis of Financial Data with R by Ruey S. Tsay. International Statistical Review = Revue Internationale de Statistique, 82(1):142–143, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Lewis:2016:BRC

[Lew16] Nando Lewis. Book reviews: Circular Statistics in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1132, October 2016. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Lin:2015:MRP

[LF15] Yan-Xia Lin and Mark James

Fielding. MaskDensity14:

an R package for the density approximant of a univariate based on noise multiplied data. SoftwareX, 3–4 (??):32–36, December 2015. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711015000163.

Liboschik:2017:TRP

[LFF17] Tobias Liboschik, Konstantinos Fokianos, and Roland Fried. tscount: An R package for analysis of count time series following generalized linear models. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i05/
v82i05.pdf.

laGrange:2009:BIB

[lGlRGL09] Anthony la Grange, Niël le Roux, and Sugnet Gardner-Lubbe. BiplotGUI: Interactive biplots in R. Journal of Statistical Software, 30(12):??, June 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v30/i12.

Langfelder:2012:FRF

[LH12] Peter Langfelder and Steve Horvath. Fast R functions for robust correlations and hierarchical clustering. Journal of Statistical Software, 46 (11):??, March 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v46/i11.

Loy:2014:HSD

[LH14] Adam Loy and Heike Hofmann. HLMdiag: A suite of diagnostics for hierarchical linear models in R. Journal of Statistical Software, 56 (5):??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/i05.

Lhachimi:2014:BRB

[Lha14] Stefan K. Lhachimi. Book reviews: R Graphics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 177(2):567, February 2014. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Liverani:2015:PRP

[LHA^{+}15] Silvia Liverani, David I. Hastie, Lamiae Azizi, Michail Papathomas, and Sylvia Richardson. PReMiuM: An R package for profile regression mixture models using Dirichlet processes. Journal of Statistical Software, 64(7):??, March 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/i07.

Luukko:2016:IPL

[LHR16] P. J. J. Luukko, J. Helske, and E. Räsänen. Introducing libeemd: a program package for performing the ensemble empirical mode decomposition. Computational Statistics, 31(2):545– 557, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-015-0603-9.

Leiva:2008:RPG

[LHS08] Víctor Leiva, Hugo Hernández, and Antonio Sanhueza. An R package for a general class of inverse Gaussian distributions. Journal of Statistical Software, 26(4): 1–21, June 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v26/i04.

Li:2011:BRI

[Li11] W. K. Li. Book review: Introductory Time Series with R, by Cowpertwait, P. S. P. and Metcalfe, A. V. Biometrics, 67(1):328–329, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ligges:2009:BRC

[Lig09]  Uwe Ligges.  Book re-

view: Crawley, M. J. (2007) The R Book. Statistical Papers, 50(2):445–446, March 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

                                        tronic). URL http://link.
                                        springer.com/accesspage/
                                        article/10.1007/s00362-
                                        008-0118-3.

Lebret:2015:PRR

[LIL^{+}15] Rémi Lebret, Serge Iovleff, Florent Langrognet, Christophe Biernacki, Gilles Celeux, and Gérard Govaert. Rmixmod: The R package of the model-based unsupervised, supervised, and semi-supervised classification mixmod library. Journal of Statistical Software, 67(6):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v067i06;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v067i06/v67i06.pdf.

Lipovetsky:2020:BRI

[Lip20a] Stan Lipovetsky. Book review: Introduction to Data Science: Data Analysis and Prediction Algorithms With R, by Rafael A. Irizarry. Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2020, xxx + 713 pp., $99.95, ISBN:

978-0-367-35798-6. Technometrics, 62(2):1–282, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2020:BRS

[Lip20b] Stan Lipovetsky. Book review: Statistical Inference via Data Science: A Modern Dive Into R and the Tidyverse, by Chester Ismay and Albert Y. Kim. Boca Raton, FL: Chapman and Hall/CRC, Taylor & Francis Group, 2020, xxx + 430 pp., $79.95, ISBN: 978-0-367-40982-1. Technometrics, 62(2):1–283, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRA

[Lip21a] Stan Lipovetsky. Book review: Advanced Statistics with Applications in R, by Eugene Demidenko, Hoboken, NJ: John Wiley & Sons, 2020, xvii + 856 pp., $135.00 (Hardback), ISBN 978-1-118-38798-6. Technometrics, 63(2):273–275, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRLa

[Lip21b] Stan Lipovetsky. Book review: Learning Microeconometrics with R, by Christopher P. Adams, Chapman and Hall/CRC, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-0-367-25538-1, 398 pp., 72 b/w illustrations, $99.95. Technometrics, 63(3):424–425, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2021:BRR

[Lip21c] Stan Lipovetsky. Book review: R for Political Data Science: A Practical Guide, edited by Francisco Urdinez and Andrés Cruz, Chapman and Hall/CRC, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-0-367-81889-0, xix + 440 pp., 209 b/w illustrations, $99.95 (Hardback). Technometrics, 63(2):277–278, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRH

[Lip22] Stan Lipovetsky. Book review: Handbook of Regression Modeling in People Analytics: With Examples in R and Python, by Keith Mc-Nulty. CRC Press, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-1-032-04174-2, xvi + 255 pp., 48 color illustrations, $63.96 (hbk). Technometrics, 64(1):143–145, 2022. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2015:BRC

[Liu15] Shuangzhe Liu. Book review: Circular Statistics in R. International Statistical Review = Revue Internationale de Statistique, 83 (1):164, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2016:BRA

[Liu16a] Shuangzhe Liu. Book review:

Analysis of Categorical Data with R. International Statistical Review = Revue Internationale de Statistique, 84(1):162–163, April 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2016:BRF

[Liu16b] Shuangzhe Liu. Book review: Fieller, N.: Basics of Matrix Algebra for Statistics with R. Statistical Papers, 57(3):857–858, September 2016. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s00362-
016-0813-4;   http://link.
springer.com/article/10.
1007/s00362-016-0813-4.

Liu:2017:BRK

[Liu17] Xulei Liu. Book review:

Kleinman, Ken and Horton, Nicholas J., SAS and R: Data Management, Statistical Analysis, and Graphics. Boca Raton: CRC Press. Biometrics, 73(4):1470, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2018:BRM

[Liu18] Shuangzhe Liu. Book review: Modern Data Science with R, Benjamin S. Baumer, Daniel T. Kaplan and Nicholas J. Horton, Chapman and Hall/CRC, 2017, 556 pages, $99.95, hardcover ISBN: 978-1-4987-2448-7. International Statistical Review = Revue Internationale de Statistique, 86 (1):162, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2023:BRM

[Liu23] Shuangzhe Liu. Book review: Modern Applied Regressions: Bayesian and Frequentist Analysis of Categorical and Limited Response Variables with R and Stan, Jun Xu, Chapman & Hall/CRC, 2023, xv + 281 pages, £80.99/ $108, hardcover ISBN: 978-0-367-17387-6 (hbk); 978-1-032-37674-5 (pbk); 978-0-429-05646-8 (ebk). International Statistical Review = Revue Internationale de Statistique, 91(2):345– 347, August 2023. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Le:2008:FRP

[LJH08] Sébastien Lê, Julie Josse, and François Husson. FactoMineR: An R package for multivariate analysis. Journal of Statistical Software, 25 (1):1–18, March 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i01.

Leuangthong:2008:SPG

[LKD08] Oy Leuangthong, K. Daniel Khan, and Clayton V. Deutsch. Solved problems in geostatistics. Wiley, New York, NY, USA, 2008. ISBN 0-470-17792-6 (paperback). vii + 207 pp. LCCN QE33.2.S82 L48 2008. URL http://

www.loc.gov/catdir/toc/
ecip0811/2008007625.html;
 http://www.wiley.com/
WileyCDA/WileyTitle/productCd-
0470177926.html.

Lawrence:2010:RGU

[LL10] Michael Lawrence and Duncan Temple Lang. RGtk2: a graphical user interface toolkit for R. Journal of Statistical Software, 37(8): ??, December 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v37/i08.

Linzer:2011:PRP

[LL11] Drew A. Linzer and Jeffrey B. Lewis. poLCA: An R package for polytomous variable latent class analysis. Journal of Statistical Software, 42(10):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v42/
                                        i10.

Luciano:2020:TRP

[LLM^{+}20] G. Luciano, K. H. Liland, M. Marsotto, R. Svoboda, and S. Berretta. takos: an R package for thermal analysis calculations. SoftwareX, 12(??):Article 100637, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711020303502.

Linn:2015:PIQ

[LLS15] Kristin A. Linn, Eric B. Laber, and Leonard A. Stefanski. iqLearn: Interactive Q-learning in R. Journal of Statistical Software, 64 (1):??, March 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v64/i01.

Ligges:2003:SRP

[LM03] Uwe Ligges and Martin

                                        Mächler. Scatterplot3d —

an R package for visualizing multivariate data. Journal of Statistical Software, 8 (11):1–20, ???? 2003. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v08/i11;
 http://www.jstatsoft.
org/v08/i11/JSSs3d.pdf.

Lawrence:2014:SGR

[LM14a] Michael Lawrence and Martin Morgan. Scalable genomics with R and Bioconductor. Statistical Science, 29(2):214–226, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.
ss/1408368572.

Liao:2014:PRP

[LM14b] Xiyue Liao and Mary C.

Meyer. coneproj: An R package for the primal or dual cone projections with routines for constrained regression. Journal of Statistical Software, 61(12): ??, November 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i12.

Lee:2018:ERP

[LM18] Sharon X. Lee and Geoffrey J. McLachlan. EMMIXcskew: An R package for the fitting of a mixture of canonical fundamental skew tdistributions. Journal of Statistical Software, 83(??): ??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.

                                        jstatsoft.org/index.php/
                                        jss/article/view/v083i03;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v083i03/v83i03.pdf.

Li:2011:TRP

[LMY^{+}11] Jiangtian Li, Xiaosong Ma, Srikanth Yoginath, Guruprasad Kora, and Nagiza F. Samatova. Transparent runtime parallelization of the R scripting language. Journal of Parallel and Distributed Computing, 71(2): 157–168, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

LoPresti:1993:IOP

[Lo 93] C. A. Lo Presti. Implementation of the object-oriented paradigm in S-Plus. In IEEE [IEE93], pages 186– 191. ISBN 0-7803-9972-2, 0-7803-9973-0. LCCN TK7868.D5N67 1993.

LoPresti:1994:IOO

[Lo 94] C. A. Lo Presti. Implementation of the object oriented paradigm in S-Plus. In IEEE [IEE94], pages 170– 177. ISBN 0-7803-9994-3, 0-7803-9995-1. LCCN TK 7801 N67 1994.

Loftus:2021:BSR

[Lof21] Stephen C. Loftus. Basic statistics with R: reaching decisions with data. Academic Press, New York, NY, USA, 2021. ISBN 0-12-820788-4, 0-12-820926-7 (ebook). xix + 283 pp. LCCN QA276.4.

Longford:2015:SSI

[Lon15] Nicholas T. Longford. Statistical studies of income, poverty and inequality in Europe: computing and graphics in R using EU-SILC. Chapman and Hall/CRC statistics in the social and behavioral sciences series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4665-6832-1 (hardcover). xxii + 354 pp. LCCN HC79.I5 L646 2015. URL

http://images.tandf.co.
uk/common/jackets/websmall/
978146656/9781466568327.
jpg.

Lortie:2018:BRA

[Lor18] Christopher Lortie. Book review: Applied Time Series Analysis with R (2nd Edition). Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085b03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085b03/
v85b03.pdf.

Lafarge:2014:RIP

[LPLPD14] Thomas Lafarge, Beatriz Pateiro-López, Antonio Possolo, and Joy Dunkers. R implementation of a polyhedral approximation to a 3D set of points using the α-shape. Journal of Statistical Software, 56(4): ??, January 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v56/i04.

Lonare:2021:ERP

[LPR21] Gunratan Lonare, Bharat Patil, and Nilesh Raut. edgar: an R package for the U.S. SEC EDGAR retrieval and parsing of corporate filings. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021001369.

Li:2012:CSR

[LQC^{+}12] Mingyu Li, Hui Quan, Joshua Chen, Yoko Tanaka, Peter Ouyang, Xiaolong Luo, and Gang Li. Code snippet: R functions for sample size and probability calculations for assessing consistency of treatment effects in multi-regional clinical trials. Journal of Statistical Software, 47(CS-1):??, April 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v47/c01.

Loecher:2015:PPU

[LR15] Markus Loecher and Karl

Ropkins. RgoogleMaps and

loa: Unleashing R graphics power on map tiles. Journal of Statistical Software, 63 (4):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i04.

Laurent:2012:GRP

[LRGTA12] Thibault Laurent, Anne Ruiz-Gazen, and Christine Thomas-Agnan. GeoXp: An R package for exploratory spatial data analysis. Journal of Statistical Software, 47(2):??, April 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v47/i02.

Lee:2018:STA

[LRN18a] Duncan Lee, Alastair Rushworth, and Gary Napier. Spatio-temporal areal unit modeling in R with conditional autoregressive priors using the CARBayesST package. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084i09/
v84i09.pdf.

Long:2018:CSS

[LRN18b] Jed Long, Colin Robertson, and Trisalyn Nelson. Code snippet: stampr:

Spatial-temporal analysis of moving polygons in R. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084c01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084c01/
v84c01.pdf.

Lopez-Raton:2014:PRP

[LRRÁCSGS14] Mónica López-Ratón, María Xosé Rodríguez-Álvarez, Carmen Cadarso-Suárez, and Francisco Gude-Sampedro.

OptimalCutpoints: An R

package for selecting optimal cutpoints in diagnostic tests. Journal of Statistical Software, 61(8): ??, November 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i08.

Lemmon:2005:DSS

[LS05] David R. Lemmon and J. L. (Joseph L.) Schafer. Developing statistical software in Fortran 95. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-28123-1, 0-387-23817-4 (paperback). xv + 323 pp. LCCN QA76.5 .L453 2005.

URL http://www.loc.gov/
catdir/enhancements/fy0662/
2004061447-d.html;   http:

//www.loc.gov/catdir/enhancements/
fy0818/2004061447-t.html.

Liland:2016:FRP

[LS16] Kristian Hovde Liland and

Lars Snipen. fixedTimeEvents:

an R package for the distribution of distances between discrete events in fixed time. SoftwareX, 5 (??):216–226, ???? 2016. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711016300310.

Leps:2020:BRI

[LŠ20] Jan Lepš and Petr Šmilauer.

Biostatistics with R: an introductory guide for field biologists. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-48038-1 (hardcover), 1-108-72734-4 (paperback), 1-108-61604-6 (e-pub). ???? pp. LCCN QH323.5 .L46 2020.

Lasinio:2020:CRP

[LSM20] Giovanna Jona Lasinio, Mario Santoro, and Gianluca Mastrantonio. CircSpaceTime: an R package for spatial and spatio-temporal modelling of circular data. Journal of Statistical Computation and Simulation, 90(7): 1315–1345, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lendle:2017:LRP

[LSPvdL17] Samuel D. Lendle, Joshua Schwab, Maya L. Petersen, and Mark J. van der Laan. ltmle: An R package implementing targeted minimum loss-based estimation for longitudinal data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081i01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081i01/
v81i01.pdf.

LafayedeMicheaux:2016:PPR

[LT16] Pierre Lafaye de Micheaux and Viet Anh Tran. PoweR: A reproducible research tool to ease Monte Carlo power simulation studies for goodness-of-fit tests in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069i03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069i03/
v69i03.pdf.

Long:2019:RCP

[LT19] J. D. Long and Paul Teetor. R Cookbook: Proven Recipes for Data Analysis, Statistics, and Graphics. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2019. ISBN 1-4920-4068-1 (paperback), 1-4920-4067-3 (e-book). xvii + 579 pp. LCCN QA276.45.R3 L66 2019.

Lu:2018:BRH

[Lu18] Zudi Lu. Book review:

Hidden Markov Models for Time Series: An Introduction Using R, 2nd Edition, by Walter Zucchini, Iain L. Macdonald, and Roland Langrock. Monographs on Statistics and Applied Probability 150, Published by CRC Press, 2016. Total number of pages: 28 + 370. ISBN: 978-1-4822-5383-2 (Hardback). Journal of Time Series Analysis, 39 (1):105–106, January 2018. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Lubinsky:1991:CTF

[Lub91] David J. Lubinsky. Comment: Two functional programming environments for statistics — Lisp-Stat and S. Statistical Science, 6 (4):352–360, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.
ss/1177011575.

Ludlam:2013:BEN

[Lud13] John Ludlam. R by Example: Numerical Solutions to Differential Equations (Soetaert, K., et al; 2012) [book review]. Computing in Science and Engineering, 15 (2):6–7, March/April 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Luhmann:2013:REEa

[Luh13] Maike Luhmann. R für Einsteiger: Einführung in die Statistiksoftware für die Sozialwissenschaften. Beltz, Weinheim, Germany, third edition, 2013. ISBN 3-621-28090-1. 320 pp. LCCN

                                        ???? URL http://d-nb.
                                        info/1034153447/04.

Lumley:2001:BRB

[Lum01] Thomas Lumley. Book review: Environmental Statistics with S-PLUS by Steven P. Millard; Nagaraj K. Neerchal. Technometrics, 43 (4):495, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1270836.

Lumley:2008:BRA

[Lum08] Thomas Lumley. Book review: Analysis of correlated data with SAS and R (3rd edn). Mohamed M. Shoukri and Mohammad A. Chaudhary, Chapman and Hall/ CRC, Boca Raton, 2007. No. of pages: 295. Price: $89.95. ISBN: 1-58488-619-6. Statistics in Medicine, 27(22):4610–4611, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lund:2002:BRBa

[Lun02] Ulric J. Lund. Book review: Understanding Statistical Concepts Using S-Plus. The American Statistician, 56(3):249–??, August 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//oberon.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/00031305/
v56n3/s18/p249. Review of

[SA01].

Lund:2006:BRB

[Lun06] Ulric J. Lund. Book review: An R and S-Plus(R) Companion to Multivariate Analysis, by Brian S. Everitt. Journal of the American Statistical Association, 101(474):857–858, June 2006. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/www.jstor.org/stable/
27590760.

Lund:2007:BRBb

[Lun07a] Robert Lund. Book review: Time Series Analysis and Its Applications: With R Examples. Journal of the American Statistical Association, 102(479):1079, September 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lund:2007:BRBa

[Lun07b] Ulric J. Lund. Book review: A Handbook of Statistical Analyses Using R. The American Statistician, 61(3):277–278, August 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Luoma:2009:SBR

[Luo09] Arto Luoma. Short book review: Simulation and Inference for Stochastic Differential Equations: With R Examples by Stefano M. Iacus. International Statistical Review = Revue Internationale de Statistique, 77(1):163, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27919714.

Lupi:2009:URC

[Lup09] Claudio Lupi. Unit root CADF testing with R. Journal of Statistical Software, 32 (2):??, October 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/i02.

Lutkepohl:2011:BRB

[Lüt11] Helmut Lütkepohl. Book review: Bernhard Pfaff (2006): Analysis of Integrated and Cointegrated Time Series with R. Statistical Papers, 52(2):495– 496, May 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s00362-
009-0278-9.

Lawson:2016:CSM

[LW16] John Lawson and Cameron Willden. Code snippet: Mixture experiments in R using mixexp. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v072c02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v072c02/
v72c02.pdf.

Lawrence:2009:EGP

[LWC^{+}09] Michael Lawrence, Hadley Wickham, Dianne Cook, Heike Hofmann, and Deborah F. Swayne. Extending the GGobi pipeline from R. Computational Statistics, 24(2):195–205, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/s00180-008-0115-
y.

Luo:2012:DRP

[LX12] Jingqin Luo and Chengjie Xiong. DiagTest3Grp: An R package for analyzing diagnostic tests with three ordinal groups. Journal of Statistical Software, 51(3): ??, October 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v51/i03.

Lyubchich:2021:ECM

[Lyu21] Vyacheslav Lyubchich. Elements of copula modeling with R. Technometrics, 62(3):416, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lin:2010:HAH

[LZ10] Shili Lin and Hongyu Zhao, editors. Handbook on Analyzing Human Genetic Data: Computational Approaches and Software. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-69263-0 (hardcover), 3-540-69264-9 (e-book). xiv + 333 pp. LCCN QH431 .H2475 2010.

Lin:2017:PAB

[LZHC17] Lifeng Lin, Jing Zhang, James S. Hodges, and Haitao Chu. Performing arm-based network meta-analysis in R with the pcnetmeta package. Journal of Statistical Software, 80(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v080i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v080i05/
v80i05.pdf.

Marchi:2014:ABD

[MA14] Max Marchi and Jim Albert. Analyzing baseball data with R. R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-7022-9 (paperback). xvii + 333 pp. LCCN GV877 .M353 2014. URL

http://images.tandf.co.
uk/common/jackets/websmall/
978146657/9781466570221.
jpg.

Machado:1998:BRBb

[Mac98] Orlando Machado. Book review: The Basics of S and S-PLUS, by A. Krause; M. Olson. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):550–551, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
2988648.

Machler:2007:SIU

[Mäc07] Martin Mächler. Statistics: An introduction using R. The American Statistician, 61(1):100–101, February 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Mullen:2011:DRP

[MAG^{+}11] Katharine M. Mullen, David Ardia, David L. Gil, Donald Windover, and James Cline. DEoptim: An R package for global optimization by differential evolution. Journal of Statistical Software, 40(6):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v40/i06.

Maindonald:2006:BRB

[Mai06] John Maindonald. Book review: Generalized Additive Models: An Introduction with R. Journal of Statistical Software, 16(BR-3):1–2, July 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/b03.

Maindonald:2008:SBR

[Mai08] John H. Maindonald. Short book review: Interactive and Dynamic Graphics for Data Analysis — With R and GGobi by Dianne Cook; Deborah F. Swayne. International Statistical Review = Revue Internationale de Statistique, 76 (3):437–438, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27919657.

Maindonald:2009:SBRb

[Mai09] John H. Maindonald. Short book review: Morphometrics with R by Julien Claude; Michel Baylac; Emmanuel Paradis; Tristan Stayton. International Statistical Review = Revue Internationale de Statistique, 77 (2):302–303, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919731.

Maindonald:2010:SBRc

[Mai10] John H. Maindonald. Short book review: Introductory Time Series with R by Paul S. P. Cowpertwait, Andrew V. Metcalfe. International Statistical Review = Revue Internationale de Statistique, 78 (3):452–453, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919869.

Maindonald:2011:SBRd

[Mai11] John H. Maindonald. Short book review: Exploratory Multivariate Analysis by Example Using R by François Husson, Sébastien Lê, Jérôme Pagès. International Statistical Review = Revue Internationale de Statistique, 79(3):498, December 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        41305079.

Maindonald:2012:SBRb

[Mai12] John H. Maindonald. Short book review: R for SAS and SPSS Users, Second Edition, by Robert A. Muenchen. International Statistical Review = Revue Internationale de Statistique, 80(1):200– 201, April 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        23257195.

Maindonald:2014:BRBa

[Mai14a] John H. Maindonald. Book review: Behavioral Research Data Analysis with R by Yuelin Li, Jonathan Baron. International Statistical Review = Revue Internationale de Statistique, 82(1):147– 148, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2014:BRL

[Mai14b] John H. Maindonald. Book review: Linear Mixed-Effects Models Using R by Andrzej Galecki, Tomasz Burzykowski. International Statistical Review = Revue Internationale de Statistique, 82(2):315–317, August 2014. CODEN ISTRDP.

ISSN 0306-7734 (print), 1751-5823 (electronic).

Maindonald:2021:BRR

[Mai21] John H. Maindonald. Book review: Random Forests with R, Robin Genuer and Jean-Michel Poggi Springer, 2019, x + 99 pp, £49.99, paperback. ISBN: 978-3-030-56484-1. International Statistical Review = Revue Internationale de Statistique, 89(2):422–423, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Maloney:2009:BRB

[Mal09] Laurence T. Maloney. Book review: Introduction to Probability with R. Journal of Statistical Software, 30 (BR-1):??, April 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v30/b01.

Maller:2013:BRC

[Mal13] Ross Maller. Book review:

Competing Risks and Multistate Models with R. Australian & New Zealand Journal of Statistics, 55(1):57– 60, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mallik:2021:BRS

[Mal21] Abhirup Mallik. Book review: Statistical Rethinking: A Bayesian Course with Examples in R and Stan, edited by Richard McElreath, Boca Raton, FL: CRC Taylor and Francis, 2020, xvi + 589 pp., $94.94 (Hardback), ISBN: 978-0-367-13991-9. Technometrics, 63(3):440–441, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Manola:2003:BRB

[Man03] Judith Manola. Book review: A Handbook of Statistical Analyses using S-Plus (2nd ed.). The American Statistician, 57(2):146– ??, May 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//oberon.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/00031305/
v57n2/s19/p146. Review of

[Eve02].

Marchetti:2006:IIG

[Mar06] Giovanni M. Marchetti. Independencies induced from a graphical Markov model after marginalization and conditioning: The R package ggm. Journal of Statistical Software, 15(6):1– 15, February 2006. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i06.

MarquesdeSa:2007:ASU

[Mar07] Joaquim P. Marques de Sá. Applied Statistics Using SPSS, STATISTICA, MAT-LAB and R. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-71972-5. xxiv + 506 + 196 pp. LCCN QA276.4 S2232 2007. Marasinghe:2011:BRM

[Mar11] Mervyn G. Marasinghe.

Book review: A Modern Approach to Regression with R, by Sheather, S. J. Biometrics, 67(2):675–677, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martin:2012:BRI

[Mar12] Michael A. Martin. Book review: An Introduction to Bootstrap Methods with Applications to R, by M. R. Chernick and R. A. LaBudde. Australian & New Zealand Journal of Statistics, 54(2):253–255, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

MathSoft:1985:PUG

[Mat85] MathSoft, Inc. S-PLUS User’s Guide: Version 3.3 for Windows. StatSci, a Division of MathSoft, Inc., Seattle, WA, USA, September 1985. various pp.

MathSoft:1994:PAI

[Mat94] MathSoft, Inc. S-PLUS for ARC/INFO: your link to a powerful statistical view. Video recording (12 min.) plus text (48 p.)., 1994.

Mathsoft:1998:GSS

[Mat98]  Mathsoft, Inc.  Getting

started with S-Plus 4.5. Data Analysis Products Division, MathSoft, Inc., Seattle, WA, USA, April 1998. 48 pp.

Matloff:2011:ARP

[Mat11] Norman S. Matloff. The Art of R Programming: Tour of Statistical Software Design. No Starch Press, San Francisco, CA, USA, 2011. ISBN 1-59327-384-3. xxiii + 373 pp. LCCN QA276.4 .M2925

                                        2011. URL http://www.
                                        loc.gov/catdir/enhancements/
                                        fy1205/2011025598-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1205/2011025598-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1205/2011025598-t.html.

Matloff:2013:BRB

[Mat13] Norman S. Matloff. Book review: The R Student Companion. Journal of Statistical Software, 52(BR-4): ??, February 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v52/b04.

Matloff:2015:BRR

[Mat15] Norman S. Matloff. Book review: R for Cloud Computing, Ajay Ohri, Springer-Verlag, New York, 2014.

ISBN 978-1-4939-1701-3. 267 pp. USD 99.00 (P).

http://www.springer.com/book/9781493917013.

Journal of Statistical Software, 66(BR-4):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/index.
php/jss/article/view/v066b04;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v066b04/v66b04.pdf.

Matloff:2016:BRU

[Mat16a] Norm Matloff. Book review: A User’s Guide to Network Analysis in R, Douglas A. Luke, Springer-Verlag, New York, 2015. ISBN 978-3-319-23882-1. 238 pp. USD 59.99 (P).

http://www.springer.com/9783319238821.

Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

www.jstatsoft.org/index.
php/jss/article/view/v072b03;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v072b03/v72b03.pdf.

Matloff:2016:PCD

[Mat16b] Norman S. Matloff. Parallel Computing for Data Science: with Examples in R, C++ and CUDA, volume 28 of Chapman and Hall/CRC: The R Series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2016. ISBN 1-4665-8701-6 (hardcover). xxiii + 324 pp. LCCN QA76.642 M37 2016. URL http:

                                        //www.tandf.net/books/
                                        details/9781466587014.

Mousavi:2023:RPA

[MAY23] Niloufar Dousti Mousavi, Hani Aldirawi, and Jie Yang. An R package AZIAD for analysing zero-inflated and zero-altered data. Journal of Statistical Computation and Simulation, 93(16): 2801–2827, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Maindonald:2003:DAG

[MB03] J. H. (John Hilary) Maindonald and John Braun. Data analysis and graphics using R: an examplebased approach. Cambridge series in statistical and probabilistic mathematics. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81336-0. xxiii + 362 pp. LCCN QA276.4 .M245 2003. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/cam031/2002031560.
                                        html;   http://www.loc.
                                        gov/catdir/toc/cam031/
                                        2002031560.html.

Marcum:2015:CMS

[MB15] Christopher Steven Marcum and Carter T. Butts. Constructing and modifying sequence statistics for

relevent: Using informR

in R. Journal of Statistical Software, 64(5):??, March 2015. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/i05.

Marsh:2018:CSD

[MBGK18] Charles J. Marsh, Louise J.

Barwell, Yoni Gavish, and William E. Kunin. Code snippet: downscale: An R package for downscaling species occupancy from coarse-grain data to predict occupancy at fine-grain sizes. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086c03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086c03/
v86c03.pdf.

Marwick:2018:PDA

[MBM18] Ben Marwick, Carl Boettiger, and Lincoln Mullen. Packaging data analytical work reproducibly using R (and friends). The American Statistician, 72(1):80– 88, 2018. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

amstat.tandfonline.com/
doi/full/10.1080/00031305.
2017.1375986.

Monteiro:2011:BRP

[MBR11] João Vitor Dias Monteiro, Sudipto Banerjee, and Gurumurthy Ramachandran. B2Z: R package for Bayesian two-zone models. Journal of Statistical Software, 43 (2):??, July 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/i02.

Moy:2020:JTV

[MBT^{+}20] C. Moy, J. Belyakova, A. Turcotte, S. D. Bartolomeo, and C. Dunne. Just TYPEical: Visualizing common function type signatures in R. In 2020 IEEE Visualization Conference (VIS), pages 121–125. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020.

Martin:1997:PDI

[MC97] R. Douglas Martin and Vikram Chalana. The S-PLUS DataBlade for INFORMIX-Universal server: The natural wedding of an object relational database with an object-oriented data analysis engine. In Hansen and Ioannidis [HI97], pages 126–130. ISBN 0-8186-7952-2. LCCN QA76.9.D3 I6 1997. URL http:

                                        //dblp.uni-trier.de/db/
                                        conf/ssdbm/MartinC97.html.

Micheas:2018:PSP

[MC18] Athanasios C. Micheas and Jiaxun Chen. sppmix: Poisson point process modeling using normal mixture models. Computational Statistics, 33(4):1767–1798, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-018-0805-z.

Maus:2019:DTW

[MCAP19] Victor Maus, Gilberto Câmara, Marius Appel, and Edzer Pebesma. dtwSat: Timeweighted dynamic time warping for satellite image time series analysis in R. Journal of Statistical Software, 88(??):??, ???? 2019. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v088i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v088i05/
v88i05.pdf.

Molnar:2018:PIR

[MCB18] Christoph Molnar, Giuseppe Casalicchio, and Bernd Bischl. iml: An R package for interpretable machine learning. Journal of Open Source Software, 3 (26):786:1–786:2, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00786.

McElreath:2016:SRB

[McE16] Richard McElreath. Statistical rethinking: a Bayesian course with examples in R and Stan, volume 122 of Chapman and Hall/CRC texts in statistical science series. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, 2016. ISBN 1-4822-5344-5 (hardcover). xvii + 469 pp. LCCN QA279.5 .M3975 2016.

McGowan:2020:BRR

[McG20] Lucy D’Agostino McGowan. Book review: Reproducible Research in R and R Studio, 3rd edition. By Gandrud, Christopher. Biometrics, 76(4):1391, December 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLean:2017:PRI

[McL17] Mathew W. McLean. RefManageR: Import and manage Bib-TeX and BibLaTeX references in R. Journal of Open Source Software, 2(16): 338:1–338:2, August 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00338.

Melnykov:2012:MRP

[MCM12] Volodymyr Melnykov, Wei-Chen Chen, and Ranjan Maitra. MixSim: An R package for simulating data to study performance of clustering algorithms. Journal of Statistical Software, 51(12): ??, November 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i12.

McNeil:2004:BRB

[McN04] Alexander J. McNeil. Book review: Modeling Financial Time Series with S-Plus, by Eric Zivot; Jiahui Wang. Journal of the American Statistical Association, 99(466): 564–565, April 2004. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:/

/miranda.asa.catchword.
org/vl=1515603/cl=66/nw=
1/rpsv/cw/asa/01621459/
v99n466/s34/p564;   http:
//www.jstor.org/stable/
27590420.

McNamara:2014:BRB

[McN14] Amelia McNamara. Book review: Dynamic Documents with R and knitr. Journal of Statistical Software, 56(BR-2):??, January 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v56/b02.

McNicholl:2022:BRP

[McN22]  Colin McNicholl.  Book

review: Project-Based R Companion to Introductory Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S768, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/185/Supplement_
2/S768/7069528.

Mora-Cantallops:2020:CNA

[MCSAGB20] Marçal Mora-Cantallops, Salvador Sánchez-Alonso, and Elena García-Barriocanal. A complex network analysis of the Comprehensive R Archive Network (CRAN) package ecosystem. The Journal of Systems and Software, 170(??): ??, December 2020. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0164121220301709.

Manukyan:2014:DPM

[MÇSD14] Artür Manukyan, Erhan Çene, Ahmet Sedef, and Ibrahim Demir. Dandelion plot: a method for the visualization of R-mode exploratory factor analyses. Computational Statistics, 29(6):1769–1791, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-014-0518-x.

Mora-Cantallops:2020:EPC

[MCSGBSA20] Marçal Mora-Cantallops, Miguel-Ángel Sicilia, Elena García-Barriocanal, and Salvador Sánchez-Alonso. Evolution and prospects of the Comprehensive R Archive Network (CRAN) package ecosystem. Journal of Software: Evolution and Process, 32(11): e2270:1–e2270:??, November 2020. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Mair:2010:GFM

[MdL10] Patrick Mair and Jan de Leeuw. A general framework for multivariate analysis with optimal scaling: The R package aspect. Journal of Statistical Software, 32 (9):??, January 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/i09.

MeSilva:2022:PMR

[MdN22] W. V. M.e Silva and F. F.

do Nascimento. MCMC4Extremes: an R package for Bayesian inference for extremes and its extensions. Communications in Statistics: Simulation and Computation, 51 (2):432–442, 2022. CODEN CSSCDB. ISSN 0361-0918.

Moreira:2010:DRP

[MdUÁC10] Carla Moreira, Jacobo de Uña-Álvarez, and Rosa M. Crujeiras. DTDA: An R package to analyze randomly truncated data. Journal of Statistical Software, 37(7): ??, November 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v37/i07.

Melnykov:2016:CRP

[Mel16] Volodymyr Melnykov. ClickClust: An R package for modelbased clustering of categorical sequences. Journal of Statistical Software, 74 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v074i09;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v074i09/v74i09.pdf.

Meulders:2013:RPP

[Meu13] Michel Meulders. An R package for probabilistic latent feature analysis of two-way two-mode frequencies. Journal of Statistical Software, 54(14):??, September 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v54/
i14.

Magis:2014:CSP

[MF14] David Magis and Bruno Facon. Code snippet: deltaPlotR: An R package for differential item functioning analysis with Angoff’s delta plot. Journal of Statistical Software, 59(CS-1):??, August 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v59/c01.

Matthews:2015:CSP

[MF15] Gregory J. Matthews and Andrea S. Foulkes. Code snippet: MixMAP: An R package for mixed modeling of meta-analysis p values in genetic association studies. Journal of Statistical Software, 66(3): ??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v066c03;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v066c03/v66c03.pdf.

Murrell:2009:QCS

[MG09] Paul Murrell and Stephen Gardiner. Quality control for statistical graphics: The graphicsQC package for R. Journal of Statistical Software, 30(1):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v30/
i01.

Ma:2016:DFP

[MGHR16] Edward Ma, Vishrut Gupta, Meichun Hsu, and Indrajit Roy. dmapply: a functional primitive to express distributed machine learning algorithms in R. Proceedings of the VLDB Endowment, 9(13):1293–1304, September 2016. CODEN ???? ISSN 2150-8097.

Mair:2007:ERM

[MH07] Patrick Mair and Reinhold Hatzinger. Extended Rasch modeling: The eRm package for the application of IRT models in R. Journal of Statistical Software, 20(9): 1–20, February 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v20/i09.

Meyer:2009:GCS

[MH09] David Meyer and Kurt Hornik. Generalized and customizable sets in R. Journal of Statistical Software, 31 (2):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v31/i02.

Muenchen:2010:RSU

[MH10] Robert A. Muenchen and Joseph M. Hilbe. R for Stata users. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-1317-3 (hardcover), 1-4419-1318-1 (e-book). xxiv + 529 pp. LCCN QA276.45.R3 M85 2010.

Marcon:2015:PER

[MH15] Eric Marcon and Bruno Hérault. entropart: An R package to measure and partition diversity. Journal of Statistical Software, 67(8):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v067i08;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v067i08/v67i08.pdf.

McNamara:2018:WCD

[MH18] Amelia McNamara and Nicholas J. Horton. Wrangling categorical data in R. The American Statistician, 72(1):97–104, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

amstat.tandfonline.com/
doi/full/10.1080/00031305.
2017.1356375.

Mahani:2016:CSS

[MHJS16] Alireza S. Mahani, Asad Hasan, Marshall Jiang, and Mansour T. A. Sharabiani. Code snippet: Stochastic Newton sampler: The R package sns. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v074c02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v074c02/
v74c02.pdf.

Morandat:2012:EDR

[MHOV12] Floréal Morandat, Brandon Hill, Leo Osvald, and Jan Vitek. Evaluating the design of the R language. Lecture Notes in Computer Science, 7313: 104–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-31057-7_
                                        6/.

Migon:2010:BRD

[Mig10] Helio S. Migon. Book review: Dynamic Linear Models with R, by Petris, G., Petrone, S. and Campagnoli, P. Biometrics, 66(4): 1311–1312, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miller:1992:SNS

[Mil92] A. J. Miller. Some notes on statistical packages for PCs. Technical report, CSIRO Division of Mathematics & Statistics, Melbourne, Australia, October 1992. 55 pp.

Millard:1998:EPU

[Mil98] Steven P. Millard. EnvironmentalStats for S-PLUS: user’s manual for Windows and UNIX. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98486-0. xii + 381 pp. LCCN GE45.S73 M55 1998.

Millard:2000:EPU

[Mil00] Steven P. Millard. EnvironmentalStats for S-Plus: User’s Manual for Windows and UNIX, Versions 1.0 & 1.1. Probability, Statistics & Information, Seattle, WA, USA, March 2000. x + 381 pp. LCCN ????

Millard:2002:EPU

[Mil02] Steven P. Millard. EnvironmentalStats for S-PLUS: user’s manual for version 2.0. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 0-387-95398-1. xiii + 264 pp. LCCN GE45.S73 M52 2002.

Militino:2010:BRB

[Mil10] Ana F. Militino. Book review: Statistics and data with R. Journal of Applied Statistics, 37(2):357– 358, February 2010. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [CC08b].

Mildenberger:2012:BRG

[Mil12] Thoralf Mildenberger. Book review: Günther Sawitzki: Computational statistics. An introduction to R. Statistical Papers, 53(4):1067–1068, November 2012. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s00362-
011-0393-2.

Miller:2015:MDS

[Mil15a] Thomas W. Miller. Marketing data science: modeling techniques in predictive analytics with R and Python. Pearson Education, Old Tappan, NJ, USA, 2015. ISBN 0-13-388766-9, 0-13-388762-6. ???? pp. LCCN HF5415.

URL http://proquest.
safaribooksonline.com/?
fpi=9780133887662;   http:
//proquest.tech.safaribooksonline.
de/9780133887662.

Millo:2015:NRS

[Mil15b] Giovanni Millo. Narrow replication of ‘A Spatio– Temporal Model of House Prices in the USA’ using R. Journal of Applied Econometrics, 30(4): 703–704, June–July 2015. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic). See [BL14b].

Miller:2017:RPD

[Mil17a] Allan M. Miller. Review of Practical Data Science with R by Nina Zumel and John Mount. ACM SIGACT News, 48(1):18–22, March 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Miller:2017:RRD

[Mil17b] Allan M. Miller. Review of R for Data Science: Import, Tidy, Transform, Visualize, and Model Data by Hadley Wickham and Garrett Grolemund. ACM SIGACT News, 48(3):14–19, September 2017. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Mirman:2014:GCA

[Mir14] Daniel Mirman. Growth curve analysis and visualization using R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-8432-7 (hardcover). xviii + 170 pp. LCCN QH324.2 .M57 2014.

Merl:2010:ARP

[MJGM10] Daniel Merl, Leah R. Johnson, Robert B. Gramacy, and Marc Mangel. amei: An R package for the adaptive management of epidemiological interventions. Journal of Statistical Software, 36 (6):??, August 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v36/i06.

Marazzi:1993:ARF

[MJR93] Alfio Marazzi, Johann Joss, and Alex Randriamiharisoa. Algorithms, Routines, and S Functions for Robust Statistics: The FORTRAN Library ROBETH with an Interface to S-PLUS. Wadsworth and Brooks/Cole Publishing Co., Pacific Grove, CA, USA and Pacific Grove, CA, USA, 1993. ISBN 0-534-19698-5. xii + 436 pp. LCCN QA276.4 .M267 1993. US$69.50. URL http:/

/www.cbooks.com/sqlnut/
SP/search/gtsumt?source=
&isbn=0534196985.

Molas:2011:HGL

[ML11] Marek Molas and Emmanuel Lesaffre. Hierarchical generalized linear models: The R package HGLMMM. Journal of Statistical Software, 39 (13):??, March 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v39/i13.

McLachlan:2013:ERP

[ML13] Geoff McLachlan and Sharon X. Lee. EMMIXuskew: An R package for fitting mixtures of multivariate skew t distributions via the EM algorithm. Journal of Statistical Software, 55(12): ??, November 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v55/i12.

MeloAlbuquerque:2022:PRD

[MM22] Pedro Henrique Melo Albuquerque and Gustavo Monteiro. RMCriteria: a decision making support system package for R. Communications in Statistics: Simulation and Computation, 51 (6):3259–3271, 2022. CO-DEN CSSCDB. ISSN 0361-0918.

Mankad:2015:TVE

[MMB15] Shawn Mankad, George Michailidis, and Moulinath Banerjee. Threshold value estimation using adaptive two-stage plans in R. Journal of Statistical Software, 67(3):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v067i03;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v067i03/v67i03.pdf.

Milton:1969:SCP

[MN69] Roy C. Milton and John Ashworth Nelder, editors. Statistical Computation: proceedings of a conference held at the University of Wisconsin, Madison, April 28–30, 1969. Academic Press, New York, NY, USA, 1969. ISBN 0-12-498150-X. LCCN QA276.4 .C6 1969.

Millard:2001:ESP

[MN01] Steven P. Millard and Nagaraj K. Neerchal. Environmental statistics with S-Plus. Applied environmental statistics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-7168-6.

830 pp. LCCN GE45.S73 M55 2001.

Millard:2003:LEB

[MN03] Steven P. Millard and Nagaraj K. Neerchal. Letter to the Editor: Book Review of Environmental Statistics with S-Plus. J. H. Petersen, Statistics in Medicine 2002; 21:2607–2608. Statistics in Medicine, 22(16):2668–2670, August 30, 2003. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [Pet02].

Morina:2014:RPP

[MN14] David Moriña and Albert Navarro. The R package survsim for the simulation of simple and complex survival data. Journal of Statistical Software, 59 (2):??, August 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v59/i02.

Morina:2017:CRS

[MN17] D. Moriña and A. Navarro. Competing risks simulation with the survsim R package. Communications in Statistics: Simulation and Computation, 46(7):5712–5722, 2017. CODEN CSSCDB. ISSN 0361-0918.

Moreno:2003:BRB

[Mor03] Victor Moreno. Book review: Understanding Statistical Concepts Using S-Plus, by R. E. Schumacker; A. Akers. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):417, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
4128232.

Morris:2009:BRW

[Mor09] Jeffrey S. Morris. Book review: Wavelet Methods in Statistics with R, by Nason, G. P. Biometrics, 65(2):667–668, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mora:2018:BRB

[Mor18] Pedro M. Valero Mora. Book review: bookdown: Authoring Books and Technical Documents with R Markdown. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v087b01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v087b01/
v87b01.pdf.

Mineo:2006:URP

[MP06] Angelo Mineo and Alfredo Pontillo. Using R via PHP for teaching purposes: R-php. Journal of Statistical Software, 17(4):1– 20, October 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v17/i04.

Millo:2012:SSP

[MP12] Giovanni Millo and Gianfranco Piras. splm: Spatial panel data models in R. Journal of Statistical Software, 47(1):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v47/
                                        i01.

Mazza:2014:CSD

[MP14] Angelo Mazza and Antonio Punzo. Code snippet: DBKGrad: An R package for mortality rates graduation by discrete beta kernel techniques. Journal of Statistical Software, 57(CS-2):??, April 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v57/c02.

Mazza:2014:PRP

[MPM14] Angelo Mazza, Antonio Punzo, and Brian McGuire.

                                        KernSmoothIRT: An R pack-

age for kernel smoothing in item response theory. Journal of Statistical Software, 58(6):??, June 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v58/i06.

Montgomery:2012:ILR

[MPV12] Douglas C. Montgomery, Elizabeth A. Peck, and G. Geoffrey Vining. Introduction to linear regression analysis. Wiley series in probability and statistics. Wiley, New York, NY, USA, fifth edition, 2012. ISBN 0-470-54281-0. xvi + 645 pp. LCCN QA278.2 .M65 2012. URL http://

proquestcombo.safaribooksonline.
com/9780470542811.

Martin:2011:MMC

[MQP11] Andrew D. Martin, Kevin M.

Quinn, and Jong Hee Park. MCMCpack: Markov chain Monte Carlo in R. Journal of Statistical Software, 42(9):??, June 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v42/i09.

Meredith:2009:TRE

[MR09] Evan Meredith and Jeffrey S. Racine. Towards reproducible econometric research: the Sweave framework. Journal of Applied Econometrics, 24(2):366– 374, March 2009. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Magis:2012:RGR

[MR12] David Magis and Gilles Raîche. Random generation of response patterns under computerized adaptive testing with the R package catR. Journal of Statistical Software, 48(8):??, May 2012.

CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v48/
                                        i08.

Marin:2014:BER

[MR14] Jean-Michel Marin and Christian P. Robert. Bayesian essentials with R. Springer Texts in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2014. ISBN 1-4614-8686-6 (hardcover), 1-4614-8687-4 (e-book). ISSN 1431-875X. xiv + 296 pp. LCCN QA279.5 .M358 2014.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy1410/
                                        2013950378-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy1410/2013950378-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1410/2013950378-
                                        t.html.

MacDonald:2015:GRP

[MRC15] Blake MacDonald, Pritam Ranjan, and Hugh Chipman. GPfit: An R package for fitting a Gaussian process model to deterministic simulator outputs. Journal of Statistical Software, 64 (12):??, April 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v64/i12.

Munda:2012:PPF

[MRL12] Marco Munda, Federico Rotolo, and Catherine Legrand. parfm: Parametric frailty models in R. Journal of Statistical Software, 51(11): ??, November 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i11.

Mebane:2011:GOU

[MS11] Walter R. Mebane, Jr. and Jasjeet S. Sekhon. Genetic optimization using derivatives: The rgenoud package for R. Journal of Statistical Software, 42 (11):??, June 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v42/i11.

Mestre:2023:PLR

[MS23] F. Mestre and B. Silva.

lconnect R package: a versatile tool for evaluating landscape connectivity and prioritizing habitat patches in conservation research. Ecological Modelling, 484(??):??, October 2023. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0304380023002193.

Mollica:2020:PPR

[MT20] Cristina Mollica and Luca Tardella. PLMIX: an R package for modelling and clustering partially ranked data. Journal of Statistical Computation and Simulation, 90(5):925–959, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marcon:2015:CST

[MTPL15] Eric Marcon, Stéphane Traissac, Florence Puech, and Gabriel Lang. Code snippet: Tools to characterize point patterns: dbmss for R. Journal of Statistical Software, 67(CS-3): ??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v067c03;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v067c03/v67c03.pdf.

Molenaar:2015:FDI

[MTvdM15] Dylan Molenaar, Francis Tuerlinckx, and Han L. J. van der Maas. Fitting diffusion item response theory models for responses and response times using the R package diffIRT. Journal of Statistical Software, 66 (4):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v066i04;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v066i04/v66i04.pdf.

Muenchen:2009:RSS

[Mue09] Robert A. Muenchen. R for SAS and SPSS users. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-09417-2. xvii + 470 pp.

Muggeo:2010:ATE

[Mug10] Vito M. R. Muggeo. Analyzing temperature effects on mortality within the R environment: The constrained segmented distributed lag parameterization. Journal of Statistical Software, 32 (12):??, January 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/i12.

Mullner:2013:FFH

[Mül13] Daniel Müllner. fastcluster:

Fast hierarchical, agglomerative clustering routines for R and Python. Journal of Statistical Software, 53(9):??, May 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v53/i09.

Mullen:2014:CGO

[Mul14] Katharine M. Mullen. Continuous global optimization in R. Journal of Statistical Software, 60(6): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v60/i06.

Mulder:2015:CSR

[Mul15]  Kees Mulder.  Circular

statistics in R. Chance, 28(1):54, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Muller:2016:BRM

[Mül16] Marlene Müller. Book review: Multivariate Nonparametric Regression and Visualization With R and Applications to Finance. Journal of Statistical Software, 70(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v070b05;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v070b05/v70b05.pdf.

Mi:2016:SRP

[MUM16] Xuefei Mi, H. Friedrich Utz, and Albrecht E. Melchinger.

Selectiongain: an R pack-

age for optimizing multistage selection. Computational Statistics, 31(2):533– 543, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-015-0583-9.

Munzert:2014:ADC

[Mun14a]  Simon Munzert.  Auto-

mated Data Collection with R: a Practical Guide to Web Scraping and Text Mining. Wiley, New York, NY, USA, 2014. ISBN 1-118-83481-X (hardcover), 1-118-83478-X. xxii + 452 pp. LCCN QA76.9.D343 M865 2014. URL http:

//catalogimages.wiley.
com/images/db/jimages/
9781118834817.jpg.

Munzert:2014:BRB

[Mun14b] Simon Munzert. Book review: XML and Web Technologies for Data Sciences with R. Journal of Statistical Software, 61(BR-1): ??, October 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/b01.

Murdoch:2003:ORP

[Mur03] Duncan Murdoch. Orientlib:

An R package for orientation data. Journal of Statistical Software, 8(19): 1–11, ???? 2003. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v08/i19;
 http://www.jstatsoft.
org/v08/i19/orientlib_
paper.pdf.

Murtagh:2005:CAD

[Mur05] Fionn Murtagh. Correspondence analysis and data coding with Java and R. Computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-528-9. 230 pp. LCCN QA76.73.J38 M877 2005.

URL http://www.loc.gov/
catdir/enhancements/fy0648/
2005041353-d.html.

Murrell:2006:RG

[Mur06] Paul Murrell. R Graphics. Computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-4200-3502-9 (electronic), 0-8493-1622-7, 1-58488-486-X. xix + 301 pp. LCCN T385 .M9

                                        2006. URL http://www.
                                        netLibrary.com/urlapi.
                                        asp?action=summary&v=1&
                                        bookid=155642.

Murrell:2009:IVG

[Mur09a] Paul Murrell. Importing vector graphics: The grImport package for R. Journal of Statistical Software, 30 (4):??, April 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v30/i04.

Murrell:2009:RG

[Mur09b] Paul Murrell. R graphics. WIREs Computational Statistics, 2(5):216– 220, September/October 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Murrell:2011:RG

[Mur11] Paul Murrell. R graphics. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2011. ISBN 1-4398-3176-9 (hardcover). xxvii + 518 pp. LCCN T385 .M9 2011.

Murrell:2014:BRR

[Mur14] Paul Murrell. Book review:

The R Student Companion. Australian & New Zealand Journal of Statistics, 56(3): 304, September 2014. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Murrell:2018:RGT

[Mur18] Paul Murrell. R Graphics. Chapman and Hall/CRC the R Ser. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2018. ISBN 1-4987-8905-6 (hardcover), 0-429-42276-8 (e-book), 0-429-74993-7 (Mobipocket ebook), 0-429-74994-5, 0-429-74995-3 (PDF ebook). xvii + 423 pp. LCCN T385 .M9 2019.

Montero:2014:PRP

[MV14] Pablo Montero and José A.

Vilar. TSclust: An R package for time series clustering. Journal of Statistical Software, 62(1):??, December 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v62/i01.

Mullen:2007:TRP

[MvS07] Katharine M. Mullen and Ivo H. M. van Stokkum. TIMP: An R package for modeling multi-way spectroscopic measurements. Journal of Statistical Software, 18(3): 1–46, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v18/i03.

Miecznikowski:2013:DRP

[MVS13] Jeffrey Miecznikowski, Albert Vexler, and Lori Shep-

herd. dbEmpLikeGOF: An

R package for nonparametric likelihood ratio tests for goodness-of-fit and twosample comparisons based on sample entropy. Journal of Statistical Software, 54 (3):??, September 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v54/i03.

Mullen:2007:FRP

[MvSB^{+}07] Katharine M. Mullen, Ivo H. M. van Stokkum, Thomas W. Binsl, Jaap Heringa, and Johannes H. G. M. van Beek. FluxSimulator: An R package to simulate isotopomer distributions in metabolic networks. Journal of Statistical Software, 18(7):1– 17, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v18/i07.

Mevik:2007:PPP

[MW07] Björn-Helge Mevik and Ron Wehrens. The pls package: Principal component and partial least squares regression in R. Journal of Statistical Software, 18(2): 1–24, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v18/i02.

Mwitondi:2011:BCR

[Mwi11] Kassim S. Mwitondi. Bayesian computation with R. Journal of Applied Statistics, 38(10):2367–2368, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [Alb09].

Mwitondi:2013:DMR

[Mwi13a] Kassim S. Mwitondi. Data mining with Rattle and R. Journal of Applied Statistics, 40(2):464, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2013:SCC

[Mwi13b] Kassim S. Mwitondi. Statistical computing in C++ and R. Journal of Applied Statistics, 40(4):916, 2013. ISSN 0266-4763 (print), 1360-0532 (electronic).

Myers:2009:BRBc

[Mye09] Donald E. Myers. Book review: Applied Econometrics With R by Christian Klieber; Achim Zeileis. Technometrics, 51(4):484– 485, November 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40586665.

Myers:2012:BRBc

[Mye12a] Donald E. Myers. Book review: Forest Analytics With R by Andrew P. Robinson; Jeff D. Hamann. Technometrics, 54(3):320– 322, August 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714901.

Myers:2012:BRBb

[Mye12b] Donald E. Myers. Book review: Numerical Ecology with R by Daniel Bocard; François Gillet; Pierre Legendre. Technometrics, 54(2):200–201, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714872.

McLeod:2007:ALT

[MYK07] A. Ian McLeod, Hao Yu, and Zinovi L. Krougly. Algorithms for linear time series analysis: With R package. Journal of Statistical Software, 23(5):1– 26, December 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i05.

McLeod:2008:ISA

[MZ08] A. Ian McLeod and Ying Zhang. Improved subset autoregression: With R package. Journal of Statistical Software, 28(2): ??, October 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v28/i02.

Mount:2018:PVR

[MZ18] John Mount and Nina Zumel. The vtreat R package: a statistically sound data processor for predictive modeling. Journal of Open Source Software, 3(23): 584:1–584:2, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00584.

Nyaga:2017:CSC

[NAA17] Victoria N. Nyaga, Marc Arbyn, and Marc Aerts. Code snippet: CopulaDTA: An R package for copula-based bivariate beta-binomial models for diagnostic test accuracy studies in a Bayesian framework. Journal of Statistical Software, 82(??): ??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v082c01;
https://www.jstatsoft.
org/index.php/jss/article/
view/v082c01/v82c01.pdf.

Nagler:2018:KRP

[Nag18] Thomas Nagler. kdecopula:

An R package for the kernel estimation of bivariate copula densities. Journal of Statistical Software, 84 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v084i07;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v084i07/v84i07.pdf.

Narasimhan:2005:LSJ

[Nar05] Balasubramanian Narasimhan. Lisp–Stat to Java to R. Journal of Statistical Software, 13 (4):1–10, ???? 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=118&url=v13/i04&
ct=2;   http://www.jstatsoft.
org/counter.php?id=118&
url=v13/i04/v13i04.pdf&
ct=1.

Nason:2008:WMS

[Nas08] G. P. Nason. Wavelet Methods in Statistics with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-75960-3 (paperback), 0-387-75961-1 (e-book). x + 257 pp. LCCN QA403.3 .N37 2008.

Nash:2014:BPO

[Nas14] John C. Nash. On best practice optimization methods in R. Journal of Statistical Software, 60(2): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v60/i02.

Nadarajah:2013:NRP

[NB13] S. Nadarajah and S. A. A.

Bakar. A new R package for actuarial survival models. Computational Statistics, 28(5):2139–2160, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-013-0400-2.

Nia:2012:HDB

[ND12] Vahid Partovi Nia and Anthony C. Davison. Highdimensional Bayesian clustering with variable selection: The R package bclust. Journal of Statistical Software, 47(5):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v47/
i05.

Noorian:2016:GGE

[NdSL16] Farzad Noorian, Anthony M.

de Silva, and Philip H. W. Leong. gramEvol: Grammatical evolution in R. Journal of Statistical Software, 71(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v071i01;
https://www.jstatsoft.
org/index.php/jss/article/
view/v071i01/v71i01.pdf.

Neubauer:2011:BRB

[Neu11] Dean V. Neubauer. Book review: Introducing Monte Carlo Methods With R by Christian P. Robert; George Casella. Technometrics, 53(2):217–218, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
23209743.

Neubauer:2012:BRBa

[Neu12] Dean V. Neubauer. Book review: The R Primer by Claus Thorn Ekstrøm. Technometrics, 54(3):326– 327, August 2012. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714907.

Ng:2006:BRB

[Ng06] H. K. Tony Ng. Book review: Statistics: An Introduction Using R by Michael J. Crawley. Technometrics, 48(2):308, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471175.

Ng:2009:BRB

[Ng09] Hon Keung Tony Ng. Book review: Probability and Statistics With R by Maria Dolores Ugarte; Ana F. Militino; Alan T. Arnholt. Technometrics, 51 (2):220, May 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40586606.

Ng:2011:BRB

[Ng11] Hon Keung Tony Ng. Book review: Introduction to Scientific Programming and Simulation Using R by Owen Jones; Robert Maillardet; Andrew Robinson. Technometrics, 53(2):213– 214, May 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
23209737.

Noguchi:2012:NRS

[NGBK12] Kimihiro Noguchi, Yulia R.

Gel, Edgar Brunner, and Frank Konietschke. nparLD: An R software package for the nonparametric analysis of longitudinal data in factorial experiments. Journal of Statistical Software, 50(12): ??, September 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v50/i12.

Nielsen:2011:BRB

[Nie11] Søren Feodor Nielsen. Book review: Introductory time series with R. Journal of Applied Statistics, 38(10): 2370–2371, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [CM09].

Nielsen:2014:BRB

[Nie14] Søren Feodor Nielsen. Book review: An introduction to analysis of financial data with R. Journal of Applied Statistics, 41(12):2777– 2778, December 2014. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nadarajah:2006:CSR

[NK06] Saralees Nadarajah and Samuel Kotz. Code snippet: R programs for truncated distributions. Journal of Statistical Software, 16(CS-2):1–8, August 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v16/c02.

Narajewski:2021:TRP

[NKHZ21] Michał Narajewski, Jens Kley-Holsteg, and Florian

                                        Ziel.  tsrobprep — an

R package for robust preprocessing of time series data. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S2352711021001084.

Nolan:2012:IAS

[NL12] Deborah Nolan and Duncan Temple Lang. Interactive and animated scalable vector graphics and R data displays. Journal of Statistical Software, 46(1):

??, January 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/i01.

Nolan:2014:XWT

[NL14] Deborah Ann Nolan and Duncan Temple Lang. XML and Web technologies for data sciences with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 1-4614-7899-

5. xxiv + 663 pp. LCCN QA76.76.H94 N65 2014.

URL http://www.loc.gov/
catdir/enhancements/fy1503/
2013954669-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1503/2013954669-d.html;
http://www.loc.gov/catdir/
enhancements/fy1503/2013954669-
t.html.

Nordhausen:2008:SBR

[Nor08] Klaus Nordhausen. Short book review: Lattice: Multivariate Data Visualization with R by Deepayan Sarkar. International Statistical Review = Revue Internationale de Statistique, 76(3):440, December 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919660.

Nordhausen:2009:SBRb

[Nor09] Klaus Nordhausen. Short book review: A First Course in Statistical Programming with R by W. John Braun, Duncan J. Murdoch. International Statistical Review = Revue Internationale de Statistique, 77(2):317, August 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919747.

Nordhausen:2014:BRBb

[Nor14] Klaus Nordhausen. Book review: An Introduction to Statistical Learning — with Applications in R by Gareth James, Daniela Witten, Trevor Hastie & Robert Tibshirani. International Statistical Review = Revue Internationale de Statistique, 82(1):156–157, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nordhausen:2015:BRS

[Nor15] Klaus Nordhausen. Book review: Statistical Analysis of Network Data with R. International Statistical Review = Revue Internationale de Statistique, 83(1):171–172, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nolan:2017:PFR

[NPP17] Rory Nolan and Sergi Padilla-Parra. filesstrings:

An R package for file and string manipulation. Journal of Open Source Software, 2 (14):260:1–260:2, June 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00260.

Nolan:2018:PIR

[NPP18] Rory Nolan and Sergi Padilla-Parra. ijtiff: An R package providing TIFF I/O for ImageJ users. Journal of Open Source Software, 3(23):633:1–633:2, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00633.

Nadarajah:2013:CRP

[NPR13] Saralees Nadarajah, Bozidar V. Popović, and Miroslav M.

Ristić. Compounding: an

R package for computing continuous distributions obtained by compounding a continuous and a discrete distribution. Computational Statistics, 28(3):977– 992, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-012-0336-y.

Nadarajah:2016:NRP

[NR16] Saralees Nadarajah and Ricardo Rocha. Newdistns: An R package for new families of distributions. Journal of Statistical Software, 69(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v069i10;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v069i10/v69i10.pdf.

Nowok:2016:SBC

[NRD16] Beata Nowok, Gillian M. Raab, and Chris Dibben. synthpop: Bespoke creation of synthetic data in R. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v074i11;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v074i11/
                                        v74i11.pdf.

Nason:1994:DWT

[NS94] G. P. Nason and B. W. Silverman. The discrete wavelet transform in S. Journal of Computational and Graphical Statistics, 3 (2):163–191, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474637.

Nummi:2013:SBRa

[Num13] Tapio Nummi. Short book review: Clinical Trial Data Analysis Using R by Ding-Geng (Din) Chen, Karl E. Peace. International Statistical Review = Revue Internationale de Statistique, 81 (1):151, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nummi:2020:BRS

[Num20] Tapio Nummi. Book review:

Semiparametric Regression with R. Jaroslaw Harezlak, David Ruppert and Matt P. Wand Springer, 2018, xi + 331 pages, EUR85.39, ebook ISBN: 978-1-4939-8853-2. International Statistical Review = Revue Internationale de Statistique, 88(2):516, August 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Nunes:2020:BRT

[Nun20] Matthew Nunes. Book review: Time Series: a Data Analysis Approach Using R By Robert H. Shumway and David S. Stoffer. Published by Taylor and Francis Group, LLC, Boca Raton, London, New York, 2019. ISBN: 978-0-367-22109-6 (Hardback). Journal of Time Series Analysis, 41(3):485– 486, May 2020. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Nash:2011:UOA

[NV11] John C. Nash and Ravi Varadhan. Unifying optimization algorithms to aid software system users: optimx for R. Journal of Statistical Software, 43(9): ??, August 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/i09.

Niu:2021:RPS

[NWH21] Mu Niu, Joe Wandy, and Dirk Husmeier. R package for statistical inference in dynamical systems using kernel based gradient matching: KGode. Computational Statistics, 36(1): 715–747, March 2021. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s00180-
                                        020-01014-x.

OBrien:2008:SBRa

[O’B08] C. M. O’Brien. Short book review: Adaptive Design Theory and Implementation Using SAS and R by Mark Chang. International Statistical Review = Revue Internationale de Statistique, 76(2):300, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27919619.

OBrien:2009:SBRe

[O’B09a]  C. M. O’Brien.  Short

book review: Introduction to Scientific Programming and Simulation Using R by [E. Stuart] Owen-Jones, Robert Maillardet, Andrew Robinson. International Statistical Review = Revue Internationale de Statistique, 77(3):471, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919776.

OBrien:2009:SBRg

[O’B09b] C. M. O’Brien. Short book review: Nonlinear Regression with R by Christian Ritz, Jens Carl Streibig. International Statistical Review = Revue Internationale de Statistique, 77 (3):473, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919778.

OBrien:2009:SBRd

[O’B09c] C. M. O’Brien. Short book review: Wavelet Methods in Statistics with R by G. P. Nason. International Statistical Review = Revue Internationale de Statistique, 77(1):159, April 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27919710.

OBrien:2010:SBRd

[O’B10] Carl M. O’Brien. Short book review: Dynamic Linear Models with R by Giovanni Petris, Sonia Petrone, Patrizia Campagnoli. International Statistical Review = Revue Internationale de Statistique, 78 (1):157, April 2010. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        27919824.

OBrien:2012:SBRg

[O’B12a] Carl M. O’Brien. Short book review: Analysis of Questionnaire Data with R by Bruno Falissard. International Statistical Review = Revue Internationale de Statistique, 80(3):474–475, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2012:SBRa

[O’B12b] Carl M. O’Brien. Short book review: Numerical Ecology with R by Daniel Borcard, François Gillet, Pierre Legendre. International Statistical Review = Revue Internationale de Statistique, 80(1):201, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
23257196.

OBrien:2013:SBRa

[O’B13] Carl M. O’Brien. Short book review: An Introduction to Bootstrap Methods with Applications to R by Michael R. Chernick, Robert A. LaBudde. International Statistical Review = Revue Internationale de Statistique, 81(1):157–158, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

OBrien:2014:BRF

[O’B14] Carl M. O’Brien. Book review: Functional and Phylogenetic Ecology in R by Nathan G. Swenson. International Statistical Review = Revue Internationale de Statistique, 82(3):486–487, December 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Oberski:2014:LSR

[Obe14] Daniel Oberski. lavaan.survey:

An R package for complex survey analysis of structural equation models. Journal of Statistical Software, 57 (1):??, March 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v57/i01.

Oliveira:2014:PRP

[OCRC14] María Oliveira, Rosa M. Crujeiras, and Alberto Rodríguez-Casal. NPCirc: An R package for nonparametric circular methods. Journal of Statistical Software, 61(9):??, November 2014. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i09.

OConnell:2011:HSM

[OH11] Jared O’Connell and Søren Højsgaard. Hidden semi Markov models for multiple observation sequences: The mhsmm package for R. Journal of Statistical Software, 39 (4):??, March 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v39/i04.

OConnell:2017:CVS

[OHD17] Mark O’Connell, Catherine B. Hurley, and Katarina Domijan. Conditional visualization for statistical models: An introduction to the condvis package in R. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081i05/
v81i05.pdf.

Ohri:2014:RCC

[Ohr14] Ajay Ohri. R for cloud computing: an approach for data scientists. Springer Science+Business Media, LLC, New York, NY, USA, 2014. ISBN 1-4939-1701-3. xvii + 267 pp. LCCN ????

URL http://www.loc.gov/
catdir/enhancements/fy1413/
2014950411-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1413/2014950411-d.html;
http://www.loc.gov/catdir/
enhancements/fy1413/2014950411-
t.html.

Oja:2010:MNM

[Oja10] Hannu Oja. Multivariate Nonparametric Methods with R: an Approach Based on Spatial Signs and Ranks, volume 199 of Lecture notes in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-0467-0 (paperback). ISSN 0930-0325. xiii + 232 pp. LCCN QA276.45.R3 O33 2010.

Owen-Jones:2009:ISP

[OJMR09] E. Stuart Owen-Jones, Robert Maillardet, and Andrew (Andrew P.) Robinson. Introduction to scientific programming and simulation with using R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6872-5. xix + 453 pp. LCCN Q183.9 .J66 2009.

Overstall:2014:PRP

[OK14] Antony M. Overstall and Ruth King. conting: An R package for Bayesian analysis of complete and incomplete contingency tables. Journal of Statistical Software, 58(7):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v58/
i07.

Oller:2017:FRP

[OL17] Ramon Oller and Klaus Langohr. FHtest: An R package for the comparison of survival curves with censored data. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081i15;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081i15/
v81i15.pdf.

Olive:2007:BRB

[Oli07] David J. Olive. Book review: Robust Statistical Methods with R by J. Jureä ová; J. Picek. Technometrics, 49(4):496, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471404.

Olive:2010:BRBc

[Oli10a] David J. Olive. Book review: Mixed Effects Models and Extensions in Ecology With R by A. F. Zuur; E. N. Ieno; N. J. Walker; A. A. Saveliev; G. M. Smith. Technometrics, 52(4):464– 465, November 2010. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40997258.

Olive:2010:BRBb

[Oli10b] David J. Olive. Book review:

Software for Data Analysis: Programming With R by J. M. Chambers. Technometrics, 52(2):261, May 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
27867236.

Olivier:2017:BRA

[Oli17] Jake Olivier. Book review:

Applied surrogate endpoint evaluation methods with SAS and R. Ariel Alonso, Theophile Bigirumurame, Tomasz Burzykowski, Marc Buyse, Geert Molenberghs, Leacky Muchene, Nolen Joy Perualila, Ziv Shkedy, Wim Van der Elst, Chapman and Hall/CRC Biostatics Series, 2017, xxi + 373, 57.99 GBP, ISBN 13: 978-1-4822-4936-1 (Hardback). Statistics in Medicine, 36(24):3919– 3920, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ooms:2010:BRB

[Oom10] Jeroen Ooms. Book review: SAS and R. Journal of Statistical Software, 37(BR-3): ??, December 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v37/b03.

Ooms:2013:RPE

[Oom13] Jeroen Ooms. The RAppArmor package: Enforcing security policies in R using dynamic sandboxing on Linux. Journal of Statistical Software, 55(7): ??, November 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v55/i07.

Oliveira:2018:UDP

[OOSM18] Amílcar Oliveira, Teresa Oliveira, and Antonio Seijas-Macías. The uniform distribution product: an approach to the (Q, r) inventory model using R. Journal of Applied Statistics, 45(2):284– 297, February 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Ormerod:2017:BRE

[Orm17] John T. Ormerod. Book review: Extending the linear model with R: generalized linear, mixed effects and nonparametric regression models, second edition by Julian J. Faraway, Chapman and Hall/ CRC, Boca Raton, 2016. No. of pages: 399. Price: £63.99 (book + eBook); £44.79 (eBook). ISBN 978-1-4987-2096-0. Statistics in Medicine, 36(14):2318– 2319, June 30, 2017. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olsen:1995:SCH

[OS95] N. H. Olsen and Pacific Biological Station. Software to complement HTI’s model 240 split-beam echosounder: a user’s guide to HAFU (hydro-acoustic file utilities) and QTS (Qualark Tools for S-Plus). Technical report 2037, Fisheries and Oceans, Canada, Nanaimo, BC, Canada, 1995. vii + 45 pp.

Olusoji:2021:PCR

[OSH^{+}21] Oluwafemi D. Olusoji, Jurg W.

Spaak, Mark Holmes, Thomas Neyens, Marc Aerts, and Frederik De Laender. cyanoFilter: an R package to identify phytoplankton populations from flow cytometry data using cell pigmentation and granularity. Ecological Modelling, 460(??):Article 109743, November 15, 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S030438002100291X.

Otneim:2017:BRB

[Otn17] Håkon Otneim. Book review: Business Analytics Using R — A Practical Approach. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v081b01;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v081b01/
                                        v81b01.pdf.

Pallmann:2015:BRA

[Pal15a] Philip Pallmann. Book reviews: Analyzing Baseball Data with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 178(4):1099, October 2015. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Pallmann:2015:AMA

[Pal15b] Philip Pallmann. Applied meta-analysis with R. Journal of Applied Statistics, 42(4):914–915, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Pandolfo:2015:CSR

[Pan15] Giuseppe Pandolfo. Book review: Circular statistics in R. Journal of Applied Statistics, 42(4):918–919, April 2015. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Panse:2018:CSR

[Pan18] Christian Panse. Code snippet: Rectangular statistical cartograms in R: The recmap package. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086c01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086c01/
v86c01.pdf.

Papastamoulis:2016:CSL

[Pap16] Panagiotis Papastamoulis.

Code snippet: label.switching:

An R package for dealing with the label switching problem in MCMC outputs. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069c01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069c01/
v69c01.pdf.

Paradis:2006:APE

[Par06a] Emmanuel Paradis. Analysis of phylogenetics and evolution with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-32914-5 (paperback), 1-4614-1743-0 (e-book). xii + 211 pp. LCCN QH83 .P37 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0824/
2006923823-d.html;   http:
//www.loc.gov/catdir/toc/
fy0703/2006923823.html.

Pardoe:2006:ARM

[Par06b] Iain Pardoe. Applied regression modeling: a business approach. Wiley-In-terscience, New York, NY, USA, 2006. ISBN 0-471-97033-6. xx + 303 pp. LCCN QA278.2 .P363 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0741/
2006044262-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2006044262-d.html;
 http://www.loc.gov/catdir/
toc/fy0702/2006044262.
html;   http://www.wiley.
com/WileyCDA/WileyTitle/
productCd-0471970336.html.

Paradis:2012:APEn

[Par12] Emmanuel Paradis. Analysis of Phylogenetics and Evolution with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-1743-0. xiv + 386 + 89 pp. LCCN QH359-425.

Park:2015:BRD

[Par15a] Mira Park. Book review:

Ding-Geng, Chen, and Karl, E. Peace, Applied metaanalysis with R. CRC Press. Biometrics, 71(3):864–865, September 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2015:BRBb

[Par15b] Taesung Park. Book review: Bilder, Christopher R. and Loughin, Thomas M., Analysis of Categorical Data with R. Boca Raton, FL: Chapman and Hall/ CRC. Biometrics, 71(4): 1198–1199, December 2015. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pavia:2016:BRH

[Pav16] Jose M. Pavía. Book review: Humanities Data in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069b02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069b02/
v69b02.pdf.

Pinheiro:2000:MEM

[PB00] Jose C. Pinheiro and Douglas M. Bates. Mixedeffects models in S and S-PLUS. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98957-9. xvi + 528 pp. LCCN QA76.73.S15 P56 2000. US$69.95.

Padoan:2015:ARF

[PB15] Simone A. Padoan and Moreno Bevilacqua. Analysis of random fields using

CompRandFld. Journal of

Statistical Software, 63(9): ??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i09.

Plummer:2011:LRC

[PC11] Martyn Plummer and Bendix Carstensen. Lexis: An R class for epidemiological studies with long-term follow-up. Journal of Statistical Software, 38(5):??, January 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v38/i05.

Parsons:2009:RMP

[PCAS09] Nick R. Parsons, Matthew L. Costa, Juul Achten, and Nigel Stallard. Repeated measures proportional odds logistic regression analysis of ordinal score data in the statistical software package R. Computational Statistics & Data Analysis, 53(3): 632–641, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947308003939.

Peng:2008:SME

[PD08] Roger D. Peng and Francesca Dominici. Statistical Methods for Environmental Epidemiology with R: a Case Study in Air Pollution and Health. Use R! Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-78166-8 (paperback), 0-387-78167-6 (e-book). ix + 144 pp. LCCN TD883.1 .P46 2008.

Peng:2016:TRP

[PDH16] Jun Peng, ZhiBao Dong, and FengQing Han. tgcd: an R package for analyzing thermoluminescence glow curves. SoftwareX, 5 (??):107–111, ???? 2016. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711016300140.

Pebesma:2012:SST

[Peb12] Edzer Pebesma. spacetime:

Spatio-temporal data in R. Journal of Statistical Software, 51(7):??, November 2012. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v51/i07.

Pebesma:2021:BRC

[Peb21] Edzer Pebesma. Book review: Christopher K. Wikle, Andrew Zammit-Mangion and Noel Cressie (2019): Spatio-temporal Statistics with R. Chapman and Hall/ CRC, 396 pp. $47.96 (Hardcover), ISBN 978-1-138-71113-6. Statistical Papers, 62(2):1093–1094, April 2021. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s00362-
                                        021-01224-5.

Peng:2003:MDP

[Pen03] Roger D. Peng. Multidimensional point process models in R. Journal of Statistical Software, 8(16): 1–27, ???? 2003. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v08/i16;
                                         http://www.jstatsoft.
                                        org/v08/i16/ptprocR102.
                                        pdf.

Peng:2008:CDS

[Pen08] Roger Peng. Caching and distributing statistical analyses in R. Journal of Statistical Software, 26(7): ??, July 2008. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v26/i07.

Peng:2017:CSY

[Pen17] Roger D. Peng. Comment on “50 Years of Data Science”. Journal of Computational and Graphical Statistics, 26(4):767, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Don17].

PerpinanLamigueiro:2014:DTS

[Per14] Oscar Perpiñán Lamigueiro.

Displaying time series, spatial, and space-time data with R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4665-6520-9 (hardcover). vii + 200 pp. LCCN QA280 .P475 2014.

Petersen:2002:BRE

[Pet02] Jørgen Holm Petersen. Book review: Environmental Statistics with S-Plus. Steven P. Millard and Nagaraj K. Neerchal. CRC Press LLC, London, 2001. No. of pages:

830. Price: $79.95. ISBN 0-8493-7168-6. Statistics in Medicine, 21(17):2607– 2608, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [MN03].

Petris:2010:RPD

[Pet10]  Giovanni Petris.  An R

package for dynamic linear models. Journal of Statistical Software, 36(12): ??, October 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v36/i12.

Peterson:2019:BRS

[Pet19] Adam Peterson. Book review: Spatio-temporal statistics with R, Wikle, Christopher K., Zammit-Mangion, Andrew, Cressie, Noel, Boca Raton, FL: CRC Press, 2019. Biometrics, 75(4):1414, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pfaff:2006:AIC

[Pfa06]  Bernhard Pfaff.  Analy-

sis of integrated and cointegrated time series with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-27959-8, 0-387-27960-1 (e-book). xi + 139 pp. LCCN QA280 .P45 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0663/
2005935027-d.html;   http:
//www.loc.gov/catdir/toc/
fy0610/2005935027.html.

Pfaff:2008:AIC

[Pfa08a] Bernhard Pfaff. Analysis of integrated and cointegrated time series with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 0-387-75966-2 (paperback). xx + 188 pp. LCCN QA280 .P45 2008. URL http://

www.loc.gov/catdir/toc/
fy0906/2008930126.html.

Pfaff:2008:VSS

[Pfa08b]  Bernhard Pfaff.  VAR,

SVAR and SVEC models: Implementation within R package vars. Journal of Statistical Software, 27(4):??, July 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v27/i04.

Pfahler:2012:BRB

[Pfa12] Lori Pfahler. Book review: A Handbook of Statistical Analyses Using R (2nd ed.) by Brian S. Everitt; Torsten Hothorn. Technometrics, 54(1):100, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714839.

Pfaff:2013:FRM

[Pfa13] Bernhard Pfaff. Financial risk modelling and portfolio optimization with R. Statistics in practice. Wiley, New York, NY, USA, 2013. ISBN 0-470-97870-8 (hardcover), 1-118-47713-8 (e-book), 1-118-47712-X (ebook). xvi + 356 pp. LCCN HG106 .P484 2013eb. URL http:

//librarytitles.ebrary.
com/Doc?id=10777209;

                                        http://rbdigital.oneclickdigital.
                                        com.

Parsons:2012:RPI

[PFT^{+}12] Nick Parsons, Tim Friede, Susan Todd, Elsa Valdes Marquez, Jeremy Chataway, Richard Nicholas, and Nigel Stallard. An R package for implementing simulations for seamless phase II/III clinical trials using early outcomes for treatment selection. Computational Statistics & Data Analysis, 56 (5):1150–1160, May 1, 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167947310004159.

Pelissier:2015:PPR

[PG15] Raphaël Pélissier and François Goreaud. ads package for R: A fast unbiased implementation of the K-function family for studying spatial point patterns in irregular-shaped sampling windows. Journal of Statistical Software, 63 (6):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v63/i06.

Pawlowsky-Glahn:2020:BRP

[PG20] Vera Pawlowsky-Glahn. Book review: Peter Filzmoser, Karel Hron, Matthias Templ: Applied compositional data analysis, with worked examples in R. Statistical Papers, 61(2):921–922, April 2020. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00362-020-01163-7.

Pallmann:2016:AMG

[PH16] Philip Pallmann and Ludwig A. Hothorn. Analysis of means: a generalized approach using R. Journal of Applied Statistics, 43 (8):1541–1560, August 2016. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Phillips:2010:CSR

[Phi10] Kem Phillips. Code snippet: R functions to symbolically compute the central moments of the multivariate normal distribution. Journal of Statistical Software, 33(CS-1):??, February 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v33/
c01.

Picka:2009:BRB

[Pic09] Jeffrey Picka. Book review: Introduction to Probability With R by K. Baclawski. Technometrics, 51 (2):221, May 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40586607.

Pienaar:2009:BRB

[Pie09] Jacques A. Pienaar. Book review: Probability and statistics with R. Journal of Applied Statistics, 36(6):703– 704, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [UMA08].

Pierson:2015:DSD

[Pie15] Lillian Pierson. Data science for dummies. For dummies. For Dummies, Hoboken, NJ, USA, 2015. ISBN 1-118-84155-7 (paperback), 1-118-84145-X (e-book), 1-118-84152-2 (e-book). xvi + 384 pp. LCCN T58.5 .P488 2015. Foreword by Jake Porway, Founder and Executive Director of DataKind.

Pinon:2024:BRH

[Piñ24] Adelson Piñón. Book review: Handbook of Regression Modeling in People Analytics, With Examples in R and Python, by Keith McNulty, 1st edition, CRC Press. 2021. 255 pp. ISBN: 978-1-032-04174-2, £59.99 (hbk). Journal of the Royal Statistical Society. Series A (Statistics in Society), 187 (2):544–545, April 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        academic.oup.com/jrsssa/
                                        article/187/2/544/7256154.

Piras:2010:SSM

[Pir10] Gianfranco Piras. sphet:

Spatial models with heteroskedastic innovations in R. Journal of Statistical Software, 35(1):??, June 2010. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/i01.

Perez-Jaume:2017:TOT

[PJSPC17] Sara Perez-Jaume, Konstantina Skaltsa, Natàlia Pallarès, and Josep L.

Carrasco. ThresholdROC:

Optimum threshold estimation tools for continuous diagnostic tests in R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i04/
v82i04.pdf.

Pineda-Krch:2008:GIG

[PK08] Mario Pineda-Krch. GillespieSSA:

Implementing the Gillespie stochastic simulation algorithm in R. Journal of Statistical Software, 25(12): 1–18, April 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i12.

Putler:2012:CBA

[PK12] Daniel S. Putler and Robert E. Krider. Customer and business analytics: applied data mining for business decision making using R. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0396-3. xxvi + 289 pp. LCCN HF5415.126 .P88 2012.

Palaseanu-Lovejoy:2023:IOS

[PL23] Monica Palaseanu-Lovejoy. iBluff: an open-source R package for geomorphic analysis of coastal bluffs/cliffs. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023000213.

Plant:2012:SDA

[Pla12] Richard E. Plant. Spatial Data Analysis in Ecology and Agriculture Using R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-1913-0 (hardcover). xv + 631 pp. LCCN S566.55 .P53 2012.

Plant:2019:SDA

[Pla19] Richard E. Plant. Spatial Data Analysis in Ecology and Agriculture Using R. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, second edition, 2019. ISBN 0-8153-9275-3 (hardcover), 1-351-18991-3 (e-book). xvii + 666 pp. LCCN S566.55 .P53 2019. Platas-Lopez:2023:DRP

[PLGHG^{+}23] Alejandro Platas-López, Alejandro Guerra-Hernández, Francisco Grimaldo, Nicandro Cruz-Ramírez, Efrén Mezura-Montes, and Marcela Quiroz-Castellanos. dplbnDE: an R package for discriminative parameter learning of Bayesian Networks by Differential Evolution. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023001383.

Poole:2011:SRC

[PLLC11] Keith Poole, Jeffrey B.

Lewis, James Lo, and Royce Carroll. Scaling roll call votes with wnominate in R. Journal of Statistical Software, 42(14):??, June 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v42/
i14.

Poole:2016:RBS

[PLR^{+}16] Keith T. Poole, Jeffrey B.

Lewis, Howard Rosenthal, James Lo, and Royce Carroll. Recovering a basic space from issue scales in R. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK.

ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v069i07;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v069i07/
                                        v69i07.pdf.

Pateiro-Lopez:2010:GCH

[PLRC10] Beatriz Pateiro-López and Alberto Rodríguez-Casal. Generalizing the convex hull of a sample: The R package alphahull. Journal of Statistical Software, 34 (5):??, April 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v34/i05.

Plummer:2001:BRP

[Plu01] Martyn Plummer. Book review: Programming with Data: a Guide to the S Language. John M. Chambers, Springer, Godalming, U.K., 1998. No. of pages:xv + 469. Price: £32.50. ISBN 0-387-98503-4. Statistics in Medicine, 20(6):985–986, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paciorek:2015:PGP

[PLZ^{+}15] Christopher J. Paciorek, Benjamin Lipshitz, Wei Zhuo, Prabhat, Cari G. G. Kaufman, and Rollin C. Thomas. Parallelizing Gaussian process calculations in R. Journal of Statistical Software, 63(10): ??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i10.

Polanco-Martinez:2023:NRP

[PMLM23] Josué M. Polanco-Martínez and José L. López-Martínez.

NonParRolCor: an R pack-

age for estimating rolling correlation for two regular time series. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000493.

Punzo:2018:CRP

[PMM18] Antonio Punzo, Angelo Mazza, and Paul D. McNi-

cholas. ContaminatedMixt:

An R package for fitting parsimonious mixtures of multivariate contaminated normal distributions. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i10;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085i10/
v85i10.pdf.

Payton:2015:PPL

[PMW^{+}15] Quinn C. Payton, Michael G.

McManus, Marc H. Weber, Anthony R. Olsen, and Thomas M. Kincaid. micromap: A package for linked micromaps. Journal of Statistical Software, 63 (2):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v63/i02.

Pillai:1997:MSD

[PN97] Krishnakumar Pillai and V. S. Sukumaran Nair. A model for software development effort and cost estimation. IEEE Transactions on Software Engineering, 23 (8):485–497, August 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Pewsey:2013:CSR

[PNR13] Arthur Pewsey, Markus Neuhäuser, and Graeme D. Ruxton. Circular statistics in R. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-967113-3. xiv + 183 pp. LCCN QA276 .P45 2013.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy1410/
                                        2013940576-b.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy1410/2013940576-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1410/2013940576-
                                        t.html.

Podgorski:2015:BRD

[Pod15] Krzysztof Podgorski. Book review: Displaying Time Series, Spatial and Space-Time Data with R. International Statistical Review = Revue Internationale de Statistique, 83(1):165, April 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2018:BRF

[Pod18] Krzysztof Podgorski. Book review: Financial Analytics with R, Mark J. Bennet and Dirk L. Hugen, Cambridge University Press, 2016, 377 pages, EUR55.99, hardcover ISBN: 978-1-107-15075-1. International Statistical Review = Revue Internationale de Statistique, 86(1):163–165, April 2018. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Podgorski:2021:BRC

[Pod21] Krzysztof Podgórski. Book review: Computational Genomics with R, Altuna Akalin Chapman and Hall/ CRC, 2021, xxii + 440 pages, £99.99, hardcover ISBN: 978-1-4987-8185-5. International Statistical Review = Revue Internationale de Statistique, 89(2):420–421, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Polasek:2009:BRP

[Pol09] Wolfgang Polasek. Book review: Phillip I. Good (2005): Introduction to Statistics Through Resampling Methods and R/S-PLUS (Paperback); Phillip I. Good (2005): Introduction to Statistics Through Resampling Methods and Microsoft Office Excel (Paperback). Statistical Papers, 50(3):673–675, June 2009. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://

link.springer.com/article/
10.1007/s00362-008-0131-
6.

Polasek:2011:BRJ

[Pol11a] Wolfgang Polasek. Book review: Jim Albert (2007): Bayesian computation with R. Statistical Papers, 52 (1):239–240, February 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://

link.springer.com/accesspage/
article/10.1007/s00362-
009-0206-z.

Polasek:2011:NPR

[Pol11b] Wolfgang Polasek. New publications on R. Statistical Papers, 52(4):987–988, November 2011. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s00362-
010-0323-8.

Polasek:2011:SBRa

[Pol11c] Wolfgang Polasek. Short book review: Using R for Data Management, Statistical Analysis, and Graphics by Nicholas J. Horton, Ken Kleinman. International Statistical Review = Revue Internationale de Statistique, 79 (2):284–285, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
41305035.

Polasek:2013:SBRc

[Pol13] Wolfgang Polasek. Short book review: Time Series Analysis and Its Applications: With R Examples, Third Edition by Robert H. Shumway, David S. Stoffer. International Statistical Review = Revue Internationale de Statistique, 81 (2):323–325, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Petris:2011:SSM

[PP11] Giovanni Petris and Sonia Petrone. State space models in R. Journal of Statistical Software, 41 (4):??, May 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v41/i04.

Pan:2017:JRP

[PP17] Jianxin Pan and Yi Pan.

jmcm: An R package for joint mean-covariance modeling of longitudinal data. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i09/
v82i09.pdf.

Perme:2018:NRS

[PP18] Maja Pohar Perme and Klemen Pavlic. Nonparametric relative survival analysis with the R package relsurv. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v087i08;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v087i08/
v87i08.pdf.

Petris:2009:DLM

[PPC09] Giovanni Petris, Sonia Petrone, and Patrizia Campagnoli. Dynamic Linear Models with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-77237-5, 0-387-77238-

3. xiii + 251 pp. LCCN QA276.45.R3 P48 2009.

URL http://www.loc.gov/
catdir/enhancements/fy1108/
2009926480-d.htm.

Palomo:2015:SRP

[PPGD15] Jesús Palomo, Rui Paulo, and Gonzalo García-Donato. SAVE: An R package for the statistical analysis of computer models. Journal of Statistical Software, 64 (13):??, April 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v64/i13.

Petzoldt:2007:SOO

[PR07] Thomas Petzoldt and Karsten Rinke. simecol: An objectoriented framework for ecological modeling in R. Journal of Statistical Software, 22(9):1–31, September 2007. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v22/
i09.

Prins:2005:LMR

[Pri05a] Samantha C. Bates Prins.

Linear models with R. The American Statistician, 59(3): 276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prins:2005:SAD

[Pri05b] Samantha C. Bates Prins.

Statistical analysis and data display: An intermediate course with examples in S-Plus, R, and SAS. The American Statistician, 59 (3):275–276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Pagano:2022:EEM

[PS22] Davide Pagano and Lorenzo Sostero. EmiR: Evolutionary minimization for R. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022000590.

Petrou:2011:OPP

[PSM^{+}11] Savvas Petrou, Terence M. Sloan, Muriel Mewissen, Thorsten Forster, Michal Piotrowski, Bartosz Dobrzelecki, Peter Ghazal, Arthur Trew, and Jon Hill. Optimization of a parallel permutation testing function for the SPRINT R package. Concurrency and Computation: Practice and Experience, 23(17):2258–2268, December 10, 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pearce:2017:HRP

[PSS^{+}17] Robert G. Pearce, R. Woodrow Setzer, Cory L. Strope, Nisha S. Sipes, and John F. Wambaugh. httk: R package for high-throughput toxicokinetics. Journal of Statistical Software, 79(??): ??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/

jss/article/view/v079i04;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v079i04/v79i04.pdf.

Plakidas:2017:ERS

[PSZ17] Konstantinos Plakidas, Daniel Schall, and Uwe Zdun. Evolution of the R software ecosystem: Metrics, relationships, and their impact on qualities. The Journal of Systems and Software, 132 (??):119–146, October 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0164121217301371.

Polzehl:2007:ASD

[PT07] Joerg Polzehl and Karsten Tabelow. Adaptive smoothing of digital images: The R package adimpro. Journal of Statistical Software, 19 (1):1–17, March 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v19/i01.

Polzehl:2009:SAS

[PT09] Jörg Polzehl and Karsten Tabelow. Structural adaptive smoothing in diffusion tensor imaging: The R package dti. Journal of Statistical Software, 31(9): ??, September 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v31/i09.

Pilhofer:2013:NAV

[PU13] Alexander Pilhöfer and Antony Unwin. New approaches in visualization of categorical data: R package extracat. Journal of Statistical Software, 53 (7):??, May 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v53/i07.

Putt:2006:BRL

[Put06] Mary E. Putt. Book review: 2. Linear models with R. Julian J. Faraway, Chapman and Hall/CRC, Boca Raton, 2005. No of pages: x + 229. Price: $69.95. ISBN: 1-58488-425-8. Statistics in Medicine, 25(8):1440– 1441, April 30, 2006. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pham:2018:GAM

[PW18] Tung H. Pham and Matt P. Wand. Generalised additive mixed models analysis via gammSlice. Australian & New Zealand Journal of Statistics, 60(3):279– 300, September 2018. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Perez:2012:CSI

[PZK^{+}12] Adriana Pérez, Saijuan Zhang, Victor Kipniss, Laurence S. Freedman, and Raymond J. Carroll. Code snip-

pet: Intake epis food():

An R function for fitting a bivariate nonlinear measurement error model to estimate usual and energy intake for episodically consumed foods. Journal of Statistical Software, 46(CS-3):??, March 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/c03.

Qian:2010:BRBa

[Qia10a] Lianfen Qian. Book review: Time Series Analysis With Applications in R (2nd ed.) by Jonathan D. Cryer; Kung-Sik Chan. Technometrics, 52(3):365, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
27867262.

Qian:2010:EES

[Qia10b] Song S. Qian. Environmental and Ecological Statistics with R. Chapman and Hall/CRC applied environmental statistics. Chapman and Hall/CRC, Boca Raton, FL, USA, 2010. ISBN 1-4200-6206-9 (paperback). xix + 421 pp. LCCN GE45.S73 Q25 2010.

Qian:2016:EES

[Qia16] Song S. Qian. Environmental and ecological statistics with R. Chapman and Hall/CRC Press applied environmental statistics. Taylor and Francis, Boca Raton, FL, USA, second edition, 2016. ISBN 1-4987-2872-3 (hardcover), 1-4987-2873-1 (e-book), 1-4987-2875-8 (e-book), 1-4987-2874-X (ebook). ???? pp. LCCN GE45.S73 Q25 2016.

Qin:2021:SLI

[QZLP21] Lu Qin, Jihong Zhang, Xinya Liang, and Qianqian Pan. R-squared of a latent interaction in structural equation model: a tutorial of using R. International Journal of Statistics and Probability, 10(3):69–??, ???? 2021. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL

                                        http://www.ccsenet.org/
                                        journal/index.php/ijsp/
                                        article/view/0/45113.

RDCT:2004:RRM

[R D04] R Development Core Team. The R reference manual: base package. Network Theory Ltd., Bristol, UK, 2004. ISBN 0-9546120-0-0 (vol. 1), 0-9546120-1-9 (vol. 2). x + 723 (vol. 1), viii + 695 (vol.

2) pp. LCCN QA76.7 RRE.

RDCT:2011:IRN

[R D11a] R Development Core Team. An Introduction to R: Notes on R: a Programming Environment for Data Analysis and Graphics: Version 2.13.0 (2011-04-13). ????, ????, 2011. ISBN 3-900051-12-7. iv + 103 pp. LCCN ???? URL http://

cran.r-project.org/doc/
manuals/R-intro.pdf.

RDCT:2011:RLE

[R D11b] R Development Core Team.

R: a Language and Environment for Statistical Computing: Reference Index. ????, ????, 2011. ISBN 3-900051-07-0. xlix + 3396 pp. LCCN ???? URL http://

cran.r-project.org/doc/
manuals/fullrefman.pdf.

RDCT:2011:RDI

[R D11c] R Development Core Team.

R Data Import/Export: Version 2.13.0 (2011-04-13). ????, ????, 2011. ISBN 3-900051-10-0. ii + 35 pp. LCCN ???? URL http://

cran.r-project.org/doc/
manuals/R-data.pdf.

RDCT:2011:RIA

[R D11d] R Development Core Team.

R Installation and Administration: Version 2.13.0 (2011-04-13). ????, ????, 2011. ISBN 3-900051-09-7. iii + 66 pp. LCCN

????  URL http://
cran.r-project.org/doc/
manuals/R-admin.pdf.

RDCT:2011:RIV

[R D11e] R Development Core Team.

R Internals: Version 2.13.0 (2011-04-13). ????, ????, 2011. ISBN 3-900051-14-3. ii + 58 pp. LCCN

                                        ????  URL http://
                                        cran.r-project.org/doc/
                                        manuals/R-ints.pdf.

RDCT:2011:RLD

[R D11f] R Development Core Team. R Language Definition: Version 2.13.0 (2011-04-13). ????, ????, 2011. ISBN 3-900051-13-5. iii + 62 pp. LCCN ???? URL http://

                                        cran.r-project.org/doc/
                                        manuals/R-lang.pdf.

RDCT:2011:WRE

[R D11g] R Development Core Team. Writing R Extensions: Version 2.13.0 (2011-04-13). ????, ????, 2011. ISBN 3-900051-11-9. iv + 167 pp. LCCN ???? URL http://

                                        cran.r-project.org/doc/
                                        manuals/R-exts.pdf.

Racine:2012:RPI

[Rac12] Jeffrey S. Racine. RStudio: A platform-independent IDE for R and Sweave. Journal of Applied Econometrics, 27(1): 167–172, January–February 2012. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Rahlf:2014:DRV

[Rah14] Thomas Rahlf. Datendesign mit R: 100 Visualisierungsbeispiele. Open Source Press, München, Germany, 2014. ISBN 3-95539-094-2. 426 pp. LCCN ????

Rahlf:2017:DVR

[Rah17] Thomas Rahlf. Data Visualisation with R: 100 Examples. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-49750-2 (hardcover), 3-319-49751-0 (e-book). xv + 385 pp. LCCN QA76.9.I52 R34 2017. URL http:

//link.springer.com/10.
1007/978-3-319-49751-8.

Rossi:2005:BSM

[RAM05] Peter E. (Peter Eric) Rossi, Greg M. (Greg Martin) Allenby, and Robert E. (Robert Edward) McCulloch. Bayesian statistics and marketing. Wiley, New York, NY, USA, 2005. ISBN 0-470-86367-6. x + 348 pp. LCCN HF5415.2 .R675 2005.

URL http://www.loc.gov/
catdir/enhancements/fy0622/
2005016418-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0514/2005016418.html;
 http://www.wiley.com/
WileyCDA/WileyTitle/productCd-
0470863676.html.

Rao:2014:BRR

[Rao14] T Subba Rao. Book review: Randall Douc, Eric Moulines and David S. Stoffer (2014) Nonlinear Time Series-Theory, Methods and Applications with R Examples. CRC Press, UK (A Chapman and Hall Book). Texts in Statistical Science. ISBN: 978-1-4665-0225-3 pages 531. Journal of Time Series Analysis, 35 (6):640–641, November 2014. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Robinson:2011:IBO

[RB11] Peter N. (Peter Nicholas) Robinson and Sebastian Bauer. Introduction to bioontologies. Chapman and Hall/CRC mathematical and computational biology series. Taylor and Francis, Boca Raton, FL, USA, 2011. ISBN 1-4398-3665-5 (hardcover). xxvii + 488 pp. LCCN QH324.2 .R618 2011.

Rigdon:2018:CSN

[RBB18] Joseph Rigdon, Michael Baiocchi, and Sanjay Basu. Code snippet: Near-far matching in R: The nearfar package. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v086c05;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v086c05/
                                        v86c05.pdf.

Riza:2015:FFR

[RBHB15] Lala Septem Riza, Christoph Bergmeir, Francisco Herrera, and José M. Benítez. frbs: Fuzzy rule-based systems for classification and regression in R. Journal of Statistical Software, 65 (6):??, June 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v65/i06.

Robert:2010:IMC

[RC10] Christian P. Robert and George Casella. Introducing Monte Carlo methods with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-1576-1 (paperback), 1-4419-1575-3 (paperback). xix + 283 pp. LCCN QA298 .R63 2010; QA298; Internet; QA298 .R63 2010eb.

Risser:2017:LLE

[RC17] Mark D. Risser and Catherine A. Calder. Local likelihood estimation for covariance functions with spatially-varying parameters: The convoSPAT package for R. Journal of Statistical Software, 81(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v081i14;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v081i14/
v81i14.pdf.

Rowlingson:1992:SSP

[RD92] B. S. Rowlingson and P. J.

Diggle. Splancs: Spatial point pattern analysis code in S-Plus. In P. Bresnahan, E. Corwin, and D. Cowen, editors, Proceedings: 5th International Symposium on Spatial Data Handling: August 3–7, 1992, Charleston, SC, USA, volume 2, pages 713–722. Humanities and Social Sciences Computing Lab, University of South Carolina, Charleston, SC, USA, 1992. ISBN 0-9633532-0-9 (vol. 1), 0-9633532-1-7 (vol. 2), 0-9633532-2-5 (set). Two volumes.

Rowlingson:1993:SSP

[RD93] B. S. Rowlingson and P. J. Diggle. Splancs: spatial point pattern analysis code in S-Plus. Computers and Geosciences, 19(5): 627–655, May 1993. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Recchia:2010:CSR

[Rec10] Anthony Recchia. Code snippet: R-squared measures for two-level hierarchical linear models using SAS. Journal of Statistical Software, 32(CS-2): ??, January 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/c02.

Reese:2009:BRBa

[Ree09] R. Allan Reese. Book review: R for SAS and SPSS Users, by R. A. Muenchen.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):697–698, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
20622537.

Reese:2018:BRG

[Ree18] R. Allan Reese. Book review: Graphics for Statistics and Data Analysis with R, 2nd edn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 181(4):1261–1262, October 2018. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Reese:2019:RGS

[Ree19] R. Allan Reese. R guides to statistics and graphics: [Getting Started with R: An Introduction for Biologists, Beckerman, Andrew, Childs, Dylan and Petchey, Owen; Oxford University Press; 256pp; February 2017; £22.99, $34.95; ISBN 978-0-19-878784-6. The New Statistics with R, Hector, Andy; Oxford University Press; 224pp; January 2015; £31.49, $49.95; ISBN 978-0-19-872906-8. R Graphics, third edition, Murrell, Paul; Chapman and Hall/CRC; 423pp; November 2018; £74.00, $99.95; ISBN 978-1-4987-8905-9]. Significance (Oxford, England), 16(2): 45, April 2019. CODEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Reimann:2008:SDA

[RFGD08] Clemens Reimann, Peter Filzmoser, Robert Garrett, and Rudolf Dutter, editors. Statistical data analysis explained: applied environmental statistics with R. Wiley, New York, NY, USA, 2008. ISBN 0-470-98581-X (hardcover), 0-470-98759-6 (e-book). xviii + 343 + 8 pp. LCCN GE45.S73 S8324 2008. URL http://

                                        www.wiley.com/WileyCDA/
                                        WileyTitle/productCd-0470987596.
                                        html.

Rodiger:2012:RCG

[RFKM12] Stefan Rödiger, Thomas Friedrichsmeier, Prasenjit Kapat, and Meik Michalke. RKWard: a comprehensive graphical user interface and integrated development environment for statistical analysis with R. Journal of Statistical Software, 49(9):??, June 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v49/i09.

Read:1996:BMU

[RG96] Edith Read and Jennifer Gough. Beyond mapping using GIS for national resource assessment and analysis. In Geography connects our world. Title: 1996 ESRI User Conference: proceedings, May 20–24, 1996, Palm Springs, CA. Environmental Systems Research Institute, Redlands, CA, USA, 1996. LCCN G70.212 .E87 1996.

Ritter:2007:ICD

[RG07] Christian Ritter and Jean Gilliard. ImpuR: a collection of diagnostic tools developed in R in the context of peak impurity detection in HPLC-DAD but potentially useful with other types of timeintensity matrices. Journal of Statistical Software, 18(9): 1–18, January 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v18/i09.

Roustant:2012:DDT

[RGD12] Olivier Roustant, David Ginsbourger, and Yves Deville. DiceKriging, DiceOptim: Two R packages for the analysis of computer experiments by kriging-based metamodeling and optimization. Journal of Statistical Software, 51(1):??, October 2012. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i01.

Racine:2002:URT

[RH02] Jeff Racine and Rob Hyndman. Using R to teach econometrics. Journal of Applied Econometrics, 17(2): 175–189, March–April 2002. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Ramsay:2009:FDA

[RHG09] J. O. (James O.) Ramsay, Giles Hooker, and Spencer Graves. Functional Data Analysis with R and MAT-LAB. Use R! Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-98184-5 (paperback), 0-387-98185-3 (e-book). xi + 207 pp. LCCN QA278 .R365 2009.

Raim:2021:STC

[RHW21] Andrew M. Raim, Scott H. Holan, and Christopher K. Wikle. Spatio-temporal change of support modeling with R. Computational Statistics, 36(1):749– 780, March 2021. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
020-01029-4.

Richardson:2011:SBRc

[Ric11] Alice Richardson. Short book review: Multiple Comparisons Using R by Frank Bretz, Torsten Hothorn, Peter Westfall. International Statistical Review = Revue Internationale de Statistique, 79(2):297, August 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
41305047.

Rieger:2019:BRM

[Rie19] Jonas Rieger. Book review:

Mónica Bécue–Bertaut (2019): Textual Data Science with R. Statistical Papers, 60(5): 1797–1798, October 2019. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://

link.springer.com/article/
10.1007/s00362-019-01126-
7.

Ripley:1987:SS

[Rip87] B. D. Ripley. Stochastic Simulation. Wiley series in probability and mathematical statistics. Applied probability and statistics. Wiley, New York, NY, USA, 1987. ISBN 0-471-81884-4. ISSN 0271-6356. xi + 237 pp. LCCN QA76.9.C65 R571 1987. US$29.95.

Ripley:2011:BRI

[Rip11] Brian D. Ripley. Book review: Introducing Monte Carlo Methods with R, by Robert, C. P. and Casella, G. Biometrics, 67(1): 328, March 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ritz:2009:NRR

[Rit09] Christian Ritz. Nonlinear regression with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-09615-9. xxiii + 252 pp. LCCN ????

Rizopoulos:2006:LRP

[Riz06] Dimitris Rizopoulos. ltm: An R package for latent variable modeling and item response analysis. Journal of Statistical Software, 17(5):1– 25, November 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v17/i05.

Rizzo:2008:SCR

[Riz08] Maria L. Rizzo. Statistical computing with R. Chapman and Hall/CRC computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-545-9. xvi + 399 pp. LCCN QA276.45.R3 R59 2008.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0740/
                                        2007034218-d.html.

Rizopoulos:2010:JRP

[Riz10] Dimitris Rizopoulos. JM: An R package for the joint modelling of longitudinal and time-to-event data. Journal of Statistical Software, 35(9):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/i09.

Rizopoulos:2012:JML

[Riz12] Dimitris Rizopoulos. Joint models for longitudinal and time-to-event data: with applications in R, volume 6 of Chapman and Hall/CRC biostatistics series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-7286-4 (hardcover). xiv + 261 pp. LCCN QA279 .R59 2012.

Rizopoulos:2016:RPJ

[Riz16] Dimitris Rizopoulos. The R package JMbayes for fitting joint models for longitudinal and time-to-event data using MCMC. Journal of Statistical Software, 72(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v072i07;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v072i07/
v72i07.pdf.

Ritter:2014:RPM

[RJH14] Stephan J. Ritter, Nicholas P.

Jewell, and Alan E. Hubbard. R package multiPIM: A causal inference approach to variable importance analysis. Journal of Statistical Software, 57 (8):??, April 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v57/i08.

Reddy:2020:SRP

[RK20] K. G. Reddy and M. G. M.

                                        Khan. stratifyR: An R

package for optimal stratification and sample allocation for univariate populations. Australian & New Zealand Journal of Statistics, 62(3): 383–405, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rasch:2011:SPU

[RKY11] Dieter Rasch, Klaus D. Kubinger, and Takuya Yanagida. Statistics in psychology using R and SPSS. Wiley, New York, NY, USA, 2011. ISBN 0-470-97124-X. ???? pp. LCCN BF39 .S7863 2011.

Rotolo:2015:BCM

[RL15] Daniele Rotolo and Loet Leydesdorff. Brief communications: Matching Medline/PubMed data with Web of Science: a routine in R language. Journal of the Association for Information Science and Technology, 66(10):2155–2159, October 2015. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Rinn:2016:LAR

[RLWP16] Philip Rinn, Pedro Lind, Matthias Wächter, and Joachim Peinke. The Langevin approach: An R package for modeling Markov processes. Journal of Open Research Software, 4 (1):e34–??, August 23, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.123/.

Rondeau:2012:FRP

[RMG12] Virginie Rondeau, Yassin Marzroui, and Juan R. Gon-

zalez. frailtypack: An

R package for the analysis of correlated survival data with frailty models using penalized likelihood estimation or parametrical estimation. Journal of Statistical Software, 47(4):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v47/
i04.

Rao:2017:BAU

[RN17] Umesh Hodeghatta Rao and Umesh Nayak. Business analytics using R — A practical approach. Apress, Berkeley, CA, USA, 2017. ISBN 1-4842-2513-9 (paperback), 1-4842-2514-7 (e-book). xvii + 280 pp. LCCN HF5548.2 .R284 2017.

Robinson:2005:BRB

[Rob05] Timothy J. Robinson. Book review: Linear Models with R by Julian J. Faraway. Technometrics, 47(3):372– 373, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        25471031.

Robinson:2007:BRG

[Rob07] Andrew Robinson. Book review: Generalized Additive Models: an Introduction in R, by Simon N. Wood. Australian & New Zealand Journal of Statistics, 49(4):455– 456, December 2007. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Robinson:2008:RB

[Rob08] Timothy J. Robinson. The R book. The American Statistician, 62(3):272–273, August 2008. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Robinson:2009:BRB

[Rob09] Andrew Robinson. Book review: Adaptive design theory and implementation using SAS and R. Journal of Applied Statistics, 36(6):701–702, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Robert:2012:SBRa

[Rob12] Christian P. Robert. Short book review: Handbook of Fitting Statistical Distributions with R by Zaven A. Karian, Edward J. Dudewicz. International Statistical Review = Revue Internationale de Statistique, 80(1):177– 178, April 2012. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
23257173.

Robert:2013:BRB

[Rob13] Christian Robert. Book review: R For Dummies. Chance, 26(4):61, 2013. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robinson:2017:BRE

[Rob17] Andrew Robinson. Book review: Extending R. John Chambers, Chapman and Hall/CRC, Boca Raton, 2016. No. of pages: 364. Price: 44.99 GBP. ISBN: 978-1-4987-7571-7. Statistics in Medicine, 36(11): 1839–1840, May 20, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robert:2018:TRC

[Rob18] Christian Robert. Testing R code. Chance, 31(1):60, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BRC

[Rob19] Christian Robert. Book review: Computational Methods for Numerical Analysis with R. Chance, 32(1):63, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2022:LBR

[Rob22] Christian Robert. Learning Base R (2nd edition). Chance, 35(2):46–47, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/full/10.1080/09332480.
                                        2022.2066419.

Robert:2023:BFF

[Rob23] Christian P. Robert. Bayes factors for forensic decision analyses with R. Chance, 36(2):32–33, 2023. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //www.tandfonline.com/
                                        doi/full/10.1080/09332480.
                                        2023.2203651.

Rohrl:2000:PPS

[Röh00] Armin Röhrl. Parallel processing in statistical computation: BSP, FPGas and MPI for the S-language. Thèse sciences, EPF Lausanne, Lausanne, Switzerland, 2000. 137 pp.

Romero:2007:IST

[Rom07] Martín Romero. Introduction to statistics through resampling methods and R/ S-Plus. Bulletin of Mathematical Biology, 69(2):795– 796, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11538-006-9120-x;
http://link.springer.
com/content/pdf/10.1007/
s11538-006-9120-x.pdf.

Rossini:2000:BRA

[Ros00] A. J. Rossini. Book review: Applied Smoothing Techniques for Data Analysis: The Kernel Approach with S-Plus Illustrations by Adrian W. Bowman and Adelchi Azzalini. Computational Statistics, 15(2):301– 302, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s001800000033.

Rosenthal:2007:ARI

[Ros07] Jeffrey S. Rosenthal. AM-CMC: an R interface for adaptive MCMC. Computational Statistics & Data Analysis, 51(12):5467–5470, August 15, 2007. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947307000916.

Rosenfeld:2008:ABG

[Ros08] Simon Rosenfeld. Approximate bivariate gamma generator with prespecified correlation and different marginal shapes. ACM Transactions on Modeling and Computer Simulation, 18(4): 16:1–16:??, September 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rosenblad:2009:BRJ

[Ros09] Andreas Rosenblad. Book review: J. J. Faraway: Extending the linear model with R: generalized linear, mixed effects and nonparametric regression models. Computational Statistics, 24(2):369– 370, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-009-0152-1.

Rosenbaum:2010:DOS

[Ros10a] Paul R. Rosenbaum. Design of Observational Studies, volume ?? of Springer Series in Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CO-DEN ???? ISBN 1-4419-1212-6 (hardcover), 1-4419-1213-4 (e-book). ISSN 0172-7397. xviii + 384 pp. LCCN QA279.R669 2010. URL http://

link.springer.com/book/
10.1007/978-1-4419-1213-
8.

Rosenblad:2010:SBRa

[Ros10b] Andreas Rosenblad. Short book review: Permutation Tests for Stochastic Ordering and ANOVA: Theory and Applications with R by Dario Basso, Fortunato Pesarin, Luigi Salmaso, Aldo Solari. International Statistical Review = Revue Internationale de Statistique, 78(3):446, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919864.

Rosseel:2012:LRP

[Ros12] Yves Rosseel. lavaan: An R package for structural equation modeling. Journal of Statistical Software, 48(2):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v48/i02.

Rovigatti:2017:PFE

[Rov17] Gabriele Rovigatti. Production function estimation in R: The prodest package. Journal of Open Source Software, 2(18):371:1, October 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00371.

Rasch:2011:OED

[RPVG11] Dieter Rasch, Jürgen Pilz, Rob Verdooren, and Albrecht Gebhardt. Optimal experimental design with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1697-2 (hardcover), 1-4398-1698-0 (e-book). xix + 325 pp. LCCN QA279 .O668 2011.

Roelstraete:2011:FRP

[RR11] Bjorn Roelstraete and Yves Rosseel. FIAR: An R package for analyzing functional integration in the brain. Journal of Statistical Software, 44(13): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v44/i13.

Roman-Roman:2012:RPE

[RRSPTR12] P. Román-Román, J. J. Serrano-Pérez, and F. Torres-Ruiz. An R package for an efficient approximation of first-passage-time densities for diffusion processes based on the FPTL function. Applied Mathematics and Computation, 218(17): 8408–8428, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300312001154.

Roman-Roman:2014:MGP

[RRSPTR14] P. Román-Román, J. J. Serrano-Pérez, and F. Torres-Ruiz. More general problems on first-passage times for diffusion processes: a new version of the fptdApprox R package. Applied Mathematics and Computation, 244(??):432–446, October 1, 2014. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300314009552.

Ritz:2005:BAU

[RS05] Christian Ritz and Jens C.

Streibig. Bioassay analysis using R. Journal of Statistical Software, 12(5): 1–22, ???? 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=112&url=v12/i05&
ct=2;   http://www.jstatsoft.
org/counter.php?id=112&
url=v12/i05/v12i05.pdf&
ct=1.

Ritz:2008:NRR

[RS08] Christian Ritz and Jens C.

Streibig. Nonlinear regression with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-09615-9 (paperback), 0-387-09616-7 (e-book). xi + 144 pp. LCCN QA278.2 .R577 2008.

Rossini:2007:SPS

[RTL07] A. J. Rossini, Luke Tierney, and Na Li. Simple parallel statistical computing in R. Journal of Computational and Graphical Statistics, 16(2):399–420, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rufibach:2009:CSR

[Ruf09] Kaspar Rufibach. Code snippet: reporttools: R functions to generate LaTeX tables of descriptive statistics. Journal of Statistical Software, 31 (CS-1):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v31/c01.

Ruiz:2016:BRM

[Rui16] Diego Andres Perez Ruiz. Book review: Modern Optimization with R. Journal of Statistical Software, 70(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v070b03;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v070b03/v70b03.pdf.

Ruiz:2017:BRS

[Rui17] Diego Andrés Pérez Ruiz. Book review: Statistical Rethinking: A Bayesian Course with Examples in R and Stan, Richard McElreath CRC Press, 2015, 469 pages, £67.99, hardcover ISBN: 978-1-4822-5344-3. International Statistical Review = Revue Internationale de Statistique, 85(2):379–380, August 2017. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Runnalls:2013:CER

[Run13] Andrew Runnalls. CXXR:

an extensible R interpreter. WIREs Computational Statistics, 5(3):181– 189, May/June 2013. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Ruppert:2004:SFI

[Rup04] David Ruppert. Statistics and finance: an introduction. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-20270-6. xx + 473 pp. LCCN HG176.5 .R87 2004.

URL http://www.loc.gov/
catdir/enhancements/fy0818/
2003063814-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0818/2003063814-t.html.

Ruppert:2011:SDA

[Rup11] David Ruppert. Statistics and data analysis for financial engineering. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-7786-4. xxii + 638 pp. LCCN HG176.7 .R87 2011.

URL http://www.loc.gov/
catdir/enhancements/fy1201/
2011282136-b.html;   http:

                                        //www.loc.gov/catdir/enhancements/
                                        fy1201/2011282136-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1201/2011282136-
                                        t.html.

Rusch:2015:BRR

[Rus15]  Thomas Rusch.  Book

review: R for Marketing Research and Analytics, Chris Chapman, Elea McDonnell Feit, Springer-Verlag, Switzerland, 2015. ISBN 978-3-319-14435-1. 454 pp. USD 64.99 (P).

                                        http://www.springer.com/book/9783319144351.

Journal of Statistical Software, 67(BR-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:

                                        //www.jstatsoft.org/index.
                                        php/jss/article/view/v067b02;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v067b02/v67b02.pdf.

Rasch:2020:DMM

[RV20] Dieter Rasch and Rob Verdooren. Determination of minimum and maximum experimental size in one-, two- and three-way ANOVA with fixed and mixed models by R. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL

                                        https://link.springer.
                                        com/article/10.1007/s42519-
                                        020-00088-6.

Roy:2023:ENW

[RVHG23] Dooti Roy, Gregory Vaughan, Jianan Hui, and Junxian Geng. An exploration of National Weather Service daily forecasts using R Shiny. Computational Statistics, 38 (3):1173–1191, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
023-01341-9.

Schumacker:2001:USC

[SA01] Randall E. Schumacker and Allen Akers. Understanding statistical concepts using Splus. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 2001. ISBN 0-8058-3623-3. xiii + 357 pp. LCCN QA276.4 .S366 2001.

Scalise:2015:UOS

[SA15] Kathleen Scalise and Diane D. Allen. Use of opensource software for adaptive measurement: Concerto as an R-based computer adaptive development and delivery platform. British Journal of Mathematical and Statistical Psychology, 68 (3):478–496, 2015. ISSN 0007-1102 (print), 2044-8317 (electronic).

Siddharth:2020:RPI

[SA20] R. Siddharth and G. Aghila.

RandPro — a practical implementation of random projection-based feature extraction for high dimensional multivariate data analysis in R. SoftwareX, 12 (??):Article 100629, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711020303423.

Sabbag:2019:BRH

[Sab19] Anelise G. Sabbag. Book review: Handbook of Educational Measurement and Psychometrics Using R, by Christopher D. Desjardins and Okan Bulut. Boca Raton, FL: Chapman & Hall/ CRC Press, 2018, xxiii + 302 pp., $79.95 (H), ISBN: 978-1-4987-7013-2. The American Statistician, 73 (4):415–416, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/00031305.
2019.1676110.

Soto:2023:RRP

[SADU23] Jonatan Soto, Johan Aparicio, Aquiles Darghan, and Milan Oldrich Urban. RankspeQ: an R package platform for genotype characterization and performance-ranking based on MultispeQ measurements. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002406.

Sahu:2022:BMS

[Sah22] Sujit K. Sahu. Bayesian modeling of spatio-temporal data with R. Chapman and Hall/CRC interdisciplinary statistics series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2022. ISBN 0-367-27798-0 (hardcover), 1-032-20957-7 (paperback), 1-000-54369-2 (e-book). xxii + 411 pp. LCCN QA278.2; QA279.5 .S34 2022; HA30.6 .S24 2022.

Sandry:2003:BRBb

[San03] Thomas D. Sandry. Book review: Introductory Statistics with R by Peter Dalgaard. Technometrics, 45 (3):274–275, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25047066.

Sanchez:2010:BRBb

[San10a] Juana Sanchez. Book review:

Data Analysis and Graphics Using R – An Example-Based Approach (3rd Edition). Journal of Statistical Software, 36(BR-6): ??, September 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v36/b06.

Sanchez:2010:BRBa

[San10b] Juana Sanchez. Book review: Introductory Time Series with R. Journal of Statistical Software, 32(BR-4):??, January 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v32/b04.

Sanchez:2011:BRB

[San11] Juana Sanchez. Book review: Hidden Markov Models for Time Series — An Introduction Using R. Journal of Statistical Software, 43(BR-4):??, October 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/b04.

Sanchez:2019:BRT

[Sán19] Brisa N. Sánchez. Book review: Textual data science with R, Bécue–Bertaut, Mónica, Boca Raton: CRC Press. Biometrics, 75(4): 1415–1416, December 2019. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sartori:2006:SRL

[Sar06]  Nicola Sartori.  Sweave

= R · L^{A}TEX^{2}. Technical report, Università “Ca’ Foscari” Venezia, Venice, Italy, November 23, 2006. 17 pp. URL http:

                                        //stat.epfl.ch/webdav/
                                        site/stat/shared/Regression/
                                        EPFL-Sweave-powerdot.pdf.

Sarkar:2008:LMD

[Sar08] Deepayan Sarkar. Lattice:

multivariate data visualization with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-75968-9 (paperback), 0-387-75969-7 (e-book). xvii + 265 pp. LCCN QA171.5 .S275 2008. URL http://

www.loc.gov/catdir/toc/
fy0805/2008920682.html.

Sevcikova:2011:BRP

[SAR11] Hana Sevcíková, Leontine Alkema, and Adrian Raftery. bayesTFR: An R package for probabilistic projections of the total fertility rate. Journal of Statistical Software, 43(1):??, July 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v43/i01.

Sarrias:2016:DCM

[Sar16] Mauricio Sarrias. Discrete choice models with random parameters in R: The Rchoice package. Journal of Statistical Software, 74 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v074i10;
https://www.jstatsoft.
org/index.php/jss/article/
view/v074i10/v74i10.pdf.

Sauter:2010:BRB

[Sau10] Roger M. Sauter. Book review: Data Manipulation With R (1st ed.) by Phil Spector. Technometrics, 52(3):365, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
27867263.

Sauter:2011:BRB

[Sau11] Roger M. Sauter. Book review: R for SAS and SPSS Users (1st ed) by Robert A. Muenchen. Technometrics, 53(1):101, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
40997301.

Sauter:2012:BRB

[Sau12] Roger M. Sauter. Book review: Data Mining With Rattle and R (1st ed.) by Graham Williams. Technometrics, 54(3):323–324, August 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714903.

Sauter:2021:BRA

[Sau21] Roger M. Sauter. Book review: Advanced R (2nd ed.). Technometrics, 62 (3):417, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Savitsky:2016:BNM

[Sav16] Terrance D. Savitsky. Bayesian nonparametric mixture estimation for time-indexed functional data in R. Journal of Statistical Software, 72(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v072i02;
https://www.jstatsoft.
org/index.php/jss/article/
view/v072i02/v72i02.pdf.

Sawitzki:2009:CSIa

[Saw09] Günther Sawitzki. Computational statistics: an introduction to R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-8678-2 (hardcover). xiv + 251 + 8 pp. LCCN QA276.45.R3 S29 2009. URL http:/

/catdir.loc.gov/catdir/
toc/fy0904/2008042298.
html.

Shin:2006:CSL

[SBMG06] Ji-Hyung Shin, Sigal Blay, Brad McNeney, and Jinko Graham. Code snippet: LDheatmap: An R function for graphical display of pairwise linkage disequilibria between single nucleotide polymorphisms. Journal of Statistical Software, 16(CS-3):

1–9, September 2006. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v16/c03.

Shoukri:2007:ACD

[SC07] M. M. (Mohamed M.) Shoukri and Mohammad A. Chaudhary. Analysis of correlated data with SAS and R. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2007. ISBN 1-58488-619-6. 295 pp. LCCN RA652.2.M3 S53

2007. URL http://www.
loc.gov/catdir/enhancements/
fy0733/2007000593-d.htm;
 http://www.loc.gov/catdir/
toc/ecip078/2007000593.
htm.

Sardy:2007:BRB

[SCD07] Sylvain Sardy and Valerie Chavez-Demoulin. Book review: Modeling Financial Time Series with S-PLUS (2nd Edition). Journal of Statistical Software, 17(BR-6):1–2, February 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v17/b06.

Schlittgen:2008:BRR

[Sch08] Rainer Schlittgen. Book review: Robert H. Shumway and David S. Stoffer: Time series analysis and its applications with R examples, 2nd edn. AStA. Advances in Statistical Analysis, 92(2): 233–234, May 2008. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10182-
008-0064-3.

Schwarz:2009:BRM

[Sch09] Carl James Schwarz. Book review: Mixed Effects Models and Extensions in Ecology with R, by Zuur, A. F., Ieno, E. N., Walker, N. J., Saveliev, A. A., and Smith, G. M. Biometrics, 65(3): 992–993, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scheipl:2011:SBV

[Sch11] Fabian Scheipl. spikeSlabGAM:

Bayesian variable selection, model choice and regularization for generalized additive mixed models in R. Journal of Statistical Software, 43(14):??, September 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v43/
i14.

Schiesser:2014:ICM

[Sch14] William E. Schiesser. An introductory comparison of Matlab and R: Applications to ordinary and partial differential equations. Report, Lehigh University, Bethlehem, PA 18105, USA, January 15, 2014. 48

pp.  URL http://www.
lehigh.edu/~wes1/R_PDE/
R-Matlab_compare.pdf.

Scholz:2016:RPC

[Sch16] Michael Scholz. R pack-

                                        age clickstream: Analyz-

ing clickstream data with Markov chains. Journal of Statistical Software, 74 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v074i04;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v074i04/v74i04.pdf.

Schiesser:2017:SCM

[Sch17a] William E. Schiesser. Spline Collocation Methods for Partial Differential Equations: with Applications in R. Wiley, New York, NY, USA, 2017. ISBN 1-119-30103-3 (hardcover), 1-119-30105-X (PDF), 1-119-30104-1 (ePub), 1-119-30106-8 (online). xv + 549 pp. LCCN QA377 .S355 2017.

Schmuller:2017:SAR

[Sch17b] Joseph Schmuller. Statistical analysis with R for dummies. For dummies. Wiley, New York, NY, USA, 2017. ISBN 1-119-33726-7 (paperback), 1-119-33709-7 (e-book), 1-119-33726-7 (ebook). xiii + 438 pp. LCCN QA276.45.R3 S386 2017.

                                        URL http://proquest.
                                        safaribooksonline.com/
                                        9781119337065.

Schwabe:2017:PBR

[Sch17c] Inga Schwabe. BayesTwin:

An R package for Bayesian inference of item-level twin data. Journal of Open Research Software, 5(1):33– ??, November 14, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.185/.

Schulman:1995:DAU

[SCK95] Daniel A. Schulman, Alec D.

Campbell, and Eric C. Kostello. Data analysis using S-Plus. Sociological methods and research, 23(3):392– ??, February 1995. ISSN 0049-1241 (print), 1552-8294 (electronic).

Skinner:2023:DDP

[SCK^{+}23] Grace Skinner, Rob Cooke, Junghyuk Keum, Andy Purvis, Cristina Raw, Ben A. Woodcock, and Joseph Millard. Dynameta: a dynamic platform for ecological meta-analyses in R Shiny. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023001358.

Soetaert:2012:SDE

[SCM12] Karline Soetaert, Jeff Cash, and Francesca Mazzia. Solving Differential Equations in R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-28069-2 (hardcover), 3-642-28070-6 (e-book). xvi + 248 pp. LCCN QA371.5.D37 S64 2012.

Scott:2009:SBRa

[Sco09] David J. Scott. Short book review: Applied Econometrics with R by Christian Kleiber, Achim Zeileis. International Statistical Review = Revue Internationale de Statistique, 77 (1):164, April 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919715.

Scott:2010:SBRb

[Sco10a] David Scott. Short book review: Introducing Monte Carlo Methods with R by Christian P. Robert, George Casella. International Statistical Review = Revue Internationale de Statistique, 78 (3):476–477, December 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919891.

Scott:2010:SBRa

[Sco10b] David J. Scott. Short book review: R Through Excel: A Spreadsheet Interface for Statistics, Data Analysis, and Graphics by Richard M. Heiberger, Erich Neuwirth. International Statistical Review = Revue Internationale de Statistique, 78(2):324, August 2010. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919849.

Scott:2011:BRI

[Sco11] David Scott. Book review:

Introduction to Scientific Programming and Simulation Using R, by O. Jones, R. Maillardet and A. Robinson. Australian & New Zealand Journal of Statistics, 53(2): 264–265, June 2011. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Scott:2012:SBR

[Sco12] David Scott. Short book review: Mathematical Statistics with Resampling and R by Laura M. Chihara, Tim C. Hesterberg. International Statistical Review = Revue Internationale de Statistique, 80(3):472, December 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:2013:SBRc

[Sco13a] David Scott. Short book review: Customer and Business Analytics: Applied Data Mining for Business Decision Making Using R by Daniel S. Putler, Robert E. Krider. International Statistical Review = Revue Internationale de Statistique, 81(2):328, August 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scott:2013:SBRa

[Sco13b] David Scott. Short book review: Statistical Computing in C++ and R by Randall L. Eubank, Ana Kupresanin. International Statistical Review = Revue Internationale de Statistique, 81(1):159–160, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Scrucca:2013:GPG

[Scr13] Luca Scrucca. GA: a package for genetic algorithms in R. Journal of Statistical Software, 53(4):??, April 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v53/
                                        i04.

Slaets:2013:WFD

[SCS13] Leen Slaets, Gerda Claeskens, and Bernard W. Silverman. Warping functional data in R and C via a Bayesian multiresolution approach. Journal of Statistical Software, 55 (3):??, October 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v55/i03.

Scutari:2010:LBN

[Scu10] Marco Scutari. Learning Bayesian networks with the bnlearn R package. Journal of Statistical Software, 35(3):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/i03.

Scutari:2015:BNE

[SD15] Marco Scutari and Jean-Baptiste Denis. Bayesian Networks: with Examples in R. Texts in statistical science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2015. ISBN 1-4822-2558-1. xv + 225 pp. LCCN QA279.5 .S38 2015. URL http:/

/catalogue.bnf.fr/ark:/
12148/cb43850800h.

Sousa:2018:PBR

[SD18] David N. Sousa and João R.

Daniel. bor: An R package for transforming behavioral observation records into data matrices. Journal of Open Source Software, 3(28):909:1–909:4, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00909.

Sax:2018:SAX

[SE18] Christoph Sax and Dirk Eddelbuettel. Seasonal adjust-

ment by X-13ARIMA-SEATS

in R. Journal of Statistical Software, 87(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v087i11;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v087i11/
v87i11.pdf.

Sekhon:2011:MPS

[Sek11] Jasjeet S. Sekhon. Multivariate and propensity score matching software with automated balance optimization: The matching package for R. Journal of Statistical Software, 42(7): ??, June 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v42/i07.

Selvin:1998:MAB

[Sel98] Steve Selvin. Modern applied biostatistical methods using S-Plus, volume 28 of Monographs in epidemiology and biostatistics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-512025-6. xiv + 461 pp. LCCN QH323.5 .S448 1998.

Selig:2017:BRM

[Sel17]  Katharina Selig.  Book

review: Munzert, Simon, Rubba, Christian, Meissner, Peter, and Nyhuis, Dominic. Automated Data Collection with R: a Practical Guide to Web Scraping and Text Mining. Hoboken: Wiley. Biometrics, 73(4):1469, December 2017. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sengupta:2014:BRBa

[Sen14] Arindam Sengupta. Book review: A Tiny Handbook on R by Mike Allerhand. International Statistical Review = Revue Internationale de Statistique, 82 (1):156, April 2014. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Saldana:2018:SRP

[SF18] Diego Franco Saldana and Yang Feng. SIS: An R package for sure independence screening in ultrahighdimensional statistical models. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v083i02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v083i02/
v83i02.pdf.

Santos-Fernandez:2012:MRP

[SFS12] Edgar Santos-Fernández and Michele Scagliarini. MPCI: An R package for computing multivariate process capability indices. Journal of Statistical Software, 47 (7):??, April 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v47/i07.

Su:2011:MID

[SGHY11] Yu-Sung Su, Andrew Gelman, Jennifer Hill, and Masanao Yajima. Multiple imputation with diagnostics (mi) in R: Opening windows into the black box. Journal of Statistical Software, 45 (2):??, December 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v45/i02.

Schrenk:2022:QRP

[SGPSzC22] Hannah Schrenk, Carlos Garcia-Perez, Nico Schreiber, and Wolfgang zu Castell. QtAC: an R-package for analyzing complex systems development in the framework of the adaptive cycle metaphor. Ecological Modelling, 466(??):Article 109860, April 2022. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0304380021003987.

Saul:2017:RIS

[SH17] Bradley C. Saul and Michael G. Hudgens. A recipe for inferference: Start with causal inference. Add interference. Mix well with R. Journal of Statistical Software, 82(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v082i02;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v082i02/
v82i02.pdf.

Shang:2011:BRB

[Sha11] Han Lin Shang. Book review: Dynamic linear models with R. Journal of Applied Statistics, 38(10): 2369–2370, October 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [PPC09].

Shang:2012:GSD

[Sha12] Han Lin Shang. Graphics for statistics and data analysis with R. Journal of Applied Statistics, 39 (8):1843–1844, August 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Shalabh:2016:BRS

[Sha16] Shalabh. Book reviews: Statistical Analysis and Data Display: an Intermediate Course with Examples in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 179(4):1129– 1130, October 2016. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shanmugam:2019:BRAa

[Sha19] Ramalingam Shanmugam. Book review: Applied compositional data analysis: with worked examples in R. Journal of Statistical Computation and Simulation, 89(16):3165, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:MAE

[Sha20] Ramalingam Shanmugam. Multivariate analysis of ecological data with ade4. Journal of Statistical Computation and Simulation, 90(6): 1150–1151, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shalabh:2021:BRI

[Sha21a] Shalabh. Book review: Interactive Web-Based Data Visualization with R, plotly, and shiny. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(3):1150, July 2021. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2021:BRR

[Sha21b] Shalabh. Book review: R Markdown Cookbook. Journal of the Royal Statistical Society. Series A (Statistics in Society), 184(4):1613, October 2021. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shanmugam:2021:BRM

[Sha21c] Ramalingam Shanmugam.

Book review: Multilevel modeling using R, by W. Holmes Finch, Jocelyn E. Bolin and Ken Kelley, 6000 Broken Sound Parkway NW, Suite 300, Boca Raton, FL 33487, CRC Press, 2014, xiii + 216 pp., $50.99 (paperback), ISBN 13: 978-1-4665-1585-7. Journal of Statistical Computation and Simulation, 91(1):216–217, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2021:OAS

[Sha21d] Ramalingam Shanmugam.

Omic association studies with R and Bioconductor. Journal of Statistical Computation and Simulation, 91 (3):648–649, 2021. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shalabh:2022:BRP

[Sha22a]  Shalabh.  Book review:

Probability, Statistics, and Data: a Fresh Approach Using R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S776–S777, December 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/185/Supplement_
2/S776/7069519.

Shalabh:2022:MDS

[Sha22b] Shalabh. Modern data science with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):735–736, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shalabh:2022:UBM

[Sha22c] Shalabh. Univariate, bivariate and multivariate statistics using R: Quantitative tools for data analysis and data science. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(2):736–737, April 2022. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Shanmugam:2023:BRBa

[Sha23a] Ramalingam Shanmugam. Book review: Bayesian modeling of spatio-temporal data with R by Sujit K. Sahu, 2022, Chapman and Hall, CRC Press, 411 + xxii pp., $74.21 (hardback), ISBN: 978-0-367-27798-7, 6000 Broken Sound Parkway, NW, (Suite 300), Boca Raton, FL 33487. Journal of Statistical Computation and Simulation, 93 (7):1224, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2023:BRBc

[Sha23b] Ramalingam Shanmugam.

Book review: Book on stochastic processes with R: an introduction by Olga Korosteleva. (2022). Stochastic Processes with R: an Introduction, Boca Raton, FL: Chapman and Hall, CRC Press, 2022, x + 190 pp., $99.95 (hardback), ISBN 978-1-032-15373-5., pp. 190 + ix, $ 99.95 (hardback). Journal of Statistical Computation and Simulation, 93 (10):1667, 2023. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheather:2009:MAR

[She09] Simon J. Sheather. A Modern Approach to Regression with R. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-09607-8 (hardcover). xiv + 392 pp. LCCN QA278.2 .S496 2009.

Shentu:2011:BRB

[She11] Yue Shentu. Book review:

Clinical Trial Data Analysis Using R. Journal of Statistical Software, 43(BR-1):??, July 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v43/b01.

Shipley:2016:CCB

[Shi16] Bill Shipley. Cause and correlation in biology: a user’s guide to path analysis, structural equations, and causal inference with R. Cambridge University Press, Cambridge, UK, second edition, 2016. ISBN 1-107-44259-1 (paperback), 1-139-97957-4 (e-book). xiii + 299 pp. LCCN QH323.5 .S477 2016.

Sparks:2017:PGG

[SHN17] Adam H. Sparks, Tomislav Hengl, and Andrew Nelson. GSODR: Global summary daily weather data in R. Journal of Open Source Software, 2(10): 177:1–177:2, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00177.

Shotwell:2013:PRP

[Sho13] Matthew S. Shotwell. profdpm: An R package for MAP estimation in a class of conjugate product partition models. Journal of Statistical Software, 53(8): ??, May 2013. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v53/i08.

Shoukri:2018:ACD

[Sho18] M. M. (Mohamed M.)

Shoukri. Analysis of Correlated Data with SAS and R. Chapman and Hall/CRC, Boca Raton, FL, USA, fourth edition edition, 2018. ISBN 1-138-19745-9 (hardcover), 1-315-27770-0 (Mobipocket e-book), 1-315-27771-9 (ePub ebook), 1-315-27772-7 (PDF e-book), 1-315-27773-5 (ebook). xv + 497 pp. LCCN RA652.2.M3 .S53 2018.

Struyf:1997:IRC

[SHR97] Anja Struyf, Mia Hubert, and Peter J. Rousseeuw. Integrating robust clustering techniques in S-PLUS. Computational Statistics & Data Analysis, 26(1):17–37, November 6, 1997. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947397000200.

Scharpf:2011:URP

[SIR^{+}11] Robert B. Scharpf, Rafael A.

Irizarry, Matthew E. Ritchie, Benilton Carvalho, and Ingo Ruczinski. Using the R package crlmm for genotyping and copy number estimation. Journal of Statistical Software, 40(12):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v40/
i12.

Sisson:2003:BRBc

[Sis03] Scott Sisson. Book review: The Basics of S-PLUS, by A. Krause; M. Olsen. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):413–414, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        4128227.

Siu:2017:PRP

[Siu17]  Chapman Siu.  Ramble:

Parser combinator for R. Journal of Open Source Software, 2(11):209:1, March 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00209.

Schep:2017:PII

[SK17] Alicia N. Schep and Sarah K. Kummerfeld. iheatmapr: Interactive complex heatmaps in R. Journal of Open Source Software, 2(16): 359:1, August 2017. CO-DEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/
                                        10.21105/joss.00359.

Sharma:2022:DRP

[SKD22] Parichit Sharma, Hasan Kurban, and Mehmet Dalkilic. DCEM: an R package for clustering big data via data-centric modification of Expectation Maximization. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001771.

Sigrist:2015:PRP

[SKS15] Fabio Sigrist, Hans R.

Künsch, and Werner A. Stahel. spate: An R package for spatio-temporal modeling with a stochastic advectiondiffusion process. Journal of Statistical Software, 63 (14):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i14.

Stegmaier:2017:PGR

[SKW17] Philip Stegmaier, Alexander Kel, and Edgar Win-

gender. geneXplainR: An

R interface for the geneXplain platform. Journal of Open Source Software, 2(18):412:1–412:3, October 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00412.

Stoll:2005:WRS

[SKZ05] Heather Stoll, Gary King, and Langche Zeng. WhatIf: R software for evaluating counterfactuals. Journal of Statistical Software, 15(4):1– 18, December 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i04.

Seefeld:2005:RB

[SL05] Kim Seefeld and Ernst Linder. R for Bioinformatics. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00544-X. 320 (est.) pp. LCCN ????

Sabatti:2009:BRB

[SL09] Chiara Sabatti and Christophe Lalanne. Book review: Applied Statistical Genetics with R for Population-Based Association Studies. Journal of Statistical Software, 31 (BR-2):??, September 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v31/
                                        b02.

Sturtz:2005:RPR

[SLG05] Sibylle Sturtz, Uwe Ligges, and Andrew Gelman. R2WinBUGS: a package for running Win-BUGS from R. Journal of Statistical Software, 12(3): 1–16, ???? 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.
                                        php?id=110&url=v12/i03&
                                        ct=2;   http://www.jstatsoft.
                                        org/counter.php?id=110&
                                        url=v12/i03/v12i03.pdf&
                                        ct=1.

Schoonees:2016:FGA

[SlRC16] Pieter Schoonees, Niël le Roux, and Roelof Coetzer. Flexible graphical assessment of experimental designs in R: The vdg package. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v074i03;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v074i03/
v74i03.pdf.

Snellenburg:2012:GJB

[SLS^{+}12] Joris J. Snellenburg, Sergey Laptenok, Ralf Seger, Katharine M. Mullen, and Ivo H. M. van Stokkum. Glotaran: a Javabased graphical user interface for the R package TIMP. Journal of Statistical Software, 49(3):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v49/
i03.

Scherer:2005:IMP

[SM05] Bernd Michael Scherer and R. Douglas Martin. Introduction to modern portfolio optimization with NuOPT and S-PLUS. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21016-4. xxi + 405 pp. LCCN HG4529.5 .S325 2005.

Stubben:2007:EAD

[SM07] Chris Stubben and Brook Milligan. Estimating and analyzing demographic models using the popbio package in R. Journal of Statistical Software, 22(11):1– 23, September 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v22/i11.

Skrabanek:2021:AFR

[SM21] Pavel Skrabánek and Natália Martínková. Algorithm 1017: fuzzyreg: an R package for fitting fuzzy regression models. ACM Transactions on Mathematical Software, 47(3):29:1– 29:18, June 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3451389.

Schmidberger:2009:SAP

[SME^{+}09] Markus Schmidberger, Martin Morgan, Dirk Eddelbuettel, Hao Yu, Luke Tierney, and Ulrich Mansmann. State of the art in parallel computing with R. Journal of Statistical Software, 31 (1):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v31/i01.

Sauerbrei:2006:MRM

[SMHBR06] W. Sauerbrei, C. Meier-Hirmer, A. Benner, and P. Royston. Multivariable regression model building by using fractional polynomials: Description of SAS, STATA and R programs. Computational Statistics & Data Analysis, 50(12):3464– 3485, August 2006. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947305001623.

Smith:2006:BRB

[Smi06] Eric P. Smith. Book review: Using R for Introductory Statistics by John Verzani. Technometrics, 48(2):307–308, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471174.

Smith:2007:BRP

[Smi07] Brian J. Smith. boa: An R package for MCMC output convergence assessment and posterior inference. Journal of Statistical Software, 21 (11):1–37, November 2007. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v21/
i11.

Smith:2017:PUP

[Smi17] Andrew P. Smith. UKCensusAPI:

Python and R interfaces to the nomisweb UK census data API. Journal of Open Source Software, 2 (19):408:1, November 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00408.

Smith:2018:PHC

[Smi18] Andrew P. Smith. humanleague: a C++ microsynthesis package with R and Python interfaces. Journal of Open Source Software, 3 (25):629:1, May 2018. CO-DEN ???? ISSN 2475-9066. URL http://

                                        joss.theoj.org/papers/
                                        10.21105/joss.00629.

Schlather:2015:ASP

[SMM^{+}15] Martin Schlather, Alexander Malinowski, Peter J. Menck, Marco Oesting, and Kirstin Strokorb. Analysis, simulation and prediction of multivariate random fields with package

                                        RandomFields. Journal of

Statistical Software, 63(8): ??, February 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v63/i08.

Smith:2022:ICC

[SMM^{+}22] Matthew J. Smith, Mohammad A. Mansournia, Camille Maringe, Paul N. Zivich, Stephen R. Cole, Clémence Leyrat, Aurélien Belot, Bernard Rachet, and Miguel A. Luque-Fernandez. Introduction to computational causal inference using reproducible Stata, R, and Python code: a tutorial. Statistics in Medicine, 41(2): 407–432, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simoneau:2020:ODT

[SMWP20] Gabrielle Simoneau, Erica E. M. Moodie, Michael P. Wallace, and Robert W. Platt. Optimal dynamic treatment regimes with survival endpoints: introducing DWSurv in the R package DTRreg. Journal of Statistical Computation and Simulation, 90(16):2991–3008, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sheffield:2018:PSR

[SNR18] Nathan Sheffield, V. P. Nagraj, and Vince Reuter.

simpleCache: R caching

for reproducible, distributed, large-scale projects. Journal of Open Source Software, 3(21):463:1, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00463.

Schutt:2013:DDS

[SO13] Rachel Schutt and Cathy O’Neil. Doing Data Science. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2013. ISBN 1-4493-5865-9 (paperback). xxiv + 375 pp. LCCN QA76.9.D343 S27 2013.

Solvang:2023:TRP

[SO23] Hiroko Kato Solvang and Mineaki Ohishi. trec: an R package for trend estimation and classification to support integrated ecosystem assessment of the marine ecosystem and environmental factors. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000055.

Salazar:2016:SGE

[SOD^{+}16] Alejandro Salazar, Begoña Ojeda, María Dueñas, Fernando Fernández, and Inmaculada Failde. Simple generalized estimating equations (GEEs) and weighted generalized estimating equations (WGEEs) in longitudinal studies with dropouts: guidelines and implementation in R. Statistics in Medicine, 35(19):3424–3448, August 30, 2016. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Soetaert:2010:BRB

[Soe10] Karline Soetaert. Book review: A Primer of Ecology with R. Journal of Statistical Software, 32(BR-3):??, January 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/b03.

Solymos:2009:PED

[Sól09] Péter Sólymos. Processing ecological data in R with the mefa package. Journal of Statistical Software, 29 (8):??, February 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/i08.

Solymos:2010:BRP

[Sól10] Péter Sólymos. Book review: A Primer of Ecology with R, by Stevens, M. H. H. Biometrics, 66(4): 1309–1310, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:2005:JES

[SP05] Christina D. Smith and Jeffrey S. Pontius. Jackknife estimator of species richness with S-PLUS. Journal of Statistical Software, 15(3):1– 12, December 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v15/i03.

Soetaert:2010:IMS

[SP10] Karline Soetaert and Thomas Petzoldt. Inverse modelling, sensitivity and Monte Carlo analysis in R using package FME. Journal of Statistical Software, 33(3): ??, February 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v33/i03.

Savitsky:2014:BSN

[SP14a] Terrance Savitsky and Susan Paddock. Bayesian semi- and non-parametric models for longitudinal data with multiple membership effects in R. Journal of Statistical Software, 57(3): ??, March 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v57/i03.

Sridharan:2014:PRC

[SP14b] Shriram Sridharan and Jignesh M. Patel. Profiling R on a contemporary processor. Proceedings of the VLDB Endowment, 8(2): 173–184, October 2014. CO-DEN ???? ISSN 2150-8097.

Sparks:2017:PGU

[Spa17] Adam H. Sparks. getCRUCLdata: Use and explore CRU CL v. 2.0 climatology elements in R. Journal of Open Source Software, 2(12):230:1, April 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00230.

Spector:1994:IP

[Spe94] Phil Spector. An Introduction to S and S-Plus. Duxbury Press, Belmont, CA, USA, 1994. ISBN 0-534-19866-X. xiii + 286 pp. LCCN QA76.73.S15 S63 1994. US$28.95.

Spector:2008:DMR

[Spe08] Phil Spector. Data Manipulation With R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-74730-3. ix + 152 pp. LCCN QA276.45.R3.S64; QA276.45.R3.S64 2008.

Spedicato:2013:LPP

[Spe13] Giorgio Alfredo Spedicato.

The lifecontingencies

package: Performing financial and actuarial mathematics calculations in R. Journal of Statistical Software, 55(10):??, November 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v55/
i10.

Sparks:2017:PBF

[SPPP17] Adam H. Sparks, Mark Padgham, Hugh Parsonage, and Keith Pembleton. bomrang: Fetch Australian Government Bureau of Meteorology data in R. Journal of Open Source Software, 2 (17):411:1–411:2, September 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00411.

Soetaert:2010:SDE

[SPS10] Karline Soetaert, Thomas Petzoldt, and R. Woodrow Setzer. Solving differential equations in R: Package deSolve. Journal of Statistical Software, 33(9): ??, February 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v33/i09.

Stasinopoulos:2007:GAM

[SR07] D. Mikis Stasinopoulos and Robert A. Rigby. Generalized additive models for location scale and shape (GAMLSS) in R. Journal of Statistical Software, 23(7): 1–46, December 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v23/i07.

Silge:2016:PTT

[SR16] Julia Silge and David Robinson. tidytext: Text mining and analysis using tidy data principles in R. Journal of Open Source Software, 1(3):37:1–37:3, July 2016. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00037.

Subbiah:2017:PCR

[SR17] M. Subbiah and V. Rajeswaran. proportion: a comprehensive R package for inference on single binomial proportion and Bayesian computations. SoftwareX, 6(??):30–35, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300018.

Scrucca:2018:CPI

[SR18] Luca Scrucca and Adrian E.

Raftery. clustvarsel: A

package implementing variable selection for Gaussian model-based clustering in R. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v084i01;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v084i01/
v84i01.pdf.

Sibuya:1992:DAU

[SS92] M. Sibuya and R. Shibata.

Data Analysis by Using S. Kyouritsu Syppan, ????, Japan, 1992. ISBN ???? ???? pp. LCCN ????

Shumway:2006:TSA

[SS06] Robert H. Shumway and David S. Stoffer. Time series analysis and its applications: with R examples. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 0-387-29317-5 (hardcover). xiii + 575 pp. LCCN QA280 .S585 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0663/
2005935284-d.html;   http:
//www.loc.gov/catdir/toc/
fy0704/2005935284.html.

Shumway:2011:TSA

[SS11] Robert H. Shumway and David S. Stoffer. Time series analysis and its applications: with R examples. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2011. ISBN 1-4419-7865-8 (e-book). xi + 596 pp. LCCN QA280 .S58

2011. URL http://rave.
ohiolink.edu/ebooks/ebc/
978144197865.

SouzadeCursi:2015:UQS

[SS15] Eduardo Souza de Cursi and Rubens Sampaio. Uncertainty Quantification and Stochastic Modeling with Matlab. ISTE Press Ltd, London, UK, 2015. ISBN 0-08-100471-0 (e-book), 1-78548-005-7. LCCN QA274.2.

URL http://alltitles.
ebrary.com/Doc?id=11040161;
 http://lib.myilibrary.
com?id=762946;   http:/
/public.eblib.com/choice/
PublicFullRecord.aspx?
p=2007484;   http://www.
sciencedirect.com/science/
book/9781785480058.

Salehi:2018:NISa

[SS18a] Younes Salehi and William E. Schiesser. Numerical integration of space fractional partial differential equations. Volume 1, Introduction to algorithms and computer coding in R, volume 19 of Synthesis lectures on mathematics and statistics. Morgan and Claypool Publishers, San Rafael, CA, USA, 2018. ISBN 1-68173-207-6 (paperback), 1-68173-208-4 (e-book). ISSN 1938-1743 (print), 1938-1751 (electronic). xii + 189 pp. LCCN QA372 .S266 2018.

Salehi:2018:NISb

[SS18b] Younes Salehi and William E.

Schiesser. Numerical integration of space fractional partial differential equations. Volume 2, Applications from classical integer PDEs, volume 20 of Synthesis lectures on mathematics and statistics. Morgan and Claypool Publishers, San Rafael, CA, USA, 2018. ISBN 1-68173-209-2 (hardcover), 1-68173-210-6 (e-book). ISSN 1938-1743 (print), 1938-1751 (electronic). xii + 183–375 pp. LCCN QA372 .S2662 2018.

Shou:2019:CRP

[SS19a] Yiyun Shou and Michael

Smithson.  cdfquantreg:

An R package for CDFquantile regression. Journal of Statistical Software, 88(??):??, ???? 2019. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v088i01;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v088i01/v88i01.pdf.

Shumway:2019:TSD

[SS19b] Robert Shumway and David Stoffer. Time Series: a Data Analysis Approach Using R. Chapman and Hall/CRC, Boca Raton, FL, USA, 2019. ISBN 0-367-22109-8 (hardcover), 0-429-27328-2 (e-book), 1-000-00156-3 (ebook). xii + 259 pp. LCCN QA280 .S586 2019.

Schon:2020:DRP

[SSH^{+}20] Marian Schön, Jakob Simeth, Paul Heinrich, Franziska Görtler, Stefan Solbrig, Tilo Wettig, Peter J. Oefner, Michael Altenbuchinger, and Rainer Spang. DTD: An R package for digital tissue deconvolution. Journal of Computational Biology, 27(3):386–389, March 2020. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL https://

                                        www.liebertpub.com/doi/
                                        abs/10.1089/cmb.2019.0469;
                                         https://www.liebertpub.
                                        com/doi/pdf/10.1089/cmb.
                                        2019.0469.

Subirana:2014:BBT

[SSV14] Isaac Subirana, Héctor Sanz, and Joan Vila. Building bivariate tables: The

                                        compareGroups package for

R. Journal of Statistical Software, 57(12):??, May 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:

                                        //www.jstatsoft.org/v57/
                                        i12.

Suess:2010:IPS

[ST10] Eric A. Suess and Bruce E.

Trumbo. Introduction to probability simulation and Gibbs sampling with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-40273-X (paperback). xiii + 307 pp. LCCN QA276.6 .S84 2010.

MathSoft:1992:PPM

[Sta92a] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Programmer’s Manual, version 3.1 edition, October 1992. various pp.

MathSoft:1992:PUM

[Sta92b] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS User’s Manual Supplement, version 3.1 edition, October 1992. various pp.

MathSoft:1993:CCP

[Sta93a] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. A Crash Course in S-PLUS, version 3.2 edition, December 1993. various pp.

StatSci:1993:CCP

[Sta93b] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. A Crash Course in S-PLUS, December 1993. various pp.

MathSoft:1993:GIP

[Sta93c] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA.

A Gentle Introduction to S-PLUS, version 3.2 edition, December 1993. various pp.

StatSci:1993:GIP

[Sta93d] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. A Gentle Introduction to S-PLUS, Version 3.2, December 1993. various pp.

MathSoft:1993:PGS

[Sta93e] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Guide to Statistical and Mathematical Analysis, version 3.2 edition, December 1993. various pp.

StatSci:1993:PGS

[Sta93f] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Guide to Statistical and Mathematical Analysis, Version 3.2, December 1993. various pp.

MathSoft:1993:IMM

[Sta93g] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Installation and Maintenance Manual, version 3.2 edition, December 1993. various pp.

MathSoft:1993:PPM

[Sta93h] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Programmer’s Manual, version 3.2 edition, December 1993. various pp.

StatSci:1993:PPM

[Sta93i] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Programmer’s Manual, Version 3.2, December 1993. various pp.

MathSoft:1993:PUM

[Sta93j] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS User’s Manual, version 3.2 edition, December 1993. various pp.

StatSci:1993:PUM

[Sta93k] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS User’s Manual, Version 3.2, December 1993. various pp.

StatSci:1995:S

[Sta95a] Statistical Sciences, Inc. S-Plus, 1995.

StatSci:1995:RMF

[Sta95b] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. Read Me First, Version 3.3, April 1995. ii + 24 pp.

StatSci:1995:PGI

[Sta95c] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Global Index, April 1995. various pp.

StatSci:1995:PIM

[Sta95d] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Installation and Maintenance Manual, Version 3.3, April 1995. various pp.

StatSci:1995:PTD

[Sta95e] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Trellis Displays User’s Manual, Version 3.3, April 1995. various pp.

StatSci:1995:SUM

[Sta95f] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS user’s manual, version 3.3. edition, September 1995. various pp.

StatSci:1995:PVS

[Sta95g] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. S-PLUS Version 3.3 Supplement, April 1995. various pp.

StatSci:1997:SGS

[Sta97a] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. SPlus-4 Guide to Statistics, July 1997. xviii + 875 pp. Order number R620203.

StatSci:1997:SPG

[Sta97b] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. SPlus-4 Programmer’s Guide, August 1997. xvi + 17–581 pp. Order number R620201.

StatSci:1997:SUG

[Sta97c] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA.

SPlus-4 User’s Guide, August 1997. xiv + 15–620 pp. Order number R620199.

StatSci:1998:SGS

[Sta98a] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. SPlus-5 Guide to Statistics, 1998. xxii + 1014 pp. Order number R620291.

StatSci:1998:SUG

[Sta98b] StatSci, a Division of Math-Soft, Inc., Seattle, WA, USA. SPlus-5 User’s Guide, 1998. xvi + 336 pp. Order number R620293.

StatSci:1999:MSP

[Sta99] StatSci. Math Soft S-PLUS 4.5. StatSci, a Division of MathSoft, Inc., Seattle, WA, USA, 4.5 (student edition) edition, 1999. ISBN 0-534-35873-X. Includes CD-ROM.

Stare:2005:BRS

[Sta05] Janez Stare. Book review:

3. Survival analysis using S: Analysis of time to event data. Mara Tableman and Jong Sung Kim with a contribution from Stephen Portnoy, Chapman and Hall/ CRC, Boca Raton, 2004. No. of pages: xv + 280. Price: $69.95. ISBN 1-58488-408-8. Statistics in Medicine, 24(2): 327–328, January 30, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Starkings:2007:SBRb

[Sta07a] Susan Starkings. Short book review: The R Book by Michael J. Crawley. International Statistical Review = Revue Internationale de Statistique, 75 (3):425–426, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
41509895.

Stasinopoulos:2007:BRG

[Sta07b] Mikis Stasinopoulos. Book review: Generalized Additive Models: an Introduction with R, by S. N. Wood. Biometrics, 63(4): 1298–1299, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staniswalis:2008:IMC

[Sta08] Joan G. Staniswalis. Incorporating marginal covariate information in a nonparametric regression model for a sample of R × C tables. Biometrics, 64(4): 1054–1061, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stapleton:2009:LSM

[Sta09] James H. Stapleton. Linear statistical models. Wiley series in probability and statistics. Wiley, New York, NY, USA, second edition, 2009.

ISBN 0-470-23146-7 (cloth). xii + 474 pp. LCCN QA279 .S695 2009. URL http://

www.wiley.com/WileyCDA/
WileyTitle/productCd-0470231467.
html.

Stadtlander:2021:BRA

[Sta21] Christian T. K.-H. Stadtländer.

Book review: Analysis of correlated data with SAS and R (4th ed.), by Mohamed M. Shoukri, Boca Raton, FL, CRC Press, 2018, 497 + xv pp., $139.95 (hardback), ISBN: 978-1-138-19745-9. Journal of Statistical Computation and Simulation, 91 (7):1478–1479, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stevens:2000:CPP

[Ste00] Al Stevens. C programming:

The S programming language. Dr. Dobb’s Journal of Software Tools, 25(2):110– 113, February 2000. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/ftp/2000/2000_
02/cprog220.txt. See com-

ment on another S language [KSC^{+}00].

Stevens:2009:PER

[Ste09] M. Henry H. (Martin Henry Hoffman) Stevens. A Primer of Ecology with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-89881-6 (softcover), 0-387-89882-4 (e-but). xvi + 401 pp. LCCN QH541.15.S5 S74 2009.

Steibel:2011:BRS

[Ste11] Juan P. Steibel. Book review: SAS and R Data Management, Statistical Analysis, and Graphics, by Kleinman, K. and Horton, N. Biometrics, 68(3):1179–1180, September 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stenborg:2023:SRP

[Ste23] Travis Stenborg. Styling R ggplot2 graphics with L^{A}TEX. TUGboat, 44 (1):108–109, 2023. CO-DEN ???? ISSN 0896-

                                        3207.  URL https://
                                        tug.org/TUGboat/tb44-1/
                                        tb136stenborg-ggplot2.
                                        pdf.

Stenborg:2024:SAT

[Ste24]  Travis Stenborg.  Semi-

automated TikZ directed acyclic graphs in R. TUGboat, 45(1):115–116, 2024. CODEN ???? ISSN 0896-3207. URL https://

                                        tug.org/TUGboat/tb45-1/
                                        tb139stenborg-dags.pdf.

Stoyanov:2011:BRBa

[Sto11] Jordan Stoyanov. Book review: Simulation and Inference for Stochastic Differential Equations: with R Examples, by S. M. Iacus. Journal of the Royal Statistical Society. Series A (Statistics in Society), 174 (2):509, April 2011. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

//www.jstor.org/stable/
23014417.

Strobl:2010:BRB

[Str10] Carolin Strobl. Book review:

Advances in Social Science Research Using R. Journal of Statistical Software, 34 (BR-2):??, April 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v34/b02.

Su:2007:FSM

[Su07]  Steve Su.  Fitting sin-

gle and mixture of generalized lambda distributions to data via discretized and maximum likelihood methods: GLDEX in R. Journal of Statistical Software, 21 (9):1–17, October 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v21/i09.

Suchard:2007:BRA

[Suc07] Marc A. Suchard. Book review: Analysis of Phylogenetics and Evolution with R, by E. Paradis. Biometrics, 63(4):1304–1305, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sund:2015:BRB

[Sun15]  Reijo Sund.  Book re-

view: Multistate Analysis of Life Histories with R. International Statistical Review = Revue Internationale de Statistique, 83 (2):324–325, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Sund:2016:BRG

[Sun16]  Reijo Sund.  Book re-

view: Graphical Data Analysis with R. International Statistical Review = Revue Internationale de Statistique, 84(2):324–325, August 2016. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Suselbeck:1993:PEP

[Süs93] Benno Süselbeck. S und S-PLUS: Eine Einführung in Programmierung und Anwendung. Fischer, Stuttgart, Germany, 1993. ISBN 3-437-40232-3. xii + 602 pp. LCCN ????

Sievert:2019:EGL

[SVC^{+}19] Carson Sievert, Susan VanderPlas, Jun Cai, Kevin Ferris, Faizan Uddin Fahad Khan, and Toby Dylan Hocking. Extending ggplot2 for linked and animated Web graphics. Journal of Computational and Graphical Statistics, 28(2):299–308, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stander:2018:BSA

[SVCB18] Julian Stander, Luciana Dalla Valle, and Mario Cortina-Borja. A Bayesian survival analysis of a historical dataset: How long do popes live? The American Statistician, 72(4):368–375, 2018. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://

amstat.tandfonline.com/
doi/full/10.1080/00031305.
2017.1328374.

Sofrygin:2017:SRP

[SvdLN17] Oleg Sofrygin, Mark J.

van der Laan, and Romain Neugebauer. simcausal R package: Conducting transparent and reproducible simulation studies of causal effect estimation with complex longitudinal data. Journal of Statistical Software, 81(??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v081i02;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v081i02/v81i02.pdf.

Salmon:2017:PRR

[SVM^{+}17] Maëlle Salmon, Sreekanth Vakacherla, Carles Milà, Julian D. Marshall, and Cathryn Tonne. rtimicropem: an R package supporting the analysis of RTI MicroPEM output files. Journal of Open Source Software, 2(16):333:1, August 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00333.

Sestelo:2017:NRP

[SVMMRP17] Marta Sestelo, Nora M. Villanueva, Luis Meira-Machado, and Javier Roca-Pardiñas. npregfast: An R package for nonparametric estimation and inference in life sciences. Journal of Statistical Software, 82 (??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v082i12;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v082i12/v82i12.pdf.

Sharaf:2020:BRP

[SWCP20] Taysseer Sharaf, Theren Williams, Abdallah Chehade, and Keshav Pokhrel. BLNN: an R package for training neural networks using Bayesian inference. SoftwareX, 11(??):Article 100432, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/S235271101930322X.

Anonymous:2013:FPE

[Swe13] Nathan G. Swenson. Functional and phylogenetic ecology in R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-9541-5. xii + 212 + 1 pp. LCCN ????

URL http://www.loc.gov/
catdir/enhancements/fy1407/
2013958447-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1407/2013958447-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1407/2013958447-
t.html.

Stadler:2017:ORL

[SWHJ17] Lukas Stadler, Adam Welc, Christian Humer, and Mick Jordan. Optimizing R language execution via aggressive speculation. ACM SIG-PLAN Notices, 52(2):84–95, February 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:2008:UGM

[SYC08] Brian J. Smith, Jun Yan, and Mary Kathryn Cowles. Unified geostatistical modeling for data fusion and spatial heteroskedasticity with R package ramps. Journal of Statistical Software, 25 (10):1–21, April 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v25/i10.

Scheike:2011:ACR

[SZ11] Thomas H. Scheike and Mei-Jie Zhang. Analyzing competing risk data using the R timereg package. Journal of Statistical Software, 38 (2):??, January 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v38/i02.

Taeger:2014:SHT

[Tae14] Dirk Taeger. Statistical hypothesis testing with SAS and R. Wiley, New York, NY, USA, 2014. ISBN 1-119-95021-X (hardcover), 1-118-76260-6 (PDF), 1-118-76258-4, 1-118-76261-4 (ePub), 1-306-32264-2. 308 (est.) pp. LCCN QA277.

Tagusari:2023:SRP

[Tag23] Junta Tagusari. shinyHugePlot: an R package for displaying the features of data with a huge sample size. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023002790.

Takezawa:2012:GRG

[Tak12] Kunio Takezawa. Guidebook to R graphics using Microsoft Windows. Wiley, New York, NY, USA, 2012. ISBN 1-118-02639-X (paperback). 263 pp. LCCN T385 .T346 2012. URL http:

                                        //catalogimages.wiley.
                                        com/images/db/jimages/
                                        9781118026397.jpg.

Tang:2018:PAR

[Tan18] Yuan Tang. autoplotly:

An R package for automatic generation of interactive visualizations for statistical results. Journal of Open Source Software, 3 (24):657:1–657:2, April 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00657.

Tattar:2018:BRS

[Tat18] Prabhanjan Tattar. Book review: Scutari, Marco and Denis, Jean-Baptiste. Bayesian Networks: With Examples in R. Boca Raton, CRC Press. Biometrics, 74(1):379–380, March 2018. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taveras:2017:REU

[Tav17] John L. Taveras. R for Excel users: an introduction to R for Excel analysts. CreateSpace Independent Publishing Platform, North Charleston, SC, USA, 2017. ISBN 1-5005-6635-7 (paperback). v + 196 pp. LCCN QA276.45.R3 T38 2017.

Taylor:2017:RPA

[TB17] Julian Taylor and David Butler. R package ASMap: Efficient genetic linkage map construction and diagnosis. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v079i06;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v079i06/
v79i06.pdf.

Thioulouse:2007:IMD

[TD07] Jean Thioulouse and Stéphane Dray. Interactive multivariate data analysis in R with

the ade4 and ade4TkGUI

packages. Journal of Statistical Software, 22(5):1– 14, September 2007. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v22/i05.

Taylor:2013:LRP

[TDRD13] Benjamin M. Taylor, Tilman M. Davies, Barry S. Rowlingson, and Peter J. Diggle. lgcp: An R package for inference with spatial and spatio-temporal log-gaussian Cox processes. Journal of Statistical Software, 52(4): ??, February 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v52/i04.

Taylor:2015:BID

[TDRD15] Benjamin M. Taylor, Tilman M. Davies, Barry S. Rowlingson, and Peter J. Diggle. Bayesian inference and data augmentation schemes for spatial, spatiotemporal and multivariate log-Gaussian Cox processes in R. Journal of Statistical Software, 63 (7):??, February 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v63/i07.

Teetor:2011:RC

[Tee11]  Paul Teetor.  R cook-

book. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 0-596-80915-

8. xviii + 413 pp. LCCN QA276.45.R3 T44 2011.

URL http://www.loc.gov/
catdir/enhancements/fy1113/
2011282768-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy1113/2011282768-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1113/2011282768-
t.html.

Teimouri:2022:BRP

[Tei22] Mahdi Teimouri. bccp: an R package for life-testing and survival analysis. Computational Statistics, 37(1): 469–489, March 2022. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
021-01129-9.

TempleLang:1997:MTE

[Tem97] Duncan Walter Temple Lang. A Multi Threaded Extension to a High Level Interactive Statistical Computing Environment. Ph.D. dissertation, University of California, Berkeley, Berkeley, CA, USA, 1997. vii +

                                        161 pp. URL https://www.
                                        proquest.com/docview/304343859/

.

Tennekes:2018:TTM

[Ten18] Martijn Tennekes. tmap: Thematic maps in R. Journal of Statistical Software, 84(??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v084i06;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v084i06/v84i06.pdf.

Turner:2012:BTM

[TF12] Heather Turner and David Firth. Bradley–Terry models in R: The BradleyTerry2 package. Journal of Statistical Software, 48(9): ??, May 2012. CODEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v48/i09.

Theussl:2012:TPD

[TFH12] Stefan Theußl, Ingo Feinerer, and Kurt Hornik. A tm plug-in for distributed text mining in R. Journal of Statistical Software, 51(5): ??, November 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v51/i05.

Thieler:2016:RRP

[TFR16] Anita M. Thieler, Roland Fried, and Jonathan Rathjens. RobPer: An R package to calculate periodograms for light curves based on robust regression. Journal of Statistical Software, 69(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v069i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v069i09/
v69i09.pdf.

Touraine:2017:SRP

[TGJ17] Célia Touraine, Thomas A.

Gerds, and Pierre Joly. SmoothHazard: An R package for fitting regression models to interval-censored observations of illness-death models. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v079i07;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v079i07/
v79i07.pdf.

Turcotte:2020:DTR

[TGKV20] Alexi Turcotte, Aviral Goel, Filip Krikava, and Jan Vitek. Designing types for R, empirically. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):181:1–181:25, November 2020. URL

                                        https://dl.acm.org/doi/
                                        10.1145/3428249.

DDJStaff:1999:NVJ

[The99] The DDJ Staff. News and views: Just say no to carpal tunnel; ACM award for S creator; cruising with Perl; you can’t judge an open Ebook …; H1-B visas in short supply; we interrupt you with this message…; great

                                        timing; www.waitandsee.

org. Dr. Dobb’s Journal of Software Tools, 24 (8):18, August 1999. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.

ddj.com/. John Chambers was awarded the ACM Software System Award for creating the S System for statistical computing.

Thiele:2014:RMN

[Thi14] Jan C. Thiele. R marries NetLogo: Introduction to the RNetLogo package. Journal of Statistical Software, 58(2):??, June 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v58/i02.

Thieme:2018:RG

[Thi18] Nick Thieme. R generation. Significance (Oxford, England), 15(4):14– 19, August 2018. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (elec-

tronic). URL https://rss.
onlinelibrary.wiley.com/
doi/full/10.1111/j.1740-
9713.2018.01169.x.

Thorley:2018:PCR

[Tho18] Joe Thorley. checkr: An R package for assertive programming. Journal of Open Source Software, 3 (23):624:1, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00624.

Tierney:2009:CAP

[Tie09] Luke Tierney. Code analysis and parallelizing vector operations in R. Computational Statistics, 24(2):217– 223, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-008-0117-9.

Tilke:1996:BRBa

[Til96] Clemens Tilke. Book review: A handbook of statistical analyses using S-Plus: Brian S. Everitt (1994): Chapman & Hall, 143 p., £14.99, ISBN 0-412-56310-X. Computational Statistics & Data Analysis, 21 (5):606, May 1996. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167947396901395.

Tanimura:2006:PSM

[TKM06] Susumu Tanimura, Chusi Kuroiwa, and Tsutomu Mizota. Proportional symbol mapping in R. Journal of Statistical Software, 15(5): 1–7, January 2006. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v15/i05.

Templ:2015:SDC

[TKM15] Matthias Templ, Alexander Kowarik, and Bernhard Meindl. Statistical disclosure control for microdata using the R package sdcMicro. Journal of Statistical Software, 67 (4):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v067i04;
                                         http://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v067i04/v67i04.pdf.

Teisseyre:2016:RSM

[TKM16] Paweł Teisseyre, Robert A. Kłopotek, and Jan Mielniczuk. Random subspace method for high-dimensional regression with the R package regRSM. Computational Statistics, 31(3):943–972, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-016-0658-2;

http://link.springer.
com/content/pdf/10.1007/
s00180-016-0658-2.pdf.

Theussl:2011:PCR

[TLH11] Stefan Theußl, Uwe Ligges, and Kurt Hornik. Prospects and challenges in R package development. Computational Statistics, 26(3):395–404, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-010-0205-5.

Terpstra:2005:RBA

[TM05] Jeff T. Terpstra and Joseph W.

McKean. Rank-based analysis of linear models using R. Journal of Statistical Software, 14(7):1– 26, July 1, 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=132&url=v14/i07&
ct=2;   http://www.jstatsoft.
org/counter.php?id=132&
url=v14/i07/v14i07.pdf&
ct=1.

Tang:2015:DRP

[TM15] Xinyu Tang and Maria Melguizo. DTR: An R package for estimation and comparison of survival outcomes of dynamic treatment. Journal of Statistical Software, 65(7):??, June 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v65/i07.

Taketomi:2022:MSR

[TMCE22] Nanami Taketomi, Hirofumi Michimae, Yuan-Tsung Chang, and Takeshi Emura.

meta.shrinkage: an R

package for meta-analyses for simultaneously estimating individual means. Algorithms (Basel), 15(1), January 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/
15/1/26.

Templ:2017:SSC

[TMKD17] Matthias Templ, Bernhard Meindl, Alexander Kowarik, and Olivier Dupriez. Simulation of synthetic complex data: The R package simPop. Journal of Statistical Software, 79(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v079i10;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v079i10/
v79i10.pdf.

Tendeiro:2016:PRP

[TMN16] Jorge N. Tendeiro, Rob R. Meijer, and A. Susan M. Niessen. PerFit: An R package for person-fit analysis in IRT. Journal of Statistical Software, 74(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/

view/v074i05;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v074i05/
v74i05.pdf.

Tarr:2018:MRP

[TMW18] Garth Tarr, Samuel Müller, and Alan H. Welsh. mplot: An R package for graphical model stability and variable selection procedures. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v083i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v083i09/
v83i09.pdf.

Teimouri:2022:MRP

[TN22] Mahdi Teimouri and Saralees Nadarajah. MPS: an R package for modelling shifted families of distributions. Australian & New Zealand Journal of Statistics, 64(1):86–108, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Totterdell:2017:BHM

[TNM17] James A. Totterdell, Darfiana Nur, and Kerrie L. Mengersen. Bayesian hidden Markov models in DNA sequence segmentation using R: the case of Simian Vacuolating virus (SV40). Journal of Statistical Computation and Simulation, 87(14): 2799–2827, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Todd:2022:BRS

[Tod22] James Todd. Book review: Supervised Machine Learning for Text Analysis in R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2): S770, December 2022. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/185/Supplement_
2/S770/7069551.

Tolvi:2003:BRBc

[Tol03] Jussi Tolvi. Book review: Modeling Financial Time Series with S-Plus, by E. Zivot; J. Wang. Journal of the Royal Statistical Society. Series D (The Statistician), 52(4):705, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
4128162.

Toller:2023:BRB

[Tol23] Owen Toller. Book review: Basic statistics with R by Stephen C. Loftus, pp 283, £57.95 (paper), ISBN 978-0-12-820788-8, Academic Press/Elsevier (2021). The Mathematical Gazette, 107 (569):379–380, July 2023.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/
journals/mathematical-
gazette/article/basic-
statistics-with-r-by-stephen-
c-loftus-pp-283-5795-paper-
isbn-9780128207888-academic-
presselsevier-2021/0BF5A0142A9544DF2FCCFB018F7415A3.

Touloumis:2015:RPP

[Tou15] Anestis Touloumis. R package multgee: A generalized estimating equations solver for multinomial responses. Journal of Statistical Software, 64(8):??, March 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v64/
i08.

Tabelow:2011:SPM

[TP11] Karsten Tabelow and Jörg Polzehl. Statistical parametric maps for functional MRI experiments in R: The package fmri. Journal of Statistical Software, 44(11): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v44/i11.

Tilahun:2007:FSM

[TPAM07] Abel Tilahun, Assam Pryseley, Ariel Alonso, and Geert Molenberghs. Flexible surrogate marker evaluation from several randomized clinical trials with continuous endpoints, using R and SAS. Computational Statistics & Data Analysis, 51(9): 4152–4163, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167947307000679.

Thomas:2014:CST

[TR14] Laine Thomas and Eric M. Reyes. Code snippet: Tutorial: Survival estimation for Cox regression models with time-varying coefficients using SAS and R. Journal of Statistical Software, 61(CS-1):??, October 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v61/c01.

Tierney:2009:SPC

[TRL09] Luke Tierney, A. J. Rossini, and Na Li. Snow: a parallel computing framework for the R system. International Journal of Parallel Programming, 37(1):78– 90, February 2009. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=
                                        37&issue=1&spage=78.

Tattar:2016:CSR

[TRM16a] Prabhanjan Tattar, Suresh Ramaiah, and B. G. Manjunath. A course in statistics with R. Wiley, New York, NY, USA, 2016. ISBN 1-119-15272-0 (hardcover). xxvi + 665 pp. LCCN QA276.45.R3 T38 2016.

Tattar:2016:AOS

[TRM16b] Prabhanjan Narayanachar Tattar, Suresh Ramaiah, and B. G. Manjunath. Appendix A: Open Source Software — An Epilogue, pages 627–630. Wiley, New York, NY, USA, 2016. ISBN 1-119-15274-7.

Trosset:2009:ISI

[Tro09] Michael W. (Michael William)

Trosset. An Introduction to Statistical Inference and Its Applications with R. Chapman and Hall/CRC texts in statistical science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-58488-947-0 (hardback). xxvii + 467 pp. LCCN QA276 .T756 2009.

Tsay:2010:AFT

[Tsa10] Ruey S. Tsay. Analysis of financial time series. Wiley series in probability and statistics. Wiley, New York, NY, USA, third edition, 2010. ISBN 0-470-41435-9 (hardcover), 0-470-64456-7 (e-book), 0-470-64455-9 (e-book). xxiii + 677 pp. LCCN HA30.3 T76 2010eb.

URL http://proquest.
safaribooksonline.com/?
fpi=9781118017098;   http:
//public.eblib.com/EBLPublic/

                                        PublicView.do?ptiID=565117;
                                         http://site.ebrary.com/
                                        id/10411605; http://www.
                                        myilibrary.com?id=270783;
                                         http://www.myilibrary.
                                        com?id=270783\%26ref=toc.

Tsay:2013:IAF

[Tsa13] Ruey S. Tsay. An introduction to analysis of financial data with R. Wiley, New York, NY, USA, 2013. ISBN 0-470-89081-9 (hardcover). xiv + 390 pp. LCCN HG106 .T76 2013.

Tsagris:2014:BRB

[Tsa14a] Michail Tsagris. Book review: Statistics through resampling methods and R, second edition. Journal of Applied Statistics, 41(12): 2780–2781, December 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Tsay:2014:MTS

[Tsa14b] Ruey S. Tsay. Multivariate time series analysis: with R and financial applications. Wiley series in probability and statistics. Wiley, New York, NY, USA, first edition, 2014. ISBN 1-118-61790-8 (hardcover), 1-118-61775-4 (e-book), 1-118-61776-2 (e-book), 1-118-61779-7 (ebook), 1-306-25450-7 (MyiLibrary). xvii + 492 pp. LCCN QA280 .T73 2013.

Thomas:2021:LHM

[TT21] Samuel Thomas and Wanzhu Tu. Learning Hamiltonian Monte Carlo in R. The American Statistician, 75 (4):403–413, 2021. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.tandfonline.com/
doi/full/10.1080/00031305.
2020.1865198.

Tu:2021:BRS

[Tu21] Yu-Kang Tu. Book review:

Structural equation modeling with partial least squares using Stata and R by Mehmet Mehmetoglu and Sergio Venturini Boca Raton, FL, USA : Chapman and Hall/CRC, 2021. Hard cover. pp. 382. £99.99. Biometrics, 77(3): 1130–1131, September 2021. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1977:EDA

[Tuk77] John W. Tukey. Exploratory data analysis. Addison-Wes-ley, Reading, MA, USA, 1977. ISBN 0-201-07616-0. xvi + 688 pp. LCCN HA29 .T916 1977.

Turner:2012:BRS

[Tur12] Robin Turner. Book review:

SAS and R: Data management, statistical analysis, and graphics. Statistics in Medicine, 31(13): 1405–1406, June 15, 2012.

CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turner:2018:PQR

[Tur18] Stephen D. Turner. qqman: an R package for visualizing GWAS results using Q-Q and Manhattan plots. Journal of Open Source Software, 3 (25):731:1–731:2, May 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00731.

Tusell:2005:BRB

[Tus05] Fernando Tusell. Book review: Linear Models with R. Journal of the Royal Statistical Society. Series A (Statistics in Society), 168 (3):632, July 2005. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http:

                                        //www.jstor.org/stable/
                                        3559846.

Tusell:2011:KFR

[Tus11] Fernando Tusell. Kalman filtering in R. Journal of Statistical Software, 39 (2):??, March 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v39/i02.

Taylor:2011:RPW

[TV11] Julian Taylor and Arunas Verbyla. R package wgaim: QTL analysis in bi-parental populations using linear mixed models. Journal of Statistical Software, 40 (7):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/i07.

Turner:2020:MRR

[TvEK20] Heather L. Turner, Jacob van Etten, and Ioannis Kosmidis. Modelling rankings

in R: the PlackettLuce

package. Computational Statistics, 35(3):1027–1057, September 2020. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL

https://link.springer.
com/article/10.1007/s00180-
020-00959-3.

Tabelow:2011:SVM

[TW11] Karsten Tabelow and Brandon Whitcher. Special volume on magnetic resonance imaging in R. Journal of Statistical Software, 44(1): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v44/i01.

Tingley:2014:PRP

[TYH^{+}14] Dustin Tingley, Teppei Yamamoto, Kentaro Hirose, Luke Keele, and Kosuke Imai. mediation: R package for causal mediation analysis. Journal of Statistical Software, 59(5):??, September 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v59/i05.

Tyler:2007:BRB

[Tyl07]  David E. Tyler.  Book

review: Robust Statistical Methods with R. Journal of the American Statistical Association, 102(478):759– 760, June 2007. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tyner:2016:URP

[Tyn16]  Sam Tyner.  Using the

R package geomnet: Visualizing trans-Atlantic slave trade of Africans, 1514–1866. Chance, 29(3):4–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Umlauf:2015:SAR

[UAK^{+}15] Nikolaus Umlauf, Daniel Adler, Thomas Kneib, Stefan Lang, and Achim Zeileis. Structured additive regression models: An R interface to BayesX. Journal of Statistical Software, 63(21): ??, February 2015. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v63/i21.

Urdinez:2021:RPD

[UC21] Francisco Urdinez and Andrýes Cruz, editors. R for Political Data Science: a Practical Guide. A Chapman et Hall book. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2021. ISBN 0-367-81889-2 (hardcover), 1-000-20447-2 (e-book), 1-000-20449-9 (Mobipocket), 1-000-20451-0 (ePub), 1-003-01062-8 (ebook). xix + 439 pp. LCCN JA71.7 .R22 2021.

Unlu:2009:FSR

[ÜKD09] Ali Ünlü, Thomas Kiefer, and Ehtibar N. Dzhafarov. Fechnerian scaling in R: The package fechner. Journal of Statistical Software, 31 (6):??, August 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v31/i06.

Ugarte:2008:PSR

[UMA08] María Dolores Ugarte, Ana F.

Militino, and Alan T. Arnholt. Probability and statistics with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2008. ISBN 1-58488-891-1. xxvi + 700 pp. LCCN QA273.19.E4 U53

2008. URL http://www.
crcpress.com/product/isbn/
9781584888918.

Unwin:2011:SBRa

[Unw11]  Antony Unwin.  Short

book review: Graphics for Statistics and Data Analysis with R by Kevin J. Keen. International Statistical Review = Revue Internationale de Statistique, 79(1):114, April 2011. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        41306167.

Unwin:2012:SBRb

[Unw12a] Antony Unwin. Short book review: An Introduction to Applied Multivariate Analysis with R by Brian Everitt and Torsten Hothorn. International Statistical Review = Revue Internationale de Statistique, 80(2):331– 332, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2012:SBRa

[Unw12b] Antony Unwin. Short book review: R Graphics, Second Edition, by Paul Murrell. International Statistical Review = Revue Internationale de Statistique, 80 (2):330–331, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2013:SBRb

[Unw13a] Antony Unwin. Short book review: Discovering Statistics Using R by Andy Field, Jeremy Miles, Zoë Field. International Statistical Review = Revue Internationale de Statistique, 81(1):169– 170, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2013:SBRa

[Unw13b] Antony Unwin. Short book review: R for Statistics by Pierre-André Cornillon, Arnaud Guyader, François Husson, Nicolas Jégou, Julie Josse, Maela Kloareg, Eric Matzner-Løber, Laurent Rouvière. International Statistical Review = Revue Internationale de Statistique, 81(1):168–169, April 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Unwin:2015:GDA

[Unw15] Antony Unwin. Graphical Data Analysis with R. Chapman and Hall/CRC the R series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4987-1523-0 (hardcover). xiii + 296 pp. LCCN QA76.9.I52 U59 2015.

Urbanek:2009:HTS

[Urb09] Simon Urbanek. How to talk to strangers: ways to leverage connectivity between R, Java and Objective C. Computational Statistics, 24(2): 303–311, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-008-0132-x.

Unlu:2010:DRP

[ÜS10] Ali Ünlü and Anatol Sargin. DAKS: An R package for data analysis methods in knowledge space theory. Journal of Statistical Software, 37(2): ??, November 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v37/i02.

Uschner:2018:RRP

[USHH18] Diane Uschner, David Schindler, Ralf-Dieter Hilgers, and Nicole Heussen. randomizeR: An R package for the assessment and implementation of randomization in clinical trials. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i08;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085i08/
v85i08.pdf.

Utlaut:2005:BRB

[Utl05] Theresa L. Utlaut. Book review: Statistical Analysis and Data Display: An Intermediate Course with Examples in S-PLUS. R, and SAS by Richard M. Heiberger; Burt Holland. Technometrics, 47(3):371, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471029.

Valiente:2009:CPM

[Val09] Gabriel Valiente. Combinatorial pattern matching algorithms in computational biology using Perl and R. Chapman and Hall/CRC mathematical and computational biology series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6973-X (hardcover), 1-4200-6974-8 (e-book). 352 pp. LCCN QH324.2 .V35 2009.

Vantas:2018:PHR

[Van18] Konstantinos Vantas. hydroscoper:

R interface to the Greek National Data Bank for Hydrological and Meteorological Information. Journal of Open Source Software, 3(23):625:1–625:2, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00625.

Varadhan:2014:NOR

[Var14] Ravi Varadhan. Numerical optimization in R: Beyond optim. Journal of Statistical Software, 60(1): ??, September 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v60/i01.

vanBuuren:2011:MMI

[vBGO11] Stef van Buuren and Karin Groothuis-Oudshoorn. mice: Multivariate Imputation by Chained Equations in R. Journal of Statistical Software, 45(3):??, December 2011. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v45/
i03.

Vidoni:2023:TTR

[VC23] Melina Vidoni and Zadia Codabux. Towards a taxonomy of Roxygen documentation in R packages. Empirical Software Engineering, 28(4):??, July 2023. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (electronic). URL

https://link.springer.
com/article/10.1007/s10664-
023-10345-4.

vanderArk:2007:MSA

[vdA07] L. Andries van der Ark. Mokken scale analysis in R. Journal of Statistical Software, 20(11):1–19, February 2007. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v20/i11.

vanderArk:2012:NDM

[vdA12] L. Andries van der Ark. New developments in Mokken scale analysis in R. Journal of Statistical Software, 48(5):??, May 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v48/i05.

Vinod:2009:MEB

[VdL09] Hrishikesh D. Vinod and Javier Lopez de Lacalle. Maximum entropy bootstrap for time series: The meboot R package. Journal of Statistical Software, 29(5): ??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/i05.

vanderWal:2011:IRP

[vdWG11] Willem M. van der Wal and Ronald B. Geskus. ipw: An R package for inverse probability weighting. Journal of Statistical Software, 43(13): ??, September 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v43/i13.

VanBerkel:2023:PRP

[VEF^{+}23] Derek Van Berkel, Thomas Estabrook, Nathan Fox, Rahul Agrawal Bejarano, Lisa Maillard, Devin Gill, Erica Akemi Goto, and Maria Carmen Lemos. PPGISr: an R package for Public Participatory GIS. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000857.

Vehkalahti:2013:SBRb

[Veh13] Kimmo Vehkalahti. Short book review: Statistics in Psychology Using R and SPSS by Dieter Rasch, Klaus Kubinger, Takuya Yanagida. International Statistical Review = Revue Internationale de Statistique, 81 (3):459–460, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Venables:2004:IRN

[Ven04] W. N. (William N.) Venables, editor. An introduction to R: notes on R: a programming environment for data analysis and graphics, version 1.9.1. Network Theory Ltd., Bristol, UK, second printing, revised and updated edition, 2004. ISBN 0-9541617-4-2 (paperback). vi + 146 pp.

Venables:2010:BRI

[Ven10] W. N. Venables. Book review: An Introduction to Statistical Inference and Its Applications with R, by Trosset, M. W. Biometrics, 66(2):656–657, June 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verzani:2005:URI

[Ver05] John Verzani. Using R for introductory statistics. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-450-9. xvi + 414 pp. LCCN QA276.4 .V47 2005.

Verzani:2012:GCI

[Ver12]  John Verzani.  gWid-

getsWWW: Creating interactive Web pages within R. Journal of Statistical Software, 49(10):??, June 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v49/
i10.

Verzani:2014:URI

[Ver14] John Verzani. Using R for introductory statistics. Chapman and Hall/CRC the R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2014. ISBN 1-4665-9073-4 (hardcover). xvii + 502 pp. LCCN QA276.4 .V47 2014.

Verde:2018:BRP

[Ver18] Pablo Emilio Verde. bamdit:

An R package for Bayesian meta-analysis of diagnostic test data. Journal of Statistical Software, 86(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v086i10;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v086i10/
v86i10.pdf.

Veytsman:2014:BRD

[Vey14] Boris Veytsman. Book review: Dynamic Documents with R and knitr, by Yihui Xie. TUGboat, 35(1):

115–119, 2014. ISSN 0896-

3207. URL http://www.
tug.org/TUGboat/tb35-1/
tb109reviews-xie.pdf.

Varmaz:2013:RAW

[VFV13] Armin Varmaz, Christian Fieberg, and Andreas Var-

wig.  RMatlab-app2web:

Web deployment of R/ MATLAB applications. Journal of Statistical Software, 54 (5):??, September 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v54/i05.

Varadhan:2009:BRP

[VG09] Ravi Varadhan and Paul Gilbert. BB: An R package for solving a large system of nonlinear equations and for optimizing a highdimensional nonlinear objective function. Journal of Statistical Software, 32(4): ??, October 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v32/i04.

Vidoni:2022:URP

[Vid22] Melina Vidoni. Understanding Roxygen package documentation in R. The Journal of Systems and Software, 188(??):??, June 2022. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0164121222000310.

Viechtbauer:2010:CMA

[Vie10] Wolfgang Viechtbauer. Conducting meta-analyses in R with the metafor package. Journal of Statistical Software, 36(3):??, August 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v36/
i03.

Viechtbauer:2011:BRG

[Vie11] Wolfgang Viechtbauer. Book review: Graphics for statistics and data analysis with R. Kevin J. Keen, Chapman and Hall/CRC, Boca Raton, 2010. No. of pages:

489. Price: $69.95. ISBN: 978-1-58488-087-5. Statistics in Medicine, 30(22):2765– 2766, September 30, 2011. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vinod:2008:HIE

[Vin08] Hrishikesh D. Vinod. Handson intermediate econometrics using R: templates for extending dozens of practical examples. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2008. ISBN 981-281-885-5, 981-4350-41-9. xxvii + 512 pp. LCCN HB141 .V56 2008.

Vinod:2010:ASS

[Vin10] Hrishikesh D. Vinod, editor.

Advances in social science research using R, volume 196 of Lecture notes in statistics - proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-1763-2 (paperback), 1-4419-1764-0 (ebook). LCCN H62.A3 A328

2010. URL http://www.
loc.gov/catdir/enhancements/
fy1011/2009942719-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1011/2009942719-
t.html.

Visscher:2010:BRH

[Vis10] Peter M. Visscher. Book review: Handbook on Analyzing Human Genetic Data: Computational Approaches and Software, edited by Lin, S. and Zhao, H. Biometrics, 66(4):1310, December 2010. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Villegas:2018:SRP

[VKM18] Andrés M. Villegas, Vladimir K. Kaishev, and Pietro Millossovich. StMoMo: An R package for stochastic mortality modeling. Journal of Statistical Software, 84 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v084i03;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v084i03/v84i03.pdf.

Vervloet:2015:PRP

[VKVC15] Marlies Vervloet, Henk A. L. Kiers, Wim Van den Noortgate, and Eva Ceulemans. PCovR: An R package for principal covariates regression. Journal of Statistical Software, 65 (8):??, June 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v65/i08.

Vishwakarma:2021:BRS

[VL21] Srishti Vishwakarma and Vyacheslav Lyubchich. Book review: Spatial Ecology and Conservation Modeling: Applications With R. Technometrics, 62(3):416–417, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Valero-Mora:2009:BRBb

[VM09a] Pedro M. Valero-Mora. Book review: Interactive and Dynamic Graphics for Data Analysis: With R and GGobi. Journal of Statistical Software, 30(BR-7):??, May 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v30/b07.

Valero-Mora:2009:BRBa

[VM09b] Pedro M. Valero-Mora. Book review: Introductory Statistics with R (Second Edition). Journal of Statistical Software, 29(BR-10):

??, January 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v29/b10.

Valero-Mora:2012:GUI

[VML12] Pedro M. Valero-Mora and Ruben Ledesma. Graphical user interfaces for R. Journal of Statistical Software, 49(1):??, June 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v49/i01.

Voelkel:2009:SBR

[Voe09] Joseph G. Voelkel. Short book review: R for SAS and SPSS Users by Robert A. Muenchen. International Statistical Review = Revue Internationale de Statistique, 77(3):465–466, December 2009. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http:

//www.jstor.org/stable/
27919771.

Vanegas:2016:ELS

[VP16] Luis Hernando Vanegas and Gilberto A. Paula. An extension of log-symmetric regression models: R codes and applications. Journal of Statistical Computation and Simulation, 86(9): 1709–1735, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Venables:1994:MAS

[VR94] William N. Venables and Brian D. Ripley. Modern Applied Statistics with S-PLUS. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94350-1 (New York), 3-540-94350-1 (Berlin). xii + 462 pp. LCCN QA276.4 .V46 1994.

Venables:1997:MAS

[VR97] W. N. (William N.) Venables and Brian D. Ripley. Modern applied statistics with S-Plus. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 0-387-98214-0. xv + 548 pp. LCCN QA276.4 .V46 1997. URL

http://www.stats.ox.ac.
uk/pub/MASS2/.

Venables:1999:MAS

[VR99] W. N. (William N.) Venables and Brian D. Ripley. Modern applied statistics with S-PLUS: Volume 1: Data Analysis. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1999. ISBN 0-387-98825-4 (hardback). xi + 501 pp. LCCN QA276.4.V46 1999. URL

http://www.stats.ox.ac.
uk/pub/MASS3/. From the

Web site: “The book is also useful with R, a freelyavailable open-source statistical system ‘not unlike S’. We have tried where possible to use code that works in all versions of S-PLUS and in R.”.

Venables:2000:MAS

[VR00a] W. N. (William N.) Venables and Brian D. Ripley. Modern applied statistics with S-PLUS: Volume 2: S Programming. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98966-8. x + 264 pp. LCCN QA76.73.S15 V46 2000. URL http://

                                        www.stats.ox.ac.uk/pub/
                                        MASS3/.

Venables:2000:P

[VR00b] W. N. (William N.) Venables and Brian D. Ripley. S programming. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98966-8. x + 264 pp. LCCN QA76.73.S15 V46 2000.

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0816/
                                        99057466-d.html;   http:
                                        //www.loc.gov/catdir/enhancements/
                                        fy0816/99057466-t.html.

Venables:2002:MAS

[VR02] W. N. (William N.) Venables and Brian D. Ripley. Modern Applied Statistics with S. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2002. ISBN 0-387-95457-0. xi + 495 pp. LCCN QA276.4 .V46 2002.

Varvia:2023:MRP

[VRP23] Petri Varvia, Janne Räty, and Petteri Packalen. mgpr: an R package for multivariate Gaussian process regression. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002595.

Venables:2002:IRN

[VS02] W. N. (William N.) Venables and D. M. (David M.) Smith. An introduction to R: notes on R: a programming environment for data analysis and graphics, version 1.4.1. Network Theory Ltd., Bristol, UK, 2002. ISBN 0-9541617-4-2. vi + 139 pp. LCCN QA276.4 .V45 2001.

Visser:2010:DRP

[VS10] Ingmar Visser and Maarten Speekenbrink. depmixS4: An R package for hidden Markov models. Journal of Statistical Software, 36 (7):??, August 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v36/i07.

Vignes:2017:BRB

[VSS^{+}17] Matthieu Vignes, Mélodie Sammarro, Anthime Schrefheere, Alex White, and Stephen Marsland. Book review: Bayesian Networks: With Examples in R. Australian & New Zealand Journal of Statistics, 59(4):545–546, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

VandenMeersche:2009:CSX

[VSV09] Karel Van den Meersche, Karline Soetaert, and Dick Van Oevelen. Code snippet: xsample(): An R function for sampling linear inverse problems. Journal of Statistical Software, 30(CS-1):??, April 2009. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v30/c01.

Villacorta:2016:FRP

[VV16] Pablo J. Villacorta and José L. Verdegay. FuzzyStatProb: An R package for the estimation of fuzzy stationary probabilities from a sequence of observations of an unknown Markov chain. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v071i08;   https:/
/www.jstatsoft.org/index.

php/jss/article/view/v071i08/
v71i08.pdf.

Vidoni:2019:RRP

[VV19] Melina Vidoni and Aldo Vecchietti. “rsppfp”: an R package for the shortest path problem with forbidden paths. SoftwareX, 9 (??):265–270, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018300955.

VanAelst:2013:FRB

[VW13] Stefan Van Aelst and Gert Willems. Fast and robust bootstrap for multivariate inference: The R package FRB. Journal of Statistical Software, 53 (3):??, April 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v53/i03.

Vitolo:2016:PFR

[VWDB16] Claudia Vitolo, Peter Wells, Martin Dobias, and Wouter Buytaert. fuse: An R package for ensemble hydrological modelling. Journal of Open Source Software, 1(8): 52:1–52:2, December 2016. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00052.

Visne:2012:SRP

[VYD^{+}12] Ilhami Visne, Ahmet Yildiz, Ekran Dilaveroglu, Klemens Vierlinger, Christa Nöhammer, Friedrich Leisch, and Albert Kriegner. speedR: An R package for interactive data import, filtering and ready-to-use code generation. Journal of Statistical Software, 51(2): ??, October 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i02.

Wang:2006:BRE

[Wan06] Haiyan Wang. Book review: Extending the Linear Model with R: Generalized Linear, Mixed Effects and Nonparametric Regression Models, Edited by Faraway J. J. Biometrics, 62 (4):1278, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2011:SRP

[Wan11] Ji-Ping Wang. SPECIES: An R package for species richness estimation. Journal of Statistical Software, 40(9):??, April 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/i09.

Wang:2013:CRP

[Wan13]  Zhu Wang.  cts: An

R package for continuous time autoregressive models via Kalman filter. Journal of Statistical Software, 53(5):??, April 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v53/i05.

Wang:2016:BRU

[Wan16] Kevin Wang. Book review:

Using R for Introductory Statistics. Australian & New Zealand Journal of Statistics, 58(1):138–139, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2022:BRO

[Wan22] Amber W. Wang. Book review: Omic association studies with R and bioconductor: Gonzalez, Juan R. and Caceres, Alejandro Boca Raton, FL: Chapman and Hall/ CRC, 2020. pp. 390. Biometrics, 78(1):415–416, March 2022. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Warasi:2023:GRP

[War23] Md S. Warasi. groupTesting:

an R package for group testing estimation. Communications in Statistics: Simulation and Computation, 52 (12):6210–6224, 2023. CO-DEN CSSCDB. ISSN 0361-0918.

Wason:2015:ORP

[Was15] James Wason. OptGS: An R package for finding nearoptimal group-sequential designs. Journal of Statistical Software, 66(2):??,

????  2015.  CODEN

JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v066i02;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v066i02/v66i02.pdf.

Wickham:2007:RG

[WC07] Hadley Wickham and Dianne Cook. R graphics. The American Statistician, 61 (1):99–100, February 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wickham:2011:TRP

[WCHB11] Hadley Wickham, Dianne Cook, Heike Hofmann, and Andreas Buja. tourr: An R package for exploring multivariate data with projections. Journal of Statistical Software, 40(2): ??, April 2011. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v40/i02.

Wilkinson:2018:PPR

[WD18] Shaun P. Wilkinson and Simon K. Davy. phylogram: an R package for phylogenetic analysis with nested lists. Journal of Open Source Software, 3(26): 790:1–790:2, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00790.

Welvaert:2011:NRP

[WDM^{+}11] Marijke Welvaert, Joke Durnez, Beatrijs Moerkerke, Geert Berdoolaege, and Yves Rosseel. neuRosim: An R package for generating fMRI data. Journal of Statistical Software, 44(10): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v44/i10.

Wallace:2012:CGB

[WDT^{+}12] Byron C. Wallace, Issa J.

Dahabreh, Thomas A. Trikalinos, Joseph Lau, Paul Trow, and Christopher H. Schmid. Closing the gap between methodologists and endusers: R as a computational back-end. Journal of Statistical Software, 49 (5):??, June 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v49/i05.

Weeks:2010:PRP

[Wee10] Jonathan P. Weeks. plink:

An R package for linking mixed-format tests using IRT-based methods. Journal of Statistical Software, 35(12):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/i12.

Weisberg:2002:DRR

[Wei02] Sanford Weisberg. Dimension reduction regression in R. Journal of Statistical Software, 7(1):1–22, 2002. CODEN JSSOBK. ISSN

????  URL http://www.
jstatsoft.org/v07/i01;
 http://www.jstatsoft.
org/v07/i01/dr_1.0.tar.
gz;   http://www.jstatsoft.
org/v07/i01/drdoc.pdf;
 http://www.jstatsoft.
org/v07/i01/updates.

Wei:2012:CSP

[Wei12] Xin Wei. Code snippet: %PROC R: a SAS macro that enables native R programming in the base SAS environment. Journal of Statistical Software, 46(CS-2):??, January 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v46/c02.

Weisberg:2014:ALR

[Wei14] Sanford Weisberg. Applied linear regression. Wiley series in probability and statistics. Wiley, New York, NY, USA, fourth edition, 2014. ISBN 1-118-38608-6 (hardcover). xvii + 340 pp. LCCN QA278.2 .W44 2014.

Weller:2018:SRP

[Wel18] Zachary D. Weller. spTest: An R package implementing nonparametric tests of isotropy. Journal of Statistical Software, 83(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.

org/index.php/jss/article/
view/v083i04;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v083i04/
v83i04.pdf.

Wehrens:2012:MSV

[WF12] Ron Wehrens and Pietro Franceschi. Meta-statistics for variable selection: The R package BioMark. Journal of Statistical Software, 51(10): ??, November 2012. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v51/i10.

Wickham:2010:BRB

[WG10] Hadley Wickham and Garrett Grolemund. Book review: Graphics for Statistics and Data Analysis with R. Journal of Statistical Software, 36(BR-3): ??, September 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v36/b03.

Wickham:2017:RDS

[WG17] Hadley Wickham and Garrett Grolemund. R for data science: import, tidy, transform, visualize, and model data. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-1039-9 (paperback), 1-4919-1033-X (e-book), 1-4919-1034-8 (e-book), 1-4919-1036-4 (e-book). LCCN QA76 .W53 2017. URL http:/

                                        /ebookcentral.proquest.
                                        com/lib/imperial/detail.
                                        action?docID=4770093.

Woodward:2017:ATS

[WGE17] Wayne A. Woodward, Henry L. Gray, and Alan C. Elliott. Applied Time Series Analysis, with R. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, second edition, 2017. ISBN 1-4987-3422-7 (hardcover), 1-4987-3431-6 (ePub e-book), 1-4987-3427-8 (PDF e-book), 1-4987-3429-4 (VitalBook ebook). xv + 618 pp. LCCN QA280 .W68 2017.

Walker:2012:MTD

[WGSL12] Steven C. Walker, Guillaume Guénard, Péter Sólymos, and Pierre Legendre. Multipletable data in R with the multitable package. Journal of Statistical Software, 51 (8):??, November 2012. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v51/i08.

Wang:2021:GRP

[WHK21] Xiao Wang, Dan Hao, and Haja N. Kadarmideen. GeneDMRs: an R package for gene-based differentially methylated regions analysis. Journal of Computational Biology, 28(3):304– 316, March 2021. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.
liebertpub.com/doi/abs/
10.1089/cmb.2020.0081;
https://www.liebertpub.
com/doi/pdf/10.1089/cmb.
2020.0081.

Wickham:2008:BRBb

[Wic08a] Hadley Wickham. Book review: A First Course in Statistical Programming with R. Journal of Statistical Software, 28(BR-3): ??, December 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v28/b03.

Wickham:2008:BRBa

[Wic08b] Hadley Wickham. Book review: Data Manipulation with R. Journal of Statistical Software, 26(BR-1): 1–2, June 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v26/b01.

Wickham:2009:GEG

[Wic09] Hadley Wickham. ggplot2:

Elegant Graphics for Data Analysis. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-98140-3 (softcover), 0-387-98141-1 (e-book). viii + 212 pp. LCCN QA90 .W53 2009.

Wickham:2011:MOR

[Wic11a] Hadley Wickham. Mutable objects in R. Computational Statistics, 26(3):405–418, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-011-0235-7.

Wickham:2011:G

[Wic11b] Hadley Wickham. ggplot2. WIREs Computational Statistics, 3(2):180–185, March/ April 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wickham:2019:AR

[Wic19] Hadley Wickham. Advanced R. The R series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2019. ISBN 0-367-25537-5 (hardcover), 0-8153-8457-2 (paperback), 1-351-20128-X (Mobipocket), 1-351-20129-8 (EPUB), 1-351-20130-1 (PDF), 1-351-20131-X (e-book), 1-4665-8697-4 (e-book). xv + 587 pp. LCCN QA276.45.R3 W5 2019.

Wiens:2004:MSP

[Wie04] Douglas P. Wiens. MATCH — a software package for robust profile matching using S-Plus. Journal of Statistical Software, 10(1): 1–35, ???? 2004. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.

php?id=91&url=v10/i01/
manual2.pdf&ct=1.

Wilkinson:2011:BRG

[Wil11a] Leland Wilkinson. Book review: ggplot2: Elegant Graphics for Data Analysis, by Wickham, H. Biometrics, 67(2):678–679, June 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:2011:DMR

[Wil11b] Graham J. Williams. Data mining with Rattle and R: the art of excavating data for knowledge discovery. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-9889-6 (paperback), 1-4419-9890-X (e-bib). xx + 374 pp. LCCN QA76.9.D343 W556 2011.

Wilson:2012:SCR

[Wil12] Jason Wilson. Statistical computing with R: selecting the right tool for the job — R Commander or something else? WIREs Computational Statistics, 4(6):518– 526, November/December 2012. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Wiley:2014:BRG

[Wil14a] Joshua F. Wiley. Book review: Growth Curve Analysis and Visualization Using R. Journal of Statistical Software, 58(BR-2):??, July 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v58/b02.

Willekens:2014:MAL

[Wil14b] Frans Willekens. Multistate analysis of life histories with R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-08383-X (e-book), 3-319-08382-1 (print). LCCN QA276.45.R3.

Wiley:2018:PRR

[Wil18] David Wiley. RISE: An R package for RISE analysis. Journal of Open Source Software, 3(28):846:1, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00846.

Wang:2018:BRP

[WLH^{+}18] Chenguang Wang, Thomas A. Louis, Nicholas C. Henderson, Carlos O. Weiss, and Ravi Varadhan. beanz: An R package for Bayesian analysis of heterogeneous treatment effects with a graphical user interface. Journal of Statistical Software, 85(??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v085i07;
 https://www.jstatsoft.

org/index.php/jss/article/
view/v085i07/v85i07.pdf.

Wallstrom:2008:IBA

[WLK08] Garrick Wallstrom, Jeffrey Liebner, and Robert E. Kass. An implementation of Bayesian adaptive regression splines (BARS) in C with S and R wrappers. Journal of Statistical Software, 26 (1):1–21, June 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v26/i01.

White:2014:PRP

[WM14] Arthur White and Thomas Brendan Murphy. BayesLCA: An R package for Bayesian latent class analysis. Journal of Statistical Software, 61(13):??, November 2014. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v61/
i13.

Wyszynski:2018:SSM

[WM18] Karol Wyszynski and Giampiero Marra. Sample selection models for count data in R. Computational Statistics, 33(3):1385–1412, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-017-0762-y;
http://link.springer.
com/content/pdf/10.1007/
s00180-017-0762-y.pdf.

Weihs:2014:FSA

[WML14] Claus Weihs, Olaf Mersmann, and Uwe Ligges. Foundations of statistical algorithms: with reference to R packages. Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2014. ISBN 1-4398-7885-4 (hardcover). xxv + 473 pp. LCCN QA276.4 .W45 2014.

Wojtys:2016:CRS

[WMR16] Magorzata Wojtyś, Giampiero Marra, and Rosalba Radice. Copula regression spline sample selection models: The R package

SemiParSampleSel. Jour-

nal of Statistical Software, 71(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v071i06;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v071i06/v71i06.pdf.

Wallace:2017:DTR

[WMS17] Michael P. Wallace, Erica E. M. Moodie, and David A. Stephens. Dynamic treatment regimen estimation via regression-based techniques: Introducing R package DTRreg. Journal of Statistical Software, 80(??): ??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v080i02;
https://www.jstatsoft.
org/index.php/jss/article/
view/v080i02/v80i02.pdf.

Wolfe:1994:GOR

[Wol94] M. A. Wolfe. On global optimization in R using interval arithmetic. In Bianco [Bia94], pages 61–76. CO-DEN OMSOE2. ISSN 1055-6788.

Wolfe:1999:DMP

[Wol99]  Michael A. Wolfe.  On

discrete minimax problems in R using interval arithmetic. Reliable Computing = Nadezhnye vychisleniia, 5(4):371–383, November 1999. CODEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A%3A1009930013647/
;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=1385-3139&
volume=5&issue=4&spage=
371;   http://www.springerlink.
com/openurl.asp?genre=
article&issn=1385-3139&
volume=5&issue=4&spage=
371-383.

Wooff:1998:BRB

[Woo98] David Wooff. Book review:

Statistical Tools for Nonlinear Regression: A Practical Guide with S-PLUS Examples, by S. Huet; A. Bouvier; M.-A. Gruet; E. Jolivet. Journal of the Royal Statistical Society. Series D (The Statistician), 47(3):545, ???? 1998. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
2988640.

Wood:2001:BRB

[Woo01] Simon Wood. Book review: Mixed-Effects Models in S and S-Plus, by J. C. Pinheiro; D. M. Bates. Journal of the Royal Statistical Society. Series D (The Statistician), 50(3):351–352, ???? 2001. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/
2680951.

Woodward:2005:BBC

[Woo05] Phil Woodward. BugsXLA: Bayes for the Common Man. Journal of Statistical Software, 14(5):1–18, January 31, 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/counter.
php?id=130&url=v14/i05&
ct=2;   http://www.jstatsoft.
org/counter.php?id=130&
url=v14/i05/v14i05.pdf&
ct=1.

Wood:2006:GAM

[Woo06] Simon N. Wood. Generalized Additive Models: an Introduction with R. Texts in statistical science. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-474-6. xvii + 391 pp. LCCN QA274.73 .W66 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0702/
2006040209-d.html.

Wood:2011:BRE

[Woo11] J. T. Wood. Book review:

Environmental and Ecological Statistics with R, by Song S. Qian. Australian & New Zealand Journal of Statistics, 53(1):127–128, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2015:VAR

[WPW15] Haichuan Wang, David Padua, and Peng Wu. Vectorization of Apply to reduce interpretation overhead of R. ACM SIGPLAN Notices, 50(10):400–415, October 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wheeler:2023:NIR

[WRD23] A. Powell Wheeler, Kyle T.

Rachels, and Kevin J. Dockendorf. NCIFD — an internal R package for a fisheries agency. Fisheries, 48 (10):411–417, October 2023. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Wijayasari:2023:MRB

[WRV23] Winda Wijayasari, Faizal Immaddudin Wira Rohmat, and Sparisoma Viridi. Malaccha: an R-based end-to-end Markov transition matrix extraction for land cover datasets. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023000110.

Whitcher:2011:QAD

[WS11] Brandon Whitcher and Volker J. Schmid. Quantitative analysis of dynamic contrast-enhanced and diffusion-weighted magnetic resonance imaging for oncology in R. Journal of Statistical Software, 44(5): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v44/i05.

Whitcher:2011:WDN

[WST11] Brandon Whitcher, Volker J. Schmid, and Andrew Thorton. Working with the DI-COM and NIfTI data standards in R. Journal of Statistical Software, 44(6): ??, October 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v44/i06.

Wu:2018:NML

[WsT18] Colin O. Wu and Xin (Mathematical statistician) Tian.

Nonparametric Models for Longitudinal Data: with Implementation in R, volume 159 of Monographs on Statistics and Applied Probability. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2018. ISBN 1-4665-1600-3 (hardcover), 0-429-93908-6 (pdf), 1-4665-1601-1 (e-book). xxix + 551 pp. LCCN QH323.5 .W82 2018.

Wickelmaier:2012:PPC

[WSZ12] Florian Wickelmaier, Carolin Strobl, and Achim Zeileis. Psychoco: Psychometric computing in R. Journal of Statistical Software, 48(1):??, May 2012. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v48/
i01.

Welty:2015:MAR

[WTB^{+}15] Ethan Z. Welty, Christian E. Torgersen, Samuel J. Brenkman, Jeffrey J. Duda, and Jonathan B. Armstrong. Multiscale analysis of river networks using the R package linbin. North American Journal of Fisheries Management, 35(4):802–809, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wu:2024:EPI

[Wu24] Haoran Wu. ecode: an R package to investigate community dynamics in ordinary differential equation systems. Ecological Modelling, 491(??):??, May 2024. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0304380024000644.

Wunschiers:2013:CBP

[Wün13] Röbbe Wünschiers. Computational biology: a practical introduction to biodata processing and analysis with Linux, MySQL, and R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2013. ISBN 3-642-34748-7, 3-642-34749-5 (e-book). xxix + 449 pp. LCCN QH324.2 .W86 2013.

Wang:2011:DEM

[WW11] Xiao-Feng Wang and Bin Wang. Deconvolution estimation in measurement error models: The R package decon. Journal of Statistical Software, 39(10): ??, March 2011. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v39/i10.

Wadhwa:2018:PTR

[WWDS18] Raoul R. Wadhwa, Drew F. K. Williamson, Andrew Dhawan, and Jacob G. Scott. TDAstats: R pipeline for computing persistent homology in topological data analysis. Journal of Open Source Software, 3(28):860:1–860:3, August 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00860.

Wikle:2019:STS

[WZMC19] Christopher K. Wikle, Andrew Zammit-Mangion, and Noel A. C. Cressie. Spatiotemporal Statistics with R. Chapman and Hall/CRC the R series. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2019. ISBN 1-138-71113-6 (hardcover), 0-429-64978-9 (PDF), 1-351-76972-3 (e-book). v + 380 pp. LCCN QA278.2 .W55 2019.

Xie:2019:RMD

[XAG19] Yihui Xie, J. J. Allaire, and Garrett Grolemund. R Markdown: the definitive guide. Taylor and Francis, CRC Press, Boca Raton, FL, USA, 2019. ISBN 1-138-35933-5. 303 pp. LCCN TK5105.8885.M39 X393 2018.

Xie:2021:RMC

[XDR21] Yihui Xie, Christophe Dervieux, and Emily Riederer. R Markdown Cookbook. The R series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2021. ISBN 0-367-56382-7 (hardcover), 0-367-56383-5 (paperback), 1-000-29080-8 (e-book), 1-000-29084-0 (Mobipocket ebook), 1-000-29088-3 (ebook), 1-003-09747-2 (ebook). xxix + 329 pp. LCCN QA276.45.R3 X54 2021.

Xie:2013:ARP

[Xie13a] Yihui Xie. animation: An R package for creating animations and demonstrating statistical methods. Journal of Statistical Software, 53(1):??, April 2013. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v53/i01.

Xie:2013:DDR

[Xie13b] Yihui Xie. Dynamic documents with R and knitr. Chapman and Hall/CRC the R series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2013. ISBN 1-4822-0353-7 (paperback). xxv + 190 pp. LCCN QA276.4 .X54 2014. URL

                                        http://images.tandf.co.
                                        uk/common/jackets/websmall/
                                        978148220/9781482203530.
                                        jpg.

Xie:2022:RRP

[Xie22] Fangzhou Xie. rethnicity: an R package for predicting ethnicity from names. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021001874.

Xu:2019:RPM

[XLX^{+}19] Cong Xu, Zheng Li, Yuan Xue, Lijun Zhang, and Ming Wang. An R package for model fitting, model selection and the simulation for longitudinal data with dropout missingness. Communications in Statistics: Simulation and Computation, 48(9):2812–2829, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xia:2010:CSP

[XMW10] Xiao-Qin Xia, Michael Mc-Clelland, and Yipeng Wang. Code snippet: PypeR, a Python package for using R in Python. Journal of Statistical Software, 35(CS-2):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/c02.

Xiao:2015:RPT

[XWHL15] Tao Xiao, G. A. Whitmore, Xin He, and Mei-Ling Ting Lee. The R package threg to implement threshold regression models. Journal of Statistical Software, 66 (8):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v066i08;
 http://www.jstatsoft.
org/index.php/jss/article/
view/v066i08/v66i08.pdf.

Xu:2022:OPO

[XYC22] Hongyan Xu, Ayten Yigiter, and Jie Chen. onlineBcp: an R package for online change point detection using a Bayesian approach. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000140.

Yalta:2010:BRB

[Yal10]  A. Talha Yalta.  Book

review: Hands-On Intermediate Econometrics Using R. Journal of Statistical Software, 33(BR-1): ??, February 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v33/b01.

Yang:1995:RMV

[Yan95] Ying Yang. Random-effects model validation: a program package using a FOR-TRAN program (REML) and S-plus. Thesis (M.S.), Graduate School of Public Health, University of Pittsburgh, Pittsburgh, PA, USA, 1995. v + 126 pp.

Yanagida:2017:SRP

[Yan17] Takuya Yanagida. seqtest: an R package for sequential triangular tests. Journal of Statistical Theory and Practice, 11(3):402–406, September 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.
springer.com/article/10.
1080/15598608.2016.1268549.

Yang:2020:ICB

[YCLL20] Chao-Tung Yang, Yu-Wei Chan, Jung-Chun Liu, and Ben-Shen Lou. An implementation of cloud-based platform with R packages for spatiotemporal analysis of air pollution. The Journal of Supercomputing, 76 (3):1416–1437, March 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yassouridis:2018:GCE

[YEL18] Christina Yassouridis, Dominik Ernst, and Friedrich Leisch. Generalization, combination and extension of functional clustering algorithms: The R package funcy. Journal of Statistical Software, 85(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL

https://www.jstatsoft.
org/index.php/jss/article/
view/v085i09;   https:/
/www.jstatsoft.org/index.
php/jss/article/view/v085i09/
v85i09.pdf.

Yee:2014:RCI

[YH14] Thomas W. Yee and Alfian F. Hadi. Rowcolumn interaction models, with an R implementation. Computational Statistics, 29(6):1427–1445, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-014-0499-9.

Yu:2017:PMR

[YL17] Qingzhao Yu and Bin Li. mma: An R package for mediation analysis with multiple mediators. Journal of Open Research Software, 5 (1):11–??, April 12, 2017. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.160/.

Young:2010:TRP

[You10] Derek S. Young. tolerance: An R package for estimating tolerance intervals. Journal of Statistical Software, 36 (5):??, August 2010. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v36/i05.

Young:2012:OED

[You12]  Derek S. Young.  Op-

timal experimental design with R. Journal of Applied Statistics, 39(8):1848– 1849, August 2012. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Yerlikaya-Ozkurt:2023:CPP

[YÖYB23] Fatma Yerlikaya-Özkurt, Ceyda Yazici, and Inci Batmaz. cmaRs: a powerful predictive data mining package in R. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002492.

Yang:2013:RPS

[YS13] Dan Yang and Dylan S.

Small. An R package and a study of methods for computing empirical likelihood. Journal of Statistical Computation and Simulation, 83(7):1363–1372, 2013. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Yu:2012:BRB

[Yu12]  Jihnhee Yu.  Book re-

view: An Introduction to Applied Multivariate Analysis With R by B. Everitt; T. Hothorn. Technometrics, 54(4):445, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
41714937.

Yuan:2007:MLM

[Yua07] Lester L. Yuan. Maximum likelihood method for predicting environmental conditions from assemblage composition: The R package bio.infer. Journal of Statistical Software, 22(3):1– 20, September 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v22/i03.

Yi:2002:IHO

[YWL02] Grace Y. Yi, Jianrong Wu, and Ying Liu. Implementation of higher-order asymptotics to S-plus. Computational Statistics & Data Analysis, 40(4):775–800, October 28, 2002. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167947302000968.

Zaghdoudi:2018:PIR

[Zag18] Taha Zaghdoudi. ivporbit: An R package to estimate the instrumental variables probit model. Journal of Open Source Software, 3 (22):523:1, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00523.

Zeileis:2010:EMF

[ZC10] Achim Zeileis and Yves Croissant. Extended model formulas in R: Multiple parts and multiple responses. Journal of Statistical Software, 34(1):??, April 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

                                        /www.jstatsoft.org/v34/
                                        i01.

Zeitler:2016:BRA

[Zei16] David W. Zeitler. Book review: Applied Multivariate Statistics with R, Daniel Zelterman, Springer-Verlag, Switzerland, 2015. ISBN 978-3-319-14092-6. 393 pp. USD 99.00 (Hardcover), USD 69.99 (eBook).

http://www.springer.com/9783319140926.

Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://

www.jstatsoft.org/index.
php/jss/article/view/v071b02;
https://www.jstatsoft.
org/index.php/jss/article/
view/v071b02/v71b02.pdf.

Zeugner:2015:BMA

[ZF15] Stefan Zeugner and Martin Feldkircher. Bayesian model averaging employing fixed and flexible priors: The BMS package for R. Journal of Statistical Software, 68(4):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068i04;
http://www.jstatsoft.
org/index.php/jss/article/
view/v068i04/v68i04.pdf.

Zang:2010:SAC

[ZFZ10] Yong Zang, Wing Kam Fung, and Gang Zheng. Simple algorithms to calculate asymptotic null distributions of robust tests in case-control genetic association studies in R. Journal of Statistical Software, 33 (8):??, February 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v33/i08.

Zagalsky:2018:HRC

[ZGS^{+}18] Alexey Zagalsky, Daniel M. German, Margaret-Anne Storey, Carlos Gómez Teshima, and Germán Poo-Caamaño. How the R community creates and curates knowledge: an extended study of stack overflow and mailing lists. Empirical Software Engineering, 23(2):953–986, April 2018. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-017-9536-y.

Zhu:2019:IRL

[ZGZvB19] Bing Zhu, Zihan Gao, Junkai Zhao, and Seppe K. L. M. vanden Broucke. IRIC: an R library for binary imbalanced classification. SoftwareX, 10 (??):Article 100341, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/S2352711019301700.

Zhang:2011:BRAb

[Zha11] Kui Zhang. Book review: Applied Statistical Genetics with R: For Populationbased Association Studies, by Foulkes, A. S. Biometrics, 67(4):1678–1679, December 2011. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zieffler:2011:CGR

[ZHL11] Andrew Zieffler, Jeffrey Harring, and Jeffrey D. Long. Comparing groups: randomization and bootstrap methods using R. Wiley, New York, NY, USA, 2011. ISBN 0-470-62169-9 (hardcover). xxxii + 298 pp. LCCN QA276.8 .Z53 2011.

Ziegel:1998:BRBd

[Zie98] Eric R. Ziegel. Book review:

The Basics of S and S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 40 (1):82–83, February 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1271413.

Ziegel:1999:BRBr

[Zie99] Eric R. Ziegel. Book review: Environmental Stats for S-Plus by Steven P. Millard. Technometrics, 41 (2):179, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1270752.

Ziegel:2000:BRBi

[Zie00] Eric R. Ziegel. Book review: Modern Applied Biostatistical Methods Using S-Plus by Steve Selvin. Technometrics, 42(2):221, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1271484.

Ziegel:2001:BRBe

[Zie01a] Eric R. Ziegel. Book review: Mixed-Effects Models in S and S-PLUS by José C. Pinheiro; Douglas M. Bates. Technometrics, 43(1):113, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1270889.

Ziegel:2001:BRBad

[Zie01b]  Eric R. Ziegel.  Book

review: Modern Applied Statistics with S-PLUS by W. N. Venables; B. D. Ripley. Technometrics, 43 (2):249, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        1271074.

Ziegel:2001:BRBac

[Zie01c] Eric R. Ziegel. Book review: The Basics of S and S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 43(2):249, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1271073.

Ziegel:2002:BRBag

[Zie02a] Eric R. Ziegel. Book review:

A Handbook of Statistical Analyses Using S-PLUS by Brian S. Everitt. Technometrics, 44(3):300, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1270517.

Ziegel:2002:BRBx

[Zie02b] Eric R. Ziegel. Book review: Analyzing Medical Data Using S-PLUS by B. S. Everitt; Sophia Rabe-Hesketh. Technometrics, 44(2):201–202, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
1271291.

Ziegel:2004:BRBg

[Zie04] Eric R. Ziegel. Book review:

An R and S-PLUS Companion to Applied Regression by John Fox. Technometrics, 46(3):371, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25470872.

Ziegel:2005:BRBb

[Zie05] Eric R. Ziegel. Book review: Statistical Analysis of Financial Data in S-PLUS by René A. Carmona. Technometrics, 47(2): 242–243, May 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

                                        //www.jstor.org/stable/
                                        25471002.

Zuur:2009:BGR

[ZIM09] Alain F. Zuur, Elena N. Ieno, and Erik H. W. G. Meesters. A Beginner’s Guide to R. Use R! Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-93836-2 (softcover), 0-387-93837-0 (e-book). xv + 218 pp. LCCN QA276.45.R3 Z88 2009.

Zivot:2005:MFT

[Ziv05] Eric Zivot. Modeling financial time series with S-Plus. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2005. ISBN 0-387-27965-2. ???? pp.

Zuur:2009:MEM

[ZIW^{+}09] Alain F. Zuur, Elena N. Ieno, Neil J. Walker, Anatoly A. Saveliev, and Graham M. Smith. Mixed Effects Models and Extensions in Ecology with R. Statistics for biology and health.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-87457-7 (hardcover). xxii + 574 pp. LCCN QH541.15.S72 M59 2009.

Zeileis:2008:RMC

[ZKJ08] Achim Zeileis, Christian Kleiber, and Simon Jackman. Regression models for count data in R. Journal of Statistical Software, 27(8):??, July 2008. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v27/i08.

Zeileis:2002:SRP

[ZLHK02] Achim Zeileis, Friedrich Leisch, Kurt Hornik, and Christian Kleiber. strucchange: an R package for testing for structural change in linear regression models. Journal of Statistical Software, 7(2):1–38, 2002. CODEN JSSOBK. ISSN

????  URL http://www.
jstatsoft.org/v07/i02;
http://www.jstatsoft.
org/v07/i02/strucchange.
pdf;   http://www.jstatsoft.
org/v07/i02/strucchange_
0.9-5.tar.gz;   http://
www.jstatsoft.org/v07/
i02/updates.

Zucchini:2009:HMM

[ZM09] W. Zucchini and Iain L.

MacDonald. Hidden Markov models for time series: an introduction using R, volume 110 of Monographs on statistics and applied probability. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-58488-573-4 (hardcover). xxii + 275 pp. LCCN QA280 .Z83 2009.

Zumel:2014:PDS

[ZM14] Nina Zumel and John Mount. Practical data science with R. Manning Publications, Greenwich, CT, USA, 2014. ISBN 1-61729-156-0 (paperback). xxv + 389 pp. LCCN QA276.45.R3 Z86 2014.

Zucchini:2016:HMM

[ZML16] Walter Zucchini, Iain L. MacDonald, and Roland Langrock. Hidden Markov Models for Time Series: an Introduction Using R, volume 150 of Monographs on statistics and applied probability. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2016. ISBN 1-4822-5383-6 (hardcover), 1-4822-5384-4 (e-book). xxviii + 370 pp. LCCN QA280 .Z83 2016. URL http:

//marc.crcnetbase.com/
isbn/9781482253849;   http:
//www.taylorfrancis.com/
books/9781482253849.

Zuccolotto:2021:AOB

[ZMS21] Paola Zuccolotto, Marica Manisera, and Marco Sandri. Alley-oop! Basketball analytics in R. Significance (Oxford, England), 18 (2):26–31, April 2021. CO-DEN ???? ISSN 1740-9705 (print), 1740-9713 (electronic).

Zheng:2018:FPS

[ZMV^{+}18] Da Zheng, Disa Mhembere, Joshua T. Vogelstein, Carey E. Priebe, and Randal Burns. FlashR: parallelize and scale R for machine learning using SSDs. ACM SIGPLAN Notices, 53 (1):183–194, January 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:2013:ERP

[ZP13] Weicheng Zhu and Changsoon Park. edcc: An R package for the economic design of the control chart. Journal of Statistical Software, 52 (9):??, February 2013. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v52/i09.

Zhang:2016:RPR

[ZPC^{+}16] Zhengzheng Zhang, Richard M. A. Parker, Christopher M. J. Charlton, George Leckie, and William J. Browne. R2MLwiN: A package to run MLwiN from within R. Journal of Statistical Software, 72(??):??, ????

2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v072i10;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v072i10/
                                        v72i10.pdf.

Zhang:2016:SRP

[ZQS16] Peng Zhang, Zhenguo Qiu, and Chengchun Shi. simplexreg: An R package for regression analysis of proportional data using the simplex distribution. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL

                                        https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v071i11;   https:/
                                        /www.jstatsoft.org/index.
                                        php/jss/article/view/v071i11/
                                        v71i11.pdf.

Zurbenko:2017:AKZ

[ZS17] Igor G. Zurbenko and Mingzeng Sun. Applying Kolmogorov–Zurbenko adaptive R-software. International Journal of Statistics and Probability, 6(5):110–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL

                                        http://www.ccsenet.org/
                                        journal/index.php/ijsp/
                                        article/view/69632.

Zeileis:2014:FGL

[ZUL14] Achim Zeileis, Nikolaus Umlauf, and Friedrich Leisch.

Flexible generation of elearning exams in R: Moodle quizzes, OLAT assessments, and beyond. Journal of Statistical Software, 58(1):??, June 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v58/i01.

Zivot:2003:MFT

[ZW03] Eric Zivot and Jiahui Wang.

Modeling financial time series with S-Plus. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95549-6 (paperback), 0-387-91624-5 (paperback). xix + 632 pp. LCCN HG106 .Z584 2003.

Zhang:2019:RPA

[ZY19] Qihuang Zhang and Grace Y.

Yi. R package for analysis of data with mixed measurement error and misclassification in covariates: augSIMEX. Journal of Statistical Computation and Simulation, 89(12):2293–2315, 2019. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-